A Complete Bibliography of Publications in The Computer Journal: 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

11 April 2024 Version 2.07

Title word cross-reference

2 [HD72]. B_{k} [Roh75]. χ^{2} [PFW74]. F (X, Y) = 0 [Sut76a]. k [Lin72b, Bay76]. K_{n}(x + iy) [Gar78]. L_{1} [OW71, SKY73]. L_{p} [FGH71, BRH70]. LR [GZ72]. N [Bel70b, Far78, Hun72, Man71b, Ull77, Var79, Har76, Woo77a]. n^{th} [Cam73, Mac73]. S [Jor73]. s^{2} [PFW74]. x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n [Sal72, Sal73b]. Y_{n}(x + iy) [Gar78].

-clusters [Lin72b]. -Formula [Bay76]. -functions [Jor73]. -gram [Ull77]. -Parallel [Woo77a]. -Space [Har76].

1 [Bar75a, Boo78, Com75a, DO72, HO78, Mac70b, Mac71b, Mer70b, Obr78b, Pit78, Ree79c, RT78]. 105 [Ano79o]. 1909’ [WR72]. 1969 [Alb71a, Web70]. 1970 [Sla71b, TCB71, WY71, YFM71]. 1971 [Gal71b, Wre71b]. 1972 [Bro74d, BS73b, Cha73b, GS73, HR73, RL73, Ull73b, WS73]. 1973 [Ano73k, IL74, NLW74, RL73]. 1974 [NLW74]. 1976 [Ano75f]. 1977 [Bra78b]. 1978 [Cor79]. 1979 [Kit74].

2 [Boo79c, Ove67, PHJ67, WHM71]. 2.1 [Ano72i]. 2.95 [Eva78]. 25 [Wil78b]. 26 [Hil69, Woo70g].

3 [Boo79d, Ran79]. 370 [Don78]. 3D [Ric71]. 3rd [Ano75f].

4 [Bea74]. 42 [Sal70b]. 43 [Woo70f]. 44 [Fie70, Red71b]. 46 [KL71, Lil71]. 47 [Mux70, Sal70c, vR70c]. 48 [GE71].

50 [BB71].

60 [BBG^{+}76, DHW76, DHW78]. 61 [Kro71]. 66 [McC72]. 67 [SM71]. 68 [Bae70, HH79a, Lin72a, ON79, Pag79a, Tan78, Val74]. 68-Rt [ON79]. 69 [Lav70].

7 [PHJ67]. 71 [Dew72a]. 74 [Hos74]. 77 [Ano76l].

Abadie [TCB71]. Abate [Dur74]. absolute [Dun73, HW74, Par74, Van73]. abstraction [Bur79]. abuse [Cha78a]. Academic [Fox74]. Accelerate [Pro74b]. Acceptance [Sha72a]. Access [Gun70, HS71, WW73, BT71, BBG^{+}71, Har70, MT70, Pez76]. according [Boo67b, Hil69, Woo70g]. accounting [Sym71]. Accuracy [Hys72]. Acm [Ano73a]. Acquisition [Var76b, Pin77]. Actual [HC78]. adaptive [DP76, Haz78, Oli72]. Addresses [Ano74e]. Addressing [CEM72]. ADI [Had71, Lew71b]. Adjustment [O’R70, Whi72b]. Administration [NT79]. Advanced [Ano76a, Ano76b, Lew71a, Eme78]. Advances [Ano78a]. advantages [Sal72]. advice [BM79]. Ai [Clo79, Fis79b]. aid [Bow73]. Aided [Wat74b, FR79]. aired [Gil73]. Albert [Gea70]. ALEC [NF76]. algebra [BBF70, BBH70, KU74, UK74, ZD71]. Algebraic [Jel79, Nas75, SG78, TD77, BF72, Fit73, Hit78, LM77a, LS78]. ALGOL [SS72b, BBG^{+}76, HH79a, Jon72, SAHS71, Bae70, CHH73, DHW76, DHW78, HP71, Hil71, Lin72a, ON79, Pag79a, Tan78, Val74, Wic72, Bro74d]. ALGOL-60 [BBG^{+}76]. Algol60 [Wil78c]. Algorithm [AHM74, Ano75a, Bel70a, BE73, Ber76, Boo67b, Boo67a, Boo79b, Boo79a, Bre71, Bro79a, Bro69, BW75b, BF78, Chr73a, Chr71a, CW70, Cra72, Day76, Day77, DA77, Del77, Dew71, EA70, Eva71, ES77, Far78, FB75, FG77, For75, Gar78, GZ72, Haz78, HP67, Hil75, Hil77, Hil79c, Hil79b, HP79, Hos70, HKA72, HK72, HM77, HGMM75, Jen71, Jon72, Kro70a, Kro70b, Kro70c, Kro70d, Kro70e, Lig73, Lil70, Lit70, Man71b, McC71a, McC71b, McC78, McL78, Obr78a, OP71, Øst79, PP79, Pik79, PFW74, PD79, Ree79b, Ree79a, Roh76, Roh73, Roh74, RL75, Sal70a, Sal71a, Smi70a, SG78, Spä69, Sut76a, Sut76b, TI75, Van73, WT79, Woo70e, Woo70d, Woo70c, Woo71c, Woo71b, Zak77, Zam70, Zam74, vR70a, vR70b, Akk73, AJ70, Alw71, Aus76, Aus71, BJS74]. algorithm [Bar78a, BE79, Bra75, Cox71, Def77, Deu69, Eva72, Eva75, EH76, GM72, Gol78a, Goo75, HD71, HL76a, HR70, Hit78, Jon70, LS78, LMS73, MR76a, MR76b, Mul75, Mux70, OW71, Ric71, Sal70c, Sal74, Sib73, Wil76b, dCS77, vR70c, vR71, Lee74, Mar74d, Ano79o, BB71, Boo78, Boo79d, Boo79c, Dew72a, Dew74, Fie70, GE71, HW74, HO78, Hos74, KL71, Lil71, McC72, Mux70, Obr78b, Ove67, Par74, PHJ67, Pit78, Ran79, Red71b, RT78, SM71, Sal70c, Sal70b, Woo70f, vR70c]. Algorithmic [BBG^{+}76]. Algorithms [Ano70a, Ano70b, Ano70c, Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, Ano73b, Ano73c, Ano74a, Ano74b, Ano75b, Ano75c, Ano76e, Ano76c, Ano76d, Ano77d, Ano77a, Ano77b, Ano77c, Ano78b, Ano78c, Ano78d, Ano78g, Ano79e, Ano79a, Ano79b, Ano79c, Ano79d, CO76, Day78a, GH75, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, HS71, Jar70, Mit76, NR79, RM78, Roy78, SNH76, Sam76, She71, She72a, She72b, She73a, She73b, She74, She79, WT79, FL79, Fle70, Gol78c, LoD71, Sab72, Sta71, Tza73, Wal77, Hil69, Kro71, TCB71, Woo70g, Wil70]. Allocating [Zmu75a, Zmu75b]. Allocation [NR79, Rob77, SZ74, Aus71, Cam71, Ree79c, Ree79d]. Alogorithms [Ano70d]. alphabetically [Far78]. Alternating [MM70, MM71]. Alternative [Par78]. Alternatives [Fry71b]. among [Kin69]. analogue [Mea73]. Analyser [Wat76a, Wat77a]. Analysing [Par75a, Tow78]. Analysis [AB78, CS77, CDZ74, Cla77a, HV78, Hai78, IL74, Jos70, Lan74b, PL74, Par75d, PB76, PB78, Phi70a, Phi74b, Sau72, Sma78a, Ste77b, Wat75a, Whe78, Wil78b, Wri74, Ano79p, Cam73, Gha72, GE71, HR70, Lit70, Mac73, MR77, Mar76, Mid77, Par75e, RT76, WW79, Bro74d]. analyst [AT70, Fel69, Fel70]. Analytic [Zve75]. analytical [Ran71b, WR72]. anatomy [LM77a]. Andrews [Ull73b]. Animated [Rep72]. annotated [SS79]. annual [Ano73k]. ANS [Eme78, Ano76j]. any [Wat77b]. APL [Ear75, Hol78, LL75, OL78]. Apology [Ano77e]. applicable [Tay70a]. Application [Cla78, Dea79, Flo76, HW77, KP79, Par75a, BF72, FB75, God75, Hoa76, Hur73, Per76, Stu71]. Applications [Ham76, LR78, Wal76c, WS73, RL73]. Applied [Par78, ZD71]. Appreciation [Gea70]. Approach [CM79, CBB^{+}76, Eat74, Jon71c, Lim74, Ben72, BD79, Bra76, DP73, Fle70, Har70, Jen74, Lle73, PW74a, Pen76, Pen77, Roh75, SL73, Wal76d, Win75]. Approaches [Bar71, Ano79p]. approximants [HGMM75]. approximating [Cox71]. Approximation [McL78, OP73a, For72, Kei79, KL71, Lil70, Lil71, MR76a, MR76b, OP73b, OW71, Rei71b, SKY73, ZL73, Sal70a, YFM71]. Approximations [BRH70, Har76, Dun73, FGH71, Jac73, OP75, Wat73]. Arbitrary [Cra72, McL74]. Architecture [Ano76b, Bro74a, Sum78, FH77, WS74]. archiving [Wig75]. Area [Smi71, Lew71b]. Areas [Bae75]. Arithmetic [Aki77, Goo77, GH73, RL72, RL75, Dew71, Dew72a, Jon72, Dew74]. ARPL [HR74]. array [Boo67b, Boo67a, BW73b, DW77, Hil69, Woo70g]. arrays [HV74, Kro70d, Smi70a, Tay76b]. Art [Sma78a, Wil74]. articulated [LP73b]. Artificial [Coo72c, BMPS71]. Arts [WCH78]. Aspects [F^{+}73, CL78, Lew71a, Pie71]. assembly [Bow73]. assembly-language [Bow73]. Assessing [SW79]. Assessment [Mae78, Par78]. assisted [WW79]. Associative [Lea78, BL72]. asynchronous [BKK78, Lew71a]. Atlas [BT71, RC70]. attitudes [Lem79]. Attribute [BW78]. Attributes [Der79]. Audit [Cha75]. Auditing [Flo74c]. authors [Ano72j]. autocode [CMRW72, RC70]. Automata [Ano75f, Coo72d, SY72]. Automated [SN75]. Automatic [AEH75, Han70, HW77, Lyn69, Mea73, WL72, BWZ71, Chr73a, CP71, GL70, JS71, LK77, McC71a, McC72, Tow77b, Ull77, Val71b, Van70, WF78]. automatically [Par75e]. availability [Mar76]. award [Ano79f]. Axiomatic [Wri74]. axis [McC71b, SM71].

B [HR73, Kit74]. Backus [CCH73]. Backus-Naur [CCH73]. Bacon [NLW74]. Balanced [WW76, ES77]. Balancing [Day76, Gra71, Wil78d]. band [Zam70]. banded [EH76]. bandwidth [Jen71]. banking [BSH70]. Barron [Eva78]. Base [BHW78, CS77, Dav78, Dea79, Eri74, Flo76, GKS79, HW77, PDY74, Str72, BJ78, CCL72]. Based [BKM78, Cam78, Cra72, KM74, Lan76b, PC78, Tur78]. Bases [CO76, CEM72, Sma78a, Wil72, Ano76k, Don78, NS77, Sch76]. BASIC [Bar72b, Sil72, BC73, Har70, Jen74, Bul72, Lee72, Wil78c]. batch [HG71, Sto72]. batched [Kol78]. Bcs [Ano79f, Dak73b]. be [BS73a]. Beaumont [Ano76f]. Behaviour [Jos70, Abb70]. Behind [Rob75]. Benchmark [Wal76d, CW76]. Bessel [Gar78]. Best [BRH70, Rob77, FGH71, Wat73]. Between [Ste77a, WCL71, Wil78f]. Bézier [For72]. bi [dCS77]. bi-directional [dCS77]. Bibliography [Sma78a, SS79]. Biharmonic [BW70a]. Binary [PB76, PB78, WW76, Wym73, Chr71b, Day76, ES77, FL79, Hur73, PW74b, Sou77, Ull77]. Binding [LM77b]. Biquinary [Bro74f]. Bird [GS73]. bisection [BE79]. bit [LMS73]. bivalent [Chr73b, Mot79b]. Blind [Fis76, Col75]. Block [Edw77b, ID73, Lee78b]. Block-length [ID73]. Blocking [MS74, NV77, Wat71a]. Board [Nor78]. Book [Alb71a, Alb70, Alb71b, Alb71c, Alb75, Alb76, Ale77a, Ale78, And75a, And77a, And77b, And78, Ano77f, Ano77g, Arm78a, Bac77, Bac78, Bal79, Ban79, Bar76a, Bar79, Bar70, Bar74c, Bar75b, Bar77b, Bar77c, Bar78b, Bar78c, Bar78d, Bar78e, Bar73b, Bea77, Bel71, Ber74, Ber73, Bes79, Bil77, Bla75, Bla79a, Bla79b, Bla79c, Boa79, Bra78a, Bra78b, Bra79a, Bra79b, Bra78c, Bro74d, Bro74b, Bro74c, Bro74e, Bro77, Bro74g, Bro75a, Bro75b, Bro76b, Bro70, Bul73, Bum78, Bum79, Bur70a, Bur70b, Bur71, BS73b, Bur73a, Bur73b, Bur75c, Cai76, Cha73a, Cha73b, Cha77a, Cha77b, Cha78b, Cha79a, Cha79b, Che71, Cla76a, Cla76b, Cla79, Cla76c, Cle75, Cle77, Clo77, Clu77a, Clu77b, Coa77, Coh70, Col78, Col77a, Col77b, Con76, Con77b, Con77a, Coo71, Coo76a, Cor79, Cra71, Cro77, Cru78, Cun77, Cur78]. Book [Dav77, Day73a, Day73b, Day75, Day78b, Day79, Den75a, Den75b, Dew75, Dic75, Dir70, Div74, Duf73, Duf74, Duf78, Duf79, Eas75, Eat72, Edm73, Edw77c, Edw77a, Edw79, Eme78, Eme79, Eme75, Eme76, Eva70, Eva78, Eve71, Eve74, Fel71, Fel78a, Fel78b, Fer74, Fis79a, Fit74, Fla75, Fla77, Fle73a, Fle73b, Flo70a, Flo70b, Flo73a, Flo73b, Flo74a, Flo74b, Flo79a, Flo79b, Flo79c, Fox74, Fox76, Fre78, Fry71a, Fry71b, Gar73, Geo70, Geo78, Gil78b, Gil79, Gir73, GS73, Gir76, Gle79a, Gle79b, Gol75, Gol77a, Gol77b, Gol77c, Gol78b, Goo70, Goo74, Gou70a, Gou70b, Gri79a, Gri79b, Gur78, Gur75, Had75, Hal71, Hal74, Hea77a, Hea73, HR73, Hea76a, Hea76b, Hea77b, Heb79, Hig77a, Hig77b, Hil73a, Hil74, Hin79, Hio70, Hio78a, Hio78b, Hoa70, Hoa77, Hol77, How75a, How76a, How76b, Hun77]. Book [Hun79, Hyn73, IL74, Ing75, Ing77a, Ing77b, Jac74, Jac77, Jac79a, Jan77, Joh79, Kem75, Kha75a, Kha75b, Kha76, Kha77, Kha78, Kha79, Kin75, Kir73, Kir74a, Kir74b, Kir75, Kir79, Kit74, Kit77a, Kit77b, Kit79, Kno75b, Kno77, Lan74a, Lan74c, Las73, Las79, Lav74, Lav78, Lea79, Lee79a, Lee79b, Lee79c, Lee75, Leh75, Leh76, Leh77, Lew73, Lew76, Lew79, Lid74a, Lid74b, Lim75, Lin73, Llo70, Llo74a, Llo74b, Los77a, Los77b, Mar74a, Mar79b, May79, May74, May76, McK79, Mee73, Mer75a, Mer75b, Mer75c, Mil79, Mit70, Mit77, Mor76, Mor79, Mor75, Moy78, Mux74, Mux76, Mux77a, Mux77b, Mux78, Nee74, NLW74, Nut78, O’C79, Oll75, OS71a, OS75, Pac77, Pag79b, Pai78, Pak73, Par79, Par75b, Par77c, Pau79, Phi70b, Pot78, Pow73, Pow74, Pow76, Pri70, Pri77a, Pro74a, Qui79, Rad73, Rad75a]. Book [Rad75b, Rad75c, Rad75d, Rad76, Rad79, Ran71a, Ray79, Ree73, Ree74, Ree76a, Ree76b, Rei71a, Rei73, RL73, Ric78, Ril77, Ril78, Ril79, Riv79, Rog74, Rog75, Rog76, Rog79a, Rog79b, Rzv79, Sam73a, Sam73b, Sam73c, Sam74a, Sam74b, Set79, Sha77a, Sha77b, Sim79a, Sim79b, Sla70, Sla71a, Sla71b, Slo78, Sma78a, Smi73, Smi75, Smi76, Smi77, SK78, Smi79, Sno78, Sta77a, Ste75, Stu73, Suf79, Sum76, Tay73a, Tay73b, Tay76a, Tho73, Tho74a, Tho74c, Tho74b, Tho79a, Tho79b, Tho79c, Tho79d, Tom71, TCB71, Tow77a, Ull73a, Ull73b, Vaj70, Vaj76a, Vaj76b, Vic79, Wal73a, Wal73b, Wal74a, Wal75a, Wal76a, Wal76b, Wal78, Wal79a, Wal79b, Wal79c, Wal73c, WS73, Wat74a, Wat71b, WY71, Web70, Wei79, Wel73, Wel75, Wel76, Wel74, Wel77, Wel79, Wes71, Wes74, Whe79, Wic78, Wil75, Wil78a, Wil70, Wil76c, Wil76d]. Book [Wis74, Wis75, Wis78, Wit79a, Wit79b, Woo74a, Woo78a, Woo78b, Woo74b, Woo76, Wre71a, Wre71b, Wyv77, You70, You71, YFM71, You72, You74, You75a, You75b, You77a, You77b, You77c, You78, You79a, You79b, dR77, Bon71]. Books [Ano75d, Ano79g]. Boolean [Edw72, LT79, Wil77e, ZD71, ZD73]. Both [SNH76]. Bounce [NI70, Osb71b]. Bounce-and-skip [NI70, Osb71b]. Boundaries [BW70a]. Boundary [BW70a, Tay70a, Cas75, Eva77, Fyf70, Sne70, Tay70b, Zaf71]. Bounds [Han70, SNH76]. Boyserf [Ree79b]. Bracketing [Sza73, Cox70, SC71]. Brain [WY71]. Brains [WY71]. Brent [SL78]. Bringing [Par75c]. British [Bon71, Ros70]. Buffer [LV79, DL76]. Building [Cor79, Sla71b, Wai70]. Bull [NLW74]. Bulletin [Ano73d]. bureau [AF76]. Business [Eat74, Ste77a, Ben70, CT74, Coa74, Fis74c, Gea73].

C [Cur78, Fry71b, IL74, Llo74b, Ull73b, Wre71b]. Cad [Clo79, Fis79b, Mar73, Mar74b]. calculate [WD79]. calculating [Abb70, Pec70, Tay70a]. Calculation [Hon71b, SH72a, FGH71]. calculations [God75]. Calculator [Wil78b]. calculators [BL77]. calculus [Tur78, WS73]. Call [Ano75e, Ano75f, Ano77j, Ano78e]. Calls [Ano79h, Ano79i]. CAM [Wat76a]. CAM02 [Wat77a]. Cambridge [Ano77j, BBH70]. Campey [Hos71]. Canadian [God75]. capabilities [SW79]. capacitated [Lig73]. Carlo [KP79]. Cartoon [Rep72]. Cascade [Sco74]. Case [Rob77, Lee78a, LP73a, Mid77, SW79, Wre71b]. cautious [Haz78]. CDC [BBG^{+}71]. Cellular [Coo72d, DW77]. central [Fos71]. Centralised [Dav78]. Certain [Bro74i, Hun70, LoD70, AHM74, Eva72, EH76, Fyf70, Goo75, HM77, Sal70b, Spä69]. Certification [Boo78, Boo79d, Boo79c, Ran79, RT78]. Chain [Mer70b, Woo70e]. Chained [Bay73, CS77]. Chains [Lor76]. Change [Tre70, Chr73a]. Changes [Ano73k]. Chapman [Eva78]. Character [Rey77, Sau72, CCL72]. characteristic [Hop71]. Characteristics [Lyn72]. charts [Dun79a, Dun79b]. Chebyshev [Fox76, Sal73b, AEH76, BS71, Cha70, KS71, OP73a, Sal72, Scr70, Sne70]. Check [Bro74f, ES75, How75b]. checker [HP71]. Checking [Dun78, Hig76a, And74, Bro74h]. checks [HK73]. chemical [KM74]. Chess [BKM78, Pro74b, Ano78a, Bel70a, BJ79, BB71, BM79, BW75b, WBB72]. chi [PFW74]. chi-squared [PFW74]. Chilton [BT71]. Chisholm [HGMM75]. Choice [Edw77b, JS71, BSH70, WCL71]. Choosing [Roh78, Wil78e]. Chow [Hur73]. Christoph [YFM71]. Chromatic [Chr71a, Mit76, SNH76]. circuit [AL72, ZD71]. Circuits [PL74, TD77, YT79, Ben72, DZW75, Wal74b]. circular [Lew71b]. circulation [Dow71]. CLAM [dRC74]. Class [Has75, LoD70, PW74a]. Classes [Bae75]. Classification [BW73a, BW75a, JW74, BW70b, JW72, JS71, Sal71b]. classifications [WCL71]. classificatory [LW71]. Clenshaw [Hun70, Oli72]. Clinical [CBB^{+}76, Gor78]. Closed [Jac73]. Closure [Weg74]. cluster [Sib73]. clustering [Mux70, Roh73, Sal70c, Sal71a, vR70a, vR70b, vR70c]. Clusters [CW70, Lin72b, Roh75]. CMSR [BG71]. COBOL [Eme78, Ano73d, LM77b, Tri74, Tri76, Tri78, Wil77a, HH79b]. CODASYL [HKT75, Sta74]. Code [AD76, Bro79b, NC76, Bem74, DP73, Daw73, Mac74b, War74, Ros70]. Codes [Bro74f, Bro74i, Gob75]. CODIL [Rey71b, Rey71a]. Coefficients [Han70]. Colin [Cha73b]. Collection [BS73b, FN78, Wil70, BPSS72, Hay79, Tem72b, Tur78]. Collector [Weg72]. College [BF71]. combination [Phi74a]. combinational [Ben72]. Combinatorial [Ham76]. Combining [Daw73]. Comm [Gol73]. command [Tes77]. Commands [Sch77b]. Comments [Deu69, Sab72, Fra70]. Commercial [GS73, Gun70, OBLP75]. Common [MJ75]. Communication [Pri78, FH77, JBD73a, JBD73b]. Compactifying [Weg72]. Compacting [FN78]. Comparative [Cla77a, Val74]. comparing [LK77]. Comparison [Cro75, SL78, AL72, LMS73, MD71, Tan78]. competition [Ano73a]. Compilation [Els79]. Compile [Bur72, Mux73]. Compile-Time [Bur72, Mux73]. Compiler [AD76, BW78, Coo76b, Cuf72, Der79, Mil70, PB74, Suf78, BL77, Cam78, CMRW72, ELR78, Per76, SA73, Ste74]. Compilers [Bur72, Mux73, Wic72]. Compiling [RL72, PR77]. complainants [Hal72b]. complete [Def77]. Complex [RL75, BM79, Jar70]. Composition [Tre70]. Comprehensive [Wat79]. Compression [GW73, Lea78, MJ75, Wel72, DP76, Tur75]. Computation [BY79, Sto73, BL72, Bra76, Cro75, ET75, HK73, Roh75, WY71]. Computational [Alb74, Coo75, MS70, Sma78b, WS73, Wil70, Wil78b, BS75, CL78, HD71, Jac71, Alb71a]. Computations [JW74, Cur78, Fox74]. Computer [Ano74c, Ano76a, Ano76b, Arm78b, Bei72, BCR72, BN78, Cha75, Cha78a, Coa74, CBB^{+}76, Dak73b, Daw74, Eas76, Eat74, Fle75, Hoa76, IL74, JP76, KTW^{+}73, KP79, Lam76, LV79, Leh73, Leh74, Nor78, O’B70, Pie74, Pro74b, RL73, Sco74, SZ74, SS75, Sum78, Var76b, WJ74, Wat74b, Wat74c, Web72, Wil74, WW79, WCH78, Zmu75a, Zmu75b, Ano77k, Ano78a, BBG^{+}71, Bel70a, BB71, Bri79, BF71, BSH70, CP79, Coa78, CW71, Com75a, Com75b, Dow71, FR79, Fox78, GP73, Lan76b, Mar73, Mar74b, Mar79a, McG77, Mea73, Pag76b, Pan70, Ric73, SS72d, SS72e, Tur79, War74, WO75, HL76b, Cor79, Cur78, Eva78, GS73, HR73, NLW74, Ree76b, WS73]. Computer-Aided [Wat74b]. Computerized [F^{+}73]. Computers [AB78, All75, BHRS70, BS79, GS73, Gol73, Ric74, Rog72, Tre70, Ben70, Hil77, HP77, Sym71, WY71, Web70]. Computing [BRH70, Bar78f, Cla77a, Fis74b, Gar78, GS78, Par75a, SB70, Ahm73, Akk73, Cox70, IWT73, Lau75, SC71, Stu71, Bel74, Bro73a, Kit74]. Concept [Eri74, SL73]. Concepts [Hou75, Pag74, Bra76]. Conceptual [BK75]. concerning [WLDC71]. concurrent [Shr75, Wal77, Sil79]. condition [BS73a, Zaf71]. conditional [Tur79]. Conditions [Goo77, Man74, San75, Wil79, Kin69, Sne70, Tay70b, Twi75, Wol71]. Conference [Ano75f, Ano76k, Ano78f, Ano79j, Dak73b, Ano73k]. Conferences [Ano79k]. Congress [Ano77j]. Congruential [Ful76]. conjugate [Cla77b]. connectivity [BP79]. Consequences [LM77b]. consideration [Hur79]. considered [Lip79]. CONSIM [Nel79]. Consistency [Flo74c, HKT75]. constituencies [HW74, Par74, Van73]. constrained [BJS74, Bra75, Tho72]. constraints [AJ70, HB70]. Constructing [GS79]. Construction [LO75, Suf78, WL72, BF71, BS79, ES77, Lin72b, SS79, WP75]. constructive [Ric73, Wal76d]. contact [Abb70]. Content [CEM72]. Content-Addressing [CEM72]. context [Gha72, Goo75, LL70, Woo70a]. context-free [Goo75, LL70, Woo70a]. contextually [EM71]. Continuation [SL78]. Continuity [Man74, San75]. Contour [Cra72]. Contouring [Sut76b]. Contours [McL74]. Control [Bar72c, BPR74, EMM72, Fra76, Ger70, HR73, KP79, Mil70, Nel79, Wal76c, Zmu75a, Zmu75b, Bar74a, Com75a, Com75b, Dak75, Gol78c, Hon74, Leh73, NI70, Osb71b]. controlled [Dow71, KK79, Pri77b]. Controversy [WLDC71]. Convergence [Bre71, Loi72, Lyn69, GL70, Scr70]. convergent [MS70, PM72]. Conversational [Cuf72, Nel79, Bar71]. Conversion [Sod53, WO78, Boo74a, Boo74b, Ham74a, KJ75]. Converting [SS75]. Convex [GS79, Cox71]. Convexity [YFM71]. Convolution [McC78]. Coordinate [Bla78, Cra72, Ter73]. Corporate [WJ74]. Correct [Bra79c]. Correcting [Bro74i]. correction [And75b, ES75, How75b, Ull77]. Correctness [LO75, Sch77b, Smi72]. Correctors [Rod70]. correlation [Jac71]. Correspondence [AAB^{+}79, AT70, All71a, All71b, And74, And75b, Ano70e, Ano74d, Art75, Bac71a, Bac71b, Bal75, Bar72a, Bar72b, Bar76b, Bar73a, Bar76c, Bar78f, Bea74, Bel74, Bem74, Boo74a, Bra75, Bra72, Bro76a, BVH74, Bro72a, Bro74h, Buc76, Cam75, Car70, Cas76, Cha72a, Cha72b, CT74, Coo72a, Coo72b, Coo72e, Coo72f, Cro75, Cro76a, Cro76b, Dak73a, Dak73b, DD70, Daw74, Den75c, Deu69, Dew72b, Dew72c, DW73, Dun78, DI78, Ecc71a, Ecc71b, ET75, Ehr72a, Ehr72b, Ell76, ES75, FT73, Fel69, Fel70, FW72, Fin72a, Fin74, Fis74a, Fis74c, Fis74b, Fla72a, Fla72b, Fow75, Fra70, Fre74a, Fre74b, Gal71a, Gal71b, Gal73, GL70, Gay72, GT72, GW73, Gla73, Gol78a, Gol78c, Gri76, Hal72a, Hal72b, Ham74a, HL70a, HL70b, HT78, Has76, HV74, Hay79, Hea74, Hil72a, Hil72b, Hil73b, HT75, Hil80, Hit78]. Correspondence [Hon71a, Hon71b, Hop78, How75b, IWT73, Jac79b, Jof76, Jon71a, Jon71b, KGC79, Kei79, Kin71a, Kin71b, KTW^{+}73, Kli75, Kra77, Lan76a, Lan71a, Lan71b, Lea75, Lee74, Lew72a, LS78, Mac71a, Mac74a, Mac74b, Mar74c, Mar74d, Mee71, Mee72a, Mee72b, Mer70a, Mer70b, MR82, Moo72, MT72, Mul75, Mux73, Mux75, Nea75, Osb71a, Osb71b, Pal72a, Pal72b, Pal73, Par76, Par77a, PH73, Phi76, Pit74, PS74, PW74b, Red71a, Red71b, Rob80, Roh79, Sab72, Sab76a, Sal73a, Sal73b, San75, SA79, SS72a, SS72b, SBS72, Scr72, Sel72a, Sel72b, SK75, Sil72, Squ71, SC71, Sta75, Str74a, Str74b, Tay80, Tem72a, Tem72b, Tem75, TW75, Tom75, Tri74, TT73, Val71a, Val71b, Vor75, Wal75b, Wal71a, Wal71b, Wan73, Wel70, Wex76, WN73, WNT73, WS74, Whi72a, Whi72b, WR72, Wil77c, WD76, Wod78, You75c]. Correspondence [Zar77, Ben70, Bow73, Bro73a, DCHR76, Dun79a, Hay77, Jar70, Lem79, Low77a, Mar79a, Nau79b, Nor78, Tes77, Tri78, WF78]. Cost [Par75c, Sym71]. costs [Chr73b]. Counting [HW74, Lun75, Par74, Van73]. coupled [AHM74]. course [Ano78f, Maz78]. courses [Ano79k, Bro73a, Fis74b]. CPM [Ste77b]. CPU [ET75, LY79]. Criteria [Lyn69, Par78, Wil78e, GL70, Van73, WLDC71, HW74, Par74]. Critical [Mae78, DCHR76]. Critique [Nau79a, Par78, Nau79b]. cross [BL77, Jac71]. cross-correlation [Jac71]. crosswords [WW79]. CRT [Rep72]. cryptic [WW79]. cubic [For75, Fyf70, HK72]. cumulative [Not70]. Current [Coa78]. Curtis [Oli72]. Curve [Sut76a, Coo72f, Ecc71b, Gro71, McC70b, McC71a, McC71b, McC72, OP71, SM71, Sab76b]. Curves [Man74, McL78, San75, Kei79, Sab72]. cutting [AO79]. Cycle [PB76, PB78]. Cyclic [Bro74h, Hil79a, HK73].

D [Eva78, IL74, NLW74, RL73, Wil70]. Dahl [Llo74b]. Dakin [Gla73]. Daniel [WY71]. Dashed [Ber76]. Data [BY79, Bla78, BHW78, CS77, CO76, CBB^{+}76, CEM72, Cro74, Dav78, Dea74b, Dea79, Dor76, EMM72, Eri74, Eth71, Flo76, GS73, Gor78, GKS79, Ham77, Har76, HW77, HC78, Jen74, JBD73a, JBD73b, KSD77, Lea77, McL74, McL76, Med70, NLW74, Pad73, Par78, PDY74, Pri78, Ros70, Sma78a, Sou76, Ste77a, Str72, Wil72, AF76, Ano75l, Ano76k, Ano79l, BJ78, BPSS72, Bur79, CS75, CPR75, CCL72, DP76, Dea74a, Don78, GM72, Hay79, Jar70, LW71, Lee78a, Mar73, MP79, Nau79a, Nau79b, NS77, OP71, Par77b, Pin77, PS74, Rob79, Sch76, SW79, SD73, Tem72b, Tur75, Tur78, Ano76l, NLW74]. data-display [Mar73]. Database [Dav76, Dea76, Dea78, HKT75, JO74, Ker76, Rob75, Sta74, Par75e]. Databases [Flo74c]. DATAFAIR [Lav70]. Datagram [Pri78]. Date [Ran74]. Date-Sensitive [Ran74]. dates [Hil77]. Davidon [KL71, Lil70, Lil71]. Davies [Hos71]. Day [Cha73b, Ano79p]. DBTG [HKT75]. deadlock [BN76]. deadly [Lle73]. Debugging [Pie74, Bul72, KK79, Sil72]. Decimal [And74, And75b, Bro74f, Bro74i]. Decision [Mae78, SS75, Ben70, Car70, Kin69, Kin71b, KJ75, PW74b, Wym73]. Decisions [Par75c]. Decomposition [DK74, Nas75, WV70, Zam70]. dedicated [CP79]. Deferred [LM77b]. Defined [Har76, Has75, Wat73]. defining [Mar74b]. definite [CP71, Jen71, Oli71, Val71b, Zam70]. Definition [Pit74, Bar72b, Bar75a, BJ78, Cam78, Gol78c, Lee72]. Definitions [Pag76a, Bur79]. Degeneracy [DW73]. degree [Cox71]. degrees [Ano75l]. delay [War74]. deletion [Ull77]. demand [Wil78d]. DeMasi [HR73]. Dendrogram [Roh74]. Denison [Wre71b]. Denotational [Pag79a]. dependence [WCL71]. DEPOT [LMP79]. Derivative [MM70, OP73b, Zaf71]. derivatives [DW73, KL71, Lil70, Lil71, MM71, OP75, Tay70b]. Description [MM78, Mot79a, Mot79b, Sau72, Zve75, BG79]. Design [BK75, CS77, Coo76b, Edw74, Hai78, Hil76, Hol75, Jel79, Mac70a, Moo79, Par75c, Rep72, Sco74, Wat72b, Wat74b, Wat74c, Wil76a, Ano79p, AL72, BKK78, CS75, Eth71, FR79, FKLM76, KL77, Lan76b, Lew71a, Nud77, Pen76, Pen77, Wat72a, ZD71, HR73, HR73, IL74]. design/drafting [CS75]. designed [Aus71]. Designs [Edw78]. Detail [Sch77a]. Detecting [YT79]. Detection [Bro74f, Ben72, BD79, Com75b]. determines [Sal74]. Determining [BF78, NV77, Boo67a, Eva75, Hil69, Lan76b, Woo70g]. Deterministic [CC74, Woo71a]. Development [BD72, Coa78, LLR73]. Developments [Smi70b, Ano73k]. device [Fos71]. devices [Pez76]. diagnosis [BL71]. Diagnostic [WL72, Pan70]. Diagnostics [Bur72, Mux73, PB74]. Diagrams [Wil77b, WO78]. dialogue [BMPS71]. Difference [CM79, BW71, Had71, KL71, Lil70, Lil71, Zaf71, Wil76e]. Different [Sch77a]. Differential [Alb71a, Lam76, Nor76, SB70, SS72c, WY71, Wat72c, BWZ71, BS71, Chr73a, Hop71, KS71, Sne70, Tay70a, Twi75, Wil76b, Wol71]. Diffusion [McC70a, Tay70b]. Digital [Coo75, Edw74, Hai78, IL74, KP79, Leh74, Lyn72, Pag74, RL73, Sma78b, Sto73, BL71, Bri79, Cro75, GP73, Wal74b, Win75, Edw78]. Digits [Bro74f, ES75, How75b, Rea70]. Digraph [Bay76]. Dijkstra [Llo74b, Gal73]. Dimensional [McL76, Kro70d, Lun71]. Dimensions [MM70, YFM71, Bel70b, Hun72, MM71]. Direct [Gun70, WD74, Pez76, SK76]. direct-access [Pez76]. Directed [Bur72, EMM72, Mux73, PR77, Wil76b]. directing [NI70, Osb71b]. Direction [MM70, MM71]. directional [dCS77]. director [Ree75, WBB72]. Dirichlet [GS78]. Disc [BK79b, HS71, Huz79, SR76, Str72, Wat75a, Wat75b, Aus71]. Discontinuities [HCC74, O’R70, Whi72b]. Discrete [Gre70, Med70, Pag74, PC78, SKY73, Coo72f, Ecc71b, McC70b, OW71, PW74a]. Discrete-Time [Med70]. Discussion [Ano79l, Bar74a, Bar78f, Ben70, Bow73, Bro73a, Coa74, Col75, DCHR76, Dun79a, Hay77, Jar70, Lau75, Lem79, Low77a, Mar79a, Nau79b, Nor78, Tes77, Tri74, Tri78, WF78]. disentangle [Woo70e]. disparate [Par75e]. display [BPSS72, Jon79, Mar73, Tem72b]. displaying [Sab76b]. Dissemination [HT78]. distance [Jac73, LoD71]. distant [HKA72, Hos74]. Distillation [Sto73, Cro75]. distinct [Far78]. distinguishing [KRK73]. Distributed [BHW78, Dav78, CP79, Mea73]. Distribution [Pag76b, McC78, Par75e, PFW74, Ree79c, Ree79d, Tur79]. Divider [PB78]. division [Ahm72, Var77]. divisive [LW71]. Document [WK77]. Documentation [Sim74]. documents [SAHS71]. DOLPHIN [Col70]. Domain [Coo75]. Domains [NW74]. Domölki [HR70]. Double [Alp71, SH72a, Hon71b]. Double-Length [SH72a, Hon71b]. Doubly [CS77, Oli72]. doubly-adaptive [Oli72]. Doubly-Chained [CS77]. down [Woo71a]. Dr [Ano75g]. drafting [CS75]. Drawing [Ber76, Gri73, McL74, Smi71, Sut76a]. Driven [Pad73]. DSTL [Edw78]. dual [Kro70e]. Dubner [Dur74]. dumps [Bro73c]. during [Aus71]. Dynamic [Bul72, GS73, Sil72, TW77, TCB71, Abb70, Aus71, Cam71, CPR75, DO72, KM74, Par75e]. Dynamically [NR79]. dynamics [Wil73].

each [Boo67a, Hil69, Woo70g]. Easter [Boo79d, Hil79c, Ran79]. econometric [BS75]. Econometrics [TCB71]. economic [BS79]. Economics [Sla71b, Web70]. Economy [Wol78, Mor78]. Edinburgh [Ano75f, BMPS71, WW73]. edited [RL73, YFM71]. edition [Sum78]. editor [Sla71b, TCB71]. Editorial [Ano74e, Ano76f, Ano78g, Ano79m, Fis74c, Gea73, Ham73, Pag73, TT73, Bar73a, Gal71b, GT72, MT72]. edits [SAHS71]. EDSIM [PC78]. EDSIM-Event [PC78]. Effect [Lyn69, GL70]. Effective [TI75]. effectiveness [Ben70]. Efficiency [Fle75]. Efficient [Dur74, Has75, Øst79, Def77, HV74, Par75e, Sal74, Sib73, Win75]. eigenfunctions [BW71]. Eigenproblem [Nas75]. Eigenproblems [Alb74]. eigenvalue [Zak77]. eigenvalues [BW71, CJ70, Eva75, Kai72, Sea71, Tay70a]. eigenvectors [CJ70, Kai72, Sea71]. EL1 [Weg74]. elastic [Sza73]. Eldon [WHM71]. electoral [Hil75]. electric [Abb70]. Electroencephalograms [Tow78]. Electronic [GS73, Lor76]. electronics [Wal74b]. element [Boo67a, DHW78, Hil69, Woo70g]. Elementary [Gou70c]. elements [Boo67b, Boo67a, Hil69, Woo70g]. Elimination [Bir79]. Ellipses [Smi71]. EMAS [MRW75, Ree75, WW73, Wig75]. embrace [Lle73]. empirical [RT76]. encapsulated [Bur79]. Encodings [GW73, Wel72]. End [BKM78]. End-Game [BKM78]. engineering [Hur79, MR77]. Engineers [Sma78b, TCB71]. Engines [Wil76e]. Enhancements [Zmu75a, Zmu75b]. entity [Ano79l, Rob79]. entity/event [Ano79l, Rob79]. entrance [Hon74]. Entrant [BW70a]. Entropy [Wil78e, Wil78f]. entry [Car70, Kin69, Kin71b]. environment [DO72]. Environmental [Dor76]. equally [OP71]. Equation [SL78, Del77, Had71, Lew71b, Man71a, Tay70b, Zaf71]. Equations [Bro72b, GH73, GH75, Lam76, Lew72b, MM70, Nor76, Phi70a, SB70, SS72c, Tew72, Wat72c, AW71, AHM74, BWZ71, BE73, BS71, Bro69, Chr73a, Cla77b, DA77, Eva72, EH76, Fie70, Hop71, HM77, Jen71, KS71, Lew72a, MM71, PD79, SK76, Sne70, Tay70a, Twi75, Wat73, Wil76b, Wol71, Zam74, Alb71a, WY71]. equi [HKA72, Hos74]. equi-distant [HKA72, Hos74]. equiangular [Sib78]. equidistant [For75, Hos70]. equifrequent [CCL72]. Eratosthenes [Hoa72]. Eric [Ano72c, Ano72d, Ano73e, Ano73f, Ano74f, Ano79p]. Errata [Ano70f, Ano71d, Ano72e, Ano72f, Ano73g, Ano73h, Ano73i, Ano73j, Ano76g, Ano76h, Ano77h, Ano78h, Cam73]. Erratum [Ano71e, Ano72g, Ano72h, Ano74g, Ano75h, Ano75i, Ano75j, Ano76i, Ano77i, Ano79o, Ano79n, Boo74b, Sav70, Sea71, Wat72a, Zmu75a]. Error [Bro74f, Bro74i, Bur72, Han70, Mux73, And75b, ES75, How75b, OP75]. error-correction [And75b]. Errors [Bel76, Bro74i, Cha70, YFM71, BD79, Jac73, Ull77]. Essays [Coo72d]. establishing [Ada76]. Estimate [LV79, Kol78]. Estimating [BK79a, Mit76, Wat75b, Pez76, Wat73]. Estimation [Cha70, CW71, HD71, Jon70, Tza73]. ETAOIN [Kli75]. Euclidean [Wri74]. Euler [Scr72, Wyn71]. European [Ano77j]. Evaluating [Hun70]. Evaluation [BPR74, Coo75, CO72, Lan76b, PB74, Roy78, Wal76c, AEH76, CJ70, CP71, GR75, GE71, HPLG79, Lit70, LP73a, Moj77, Oli71, Sea71, Val71b]. Event [Pag74, PC78, Ano79l, Rob79]. evolutionary [SL73]. examination [DCHR76]. Examples [Wil77a]. Excluding [Kam77]. Executable [Bro79b]. Execution [WD74, Wal77]. Executive [BC73]. Exercises [Wre71b]. exhibit [Par75e]. Expanding [Nor76]. Experience [ON79, Fis76, MS70]. Experiment [LK76, Lew71b]. Experimental [BC73, Mil70, Hen78, HH79a, LKK71, SS72d, SS72e]. experiments [KRK73]. Expert [Mic79]. exploit [Ros70]. Exploration [Gor78]. exponential [Lip79]. Expressions [RL72, SY72]. Extended [Wil79, HL76a]. extending [Tes77]. Extensible [NF76, FKLM76, Mol71]. Extension [Fin72b].

Extensions [Ehr72b, Fin72a, Mee72b]. extrapolated [Had71]. Extrapolation [Vit75].

F [Web70]. faces [Jar75]. Facilities [RC70]. Facility [BD72, Cam78]. Factor [TI75]. factored [Kom74, Woo70b]. Factorising [MJ75, Alw71]. Factors [MS74, NV77]. faculty [Alm69, DD70]. failure [MR77]. Fallacies [Wat72b, Wat72a]. Family [Ram70, MP79]. Fast [MH77, Tay76b, DA77, Eva77, vR70b]. Fault [BL71, Com75b, Whe78]. Faults [BKL74, Hil71, SS72b, YT79]. Features [Wil78c, FH77, LT71, Par75e]. February [Gal71b]. Federation [Ano77j]. Feedback [GK73]. Fewer [Lin72a]. Fibonacci [Boo79c, Ove67, PHJ67, PP79]. Fike [Roh76]. File [GW73, Par75c, SR76, SH72b, Wat72a, Wat72b, Wel72, Aus71, BBG^{+}71, Bon71, Wil78f]. Files [Edw77b, Huz79, ID73, NV77, Ran74, Wat75a, Kol78, Wat71a]. filing [Col70, SS72e]. film [Ano77k]. filtering [Hun72]. Filters [Han78, Lyn72]. Find [Kro70a]. Finding [Bre71, SNH76, Kai72, KL71, Lil70, Lil71]. Finite [Ham77, Hun70, Woo77a, YFM71, BW71, DL76, Lle73]. firm [BHRS70]. First [Lew72b, Rob77, AW71, Chr73a, Cox71, Lew72a, OP73b, Wol71]. Fit [Rob77, TI75, Cam71, Ree79c, Ree79d]. Fitting [Sou76, Gro71, McC71b, OP71, SM71]. Five [Wic72]. Fixed [Smi71, BP79, Sto72]. Flexibility [ID73, Rey71a]. Flexible [SZ74, Zmu75a, Zmu75b]. Flocoder [MKL71]. Flow [Wil77b, WO78, Mid77, NI70, Osb71b]. FOCUS [BBG^{+}71]. folding [EH76, Lun71]. following [Dak73b]. Form [JW74, WO78, CCH73, Jac73, LL70, Mid77, Woo70a]. Formal [Gha72, Bar72b, Lee72]. Formalisms [Sch77b]. format [Car70, Kin69, Kin71b]. Formats [Zve75]. forms [FR79]. Formula [Bay76, Ada76, Sym71]. Formulae [AEH75, Bar75a, BJ78, Oli71, Pie71]. Forsythe [Ano73a, Cur78, Hit78, LS78, SG78]. Fort [Tay70b]. FORTRAN [Cha72b, Dew72c, Ehr72b, Fla72b, Mee72b, Ree76b, SS72b, Art75, Bar71, BLY70, Bur79, Coo72b, Fin72a, Hal72b, HV74, Tay76b, Ano76j, Cha71, CPR75, CC74, DCHR76, Fin72b, HL70b, Hil71, Hou71, Mac70b, Mac71b, NC76, PC78, Rey77, RT76, Sal71b, SA73, Sta74, Wel70, Cha73b]. foundation [VG73]. Fourier [MH77]. Fourth [O’R70, Fyf70, Whi72b]. Fragmentation [Rob77, Aus71]. Francis [HR73]. Frankel [Tay70b]. fraud [Cha78a]. Fredholm [DA77]. Free [MH77, Ree79c, Ree79d, Gob75, Goo75, LL70, Woo70a]. free-text [Gob75]. French [McC71a, McC72]. French-curve [McC71a, McC72]. frequency [Mea73, Par75e]. Full [Bur75a, Cuf72]. Function [Aki77, Bre71, dRC74, CP79, Cox70, Dun73, FB75, HB70, KL71, Lil70, Lil71, OP73b, Phi74a, SC71]. Functional [Lan74b, Rin75, Stu71]. Functions [BRH70, Cra72, Has75, Kno75a, Sou76, WV70, Cox71, EM71, Gar78, Hil71, Hur73, Jac73, Jor73, McC78, SS72b, Wat73, TCB71]. Fundamentals [Wal76c]. Further [Fra70, Woo71a]. future [Bro73a, Fis74b, Nor78].

G [Cur78, Fox74, HR73, NLW74, Sla71b]. Galerkin [DA77, Del77]. Game [BKM78, Bra79c]. GAMM [WD79]. Garbage [FN78, Weg72]. Gate [Edw78]. Gates [EH78, Win75]. Gauss [Cha70]. Gaussian [Pie71]. General [Has74, PC78, Alw71, Bra74, Bri79, Dak75, EA70, Eva71, Mar74b, Sne70, SS72d]. Generalisation [Gou70c, Par75e]. Generalised [BHW78, OP73a, Weg72, Zak77, BJS74, Bra75, Pin77, PM72, Scr72, Woo70a, Wyn71]. Generated [Ful76]. Generating [CW70, Day77, Roh76, Roh78, Wil77b, FL79, Pan70]. Generation [AEH75, AD76, Far78, HGMM75, OS70, OS71b, SY72, Woo77b, YT79, Ben72, Fik75, Pag71, Tow77b, Var77, WF78]. Generative [RW74]. Generator [SS72c, MWC70]. Generators [Woo77a]. Generic [Sch77b]. geometric [FR79, FB75]. geometry [Ric73]. George [Ano73a]. Gill [Ano75g]. Girling [HR73]. given [Tur79]. giving [Woo70d]. Global [Bar74b, Pri77b]. Go [LR72]. goals [Lan76b]. Golden [Hil79b]. GOTO [Gal73]. Gozinto [DK74]. Gradient [Ram70, Cla77b]. gradients [Pec70]. graduate [Ano79q]. gram [Ull77]. Grammars [BW78, Has74, LL70, Ter73, Woo70a]. Graph [Chr71a, DK74, Mau73, Mit76, SNH76, Wil76b]. Graphic [WCH78, CS75]. Graphical [BJ78, EM71, Gor78, Mac70a, CCH73]. Graphics [CDZ74, Gro73, Mar73, Mar74b, Ric73]. Gregory [Wil70]. ground [Tur75]. Group [WCL71, Sta74]. Group-size [WCL71]. grouping [CO72, Moj77]. Guaranteed [Bre71]. Guidelines [Hil76].

H [Alb71a, Bro74d, BS73b, Cor79, HR73, NLW74, YFM71]. Hall [Cur78, Eva78]. Handling [Cro74, Dor76, Hil77, Rey77, Tri74, BBG^{+}71]. Hanoi [Wod78, Hay77]. Hardware [Edw74, HP77, Mar76]. hardware/software [HP77]. Hardwired [BK75]. Harry [Ull73b]. Hash [AK74, Bay73, HD72, SD78, Eck74, Gob75]. Hashing [Kno75a, PH73]. Having [Han70, SH72b]. Head [BK79b]. Health [Ker76]. heat [Zaf71]. held [Ano79p]. Helsinki [Ano79q]. Herbert [BS73b]. Hermitian [Gou70c]. Heuristic [Coo72c, Mic71, NV77, AO79, Wym73, dCS77]. Heuristics [Edm74]. Hidden [Jon71c, Wil77d, Ric71]. Hierarchic [BW73a, SR74, vR70b]. Hierarchical [Moj77, Roh73, CO72, Wil78d]. High [Bac71b, Eme78, ER71, Ger70, Hil72b, Wal71b, Ahm72, BLY70, BP79, Bra72, Hop71, Man71a, Oli71, Pal72b]. High-Level [Ger70, BLY70]. high-order [Hop71, Man71a, Oli71]. Himmelbett [Fre74b, Str74b, Bel73]. Hints [Bra77]. historic [Ano77k]. Hit [Wat76b, Pez76]. Hoare [Llo74b]. horizons [Ben70]. house [AF76]. HP [Wil78b]. HP-25 [Wil78b]. Hull [GS79]. Human [vO78]. Hybrid [Coo75, KP79, SB70, Sto73, Cro75, Fox78, Jac71]. Hyperbolas [Smi71]. Hyperbolic [SB70, Hop71, Twi75]. Hyvärinen [TCB71].

I/O [LV79, Sil79]. IBM [Pal73]. ICL [HP77]. Ideas [Rob75]. Identification [Hol75, CCL72]. IFIP [Ano72i]. II [HH79a, Pen77]. illustrative [Kro70b, Kro70c, Kro70d, Kro70e]. image [DW77]. Imbedding [Sco70].

IMP [Ste74]. Implementation [Gil73, Has75, Hou75, dRC74, Bar75a, Bra74, LMP79, McD77, OL78]. Implementing [KRR79]. implementors [ELR78]. Implicit [Nor76, CL78, Wol71, Zaf71]. implied [EM71]. importance [Rey71a]. Impressions [Dak73b]. Improved [AW71, CW70, Edw78, Lew72a, Lew72b, Sal71a]. Improvement [DL76, Dur74, Aus76]. Improvements [EH78, Roh76]. Improving [KSD77, LY79, Scr70, War74]. in-house [AF76]. Inaccuracy [JP76]. Inadequate [Lyn69, GL70]. income [God75]. incremental [Gri73, McC71a, McC72]. Index [NT79, NS77, WK77, Bon71, Kol78, McD77]. Indexed [Huz79]. Indexes [Ing74, LP73b]. Indexing [Smi70a, Wat75a]. India [Ano74c]. Individual [YT79, Ada76]. induction [Bra77]. Industrial [Wal76c]. Industry [Sla71b]. inequalities [Wil77e]. inequality [AJ70]. Informal [Atk76]. Informatics [Ano77j]. Informatik [Ric76]. Information [Ano73d, BHW78, BW73a, BW75a, F^{+}73, Ham74b, HM70, IL74, Lan74b, MJ75, Sch77a, Ben70, BG71, HT78, Lan76b, vR71, TCB71, NLW74]. Infotech [Sma78a]. Initial [Sco70, Wat72c, Hop71, Twi75, Wol71, Bar75a]. inner [Bro73b]. Input [Ing76, Pad73, BS75, DI78, Hop78, SS72e, Tay76b, Wil76b]. input-output [BS75]. input/output [SS72e, Tay76b]. inputs [Sal74]. inquiry [Sto72]. Inscribed [Smi71]. insertion [Ull77]. Insight [Sam72]. Institute [Ano74j, Ano76a, Ano76b]. instruction [Ibb72]. instructional [WO75]. Insuring [All75]. integer [FL79, Jon72, Wil77e, TCB71]. Integers [Bro74i, Bel70b]. Integral [Lew72b, Phi70a, AW71, GE71, Lew72a, Lit70, SY72, Wat73]. Integrals [AEH75, AEH76, CP71, McC78, Oli71, Val71b]. integrands [Sha72b]. Integrated [Col75, Wal74b, Mac70b]. Integration [Har79, HCC74, HG71, RM78, BS71, Don78, GM72]. Integrity [Dav76, Wil72]. Integrodifferential [Phi70a]. Intelligence [Coo72c, BMPS71, Bro73c]. Intelligent [KSD77]. intending [Ano72j]. Interaction [Ano76a, NC76]. Interactive [BY79, CDZ74, For72, GP73, Gor78, Hil76, Mac70a, Wil76a, CS75, God75, Gro73, Mar73, Mar74b, MR76b, PW74a, Pin77, Stu71]. Interface [HW77, Sta74, Ste77a, HP77]. interior [BW71]. Intermediate [Els79]. Internal [BN78, BF78, Woo70d, Woo70c]. International [Ano75f, Ano76k, Ano76l]. Interpolating [Hos70]. Interpolation [HKA72, Hos74, Mau73, McL76, Sal70b, Spä69, TI75, Vit75, For75, For72, HK72, Sal72, Sal73b]. Interpretation [Car70, Kin71b, Kin69]. Interpreter [Pag76a, Rey71b]. Interpreter-Oriented [Pag76a]. interpretive [Bem74, Daw73, Mac74b]. Interrogating [Ran74]. intersection [Sab76b]. Interval [GH73, Han70, RL75, Dew71, Dew72a, Dew74]. intrinsic [Coo72f, Ecc71b, McC70b]. Introduction [HR73, Cor79, Bro74d, Fox74, IL74, RL73, Ull73b, WS73]. introductory [Maz78]. Invariant [Sco70, McC71b, SM71]. inverse [AL72]. Inverses [PW70]. Inversion [Dur74, Pie71, ZL73]. Inverted [Ing74, McD77, NS77]. inverting [Goo75]. Investigation [Dea76, Phi74a]. Invited [Sma78a]. involving [McC78, Tay70b]. Irregular [Cra72]. Island [CDZ74]. ISO [Nau79a, Nau79b]. issue [Ano79g]. Issues [Nel79]. items [Ada76, Woo70d]. Iteration [Mid76, CJ70, Gor79, Lew71b, Sea71]. Iterative [Ahm72, BW71, CM79, Phi70a, Had71, PM72]. IV [Ehr72b, Mee72b, Art75, CPR75, Fin72a, Fin72b, HV74, Tay76b].

J [Fox76, GS73, HR73, IL74, Llo74b, NLW74, TCB71, WY71, Wre71b]. Jacobi [Loi72]. James [WY71]. JANUS [Lav70]. Jardine [CW70, Roh75]. Jardine-Sibson [CW70, Roh75]. JK [Kai72]. Job [Bar74a, Fra76, HR73, Moo79]. Jobs [MS74]. jobstream [HPLG79]. John [Fox76, Wil70]. joint [Ada76]. Josef [YFM71]. journal [Pag73, Gil73, Ano74c, Daw74, KTW^{+}73]. Journals [Ano75k]. Jr [NLW74]. July [Ano75f]. Jumping [ASG79, Hil80, MR82, Rob80, Tay80]. Junction [Sav69, Sav70].

Karney [Wil70]. KATE [Tes77]. Katzan [NLW74]. KDF9 [WHM71]. Kendall [Sla71b]. Key [BK79b]. Keys [SD78, Pan70]. keyword [Bon71]. Kind [Lew72b, AW71, Lew72a]. knots [Hos70, HKA72, Hos74]. Knowledge [BKM78]. Kochenburger [GS73]. Kriegspiel [BW75b, WBB72]. Kuo [WS73]. Kutta [Whi72b, CL78, O’R70].

L [Kit74, RL73, TCB71, Wil70]. Lagrangian [Sal73b, Sal72]. lags [Mea73]. Laguerre [GZ72]. Lance [Sib71]. Lang [Wil76a]. Lang-Pak [Wil76a]. Language [Atk76, BBG^{+}76, BHW78, Duf72, Ear75, Els79, Hig76a, KP79, Mil70, Moo79, NF76, PDY74, Pie74, Sta77b, Wil76a, Wol78, Bar72b, BLY70, BS73a, BG79, Bow73, Bro73c, BKK78, CMRW72, CCL72, Dak75, ELR78, Hig76b, Hol78, KL77, KM74, Lee72, LMP79, LP73a, Mol71, Mor78, Pal73, Par75e, Per76, Rey71b, Smi70b, Sta71, Ste74, UK74, WP75, HR73, HR74]. language-oriented [UK74]. Languages [Ano75f, Bar72c, Lim74, Pag76a, PC78, WD74, Bac71b, Bar74a, Bra72, ER71, Gha72, Goo75, Hil72b, Kom74, Pal72b, Tes77, Wal71b, Woo70b, Woo71a]. Laplace [AL72, BW71, Dur74, Had71, Pie71, ZL73]. Large [Bro72b, Chr73b, Ano76k, BP79, BPSS72, Cla77b, Com75a, Com75b, Sch76, SK76, Tem72b]. large-scale [BPSS72, Tem72b]. Laurent [BS73b]. Least [LR72, PW70, TI75, Del77, Jon70, PM72, ZL73]. Least-Squares [TI75, ZL73]. left [Kom74, Woo70b]. Legal [F^{+}73, Bel70a, BB71]. LEGOL [Sta77b]. Lehman [Cor79]. Lehmer [Ful76]. Length [GW73, SH72a, Wel72, Art75, CCL72, Hon71b, ID73, Tay76b, Tur79]. letter [Gla73]. Letters [KTW^{+}73]. Level [Dea79, Eme78, Ger70, GKS79, Bac71b, BLY70, Bra72, ER71, FH77, Hil72b, Pal72b, Per76, Wal71b]. Levels [Sch77a]. Lexical [HV78]. library [Not70]. Limited [Aki77, Car70, Kin69, Kin71b]. Limiting [GH75]. Line [Jon71c, Par75a, BBG^{+}71, CW71, DP76, Deu69, Fra70, HP71, Sab72, Sma78a]. line-syntax [HP71]. line-thinning [Fra70]. LINEAL [KU74]. Linear [Bro72b, Fle75, Krá71, Lyn72, SG78, Tew72, YFM71, Akk73, AHM74, BE73, Cas75, DA77, Del77, EA70, Eva71, Eva72, EH76, FB75, FGH71, Gol78c, GE71, Hit78, HM77, Jon70, KU74, Kro70b, Kro70c, Kro70d, Kro70e, LS78, Lit70, Øst79, SK76, Sne70, Tza73, UK74, Zam74]. Linear-Phase [Lyn72]. linearisation [Tho72]. Lines [Ber76, Ric71, Sal70a]. Link [BW75a, Com75a, Com75b, Def77, Sib73]. linked [Tho72]. LISP [Jor73, Der79]. List [Atk76, Ing74, Mil70, RC70, Bro79a, Gor79]. listed [Woo70f]. LL [BS73a]. local [McC71b, SM71]. Locally [Sib78, WW76]. lock [Sch76]. Lockout [Lor76]. locus [Mea73]. Logic [Edw72, Edw78, EH78, HR73, PL74, Sco74, VH72, Ben72, BVH74, BKK78, DW77, DZW75, Hur79, KL77, Lew71a, HR73]. Longo [HR73]. Look [Wil77a]. Look-Up [Wil77a]. Loop [DFO79]. Louis [RL73]. Low [FH77, Bac71b, ER71, Hil72b, Wal71b]. Ltd [Wil70]. Ludgate [WR72, Ran71b]. Lump [SNH76]. Lyapunov [Man71a].

M [BS73b, Cur78, HR73, NLW74, Sla71b]. Machine [AD76, BC73, GH75, KRK73, Ahm73, Bem74, BG79, CD77, Daw73, IWT73, Mac74b, Ran71b, Sal74, WR72]. Machines [Duf72, Tza73]. macro [Cam78, Tes77]. macro-processor [Tes77]. macroprocessor [Mac71b, Nud77]. Macros [Mid76, Bow73]. Magnetic [Edw77b, ID73, Wat75b, Wat71a]. magnitudes [Boo67b, Hil69, Woo70g]. Main [BK75, HH76]. Mainly [Mau73]. Majority [Wat77b, HW74, Par74, Van73]. make [BJ79]. Making [WCH78, Ben70]. Mal’cev [TCB71]. Malcolm [Cur78]. Man [Ano76a]. Man-Computer [Ano76a]. Management [Ano76l, BY79, BK75, CO76, Coa74, Dea74b, Dea79, Gor78, HM70, Jen74, KSD77, NW74, PDY74, Dea74a, PS74, SD73, NLW74]. manager [Fis76]. Manipulation [PDY74]. manipulative [BF72]. Mano [HR73]. Manual [Sim74]. many [Jar75]. map [Lun71]. map-folding [Lun71]. Mapping [BHW78]. Maps [Sau72]. marks [Var79]. Marquardt [TI75]. Martin [Bro74d, Wre71b]. MASCOT [SJ79]. Massachusetts [Ano74j]. matching [Sza73]. Mate [Pro74b, Man71b]. Mate-In-Two [Pro74b]. Mathematical [Bar74b, CDZ74, Cor79, Cur78, Sma78b, Ull73b, Wre72, Sla71b]. mathematics [McG77]. Matrices [Edw72, Gou70c, GZ72, LoD70, Loi72, RL73, Wil70, CJ70, Hat76, Hur73, Sea71, Zam70]. Matrix [Alb74, Coo75, Fox74, Gra71, Bra74, DW73, Eva75, Kai72, Man71a]. Matrizen [Alb71a]. Maurer [BS73b]. MAX [Nud77]. Maximum [Smi71, Wil78e]. McNeary [WS73]. Mead [Aki77]. Meadow [IL74]. mean [GE71, Lit70]. Meaning [BS73b]. means [Cox71]. Measure [BW73a, BW75a, Hun75, WD79]. Measurement [Sou76]. measurements [Jac71]. Measuring [Lem79]. Medical [Ano77j, JO74]. Medicine [Ric74]. Medizin [Ric76]. meets [Clo79, Fis79b]. member [HW74, Par74, Van73].

MEMBERS [BC73]. Memorial [Ano73e, Ano73f, Ano72c, Ano72d, Ano74f]. Memory [HH76, DO72, Stu71]. Memoryless [Fen74]. Merbach [YFM71]. merge [Woo70d, Woo70c]. Mersenne [Hil79a]. Mersey [Hon74]. Mesh [Cra72]. Meshes [Har76]. Message [Gro73]. Metalanguage [Pag79a]. Method [Aki77, CW70, Dur74, Ful76, HD72, Hun70, LV79, Loi72, Mal77, Med70, Nas75, Pro74b, Sod53, SL78, Abb70, BWZ71, Cam71, Chr73a, Cla77b, Def77, Dun79a, Dun79b, Fik75, HB70, HV74, Hil75, Kai72, KL71, Lil70, Lil71, LP73a, Man71a, Mar74b, Oli72, Pec70, PM72, Rob79, Sab76b, Scr70, SK76, Sib73, Tay70a, Tay70b, War74, Ano79l]. Methoden [Ric76]. Methodology [Wat74b, Wat74c, Coa78, ELR78, Mar76]. Methods [Alb71a, Alb74, Bro79b, Cur78, Ham76, Lew72b, MM70, NV77, O’R70, Pag74, Phi70a, Ram70, WS73, AW71, BS75, Bri79, CL78, Del77, Eck74, Eva77, Jar70, JS71, Lew72a, LB76, MD71, MM71, Moj77, MS70, Phi74a, Whi72b, Zaf71, WS73]. metric [Fle70]. Microcode [Dea79]. Microcoded [BGL^{+}78]. Microcomputers [Mar79a, Wal76c]. Microprogrammable [NR79]. Microprogrammed [BC73]. Microprogramming [Bro74a, WS74]. Microtones [O’B70]. Mikhailov [Mea73]. Milton [Fry71b]. mincing [Rea70]. Minicomputer [Gor78, Wal76c, KK79]. Minicomputers [Wal76c]. Minimax [Tew72, MR76a]. Minimisation [Aki77, Hos71, LoD71, MS70, ZD73]. Minimising [LT79, Pec70]. MINIMOP [MT70]. Minimum [FB75, KL71, Lil70, Lil71, Roh73]. Mintree [Boo78, HO78, Obr78a, Obr78b, Pit78, RT78, WT79]. MINX [Boo79b, PHJ67]. Mitchell [Alb71a]. Mix [FG77, MS74]. Mixed [MM70, NC76, DP73, LW71, MM71]. mixing [Bem74, Mac74b]. Mobile [Wai70]. Mode [Pri78]. Model [AD76, BPR74, Dea74b, GKS79, Sla71b, Dea74a, Lle73, Tur78, WO75, Cor79]. Modeling [Cor79]. Modelling [Huz79, Sch77a, Sma78b, Ano79l, EM71, Fox78, Rob79]. Models [Flo76, RW74, WJ74, BS75, BS79, Don78, Jar70]. Modern [Jen74]. modification [MR76a]. Modifications [Hit78, LS78, SG78]. Modified [BBG^{+}76, DHW78, KL71, Lil70, Lil71, MD71]. Modulo [LT79]. molecular [Bar75a, BJ78]. Moler [Cur78, Hit78, LS78, SG78]. Monitoring [Dea78, Huz79]. Monitors [Wie79]. Monte [KP79]. Moore [WY71]. Moring [HR73]. Morris [HR73]. move [Man71b]. Movements [BK79b]. moves [BJ79, Man71b]. Mp [Mac71b]. Mp/1 [Mac71b]. MU5 [CMRW72, Fra76, Ibb72, IH77, MDFS72]. Müller [YFM71]. Müller-Merbach [YFM71]. Multi [BT71, Bla78, GKS79, Ing74, Lun71, Par78, WV70, Wat77b, WW73, Com75a, Com75b, HW74, MT70, Par74, Van73]. Multi-Access [WW73, BT71, MT70]. multi-computer [Com75a, Com75b]. Multi-Coordinate [Bla78]. Multi-Criteria [Par78]. Multi-dimensional [Lun71]. Multi-Level [GKS79]. Multi-List [Ing74]. multi-member [HW74, Par74, Van73]. Multi-Processing [Wat77b]. Multi-Valued [WV70]. Multicomponent [Sto73]. Multics [Jar75]. Multidimensional [Lun75, RM78, Smi70a].

Multioutput [EH78]. Multiple [Jon72, BSH70]. Multiplier [PB76]. Multipliers [VH72, BVH74, Kro70c, Kro70d, Kro70e]. Multiprocessor [BGL^{+}78, NR79]. Multiprogramming [LY79, Sym71]. Multiway [TW77]. Music [O’B70]. Mutch [Ano72c, Ano72d, Ano73e, Ano74f, Ano73f, Ano79p].

Naive [Eas76]. name [IH77]. NAND [LT79, TD77, ZD71]. NAND-NOR [TD77]. NAND-NOR-XOR [LT79]. Nashelsky [RL73]. National [Ker76]. Natural [Wol78, CCL72, Mor78]. Nature [Wil77b]. Naur [CCH73]. necessary [BS73a]. Nelder [Aki77]. Nesting [DK74]. Network [LT79, Ste77b, Eth71, Jon79, Bar78f]. Networks [Pri78, BP79, Lip79, Mar74b]. Newton [MD71, Med70]. Nimbus [Cro74]. No [Gil78a]. Nodal [Cra72]. nodes [For75]. Non [Bay76, CC74, JW74, Lea77, SD78, BL72, Chr71b, FB75, JW72, Jon70, Smi70b, Tza73, Woo71c]. non-binary [Chr71b]. Non-Deterministic [CC74]. non-linear [FB75, Jon70, Tza73]. non-numerical [BL72]. Non-Parametric [JW74]. Non-Procedural [Lea77]. non-programmer [Smi70b]. Non-Recursive [Bay76, Woo71c]. non-statistical [JW72]. Non-Unique [SD78]. Nonlinear [BRH70, GK73, Gre70, Med70, Mot79a, Sou76, SL78, TI75, TS75, AJ70, Bro69, Fie70, HB70, MR76a, MM78, Mot79b, OW71, PM72, PD79]. Nonmorphological [Hay79, Tur78]. Nonrecursive [Bur75b, Sou77]. NOR [LT79, TD77, ZD71]. norm [SKY73]. Normal [Fry71b, JW74, Loi72, Woo70a]. Notable [LT71]. notation [DZW75]. Note [BB71, Dew72a, Dew74, Fie70, FN78, GE71, HP67, Hil69, HW74, Hos74, Hys72, KL71, Kro71, Lil71, McC72, Mux70, Ove67, Par74, PHJ67, Pit78, Red71b, RL72, SM71, Sal70c, Sal70b, Sod53, Woo70f, Woo70g, vR70c, BS75, BN76, BS71, Cam71, CPR75, Cla77b, Don78, Hay77, HM77, Jor73, Kom74, LW71, LL70, Low77a, Low77b, Pez76, Rei71b, Sou77, Van70, Wod78, Woo71a, Wyn71]. Notes [Ano72j, Ano72k, Val74, Ano78g]. notice [Ano73l, Ano73m, Ano73n, Ano74h, Ano74i, Ano75m, Ano75n, Ano75o, Ano75p, Ano77l]. November [Ano73k]. Number [Ano73d, Chr71a, Hol75, Mit76, PB76, PB78, Rog72, SNH76, SH72a, Smi71, And74, Bea74, BF78, Hon71b]. Numbers [Ful76, Sam72]. Numerical [Abb70, BW70a, CM79, Dur74, Hun72, Lew72b, OP73a, Phi70a, Phi74b, RM78, Sco70, SS72c, SL78, WS73, Wat72c, ZL73, AW71, BL72, BW70b, Cas75, CP71, Hen78, HH79a, KU74, Lew71b, Lew72a, LB76, Pie71, Twi75, UK74, Val71b, WCL71, Wol71]. Numerik [Alb71a].

O [Llo74b, LV79, Sil79]. O.-J [Llo74b]. object [DO72]. Objectives [Par75d]. objects [Far78]. Observations [LR72, PW74b, Sib71]. Obtain [Wil78e]. Occupancy [BW73b]. occurring [GE71, Lit70]. odd [HKA72, Hos74]. off [YFM71]. Office [JBD73a, JBD73b]. on-ground [Tur75]. On-Line [Par75a, CW71, BBG^{+}71, Sma78a]. One [BRH70, Bro74f, Nas75, Ano79p, Chr71b]. one-day [Ano79p]. One-sided [Nas75]. Online [Gil78a]. opaque [Gri73]. Operating [EMM72, Eva78, Fra76, Pri78, SH72b, Var76a, FW72, MT70, MDFS72, SS72d, SS72e, VG73, WHM71]. Operation [Dea74b, Str72]. Operational [Ano73k]. Operations [TCB71, Wre71b, God75]. Operator [Wil79, BW71, HL76a]. operators [ZD71]. Optimal [DO72, Sto72, Bra79c, Cam71, Dun73, Wym73]. optimal-fit [Cam71]. optimality [dCS77]. optimally [Sib73]. Optimisation [DFO79, HH76, Hys72, Sav69, TD77, MD71, OP75, PW74a, Phi74a, Pri77b, Sav70, Tho72]. Optimising [BK79b, SA73]. Optimization [Bar74b, Ram70, YFM71, HB70]. Optimizing [Coo76b]. Optimum [Had71, MS74, Akk73]. Option [AB78]. Order [Boo67b, Fry71b, Hil69, O’R70, Woo70g, Ada76, Cam73, Chr73a, Fyf70, Hop71, HKA72, Hos74, Mac73, Man71a, Oli71, Twi75, Whi72b, Wol71]. Ordered [AK74, Bro79a, Far78, Pag71]. Ordering [Pro74b, SN75]. Ordinary [Lam76, Nor76, WY71, Wat72c, BWZ71, BS71, Chr73a, Tay70a, Wil76b, Wol71]. Organisation [Bar78f, CS77, Moo79, BE79]. Organising [Dea74b, Dea74a, PS74, SD73]. Organization [Edw74]. orientated [Gro73, KL77]. Oriented [Pag76a, Alp71, ELR78, KM74, UK74]. origin [Mul75]. Orion [LT71]. OS6 [SS72d, SS72e]. Osborne [MR76a]. oscillating [Low77a, Low77b]. Oscillatory [AEH75, AEH76]. other [CD77]. outcomes [Par77b]. outer [Bro73b]. Outline [Atk76]. Output [WCH78, BS75, Sal74, SS72e, Tay76b]. Overflow [Goo77, Sam76]. overhead [Abb70]. Overlapping [CW70]. own [Hig76b].

P [Cor79, TCB71, WY71]. package [BLY70]. Packing [Coo72b]. Packs [Cha75]. pages [Cur78, Eva78, Fox74, Fox76, Wil70]. Paging [BCR72, Jos70, Hoa73, LKK71, Wil73]. Pak [Wil76a]. paper [Ano73a, Sib71]. Papers [Ano72k, Ano77j, Ano79i, Ano75e, Ano75f, Ano78e, Ano79h, Sma78a]. Parabolas [Smi71]. Parabolic [MM70, SB70, KS71, MM71]. Parallel [Arm78b, DFO79, Jel79, Lea78, Lis74, RW74, VH72, Woo77a, BE79, BL72, BVH74, ES77, War74]. parallelism [DL76]. Parameter [BRH70, Cra72, Jon70]. Parameters [Bir79, Hur73, Lew71b]. Parametric [Gro71, JW74, HK72, Kro70e, McC71b, SM71]. parity [Day77]. parser [Bar75a]. parsers [Ter73]. Parsing [Wil79, Bar75a, BJ78, HL76a]. Part [BJ78, Com75b, Rey71b, Rey71a, OS70, Ree79d, Bar75a, Com75a, JBD73a, JBD73b, OS71b, Ree79c, SS72d, SS72e, Woo70b]. Partial [Alb71a, Sch77b, SB70, SN75, DW73, Hop71, KS71, Twi75, Zam74]. Partitioning [Bel70b, Stu71]. partitions [FL79]. PASCAL [Sil79, Tan78]. passing [Coo72f, Ecc71b, McC70b]. past [Nor78]. PATCOSY [Ker76]. Patient [JO74]. Pattern [Bra78b, BKM78, Ull73b]. Pattern-Based [BKM78]. Pavlidis [Bra78b]. payroll [WF78]. Performance [BPR74, Cla77a, Dea78, KSD77, Don78, Fox78, HPLG79, Dak73b]. performances [Cro75]. Period [Ful76, Eck74]. Periodic [For75, HK72, BE73, HKA72, Hos74]. PERM [Boo79a]. Permanent [SH72b]. Permutation [LoD70, Roy78, Fik75, OS70, OS71b, Woo77b]. Permutations [Roh76, Roh78, Var79]. personal [Ano76f]. personnel [BG71]. Ph.D. [HL76b, Lau75]. Phase [Lyn72]. Phrase [Has74, Hou71]. Physical [Sch77a]. Picture [Duf72]. pipeline [Ibb72]. pivoting [Zam74]. PL [Cuf72, DO72, Lim74, Mer70b, Val74]. PL/1 [DO72, Mer70b]. PL/I [Cuf72, Lim74, Val74]. place [Hig76b]. placement [LoD71]. Planar [LR78]. Plane [GS78, GS79, Rei71b]. Planning [WJ74]. plants [KM74]. Play [O’B70, Bel70a, BB71]. playing [Bra79c]. plot [Roh74]. plotter [Gri73, LN79, McC71a, McC72]. Plotters [Mau73]. Plotting [Cra72, Mea73]. Pocket [Wil78b, BL77]. Point [SA79, McC71b, SM71]. Points [Cra72, GS79, McL74, Smi71, Coo72f, Ecc71b, McC70b, McC71b, SM71, Sal72, Sal73b]. Poisson [Lew71b]. policies [Nor78, Web72]. Policy [Ano75a, Ano76e, Ano77d, Ano79e, She73a, She73b, Bar73a, Day78a, DL76, GT72, MT72, She72b, She74, She79]. polygon [SA79]. Polynomial [GH73, GH75, BJS74, Bra75, MR76b, OP75, Var77]. Polynomials [Han70, OP73a, For72, Van70, Fox76]. Polyominoes [Lun75]. polyphase [Cam73, Mac73]. popularisation [Pen77]. portability [Tri78]. Portable [HH79b]. positions [Ano79q]. positive [Jen71, Zam70]. Post [JBD73a, JBD73b]. postage [Lun69, Sel72b]. postfix [DZW75]. postman [Bar78a, Gol78a]. Power [BY79, HD72, Nor76, BS73a, Gor79]. Practical [Eat74, ON79, Bro73b]. PRBS [Jac71]. Precedence [Has74, Wat76a, Wil79, God75, HL76a, LL70, Wat77a]. Precision [Aki77, GH75, SH72a, Hon71b, Hun75]. Predicting [Maz78]. preferential [HW74, Par74, Van73]. preliminary [CO72, Phi74a]. Prentice [Cur78]. Prentice/Hall [Cur78]. preprocessor [Bur79]. Prescheduling [FG77]. present [Nor78]. Presentation [Bla78]. Preserving [LO75, Wil72]. Press [Fox74]. pretreatment [Van70]. Previously [Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a]. Pricing [BN78, SZ74, Zmu75a, Zmu75b]. primal [Kro70b, Kro70c, Kro70d]. prime [Hil79a]. primitive [Sil79]. Princeton [Wre72]. Principles [Col76, Kit74, Leh74, SS72d]. Prior [FG77]. Priority [NT79]. Privacy [Gol73]. Prize [Ano73f, Ano73e]. Probabilistic [JW74]. Probabilities [Fry71b]. Probability [SR76, McC78, Sma78b]. probable [Par77b]. Probing [Krá71]. Problem [Jon71c, KL77, PW70, AO79, Bar78a, BN76, BP79, ELR78, Gol78a, Hay77, Hop71, Lig73, Lun69, Mot79b, PM72, Sch76, Sel72b, Wod78, Zak77]. Problems [BW70a, BS73b, Edm74, Lor76, McC70a, Pro74b, Sco70, Wat72c, Zve75, AJ70, BJS74, Bra75, Cas75, Chr73b, Eva77, Fyf70, Gil73, JW72, MM78, PW74a, Tay70a]. Procedural [Lea77]. Procedure [GE71, Lit70, Weg74, Kai72, McC71a, McC71b, McC72, PFW74, Pri77b, PD79, SM71, Woo70c]. Procedures [Dew71, Dew72a, Dew74, Gun70, Jon72, KJ75]. Proceedings [Wre72]. Process [Bar72c, BPR74, Ger70, HR74, SJ79, Wal76c, Boo74a, Boo74b, Com75a, Com75b, FH77, Hos71, Rea70, Shr75]. processable [War74]. Processed [MS74, Wat71a]. Processes [Cla78, Lis74, NW74, RW74]. Processing [Atk76, CBB^{+}76, DFO79, FG77, GS73, Ham74b, Jen74, Lan74b, Lea77, OBLP75, Par78, RC70, SR76, Ste77a, Wat77b, AF76, Alp71, Ano75l, BL72, BLY70, ES77, HN79, Hat76, Lee78a, Mer70b, Nau79a, Nau79b, Par77b]. Processor [Jel79, Lea78, DW77, Fos71, LLR73, Mac70b, Tes77]. processors [CP79]. procurement [Web72]. Product [Har76]. production [LP73b]. Products [LT79]. Professional [HL76b]. Program [Bar74b, BD72, BK79a, BKL74, HW77, Jos70, LO75, Mae78, Mot79a, Smi72, Bel70a, BB71, BW70b, BM79, Bro74h, BW75b, BSH70, CHH73, FH71, HK73, Hoa72, Jon79, Kro70b, Kro70c, Kro70d, Kro70e, LW71, LLR73, LB76, LMS73, LP73b, MWC70, MM78, Mot79b, Pan70, SAHS71, Tow77b, WD79, Wig75]. programme [Ano74j]. Programmer [Bel76, Col75, Smi70b]. programmers [Fis76]. programmes [Mar79a]. Programming [Ano75f, Arm78b, Bel76, BS73b, Col76, Coo72c, Ear75, Fle75, Ger70, Ham76, Ing76, Lim74, LK76, Mot79a, NF76, OBLP75, Pag76a, Par75d, Ree76b, Roh76, Wai70, WD74, Wil74, AJ70, BJS74, Bra75, Bra76, BM79, Bro73c, Chr71b, Coa78, DI78, Eme78, FB75, Hig76b, Hol78, Hop78, LP73a, Maz78, Mol71, MM78, Mot79b, OL78, PW74b, Wym73, Wre72, YFM71, Bro74d, Cor79, TCB71, Llo74b, Eme78]. Programs [Bir74, Cha75, HV78, Ham77, JP76, LL75, O’B70, SS75, Bar74a, BF72, Bra79c, Bra74, DO72, Dun78, LM77a, NI70, Osb71b, RT76, SAHS71, War74, WF78]. Progress [WY71]. project [Bro73a, Fis74b, Hon74, Par77b]. Proof [Der79, FH71, Hoa72, Smi72]. proofs [Bra77]. Properties [Krá71, Hil79a, SW78]. PROPLAN [Pen77]. Proposal [Ste77a]. Proposals [HKT75, Lau75]. proposed [DCHR76]. protected [FKLM76]. Protection [NW74]. Proteus [BGL^{+}78]. Prototype [Sta77b]. provide [McC71b, SM71]. Proving [SN75]. Pseudo [Ful76, PW70, Hil79a, Var77]. Pseudo-Inverses [PW70]. Pseudo-Random [Ful76, Hil79a]. pseudo-remainder [Var77]. published [Dun78, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a]. Purchase [AB78]. Purpose [ASG79, PC78, Bri79, Hil80, LMP79, MR82, Rob80, Tay80]. purposes [WO75]. put [Bro73c].

Q [Dun79a, Dun79b]. Quadratic [Bur75a, HD72, Loi72, Eck74, Tho72]. quadratically [MS70]. Quadrature [AEH75, Haz78, LK77, Oli72, Pie71]. Quadratures [Cha70]. quadric [Sab76b]. Quality [WCH78, All71b, Jon71b]. quantitative [Hun75]. quantities [Ada76]. quasi [Coo72f, Ecc71b, McC70b]. quasi-intrinsic [Coo72f, Ecc71b, McC70b]. ‘Queensway’ [Hon74]. queries [BJ78]. questions [BSH70]. Queue [NT79, Tur79]. Queueing [Par75a, Lip79]. Quicksort [FH71]. quindiagonal [BE73, Eva75]. quintic [Hos70, Sal70b, Spä69]. Quotient [Bur75a].

R [Alb71a, Bro74d, BS73b, Cor79, GS73, Llo74b, NLW74, RL73, Wre71b]. Rademacher [Hur73]. Rademacher-Walsh [Hur73]. Radford [NLW74]. Radix [Sod53, Boo74a, Boo74b, Ham74a, Woo70f]. railways [Abb70]. Ramon [WY71]. RANCOMB [Pik79]. Random [Ful76, McL76, Rea70, Bea74, Hil79a, Jac73, Kol78, Pit74, Pri77b, Ree79c, Ree79d]. randomness [SL73]. range [HP77]. rank [Boo67a, Hil69, Woo70g, Fry71b]. ranks [Woo70d]. rapidly [PM72]. rates [BF78]. Rational [Vit75, Whe78]. rationale [WCL71]. Ratios [Wat76b, Pez76]. Re [BW70a, FS75, Ada76]. Re-Entrant [BW70a]. re-order [Ada76]. Re-Set [FS75]. Reactor [McC70a]. read [BJ79]. readers [BN76]. Reading [Art75, HV74]. Real [Bar72c, BD72, BC73, BKL74, Han70, Sha72a, Wil77d, CJ70, DO72, Kai72, Sea71, Sto72]. Real-Time [BD72, BKL74, BC73, Sto72]. realistic [Ben72]. received [Ano75d]. Recoding [Mor78, Wol78]. recognition [War74, Bra78b, Ull73b]. reconsideration [Lee78b]. Record [Bae75, JO74]. Record-Classes [Bae75]. Records [SD78, Art75, Tri74]. Recovery [Gun70, Lee78b]. Recreating [Bay76]. Rectangular [Har76, AO79, BW73b]. recurrence [Not70, Pag71]. Recursion [Bir79, Bra77]. Recursive [Bay76, Lyn72, TCB71, Eva75, FH71, Woo71c, Woo71b]. Recursively [Has75]. reduce [Aus71, Har79]. reduced [Zam74]. Reducing [Mal77]. Reduction [JW74, Tho72]. redundancy [Bro74h, HK73]. referee [BW75b]. Refinements [LO75]. reflections [Lav70]. Regenerative [Aki77]. Regional [Bar78f]. Regions [Kam77, LR78, Rei71b]. Registers [GK73]. regular [Sal74, SY72]. rejection [PFW74]. related [Eck74, FL79]. Relational [CO76]. Relations [Has74, MP79, Pag71]. relationship [Wil78f]. relationships [Kin69]. Relative [Wil78e]. Reliability [BK79a]. reloading [Aus71]. remainder [Var77]. Remark [Obr78b, Sut76b]. Remarks [HO78, Var76b, SS72b]. remote [BBG^{+}71, Har70]. removal [Ric71]. removes [Par75e]. Renal [Gor78]. repetition [HN79]. Replacement [CO76, TW77]. replenishment [Ada76]. Report [BBG^{+}76, DHW76, Sma78a, Nor78, Gol73]. Reports [Ano75l]. Representation [Bay76, BKM78, Mae78, CCH73, FL79, Sab72, BS73b]. Representations [Flo76]. Reprinting [Ano78g]. reproduction [Dun78]. Requirements [PDY74, Zam74]. Research [Ano73k, TCB71, BMPS71, JBD73a, JBD73b, WY71, Wre71b]. Reservoir [TW77]. residues [Hil79a]. Resilience [Dea76]. Resources [SZ74, Zmu75a, Zmu75b]. response [GE71, Lit70]. Restart [SH72b]. resulting [Ano73k]. Results [SR76, Sto73, Bra76]. Retrieval [HR74, HM70, Mal77, WK77, Gob75, Smi70b, SW79, vR71]. Retrievals [BK79b]. return [BF78]. reversal [Ull77]. Review [Alb70, Alb71b, Alb71c, Alb75, Alb76, Ale77a, Ale77b, Ale78, And75a, And77a, And77b, And78, Ano73l, Ano73m, Ano73n, Ano74h, Ano74i, Ano75m, Ano75n, Ano75o, Ano77f, Ano77g, Arm78a, Bac77, Bac78, Bal79, Ban79, Bar76a, Bar77a, Bar79, Bar70, Bar74c, Bar75b, Bar77b, Bar77c, Bar78b, Bar78c, Bar78d, Bar78e, Bar73b, Bea77, Bel71, Ber74, Ber73, Bes79, Bil77, Bla75, Bla79a, Bla79b, Bla79c, Boa79, Bra78a, Bra79a, Bra79b, Bra78c, Bro74b, Bro74c, Bro74e, Bro77, Bro74g, Bro75a, Bro75b, Bro76b, Bro70, Bul73, Bum78, Bum79, Bur70a, Bur70b, Bur71, Bur73a, Bur73b, Bur75c, Cai76, Cha73a, Cha77a, Cha77b, Cha78b, Cha79a, Cha79b, Che71, Cla76a, Cla76b, Cla79, Cla76c, Cle75, Cle77, Clo77, Clo79, Clu77a, Clu77b, Coa77, Coh70, Col78, Col77a, Col77b, Con76, Con77b, Con77a]. Review [Coo71, Coo76a, Cra71, Cro77, Cru78, Cun77, Dav77, Day73a, Day73b, Day75, Day78b, Day79, Den75a, Den75b, Dew75, Dic75, Dir70, Div74, Duf73, Duf74, Duf78, Duf79, Eas75, Eat72, Edm73, Edw77c, Edw77a, Edw79, Eme78, Eme79, Eme75, Eme76, Eva70, Eve71, Eve74, Fel71, Fel78a, Fel78b, Fer74, Fis79a, Fis79b, Fit74, Fla75, Fla77, Fle73a, Fle73b, Flo70a, Flo70b, Flo73a, Flo73b, Flo74a, Flo74b, Flo79a, Flo79b, Flo79c, Fox74, Fox76, Fre78, Fry71a, Gar73, Geo70, Geo78, Gil78a, Gil78b, Gil79, Gir73, Gir76, Gle79a, Gle79b, Gol75, Gol77a, Gol77b, Gol77c, Gol78b, Gol73, Goo70, Gou70a, Gou70b, Gri79a, Gri79b, Gur78, Gur75, Had75, Hal71, Hal74, Hea77a, Hea73, Hea76a, Hea76b, Hea77b, Heb79, Hig77a, Hig77b, Hil73a, Hil74, Hin79, Hio70, Hio78a, Hio78b, Hoa70, Hoa77, Hol77]. Review [How75a, How76a, How76b, Hun77, Hun79, Ing75, Ing77a, Ing77b, Jac74, Jac77, Jac79a, Jan77, Joh79, Kem75, Kha75a, Kha75b, Kha76, Kha77, Kha78, Kha79, Kin75, Kir73, Kir74a, Kir74b, Kir75, Kir79, Kit77a, Kit77b, Kit79, Kno75b, Kno77, Lan74a, Lan74c, Las73, Las79, Lav74, Lav78, Lea79, Lee79a, Lee79b, Lee79c, Lee75, Leh75, Leh76, Leh77, Lew73, Lew76, Lew79, Lid74a, Lid74b, Lim75, Lin73, Llo70, Llo74a, Llo74b, Los77a, Los77b, Mar74a, Mar79b, May79, May74, May76, McK79, Mee73, Mer75a, Mer75b, Mer75c, Mil79, Mit70, Mit77, Mor76, Mor79, Mor75, Mor78, Moy78, Mux74, Mux76, Mux77a, Mux77b, Mux78, Nee74, Nut78, O’C79, Oll75, OS71a, OS75, Pac77, Pag79b, Pai78, Pak73, Par79, Par75b, Par77c, Pau79, Phi70b, Pot78, Pow73, Pow74, Pow76, Pri70, Pri77a, Pro74a, Qui79, Rad73]. Review [Rad75a, Rad75b, Rad75c, Rad75d, Rad76, Rad79, Ran71a, Ray79, Ree73, Ree74, Ree76a, Ree76b, Rei71a, Rei73, Ric78, Ril77, Ril78, Ril79, Riv79, Rog74, Rog75, Rog76, Rog79a, Rog79b, Rzv79, Sam73a, Sam73b, Sam73c, Sam74a, Sam74b, Set79, Sha77a, Sha77b, She77, Sim79a, Sim79b, Sla70, Sla71a, Slo78, Smi73, Smi75, Smi76, Smi77, SK78, Smi79, Sno78, Sta77a, Ste75, Stu73, Suf79, Sum76, Tay73a, Tay76a, Tho73, Tho74a, Tho74c, Tho74b, Tho79a, Tho79b, Tho79c, Tho79d, Tom71, Tow77a, Ull73a, Vaj70, Vaj76a, Vaj76b, Vic79, Wal73a, Wal73b, Wal74a, Wal75a, Wal76a, Wal76b, Wal78, Wal79a, Wal79b, Wal79c, Wal73c, Wat74a, Wat71b, Wei79, Wel73, Wel75, Wel76, Wel74, Wel77, Wel79, Wes71, Wes74, Whe79, Wic78, Wil75, Wil78a, Wil70, Wil76c, Wil76d, Wis74, Wis75, Wis78, Wit79a, Wit79b, Woo74a]. Review [Woo78a, Woo78b, Woo74b, Woo76, Wre71a, Wyv77, You70, You71, You72, You74, You75a, You75b, You77a, You77b, You77c, You78, You79a, You79b, dR77, BF72, BL71, Bra78b, Bri79, Cha73b, Cur78, Eva78, Fry71b, Kit74, Sla71b, Sma78a, Ull73b, Web70, Wre71b]. reviewed [Ano79g]. Reviews [Goo74, Hyn73, Tay73b, Alb71a, Bro74d, BS73b, Cor79, GS73, HR73, IL74, NLW74, RL73, TCB71, WS73, WY71, YFM71]. Revised [DHW76]. ring [BLY70]. Ringing [Tre70]. Rivlin [Fox76]. Road [Bei72, Bri79]. Robert [Wil70]. Robot [Arm78b]. robots [BMPS71]. Romberg [Sha72b]. Ronald [HR73]. Root [SH72a, Hon71b]. Roots [Han70]. Rose [RL73]. Round [YFM71, Ano79j, Cha71, Cha72b, Dew72c, Fla72b]. Round-off [YFM71]. Routines [HCC74, Lyn69, GL70, LK77]. routing [BP79]. Roy [Fry71b]. Rt [ON79]. rule [PW74b, Wym73]. rules [Moj77]. Run [NC76]. Runge [Whi72b, CL78, O’R70]. Rutishauser [Alb71a].

S [Ano75g, Cor79, WS73]. safety [Sil79]. Sample [Fry71b, LMP79]. sampling [Pag76b, PFW74]. satellite [Tur75]. Saving [ET75]. scale [BPSS72, Tem72b]. Scatter [Krá71, Mal77]. scenes [Ric71]. Schade [WY71]. Scheduling [FG77, LR72, Wal77, Chr73b]. scheme [Ada76, Ano79f, Coo72f, Ecc71b, Fra70, Lee78b, McC70b, SR74]. Schemes [CM79, Ahm72, And74, Had71]. school [Aus76]. Schultz [Bro74d]. Schwarz [Alb71a]. Science [Cla77a, Fis74b, Hoa76, Ree76b, Lau75, Mar79a, Bel74, Bro73a, HL76b]. Scientific [ET75, NF76]. Scientists [Sma78b]. Scotland [Ano75f]. SCR [Cro74]. Search [SD78, Boo79c, Chr71b, Eck74, Hil79b, Mic71, Ove67, PHJ67, PP79, Pri77b, Wil78f, dCS77]. Searching [Bur75a, Wil74, Kol78, PH73]. Second [Ano76k, Sla71b, Sum78, Twi75]. Secondary [BK79b]. section [Boo67b, Boo67a, Hil69, Hil79b, Woo70g]. Security [GKS79, NLW74]. Seek [SR76, Kol78]. Seeks [Wat75b]. segments [Sab72]. SEL [Mol71]. seldom [Gil73]. selecting [Wym73]. Selection [Bel76, TW77, Wal76c, Lew71b]. Self [Aki77, Dea74b, PS74, SD73, Wat75a, Dea74a, Kro70e, LLR73, Mol71]. self-dual [Kro70e]. self-extensible [Mol71]. Self-Indexing [Wat75a]. Self-organising [PS74, SD73, Dea74a]. self-test [LLR73]. Semantic [Der79, Gol78c, Wil78c, BD79]. Semantics [Pag79a]. Semaphores [KRR79]. Semi [Coo75, CL78]. Semi-Hybrid [Coo75]. semi-implicit [CL78]. seminar [Ano79p]. Sensitive [Ran74, Gha72]. sentential [Tur78]. Sequences [YT79, Far78, Hil79a, OS70, OS71b, Pag71, Woo77b]. sequencing [LN79]. Sequential [Hai78, Huz79, NV77, PL74, YT79, HD71, KJ75, Kol78, Sal74, Tza73]. Sequentially [MS74, Wat71a]. Series [Hun70, Nor76, Ric74, Sla71b, AEH76, BBG^{+}71, BWZ71, BS71, KS71, Scr70, Sne70, Wil76b]. servers [Lip79]. Service [AF76, Wie79, Ker76]. Services [BN78, HG71]. Servomechanism [Coo75]. Set [FS75, GS79, Ben72, Coo72f, Ecc71b, Far78, McC70b]. Sets [Bro72b, Ham77, KRR79, Nor76]. Shan [WS73]. sharing [DL76, Lip79]. Sharp [Bro74d]. Sharpe [Web70]. Shell [GR75]. Shift [Fry71b, GZ72, GK73]. Shifts [PB76, PB78]. Short [Ano73l, Ano73m, Ano73n, Ano74h, Ano74i, Ano75m, Ano75n, Ano75o, Ano77l]. Shortest [YT79]. SHRDLU [Kli75]. Sibson [CW70, Roh75]. sided [Nas75]. Sidney [Gea70]. sieve [Hoa72]. Sigmod [Ano76l]. Sigmod-77 [Ano76l]. significance [Tri76]. Siklossy [BS73b]. silent [CT74, Fis74c, Gea73]. similar [BLY70, Web72]. Simon [BS73b]. SIMPL [Pal73]. SIMPL/I [Pal73]. simple [Cam73, CS75, Mac73]. Simplex [Aki77, Kro70b, Kro70c, Kro70d, Kro70e, PD79]. simplification [Fit73, Jar70, Wil77e]. Simula [KM74]. Simula-based [KM74]. Simulation [Bei72, BPR74, BCR72, BKL74, Cor79, Eat74, GS73, KP79, Lew78, Nel79, Pag74, PC78, Pri78, Sto73, Bra74, Bri79, BPSS72, Cro75, GP73, KM74, KK79, Pag76b, Pal73, Per76, Sma78b, Tem72b]. Simultaneous [OP73b, Bro69, Chr73a, CJ70, Fie70, Jen71, PD79, Sea71]. Single [BW75a, SH72a, SL78, Hon71b, Kro70d, Mar73, Sib73]. single-link [Sib73]. Singular [Nas75, Sha72b, Tay70a]. Singularities [RM78]. Size [HH76, HD72, LV79, O’R70, Sam76, Whi72b, WCL71]. Sizes [Edw77b, Sto72]. skip [NI70, Osb71b]. Skyline [Ham74b]. SL5 [Han78]. SLINK [Sib73]. slopes [McC71b, SM71]. SLPL [Atk76]. Small [Bei72, Pie74, BHRS70, CW71, GP73, Hil77, SS72d, SS72e]. Smith [WY71]. smooth [Coo72f, Ecc71b, McC70b]. Smoothing [Med70, Hun72]. SOAP [SAHS71]. Society [Ano74c, Ano73k]. Software [Bro79b, Rin75, Coa78, Hen78, HH79a, HP77, Jac79b, LP73a, MR77, Mar76]. solids [Gri73]. Solution [Bro69, CM79, Fie70, GH73, GH75, Jen71, Lam76, McC70a, Pro74b, Sco70, SS72c, SL78, AO79, AHM74, And75b, BJS74, Bar78a, BWZ71, BW71, BE73, Bra75, Cas75, Cla77b, DA77, EA70, Eva71, Eva72, EH76, Fyf70, Gol78a, Had71, Hop71, HM77, KS71, Lew71b, Man71a, Øst79, PW74a, PM72, PD79, SK76, Sne70, Twi75, Wol71]. Solutions [Nor76, Tew72, Zam74]. solver [Del77]. Solving [Chr73a, Phi70a, SB70, SG78, TS75, AJ70, Eva77, Hit78, LS78, MM78, Wil76b]. Some [ASG79, Bay73, Bro74i, Edw78, FR79, Krá71, LM77b, LR72, Pie71, Sib71, SR76, SW78, Var76b, Zaf71, AL72, Bra76, Hil80, MR82, Rob80, Sab72, Sal72, Tay80]. Sons [Wil70]. Sort [Boo67a, Hil69, TW77, Woo70g, Cam73, Low77a, Low77b, Mac73, Woo70d, Woo70f, Woo71c, Woo71b]. Sorting [Wil74, GR75, Woo70c]. Source [Pie74, Per76]. SP [Mac70b]. SP/1 [Mac70b]. Space [Har76, MM70, Sch77a, Had71, MM71, Tay70b]. spaced [GM72, OP71, Twi75]. spanning [Roh73]. Sparse [Bro72b, Bra74, Cla77b, Hat76, SK76, RL73]. Special [Ano74j, Ano75p, LMP79, Mot79b]. specialised [HT78]. Specification [Wri74, Dun79a, Dun79b]. Specifications [Sta74, Wat79]. specified [Cra72]. Specifying [Cla78]. SPECOL [Smi70b]. spectral [Hur79]. Speed [Dea79, HH76, PB76, PB78, Ahm72, Ahm73, IWT73]. Speeding [Bir74]. speeds [Fos71]. spelling [Lee74, Mar74d]. SPG [OBLP75]. Spiral [Jon70]. Spline [Man74, San75, For75, HK72, Bro74d]. splines [Cox71, Fyf70, Hos70, HKA72, HK72, Hos74, Sal70b, Spä69]. Square [SH72a, GE71, Hon71b, Lit70]. squared [PFW74]. Squares [PW70, TI75, Del77, Jon70, Pec70, PM72, ZL73]. SST [Tow77b]. Stability [JW72, Tay70b]. Stable [Rod70, Zaf71]. Stack [Goo77]. stamp [Lun69, Sel72b]. Standard [Ano76j, Ros70, DCHR76, MRW75, Tri76]. Standardised [Coa78]. Standards [Wie79, HL76b]. State [Day70, FS75, Sma78a, Tza73, Woo77a, CW71, Lan71b, Lle73, Tow77b]. Statement [Ano75a, Ano76e, Ano77d, Day78a, She72b, She73a, She73b, She74, She79, Gal73]. Statements [Ing76, DI78, Hop78, Sal71b]. Static [Wil78c]. statistical [JW72]. Step [O’R70, Whi72b, Chr73a, DHW78]. step-until [DHW78]. stereo [Ort71]. Stewart [Fox74]. Stiefel [Alb71a]. stiff [Cas75]. stochastic [Tza73]. stock [Ada76, AO79]. Stoer [YFM71]. stopping [Moj77]. Storage [BK75, Bro79b, CO76, Krá71, Mal77, Rob77, Sch77a, Ste77b, Wol78, Cam71, CPR75, Fos71, Mor78, Pez76, SR74, Zam74]. Store [DK74, BJ79, IH77, Ree79c, Ree79d]. straight [Sab72, Sal70a]. Strategies [CS77, Rob77, Bra79c, WLDC71]. strategy [Lan76b]. Streams [Cla78]. Strema [Cla78]. String [Gre70, Hou75, Rey77, Mac70b]. Strings [MJ75, CCL72]. structural [VG73, Bra78b]. Structure [Has74, Mil70, SH72b, dRC74, BBH70, Ben70, CS75, Dak75, FKLM76, KM74, Mar73, MP79, MDFS72, SS72d, Tri74, Wil78d]. Structured [DI78, GKS79, Hop78, Ing76, LL75, LK76, Wil77b, WO78, Eme78, Hoa72, Hoa73, Kol78, Lem79, OL78, Pin77, Wat77a, WP75, Llo74b]. Structures [HC78, Ing74, Mae78, Eth71, Hou71, SR74]. Structuring [Bay73, NS77, vR71]. student [Ano73a, Lem79, Maz78]. Students [WS73, GS73]. Studies [BY79, KP79, Pri78, Lee78a]. Study [Ano76a, Lip79, Nel79, CO72, GR75, HN79, SW79, Tri78]. subarrays [Smi70a]. subject [AJ70, HB70, LP73b]. Sublist [NT79]. Submission [Ano72k]. subroutines [Dun73]. subscripted [HV74]. subscripts [Boo67b, Hil69, Woo70g]. subset [Var79]. Subsets [Wil78e]. substitution [Ull77]. Substring [WK77]. subsystem [MRW75]. Subsystems [Fen74]. success [Maz78]. such [PC78]. SUD [Ree79a]. sufficient [BS73a]. Suggested [Ehr72b, Fin72a, Fin72b, Mee72b]. Suitable [FG77]. suite [LP73b]. Sum [LT79, Pec70]. summary [Nor78]. Summation [Edw78]. Summer [Ano76b, Ano74j]. SUMT [Vit75]. Supplement [Ano70d, Ano70b, Ano70c, Ano71a, Ano71b, Ano71c, Ano72a, Ano72b, Ano74b, Ano77b, Ano77c, Ano78b, Ano70a, Ano73b, Ano73c, Ano74a, Ano75a, Ano75b, Ano75c, Ano76e, Ano76c, Ano76d, Ano77d, Ano77a, Ano78c, Ano78d, Ano79e, Ano79a, Ano79b, Ano79c, Ano79d, Day78a, DHW76, Ham70a, Ham70b, Ham70c, Ham70d, Ham71a, Ham71b, She71, She72a, She72b, She73a, She73b, She74, She79]. supplement-Statement [Ano76e, Ano77d]. support [Bur79]. supporting [FH77]. Surface [Wil77d, FR79]. Surfaces [Mac70a, Sab76b]. survey [CHH73, UK74]. Swann [Hos71]. Switching [TD77, WV70, ZD71]. Symbol [Day70, Lan71b, Tow77b]. Symbol-State [Day70, Lan71b]. Symbolic [Har79]. Symmetric [GZ72, CJ70, Eva71, Kai72, Sea71, Zam70]. Symmetrical [Has74]. Symmetrischer [Alb71a]. Symposium [Wre72]. Synchronisation [SJ79, Shr75]. Synchronous [SS72c]. Syntactic [Bar75a, BJ78, Cam78]. Syntax [Bur72, HR70, Mux73, Alp71, Bar75a, Bro73b, HP71, Wil76b]. Syntax-Directed [Bur72, Mux73]. synthesis [BJ78, Hur73, Hur79, Win75]. synthetic [CW76, HPLG79]. System [BY79, BPR74, BW78, BCR72, Bro74a, BGL^{+}78, CDZ74, Cra72, Dea74b, Dea78, Fra76, Gor78, Har79, Ker76, KSD77, LY79, LV79, OBLP75, Sha72a, SH72b, Sta77b, Ste77a, Var76a, Wai70, Wil76a, Zmu75a, Zmu75b, Abb70, BBG^{+}71, BBF70, BBH70, BL72, BG71, BPSS72, Chr73a, Cla77b, Col70, DP76, Dea74a, Dow71, Eth71, FKLM76, Fox78, Gob75, Hoa73, KU74, KK79, Lan76b, LN79, LMP79, MT70, MWC70, MDFS72, Ort71, PR77, Pin77, SS72d, SS72e, Tem72b, Tur79, VG73, WHM71, Wil76b, Bea74, WW73]. Systematic [Pag71]. Systems [Cla77a, Col76, Edw74, EMM72, F^{+}73, FS75, Gun70, Hai78, HCC74, Hil76, Leh74, Lew78, Lis74, NLW74, NR79, Par75a, Par78, PDY74, SG78, Ste77a, TCB71, TS75, WJ74, Wat74b, Wat74c, AHM74, Ano79p, BWZ71, BL71, BE73, Com75a, Com75b, EA70, Eva71, Eva72, EH76, FW72, Gha72, GE71, Hit78, HM77, Jen74, Lan76b, Lip79, LS78, Lit70, Mea73, Mic79, Øst79, Pen76, Pen77, PS74, SK76, Sne70, SD73, Sto72, Tho72, UK74, WS74, Zam74, BC73, Eva78, GS73, IL74].

T [Bra78b, IL74, Wil70]. Table [Bur75a, HD72, Sam76, Wil77a, Car70, Day77, Kin69, Kin71b, PH73]. Tables [AK74, Bay73, Day70, Mae78, SD78, SS75, WL72, KJ75, Lan71b, Sha72b, Tow77b]. Talbot [YFM71]. Tape [Edw77b]. target [CMRW72]. Task [Sta74, BM79]. Tasker [Gea70]. tax [God75]. taxonometric [WLDC71]. Taylor [BWZ71, Wil76b]. Teaching [AD76, McG77, Ano79q, Bow73, Lee78a]. Tears [Lin72a, MH77]. Technique [Bay76, Krá71, Sav69, Wil77d, Cox70, FB75, GR75, LK77, Mea73, NI70, Osb71b, Sav70, SC71, Sza73, Ull77]. Techniques [Bay73, Coo75, Hys72, Mal77, SH72a, SB70, Ull73b, AL72, Bra74, CO72, Hon71b, Jac71, PH73, Wal76d, Wym73, BS73b, Cha73b, Kit74]. Technology [Hoa76, Ano74j, RL73]. telecommunications [Ano73k]. Tensor [Har76]. term [EA70, Eva71]. Terminal [Ste77b]. Ternary [BVH74, GK73, VH72, Hur79, SW78]. Tessellations [GS78]. Test [Cha75, YT79, Ben72, LLR73]. testbed [Hen78, HH79a]. Testing [Ham77, Sam76, Wil70, Day77, KJ75, LLR73, LMS73]. Text [Lea78, Zve75, Col70, Gob75]. Their [Hou75, OP73a, BF72, Fis76, Hoa76, RL73]. Theodore [Fox76]. Theorem [SN75, Woo70a, dCS77]. Theorem-Proving [SN75]. Theories [SN75]. Theory [Bar74b, Jel79, Par75a, Rog72, SR76, TCB71, WY71, Bra76, Hig76b, JW72, Lin72b, Woo70b, YFM71]. thesis [Lau75]. thinning [Deu69, Fra70]. Three [Bro74f, Wat77b, EA70, Eva71, Had71, McC71b, MM71, SM71]. three-point [McC71b, SM71]. Threshold [Edw78, EH78, Win75]. Threshold-Logic [EH78]. throughs [Lem79]. Time [Bar72c, BD72, Bur72, BKL74, Coo75, LR72, Med70, Mux73, NC76, Sha72a, Wil77d, BC73, ET75, Kol78, Lip79, Sto72, Tur79]. Time-Domain [Coo75]. Times [Mal77, SD78, SR76, Wil78f]. timesharing [HG71]. Timetable [SS79, Akk73, Alm69, BF71, DD70]. timetabling [Aus76]. Timing [Lor76]. titles [Bon71]. Tokyo [BMPS71]. Tokyo-Edinburgh [BMPS71]. Tolerance [JP76]. Tool [Sch77b, KK79]. top [Woo71a]. top-down [Woo71a]. Topics [Phi74b]. Torquay [Ano73k]. Torso [vO78]. Total [SN75]. Towers [Wod78, Hay77]. Traffic [Bei72, Bri79]. training [Col75]. Transaction [OBLP75]. Transcription [Bro74i]. Transferral [BHW78]. Transfers [Huz79]. Transform [MH77, AL72, Pie71]. Transformation [Nas75, PR77, Scr72, Wyn71]. transformation-directed [PR77]. transforming [LL70]. Transforms [Dur74, Hur79, ZL73]. translation [Dak75]. translator [WP75]. translators [Goo75, LP73a]. Transmission [Wol78, Mor78, NLW74]. Transportable [Moo79]. Transportation [HV78, Lig73]. Transposition [Bro74i]. traversal [Sou77]. Traversals [Bur75b]. Treatment [BW70a, TI75]. Tree [CS77, Pad73, Chr71b, Day76, FL79, Roh73, Woo71c, Woo71b]. tree-search [Chr71b]. Trees [Bur75b, WW76, Bae70, ES77, Sou77, SW78]. trends [Coa78]. Trials [Sha72a]. Triangulating [vO78]. Triangulation [LR78]. triangulations [Sib78]. tribute [Ano79p]. Tridiagonal [GZ72, AHM74, Eva72, HM77, Øst79]. TRIES [NS77]. Trigonometric [OP71, OP73b]. trolley [Abb70]. truncation [OP75]. Tube [Ste77b]. Tuning [Wie79]. tunnel [Hon74]. Turing [CD77]. Two [Bro74f, Cra72, Fry71b, GH75, MM70, McL76, Pro74b, Wat77b, Woo77a, Zmu75a, Zmu75b, Sab76b, Woo70c, HP67]. Two-Parameter [Cra72]. two-way [Woo70c]. TWOBYTWO [HP79]. Type [Hig76a, Mid77]. Typeless [Hig76a].

Uncertainties [Sou76]. unconstrained [MD71]. Uncoupling [Fos71]. Underflow [Goo77]. Understandability [Var76a]. Understanding [Eas76, HC78]. unequally [GM72, OP71, Twi75]. Uniform [Mal77]. Unique [SD78]. Unit [BK75, LKK71]. universal [Jor73]. universities [Web72]. University [Ano75f, Ano79q, DD70, Akk73, Alm69, Col70]. unnoted [Sal72]. UNRAVEL [Bro73c]. Unstable [Sco70]. Unstructured [WO78]. Unstructuredness [Wil77b]. until [DHW78]. unwanted [Par75e]. updating [Bro79a]. usage [Bro73b, Leh73]. Use [Day70, DK74, Lee78a, Mid76, OP73a, AEH76, AL72, Bae70, BLY70, Bra74, Cla77b, Col70, DO72, Eva77, FT73, Fyf70, HPLG79, KK79, Lan71b, McC71a, McC72, OP75, WNT73]. User [Eas76, NF76, CT74, Fis74c, Gea73, JBD73a, JBD73b]. users [Bac71b, ER71, Hil72b, Wal71b, Web72]. Using [Aki77, Cla78, Coo75, GW73, NC76, PB76, PB78, SD78, Wel72, WK77, Art75, BS71, Bow73, Bri79, Chr73a, Chr71b, EM71, EH76, Gob75, Gri73, HV74, HW74, Hon71b, HKA72, HK72, Hos74, Jon70, KL71, Kro70c, Kro70d, Kro70e, Lil70, Lil71, Mar74b, Oli71, Pag71, Par74, PC78, Roh73, SH72a, Van73, Woo70c]. Utilisation [LY79].

Validating [Jac79b]. Validation [Lis74, Pad73, Smi72]. Value [BW70a, Nas75, Sco70, Wat72c, Cas75, Eva77, Fyf70, Hop71, Sch76, Tay70a]. Valued [WV70]. Vandermonde [TS75]. Variability [Rin75]. Variable [Bir79, GW73, PB76, PB78, Sam76, Wel72, Art75, CCL72, Fle70, Jen71, Tay76b, Tri74, War74]. variable-length [CCL72, Tay76b]. Variables [Sou76, Had71, Hig76b]. variance [PFW74]. variant [WBB72]. Variations [Woo77a]. Various [Mit76]. VDU [Bro79a]. Vector [Kei79, McL78]. Verdicts [Wat77b]. verification [LB76]. Versus [HH76]. very [Ano76k]. VI [HL70b, Wel70]. via [BHW78, Els79]. Viable [Ear75]. Vibrations [Gre70]. Vickery [Kit74]. view [Ano76f, Mar73, Shr75]. viewing [Ort71]. Viewpoint [Mal77]. Virtual [Bro79b]. Vision [Arm78b]. Visiting [Ano79q]. VM [Don78]. VM/370 [Don78]. vocabulary [Nau79a, Nau79b]. Vol [Gil78a, WY71]. Volterra [Lew72a, AW71, Lew72b]. Volume [Wil74]. votes [HW74, Par74, Van73]. vs [Wym73].

W [Bro74d, Eva78, Fox74, Llo74b, WY71]. waiting [Tur79]. walk [Lem79]. walk-throughs [Lem79]. Walsh [Hur73]. WATFIV [Ree76b]. WATFOR [Ree76b]. Watson [MR76a]. way [Woo70c]. Weber [LL70]. weekend [Ano76f]. weighted [PD79, ZL73]. WG [Ano72i]. Which [O’B70, Par75e, Sal74, SAHS71]. White [Man71b]. widely [Tay70a]. Wiley [Fox76, Wil70]. William [Web70]. Williams [BS73b, Cor79, Sib71]. Willoughby [RL73]. wire [Abb70]. Wirth [LL70]. Within [AD76, BHW78, Ada76]. without [Pec70, Wie79]. Witzgall [YFM71]. wonderland [FW72]. Wooloms [IL74]. word [FT73, Mul75, WNT73]. words [Ull77]. work [HL76b]. Workshop [Ano76l]. Worst [Rob77]. writer [BN76]. Writing [BW78, All71b, HV74, Jon71b].

XOR [LT79]. XVII [TCB71].

Younger [Gol73].

Zero [Bre71, Chr71b, Cox70, SC71]. Zero-one [Chr71b].

References

Albus:1979:C

[AAB^{+}79] J. Albus, R. H. Anderson, J. M. Brayer, R. DeMori, K. S. Fu, S. L. Horowitz, B. Moayer, T. Pavlidis, W. Stallings, and T. Vamos. Correspondence. The Computer Journal, 22(2):190–191, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

         URL http://comjnl.oxfordjournals.org/content/22/2/190.2.
         full.pdf+html.

Ahituv:1978:APO

[AB78] N. Ahituv and I. Borovits. Analysis of the purchase option of computers. The Computer Journal, 21(2):105–109, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/105.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/105.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/106.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/108.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/109.tif.

Abbott:1970:NMC

[Abb70] M. R. Abbott. Numerical method for calculating the dynamic behaviour of a trolley wire overhead contact system for electric railways. The Computer Journal, 13(4):363–368, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/363.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130363.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/365.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/367.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/368.tif.

Atkinson:1976:MCG

[AD76] L. V. Atkinson and G. J. Duffus. Machine code generation within a compiler — a teaching model. The Computer Journal, 19(2): 160–165, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/160.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/160.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/163.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/164.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/165.tif.

Adamczewski:1976:FER

[Ada76] B. F. M. Adamczewski. A formula for establishing re-order quantities for individual items within a joint stock replenishment scheme. The Computer Journal, 19(3):194–197, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/194.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/194.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/195.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/197.tif.

Alaylioglu:1975:AGQ

[AEH75] Ayse Alaylioglu, G. A. Evans, and J. Hyslop. Automatic generation of quadrature formulae for oscillatory integrals. The Computer Journal, 18(2):173–176, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/173.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/174.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/175.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/176.tif.

Alaylioglu:1976:UCS

[AEH76] A. Alaylioglu, G. A. Evans, and J. Hyslop. The use of Chebyshev series for the evaluation of oscillatory integrals. The Computer Journal, 19(3):258–267, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/3/258.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/258.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/259.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/260.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/261.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/262.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/263.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/265.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/266.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/267.tif.

Alewine:1976:SBH

[AF76] T. Alewine and R. A. Fleck, Jr. Service bureau or in-house data processing. The Computer Journal, 19(3):198–201, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/198.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/198.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/199.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/200.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/201.tif.

Ahmad:1972:ISH

[Ahm72] M. Ahmad. Iterative schemes for high speed division. The Computer Journal, 15(4):333–336, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/333.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150333.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/334.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/335.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/336.tif.

Ahmad:1973:CSN

[Ahm73] M. Ahmad. The computing speed of a new machine. The Computer Journal, 16(2):161–164, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/161.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160161.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/161.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/162.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/163.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/164.tif.

Andres:1974:ACA

[AHM74] T. Andres, W. D. Hoskins, and G. E. McMaster. Algorithm 84: a coupled algorithm for the solution of certain tridiagonal systems of linear equations. The Computer Journal, 17 (4):378–379, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_17/Issue_04/tiff/378.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/379.tif.

Allran:1970:ASN

[AJ70] R. R. Allran and S. E. J. Johnsen. An algorithm for solving nonlinear programming problems subject to nonlinear inequality constraints. The Computer Journal, 13(2):171–177, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/171.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130171.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/171.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/172.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/173.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/174.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/176.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/177.
tif.

Amble:1974:OHT

[AK74] O. Amble and D. E. Knuth. Ordered hash tables. The Computer Journal, 17(2):135–142, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/135.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/135.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/137.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/140.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/141.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/142.tif.

Akitt:1977:FMU

[Aki77] J. W. Akitt. Function minimisation using the Nelder and Mead simplex method with limited arithmetic precision: the self regenerative simplex. The Computer Journal, 20(1):84–85, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/84.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/85.
tif.

Akkoyunlu:1973:LAC

[Akk73] E. A. Akkoyunlu. A linear algorithm for computing the optimum university timetable. The Computer Journal, 16(4):347–350, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/347.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160347.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/347.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/348.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/349.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/350.tif.

Atkinson:1972:CSI

[AL72] M. P. Atkinson and S. R. Lang. A comparison of some inverse Laplace transform techniques for use in circuit design. The Computer Journal, 15(2):138–139, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/138.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150138.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/138.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/139.tif.

Albasiny:1970:BR

[Alb70] E. L. Albasiny. Book review. The Computer Journal, 13(3):262, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/3/262.full.pdf+html.

Albaseny:1971:BRB

[Alb71a] E. L. Albaseny. Book reviews: Numerik Symmetrischer Matrizen, by H. Rutishauser, E. Stiefel and H. Schwarz, 1968; Computational Methods in Partial Differential Equations, by A. R.

Mitchell, 1969. The Computer Journal, 14(4):374, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/374.tif.

Albasiny:1971:BRa

[Alb71b] E. L. Albasiny. Book review. The Computer Journal, 14(4):374, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/374.1.full.pdf+html.

Albasiny:1971:BRb

[Alb71c] E. L. Albasiny. Book review. The Computer Journal, 14(4):374, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/374.2.full.pdf+html.

Albasiny:1974:CMM

[Alb74] E. L. Albasiny. Computational methods for matrix eigenproblems. The Computer Journal, 17(4):317, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/4/317.full.pdf+html.

Albasiny:1975:BR

[Alb75] E. L. Albasiny. Book review. The Computer Journal, 18(3):230, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/230.2.full.pdf+html.

Albasiny:1976:BR

[Alb76] E. L. Albasiny. Book review. The Computer Journal, 19(4):321, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/321.2.full.pdf+html.

Aleksander:1977:BR

[Ale77a] I. Aleksander. Book review. The Computer Journal, 20(1):67, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/67.1.full.pdf+html.

Aleksander:1977:R

[Ale77b] I. Aleksander. Review. The Computer Journal, 20(3):268, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/268.2.
full.pdf+html.

Aleksander:1978:BR

[Ale78] I. Aleksander. Book review. The Computer Journal, 21(3):223, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/223.1.full.pdf+html.

Allen:1971:C

[All71a] R. L. Allen. Correspondence. The Computer Journal, 14(1):86, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/86.full.pdf+html.

Allen:1971:CQW

[All71b] R. L. Allen. Correspondence: On quality of writing. The Computer Journal, 14(1):86, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
140086.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/86.tif. See [Jon71b].

Allen:1975:IC

[All75] K. J. Allen. Insuring computers. The Computer Journal, 18(3): 194–197, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/194.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/194.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/195.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/197.tif.

Almond:1969:UFT

[Alm69] Mary Almond. A University faculty timetable. The Computer Journal, 12(3):215–217, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/215.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
120215.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/215.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/216.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_

03/tiff/217.tif. See correspondence [DD70].

Alpiar:1971:DSO

[Alp71] R. Alpiar. Double syntax oriented processing. The Computer Journal, 14(1):25–37, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/1/25.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/140025.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/32.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/35.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/37.tif.

Alway:1971:GFA

[Alw71] G. G. Alway. A general factorising algorithm. The Computer Journal, 14(2):166–168, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/2/166.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/
140166.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/166.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/167.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/168.tif.

Andrew:1974:CDN

[And74] A. M. Andrew. Correspondence: Decimal number checking schemes. The Computer Journal, 17(4):382–383, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_04/tiff/382.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_04/tiff/383.tif. See correspon-

dence [And75b, How75b, ES75].

Andrew:1975:BR

[And75a] A. M. Andrew. Book review. The Computer Journal, 18(3):233, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/233.full.pdf+html.

Andrew:1975:CDE

[And75b] A. M. Andrew. Correspondence: Decimal error-correction — a solution. The Computer Journal, 18(3):284–285, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/284.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/284.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/285.tif.

See [And74, ES75, How75b].

Andrew:1977:BRa

[And77a] A. M. Andrew. Book review. The Computer Journal, 20(2):123, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/123.1.full.pdf+html.

Andrew:1977:BRb

[And77b] A. M. Andrew. Book review. The Computer Journal, 20(2):123, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/123.2.full.pdf+html.

Andrew:1978:BR

[And78] A. M. Andrew. Book review. The Computer Journal, 21(4):315, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/4/315.full.pdf+html.

Anonymous:1970:ASa

[Ano70a] Anonymous. Algorithms supplement. The Computer Journal, 13 (1):110–115, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/1/110.full.pdf+html.

Anonymous:1970:ASc

[Ano70b] Anonymous. Algorithms supplement. The Computer Journal, 13 (3):323–326, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/323.full.pdf+html.

Anonymous:1970:ASd

[Ano70c] Anonymous. Algorithms supplement. The Computer Journal, 13(4):421–438, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/4/421.full.pdf+html.

Anonymous:1970:ASb

[Ano70d] Anonymous. Alogorithms supplement. The Computer Journal, 13 (2):206–219, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/2/206.full.pdf+html.

Anonymous:1970:C

[Ano70e] Anonymous. Correspondence. The Computer Journal, 13(4):396, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/396.full.pdf+html.

Anonymous:1970:E

[Ano70f] Anonymous. Errata. The Computer Journal, 13(1):91, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/91.full.pdf+html.

Anonymous:1971:ASa

[Ano71a] Anonymous. Algorithms supplement. The Computer Journal, 14 (1):103–107, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/1/103.full.pdf+html.

Anonymous:1971:ASb

[Ano71b] Anonymous. Algorithms supplement. The Computer Journal, 14 (2):207–215, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/207.full.pdf+html.

Anonymous:1971:ASc

[Ano71c] Anonymous. Algorithms supplement. The Computer Journal, 14 (4):444–451, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/4/444.full.pdf+html.

Anonymous:1971:Ea

[Ano71d] Anonymous. Errata. The Computer Journal, 14(3):326, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
3/326.full.pdf+html.

Anonymous:1971:Eb

[Ano71e] Anonymous. Erratum. The Computer Journal, 14(4):351, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/351.2.
full.pdf+html.

Anonymous:1972:ASa

[Ano72a] Anonymous. Algorithms supplement. The Computer Journal, 15(3): 280–285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/280.full.pdf+html.

Anonymous:1972:ASb

[Ano72b] Anonymous. Algorithms supplement. The Computer Journal, 15 (4):382–384, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/382.full.pdf+html.

Anonymous:1972:EMMa

[Ano72c] Anonymous. Eric Mutch Memorial. The Computer Journal, 15 (3):270, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/270.full.pdf+html.

Anonymous:1972:EMMb

[Ano72d] Anonymous. Eric Mutch Memorial. The Computer Journal, 15 (4):355, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/355.1.full.pdf+html.

Anonymous:1972:Ea

[Ano72e] Anonymous. Errata. The Computer Journal, 15(1):79a–79, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/79-a.full.pdf+html.

Anonymous:1972:Eb

[Ano72f] Anonymous. Errata. The Computer Journal, 15(2):175a–175, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
2/175-a.full.pdf+html.

Anonymous:1972:Ec

[Ano72g] Anonymous. Erratum. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
3/258.2.full.pdf+html.

Anonymous:1972:Ed

[Ano72h] Anonymous. Erratum. The Computer Journal, 15(4):355, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/355.2.
full.pdf+html.

Anonymous:1972:IW

[Ano72i] Anonymous. IFIP WG 2.1. The Computer Journal, 15(4):297, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
4/297.full.pdf+html.

Anonymous:1972:NIA

[Ano72j] Anonymous. Notes for intending authors. The Computer Journal, 15(4):349, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/349.full.pdf+html.

Anonymous:1972:NSP

[Ano72k] Anonymous. Notes on the submission of papers. The Computer Journal, 15(1):93, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/1/93.full.pdf+html.

Anonymous:1973:AGF

[Ano73a] Anonymous. Acm George E. forsythe student paper competition. The Computer Journal, 16(2):151, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/151.full.pdf+html.

Anonymous:1973:ASa

[Ano73b] Anonymous. Algorithms supplement. The Computer Journal, 16 (2):187–188, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/187.full.pdf+html.

Anonymous:1973:ASb

[Ano73c] Anonymous. Algorithms supplement. The Computer Journal, 16 (3):273–280, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/273.full.pdf+html.

Anonymous:1973:CIB

[Ano73d] Anonymous. COBOL information bulletin number 16. The Computer Journal, 16(1):81–92, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/81.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/83.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/84.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/86.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/87.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/89.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/90.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_16/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/92.tif.

Anonymous:1973:EMMa

[Ano73e] Anonymous. Eric Mutch Memorial Prize. The Computer Journal, 16(2):140, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/140.1.full.pdf+html.

Anonymous:1973:EMMb

[Ano73f] Anonymous. Eric Mutch memorial prize. The Computer Journal, 16 (3):244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/3/244.1.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/244.tif.

Anonymous:1973:Ea

[Ano73g] Anonymous. Errata. The Computer Journal, 16(1):51, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/51.full.pdf+html.

Anonymous:1973:Eb

[Ano73h] Anonymous. Errata. The Computer Journal, 16(2):140, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/140.2.
full.pdf+html.

Anonymous:1973:Ec

[Ano73i] Anonymous. Errata. The Computer Journal, 16(3):244, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/244.2.
full.pdf+html.

Anonymous:1973:Ed

[Ano73j] Anonymous. Errata. The Computer Journal, 16(4):382, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/382.full.pdf+html.

Anonymous:1973:ORS

[Ano73k] Anonymous. Operational Research Society annual conference: Torquay November 13–16 1973 — Changes resulting from developments in telecommunications. The Computer Journal, 16(3): 280, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/280.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/tiff/280.tif.

Anonymous:1973:RSNa

[Ano73l] Anonymous. Review: Short notice. The Computer Journal, 16 (3):204, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/204.3.full.pdf+html.

Anonymous:1973:RSNb

[Ano73m] Anonymous. Review: Short notice. The Computer Journal, 16 (4):356, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/356.2.full.pdf+html.

Anonymous:1973:RSNc

[Ano73n] Anonymous. Review: Short notice. The Computer Journal, 16 (4):356, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/356.3.full.pdf+html.

Anonymous:1974:ASa

[Ano74a] Anonymous. Algorithms supplement. The Computer Journal, 17 (1):89–92, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/1/89.full.pdf+html.

Anonymous:1974:ASb

[Ano74b] Anonymous. Algorithms supplement. The Computer Journal, 17(4):377–380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/377.full.pdf+html.

Anonymous:1974:BCS

[Ano74c] Anonymous. Journal of the Computer Society of India. The Computer Journal, 17(1):81, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/1/81.2.full.pdf+html.

Anonymous:1974:C

[Ano74d] Anonymous. Correspondence. The Computer Journal, 17(4):381– 383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/381.full.pdf+html.

Anonymous:1974:EA

[Ano74e] Anonymous. Editorial addresses. The Computer Journal, 17(4): 383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/383.full.pdf+html.

Anonymous:1974:EMM

[Ano74f] Anonymous. Eric Mutch Memorial. The Computer Journal, 17(1): 2, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/2.1.full.pdf+html.

Anonymous:1974:E

[Ano74g] Anonymous. Erratum. The Computer Journal, 17(1):81, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/81.1.full.pdf+html.

Anonymous:1974:RSNa

[Ano74h] Anonymous. Review: Short notice. The Computer Journal, 17 (2):142, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/142.3.full.pdf+html.

Anonymous:1974:RSNb

[Ano74i] Anonymous. Review: Short notice. The Computer Journal, 17 (2):147, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/147.2.full.pdf+html.

Anonymous:1974:SSP

[Ano74j] Anonymous. Special summer programme at Massachusetts Institute of Technology. The Computer Journal, 17(2):192, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/192.tif.

Anonymous:1975:ASS

[Ano75a] Anonymous. Algorithm supplement — statement of policy. The Computer Journal, 18(2):185, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/185.full.pdf+html.

Anonymous:1975:ASa

[Ano75b] Anonymous. Algorithms supplement. The Computer Journal, 18 (1):81–92, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/1/81.full.pdf+html.

Anonymous:1975:ASb

[Ano75c] Anonymous. Algorithms supplement. The Computer Journal, 18 (2):177–185, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/177.full.pdf+html.

Anonymous:1975:BR

[Ano75d] Anonymous. Books received. The Computer Journal, 18(2):156, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/156.full.pdf+html.

Anonymous:1975:CP

[Ano75e] Anonymous. Call for papers. The Computer Journal, 18(3):287, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/287.3.full.pdf+html.

Anonymous:1975:CPI

[Ano75f] Anonymous. Call for papers: 3rd International Conference on Automata, Languages and Programming, Edinburgh University, Scotland, July 20–23, 1976. The Computer Journal, 18(3):287, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/287.tif.

Anonymous:1975:DG

[Ano75g] Anonymous. Dr S. Gill. The Computer Journal, 18(2):101, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/101.full.pdf+html.

Anonymous:1975:Ea

[Ano75h] Anonymous. Erratum. The Computer Journal, 18(1):95, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/95.full.pdf+html.

Anonymous:1975:Eb

[Ano75i] Anonymous. Erratum. The Computer Journal, 18(2):191, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/191.2.
full.pdf+html.

Anonymous:1975:Ec

[Ano75j] Anonymous. Erratum. The Computer Journal, 18(3):287, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/287.2.
full.pdf+html.

Anonymous:1975:NJ

[Ano75k] Anonymous. New journals. The Computer Journal, 18(2):121, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/121.3.full.pdf+html.

Anonymous:1975:RDD

[Ano75l] Anonymous. Reports on degrees in data processing. The Computer Journal, 18(4):382–398, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/382.full.pdf+html.

Anonymous:1975:RSNa

[Ano75m] Anonymous. Review: Short notice. The Computer Journal, 18 (2):121, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/121.1.full.pdf+html.

Anonymous:1975:RSNb

[Ano75n] Anonymous. Review: Short notice. The Computer Journal, 18 (2):121, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/121.2.full.pdf+html.

Anonymous:1975:RSNc

[Ano75o] Anonymous. Review: Short notice. The Computer Journal, 18 (3):222, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/222.4.full.pdf+html.

Anonymous:1975:SN

[Ano75p] Anonymous. Special notice. The Computer Journal, 18(3): 287, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/18/3/287.1.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_18/Issue_03/tiff/287.tif.

Anonymous:1976:ASIa

[Ano76a] Anonymous. Advanced study institute on man-computer interaction. The Computer Journal, 19(2):181, May 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/2/181.full.pdf+html.

Anonymous:1976:ASIb

[Ano76b] Anonymous. Advanced summer institute on computer architecture. The Computer Journal, 19(2):181a–181, May 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/181-a.
full.pdf+html.

Anonymous:1976:ASa

[Ano76c] Anonymous. Algorithms supplement. The Computer Journal, 19(1): 95, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/95.3.full.pdf+html.

Anonymous:1976:ASb

[Ano76d] Anonymous. Algorithms supplement. The Computer Journal, 19 (4):360–363, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/360.full.pdf+html.

Anonymous:1976:ASS

[Ano76e] Anonymous. Algorithms supplement-statement of policy. The Computer Journal, 19(4):363, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/4/363.full.pdf+html.

Anonymous:1976:EPV

[Ano76f] Anonymous. Editorial: a personal view of the ‘Beaumont’ weekend. The Computer Journal, 19(1):2, February 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/2.tif.

Anonymous:1976:Eb

[Ano76g] Anonymous. Errata. The Computer Journal, 19(4):379, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/379.full.pdf+html.

Anonymous:1976:Ec

[Ano76h] Anonymous. Errata. The Computer Journal, 19(4):384, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/384.1.
full.pdf+html.

Anonymous:1976:Ea

[Ano76i] Anonymous. Erratum. The Computer Journal, 19(4):352, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/352.3.
full.pdf+html.

Anonymous:1976:NAF

[Ano76j] Anonymous. New ans Fortran standard. The Computer Journal, 19(1):7, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/1/7.full.pdf+html.

Anonymous:1976:SIC

[Ano76k] Anonymous. Second international conference on very large data bases. The Computer Journal, 19(1):78, ???? 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/comjnl.oxfordjournals.org/content/19/1/78.full.pdf+html.

Anonymous:1976:SIW

[Ano76l] Anonymous. Sigmod-77 — International Workshop on Management of Data. The Computer Journal, 19(4):384, ???? 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/384.2.
full.pdf+html.

Anonymous:1977:ASa

[Ano77a] Anonymous. Algorithms supplement. The Computer Journal, 20 (1):86–91, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/1/86.full.pdf+html.

Anonymous:1977:ASb

[Ano77b] Anonymous. Algorithms supplement. The Computer Journal, 20 (3):280, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/280.1.full.pdf+html.

Anonymous:1977:ASc

[Ano77c] Anonymous. Algorithms supplement. The Computer Journal, 20 (4):371–379, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/371.full.pdf+html.

Anonymous:1977:ASS

[Ano77d] Anonymous. Algorithms supplement-statement of policy. The Computer Journal, 20(3):280–281, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/3/280.2.full.pdf+html.

Anonymous:1977:A

[Ano77e] Anonymous. Apology. The Computer Journal, 20(2):131, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/131.2.
full.pdf+html.

Anonymous:1977:BRa

[Ano77f] Anonymous. Book review. The Computer Journal, 20(1):36, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/36.2.full.pdf+html.

Anonymous:1977:BRb

[Ano77g] Anonymous. Book review. The Computer Journal, 20(2):150, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/150.1.full.pdf+html.

Anonymous:1977:Eb

[Ano77h] Anonymous. Errata. The Computer Journal, 20(4):329, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/329.3.
full.pdf+html.

Anonymous:1977:Ea

[Ano77i] Anonymous. Erratum. The Computer Journal, 20(1):96, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/96.full.pdf+html.

Anonymous:1977:EFM

[Ano77j] Anonymous. European Federation for Medical Informatics — Cambridge Congress call for papers. The Computer Journal, 20(2):177, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/177.2.full.pdf+html.

Anonymous:1977:HCF

[Ano77k] Anonymous. A historic computer film. The Computer Journal, 20 (2):177, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/177.1.full.pdf+html.

Anonymous:1977:SN

[Ano77l] Anonymous. Short notice. The Computer Journal, 20(1):44, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/44.4.full.pdf+html.

Anonymous:1978:ACC

[Ano78a] Anonymous. Advances in computer chess. The Computer Journal, 21(1):94, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/1/94.full.pdf+html.

Anonymous:1978:ASa

[Ano78b] Anonymous. Algorithms supplement. The Computer Journal, 21 (2):178–183, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/178.full.pdf+html.

Anonymous:1978:ASb

[Ano78c] Anonymous. Algorithms supplement. The Computer Journal, 21 (3):270–277, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/270.full.pdf+html.

Anonymous:1978:ASc

[Ano78d] Anonymous. Algorithms supplement. The Computer Journal, 21 (4):373–377, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/373.full.pdf+html.

Anonymous:1978:CP

[Ano78e] Anonymous. Call for papers. The Computer Journal, 21(3):204, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/204.full.pdf+html.

Anonymous:1978:CC

[Ano78f] Anonymous. Conference and course. The Computer Journal, 21 (4):346, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/346.full.pdf+html.

Anonymous:1978:ENR

[Ano78g] Anonymous. Editorial notes: Reprinting of algorithms. The Computer Journal, 21(3):276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/276.tif.

Anonymous:1978:E

[Ano78h] Anonymous. Errata. The Computer Journal, 21(4):332, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
4/332.full.pdf+html.

Anonymous:1979:ASa

[Ano79a] Anonymous. Algorithms supplement. The Computer Journal, 22 (1):86–90, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/1/86.full.pdf+html.

Anonymous:1979:ASb

[Ano79b] Anonymous. Algorithms supplement. The Computer Journal, 22 (2):188–189, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/2/188.full.pdf+html.

Anonymous:1979:ASc

[Ano79c] Anonymous. Algorithms supplement. The Computer Journal, 22 (3):282–285, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/3/282.full.pdf+html.

Anonymous:1979:ASd

[Ano79d] Anonymous. Algorithms supplement. The Computer Journal, 22 (4):383, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/4/383.full.pdf+html.

Anonymous:1979:ASS

[Ano79e] Anonymous. Algorithms supplement — supplement of policy. The Computer Journal, 22(1):90, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/1/90.1.full.pdf+html.

Anonymous:1979:BAS

[Ano79f] Anonymous. Bcs award scheme. The Computer Journal, 22(3):208, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/3/208.1.full.pdf+html.

Anonymous:1979:BRI

[Ano79g] Anonymous. Books reviewed in this issue. The Computer Journal, 23(4):306, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/4/306.full.pdf+html.

Anonymous:1979:CPa

[Ano79h] Anonymous. Calls for papers. The Computer Journal, 22(4):295, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/295.full.pdf+html.

Anonymous:1979:CPb

[Ano79i] Anonymous. Calls for papers. The Computer Journal, 23(4):297, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/297.full.pdf+html.

Anonymous:1979:CR

[Ano79j] Anonymous. Conference round up. The Computer Journal, 22(3):208, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/3/208.2.full.pdf+html.

Anonymous:1979:CC

[Ano79k] Anonymous. Conferences and courses. The Computer Journal, 22(1): 79, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/79.full.pdf+html.

Anonymous:1979:DEE

[Ano79l] Anonymous. Discussion on “An entity/event data modelling method. The Computer Journal, 22(3):279–281, August 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/279.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/280.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/281.tif. See

[Rob79].

Anonymous:1979:Ea

[Ano79m] Anonymous. Editorial. The Computer Journal, 22(1):2, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/2.tif.

Anonymous:1979:Eb

[Ano79n] Anonymous. Erratum. The Computer Journal, 23(4):383, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/4/383.1.
full.pdf+html.

Anonymous:1979:EAa

[Ano79o] Anonymous. Erratum: Algorithm 105. The Computer Journal, 22 (1):90, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/1/90.2.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_01/tiff/90.tif. See

[BF78].

Anonymous:1979:NAS

[Ano79p] Anonymous. New approaches to systems analysis and design: a oneday seminar held as a tribute to Eric Mutch. The Computer Journal, 22(2):168, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/2/168.full.pdf+html.

Anonymous:1979:VGT

[Ano79q] Anonymous. Visiting graduate teaching positions at the university of helsinki. The Computer Journal, 23(4):343, ???? 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/23/4/343.full.pdf+html.

Albano:1979:HSR

[AO79] A. Albano and R. Orsini. A heuristic solution of the rectangular cutting stock problem. The Computer Journal, 23(4):338–343, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/338.full.pdf+html.

Armstrong:1978:BR

[Arm78a] J. L. Armstrong. Book review. The Computer Journal, 21(1):14, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/14.3.full.pdf+html.

Armstrong:1978:PPC

[Arm78b] J. L. Armstrong. Programming a parallel computer for robot vision. The Computer Journal, 21(3):215–218, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/215.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/215.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/216.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/218.tif.

Arthur:1975:CRV

[Art75] D. F. Arthur. Correspondence: Reading variable length records using FORTRAN IV. The Computer Journal, 18(3):283, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/283.1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/283.tif. See [HV74, ET75].

Arblaster:1979:JSP

[ASG79] A. T. Arblaster, M. E. Sime, and T. R. G. Green. Jumping to some purpose. The Computer Journal, 22(2):105–109, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/105.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/105.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/106.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/108.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/109.tif.

See correspondence [Hil80, Rob80, Tay80, MR82].

Allen:1970:CWA

[AT70] R. L. Allen and The Editor. Correspondence: What is an analyst? The Computer Journal, 13(1):119–120, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/120.tif. See [Fel69, Fel70].

Atkinson:1976:LPL

[Atk76] L. V. Atkinson. List processing language, SLPL — an informal outline. The Computer Journal, 19(1):32–39, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/32.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/32.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/33.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

19/Issue_01/tiff/34.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/35.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/36.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/38.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/39.
tif.

Austin:1971:DDA

[Aus71] B. J. Austin. A dynamic disc allocation algorithm designed to reduce fragmentation during file reloading. The Computer Journal, 14(4):378–381, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/4/378.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140378.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/378.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/379.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/380.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/381.tif.

Aust:1976:IAS

[Aus76] R. J. Aust. An improvement algorithm for school timetabling. The Computer Journal, 19(4):339–343, November 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/339.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/339.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/340.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/341.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/342.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/343.tif.

Anderssen:1971:INM

[AW71] A. S. Anderssen and E. T. White. Improved numerical methods for Volterra integral equations of the first kind. The Computer Journal, 14(4):442–443, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/4/442.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140442.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_04/tiff/442.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/443.tif.

See corrections [Lew72a].

Backhouse:1971:C

[Bac71a] R. C. Backhouse. Correspondence. The Computer Journal, 14(4):365, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/365.2.full.pdf+html.

Backhouse:1971:CHL

[Bac71b] R. C. Backhouse. Correspondence: On “High level languages for low level users”. The Computer Journal, 14(4):365, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/365.tif. See [ER71, Wal71b, FW72,

Hil72b].

Backhouse:1977:BR

[Bac77] R. C. Backhouse. Book review. The Computer Journal, 20(1):72, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/72.full.pdf+html.

Backhouse:1978:BR

[Bac78] R. C. Backhouse. Book review. The Computer Journal, 21(3):198, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/198.2.full.pdf+html.

Baecker:1970:UAT

[Bae70] H. D. Baecker. The use of Algol 68 for trees. The Computer Journal, 13(1):25–27, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/25.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130025.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/27.tif.

Baecker:1975:ARC

[Bae75] H. D. Baecker. Areas and record-classes. The Computer Journal, 18(3):223–226, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/3/223.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/223.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/225.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/226.tif.

Baldota:1975:C

[Bal75] S. N. Baldota. Correspondence. The Computer Journal, 18(2):190, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/190.2.full.pdf+html.

Bale:1979:BR

[Bal79] D. W. B. Bale. Book review. The Computer Journal, 22(2):109, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/2/109.2.full.pdf+html.

Band:1979:BR

[Ban79] T. A. Band. Book review. The Computer Journal, 23(4):323, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/323.2.full.pdf+html.

Barron:1970:BR

[Bar70] D. W. Barron. Book review. The Computer Journal, 13(2):135, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/135.full.pdf+html.

Barron:1971:ACF

[Bar71] D. W. Barron. Approaches to conversational FORTRAN. The Computer Journal, 14(2):123–127, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/2/123.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/

140123.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/123.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/124.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/125.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/126.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/127.tif.

Bares:1972:C

[Bar72a] P. Bares. Correspondence. The Computer Journal, 15(3):276, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/276.full.pdf+html.

Bares:1972:CFD

[Bar72b] P. Bareš. Correspondence: On “The formal definition of the BA-SIC language”. The Computer Journal, 15(3):276, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/150276.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/tiff/276.tif. See

[Lee72].

Barnes:1972:RTL

[Bar72c] J. G. P. Barnes. Real time languages for process control. The Computer Journal, 15(1):15–17, February 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/1/15.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/150015.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/15.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/16.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/17.tif.

Barron:1973:CEP

[Bar73a] D. W. Barron. Correspondence: On editorial policy. The Computer Journal, 16(3):284, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_16/Issue_03/tiff/284.tif. See

[GT72].

Barry:1973:BR

[Bar73b] J. N. Barry. Book review. The Computer Journal, 16(2):105, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/105.1.full.pdf+html.

Barron:1974:DJC

[Bar74a] D. W. Barron. Discussion: Job control languages and job control programs. The Computer Journal, 17(3):282–286, August 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/282.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/284.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/285.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/286.tif.

Barron:1974:MTG

[Bar74b] D. W. Barron. A mathematical theory of global program optimization. The Computer Journal, 17(4):336, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
4/336.full.pdf+html.

Barron:1974:BR

[Bar74c] I. M. Barron. Book review. The Computer Journal, 17(1):37, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/37.full.pdf+html.

Barker:1975:SDP

[Bar75a] P. G. Barker. Syntactic definition and parsing of molecular formulae: Part 1. Initial syntax definition and parser implementation. The Computer Journal, 18(4):355–359, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/355.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/355.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/356.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/358.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/359.tif.

Barron:1975:BR

[Bar75b] I. M. Barron. Book review. The Computer Journal, 18(4):365, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
4/365.2.full.pdf+html.

Barber:1976:BR

[Bar76a] D. L. A. Barber. Book review. The Computer Journal, 19(3):288, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/288.2.full.pdf+html.

Barker:1976:C

[Bar76b] P. G. Barker. Correspondence. The Computer Journal, 19(1):94, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/94.1.full.pdf+html.

Barron:1976:C

[Bar76c] D. W. Barron. Correspondence. The Computer Journal, 19(1):93, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/93.2.full.pdf+html.

Barber:1977:R

[Bar77a] B. Barber. Review. The Computer Journal, 20(3):268, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/268.1.
full.pdf+html.

Barron:1977:BRa

[Bar77b] D. W. Barron. Book review. The Computer Journal, 20(1):67, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/67.2.full.pdf+html.

Barron:1977:BRb

[Bar77c] I. M. Barron. Book review. The Computer Journal, 20(2):140, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/140.2.full.pdf+html.

Barlow:1978:ASP

[Bar78a] J. F. Barlow. An algorithm for the solution of the postman’s problem. The Computer Journal, 21(1):86–89, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/86.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/86.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/87.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/89.tif. See correspon-

dence [Gol78a].

Barron:1978:BRa

[Bar78b] D. W. Barron. Book review. The Computer Journal, 21(3):233, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/233.2.full.pdf+html.

Barron:1978:BRb

[Bar78c] I. M. Barron. Book review. The Computer Journal, 21(3):285, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/285.1.full.pdf+html.

Barron:1978:BRc

[Bar78d] I. M. Barron. Book review. The Computer Journal, 21(3):285, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/285.2.full.pdf+html.

Barron:1978:BRd

[Bar78e] I. M. Barron. Book review. The Computer Journal, 21(4):354, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
4/354.1.full.pdf+html.

Barry:1978:DCR

[Bar78f] P. T. Barry. Discussion and correspondence: the Regional Computing Organisation Network. The Computer Journal, 21(2):184– 187, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/184.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/184.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/185.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/186.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/187.tif.

Barber:1979:BR

[Bar79] B. Barber. Book review. The Computer Journal, 22(1):36, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/36.3.full.pdf+html.

Bays:1973:STS

[Bay73] C. Bays. Some techniques for structuring chained hash tables. The Computer Journal, 16(2):126–131, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/126.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/160126.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/126.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/127.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/128.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/129.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/131.tif.

Bays:1976:NRT

[Bay76] C. Bays. Non-recursive technique for recreating a digraph from its K-formula representation. The Computer Journal, 19(4):326– 328, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/326.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/326.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/

tiff/327.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/328.tif.

Berry:1971:NAH

[BB71] J. L. Berry and A. G. Bell. Note on Algorithm 50: How to program a computer to play legal chess. The Computer Journal, 14(1):106–107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_01/tiff/106.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/tiff/107.tif. See [Bel70a].

Barton:1970:AS

[BBF70] D. Barton, S. R. Bourne, and J. P. Fitch. An algebra system. The Computer Journal, 13(1):32–39, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
13/1/32.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130032.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/32.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/34.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/35.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/37.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/38.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/39.tif.

Ball:1971:FRA

[BBG^{+}71] D. Ball, P. M. Blackall, V. Gerard, G. R. Macleod, P. J. Marcer, and E. M. Palandri. FOCUS — a remote access file handling system on-line to a CDC 6000 series computer. The Computer Journal, 14 (2):113–118, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/113.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140113.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/115.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/116.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/117.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/118.tif.

Backus:1976:MRA

[BBG^{+}76] J. W. Backus, F. L. Bauer, J. Green, C. Katz, J. McCarthy, P. Naur, A. J. Perlis, H. Rutishauser, K. Samelson, B. Vauquois, J. H. Wegstein, A. van Wijngaarden, M. Woodger, R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified report on the algorithmic language ALGOL-60. The Computer Journal, 19(4):364–379, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/364.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/364.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/365.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/367.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/368.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/369.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/370.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/371.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/372.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/373.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/374.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/375.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/376.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/377.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/378.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/379.tif.

Barton:1970:SCA

[BBH70] D. Barton, S. R. Bourne, and J. R. Horton. The structure of the Cambridge algebra system. The Computer Journal, 13(3):243– 247, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/243.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_03/130243.sgm.abs.html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/243.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/244.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/245.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/246.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/247.tif.

Broadbent:1973:MME

[BC73] J. K. Broadbent and G. F. Coulouris. MEMBERS — a Microprogrammed Experimental Machine with a Basic Executive for Realtime Systems. The Computer Journal, 16(3):205–208, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/205.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160205.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/207.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/208.
tif.

Boote:1972:SPC

[BCR72] W. P. Boote, S. R. Clark, and T. A. Rourke. Simulation of a paging computer system. The Computer Journal, 15(1):51–57, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/51.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150051.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/51.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/52.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/53.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/54.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/55.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/56.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/57.tif.

Bennett:1972:FRT

[BD72] D. G. Bennett and R. A. Davenport. A facility for real-time program development. The Computer Journal, 15(1):5–7, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
1/5.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/150005.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/5.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/7.tif.

Boccalatte:1979:ADS

[BD79] A. Boccalatte and M. Di Manzo. An approach to the detection of semantic errors. The Computer Journal, 23(4):317–323, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/317.full.pdf+html.

Benson:1973:AAS

[BE73] A. Benson and D. J. Evans. Algorithm 80: an algorithm for the solution of periodic quindiagonal systems of linear equations. The Computer Journal, 16(3):278–279, August 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/278.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/279.tif.

Barlow:1979:POB

[BE79] R. H. Barlow and D. J. Evans. A parallel organisation of the bisection algorithm. The Computer Journal, 22(3):267–269, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/267.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/267.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/268.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/269.tif.

Beasley:1974:CSR

[Bea74] J. D. Beasley. Correspondence: System 4 random number. The Computer Journal, 17(4):381, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/381.tif.

Beale:1977:BR

[Bea77] E. M. L. Beale. Book review. The Computer Journal, 20(2):131, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/131.1.full.pdf+html.

Beilby:1972:RTS

[Bei72] M. H. Beilby. Road traffic simulation on a small computer. The Computer Journal, 15(2):134–137, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/134.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150134.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/134.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/135.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/136.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/137.tif.

Bell:1970:AHP

[Bel70a] A. G. Bell. Algorithm 50: How to program a computer to play legal chess. The Computer Journal, 13(2):208–219, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/208.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/209.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/210.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/211.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/212.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/213.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/214.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/215.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/216.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/218.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/219.

tif. See note [BB71].

Bell:1970:PID

[Bel70b] A. G. Bell. Partitioning integers in N dimensions. The Computer Journal, 13(3):278–283, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/3/278.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/
130278.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/278.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/279.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/280.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/281.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/282.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/283.tif.

Bell:1971:BR

[Bel71] D. A. Bell. Book review. The Computer Journal, 14(4):343, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/343.3.full.pdf+html.

Bell:1973:H

[Bel73] A. G. Bell. Himmelbett. The Computer Journal, 16(3):269– 272, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/269.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/160269.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/270.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/271.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_

03/tiff/272.tif. See correspondence [Fre74b, Str74b].

Bell:1974:CCS

[Bel74] D. A. Bell. Correspondence: Computing Science. The Computer Journal, 17(4):383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_17/Issue_04/tiff/383.tif. See

[Bro73a, Fis74b].

Bell:1976:PSP

[Bel76] D. Bell. Programmer selection and programming errors. The Computer Journal, 19(3):202–206, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/3/202.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/203.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/204.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/205.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/206.tif.

Bemer:1974:CMI

[Bem74] R. W. Bemer. Correspondence: On mixing interpretive and machine code. The Computer Journal, 17(2):188–189, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/188.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/188.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/189.tif.

See [Daw73, DP73, Mac74b].

Bennett:1970:DCC

[Ben70] R. D. Bennett. Discussion and correspondence: computers and the business information structure: new horizons in decision making effectiveness. The Computer Journal, 13(3):327–329, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/327.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130327.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/327.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/329.tif.

Bennetts:1972:RAD

[Ben72] R. G. Bennetts. A realistic approach to detection test set generation for combinational logic circuits. The Computer Journal, 15(3): 238–246, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/238.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/150238.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/238.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/239.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_15/Issue_03/tiff/240.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/241.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/242.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/243.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/244.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/245.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/246.tif.

Berry:1973:BR

[Ber73] J. L. Berry. Book review. The Computer Journal, 16(2):105, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/105.3.full.pdf+html.

Bernal:1974:BR

[Ber74] M. J. M. Bernal. Book review. The Computer Journal, 17(3):204, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/204.2.full.pdf+html.

Berry:1976:ADD

[Ber76] G. Berry. Algorithm 92: The drawing of dashed lines. The Computer Journal, 19(4):361–363, November 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/361.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/363.tif.

Best:1979:BR

[Bes79] E. Best. Book review. The Computer Journal, 22(3):266, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/266.1.
full.pdf+html.

Brittan:1971:CTC

[BF71] J. N. G. Brittan and F. J. M. Farley. College timetable construction by computer. The Computer Journal, 14(4):361–365, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/361.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_04/140361.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/361.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/363.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/364.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/365.tif.

Barton:1972:RAM

[BF72] D. Barton and J. P. Fitch. A review of algebraic manipulative programs and their application. The Computer Journal, 15(4):362– 381, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/362.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_04/150362.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/362.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/363.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/364.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/366.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/367.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/370.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/372.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/373.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/374.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/375.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/376.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/378.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/379.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/380.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/381.tif.

Buckley:1978:ADN

[BF78] James J. Buckley and Robert A. Fleck, Jr. Algorithm 105: Determining the number of internal rates of return. The Computer Journal, 21(4):373–377, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_21/Issue_04/tiff/373.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/374.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/375.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/376.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_

04/tiff/377.tif. See erratum [Ano79o].

Bridle:1971:CPI

[BG71] J. W. Bridle and R. J. Gregersen. CMSR — a personnel information system. The Computer Journal, 14(4):338–343, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/338.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140338.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/338.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/339.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/340.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/341.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/342.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/343.tif.

Boulton:1979:MDL

[BG79] P. I. P. Boulton and J. R. Goguen. A machine description language. The Computer Journal, 22(2):132–135, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/132.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/133.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/135.tif.

Bull:1978:PMM

[BGL^{+}78] G. M. Bull, S. Gilbey, Kathleen Levine, A. G. Lippiatt, and C. H. C. Machin. Proteus — a microcoded multiprocessor system. The Computer Journal, 21(2):154–160, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/154.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/154.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/155.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/156.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/157.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/158.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/160.tif.

Buckle:1970:CSF

[BHRS70] D. G. R. Buckle, C. H. Hunt, M. M. Ryder, and S. H. Storey. Computers and the small firm: 3. The Computer Journal, 13(1):7–9, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/1/7.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
130007.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/7.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/9.tif.

Bonczek:1978:ITW

[BHW78] R. H. Bonczek, C. W. Holsapple, and A. B. Whinston. Information transferral within a distributed data base via a generalised mapping language. The Computer Journal, 21(2):110–116, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/110.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/110.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/111.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/113.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/114.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/115.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/116.tif.

Billett:1977:BR

[Bil77] R. W. Billett. Book review. The Computer Journal, 20(4):295, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/295.1.full.pdf+html.

Bird:1974:SP

[Bir74] R. Bird. Speeding up programs. The Computer Journal, 17(4):337– 339, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/337.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/337.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/338.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/339.tif.

Bird:1979:REV

[Bir79] R. S. Bird. Recursion elimination with variable parameters. The Computer Journal, 22(2):151–154, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/22/2/151.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/151.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/152.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/153.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/154.tif.

Barker:1978:SDP

[BJ78] P. G. Barker and P. S. Jones. Syntactic definition and parsing of molecular formulae. part 2: Graphical synthesis of molecular formulae for data base queries. The Computer Journal, 21(3):224– 233, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/224.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/224.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/226.tif;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_21/Issue_03/tiff/227.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/228.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/229.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/230.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/231.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/233.tif.

Bell:1979:HRM

[BJ79] A. G. Bell and N. Jacobi. How to read, make and store chess moves. The Computer Journal, 22(1):71–75, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/71.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/72.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/74.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/75.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/76.tif.

Ballard:1974:ASC

[BJS74] Dana H. Ballard, C. O. Jelinek, and R. Schinzinger. An algorithm for the solution of constrained generalised polynomial programming problems. The Computer Journal, 17(3):261–266, August 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/261.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/261.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/262.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/264.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/265.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_

03/tiff/266.tif. See correspondence [Bra75].

Berkling:1975:CDH

[BK75] K. Berkling and W. Kluge. Conceptual design of a hardwired main storage management unit. The Computer Journal, 18(4):360– 365, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/4/360.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/360.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/361.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/362.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/363.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/365.tif.

Boulton:1979:EPR

[BK79a] P. I. P. Boulton and M. A. R. Kittler. Estimating program reliability. The Computer Journal, 22(4):328–331, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/328.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/328.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/329.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/331.tif.

Burton:1979:ODH

[BK79b] F. W. Burton and J. Kollias. Optimising disc head movements in secondary key retrievals. The Computer Journal, 22(3):206– 208, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/3/206.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/206.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/207.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/208.tif.

Burston:1978:DLA

[BKK78] A. K. Burston, D. J. Kinniment, and Hilary Kahn. A design language for asynchronous logic. The Computer Journal, 21(4):347– 354, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/347.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/347.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/

tiff/348.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/349.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/350.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/351.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/352.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/353.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/354.tif.

Burnett:1974:SRT

[BKL74] P. Burnett, P. A. Kidd, and A. M. Lister. Simulation of realtime program faults. The Computer Journal, 17(1):25–27, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/25.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/27.tif.

Bratko:1978:PBR

[BKM78] I. Bratko, D. Kopec, and D. Michie. Pattern-based representation of chess end-game knowledge. The Computer Journal, 21(2): 149–153, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/149.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/149.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/150.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/151.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/152.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/153.tif.

Bennetts:1971:FDD

[BL71] R. G. Bennetts and D. W. Lewin. Fault diagnosis of digital systems — a review. The Computer Journal, 14(2):199–206, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/199.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140199.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/199.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

14/Issue_02/tiff/200.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/201.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/202.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/203.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/204.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/205.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/206.tif.

Beaven:1972:APP

[BL72] P. A. Beaven and D. W. Lewin. An associative parallel processing system for non-numerical computation. The Computer Journal, 15(4):343–349, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/343.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150343.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/343.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/344.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/345.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/346.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/347.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/348.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/349.tif.

Blazie:1977:CCP

[BL77] Deane B. Blazie and L. S. Levy. A cross compiler for pocket calculators. The Computer Journal, 20(3):213–221, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/213.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/215.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/216.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/217.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/218.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/219.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/220.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/221.tif.

Black:1975:BR

[Bla75] J. W. Black. Book review. The Computer Journal, 18(1):33, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/33.2.full.pdf+html.

Black:1978:MCD

[Bla78] D. D. Black. Multi-coordinate data presentation. The Computer Journal, 21(2):148a–148, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/148-a.full.pdf+html.

Black:1979:BRa

[Bla79a] A. P. Black. Book review. The Computer Journal, 22(1):21, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/21.2.full.pdf+html.

Black:1979:BRb

[Bla79b] D. D. Black. Book review. The Computer Journal, 22(1):62, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/62.1.full.pdf+html.

Blanin:1979:BR

[Bla79c] A. Blanin. Book review. The Computer Journal, 23(4):352, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/352.1.full.pdf+html.

Blake:1970:RPP

[BLY70] L. F. Blake, R. E. Lawson, and I. M. Yuille. A ring processing package for use with FORTRAN or a similar high-level language. The Computer Journal, 13(1):40–47, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
13/1/40.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130040.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/40.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_13/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/42.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/43.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/45.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/46.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/47.tif.

Bratko:1979:APC

[BM79] I. Bratko and D. Michie. An advice program for a complex chess programming task. The Computer Journal, 23(4):353–359, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/353.full.pdf+html.

Barrow:1971:TED

[BMPS71] H. G. Barrow, D. Michie, R. J. Popplestone, and S. H. Salter. Tokyo-Edinburgh dialogue on robots in artificial intelligence research. The Computer Journal, 14(1):91–95, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
14/1/91.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140091.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/91.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/92.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/93.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/94.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/95.tif.

Boari:1976:NDR

[BN76] M. Boari and G. Neri. A note on deadlock in the ‘readers and writers’ problem. The Computer Journal, 19(3):275, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/275.tif.

Borovits:1978:IPC

[BN78] I. Borovits and S. Neumann. Internal pricing for computer services. The Computer Journal, 21(3):199–204, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/199.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/199.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/200.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/201.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/203.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/204.tif.

Boardman:1979:BR

[Boa79] K. Boardman. Book review. The Computer Journal, 22(1):66, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/66.full.pdf+html.

Bonney:1971:KIF

[Bon71] R. W. Bonney. A keyword index file of British book titles. The Computer Journal, 14(4):333–337, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/333.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140333.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/333.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/335.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/336.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/337.tif.

Boothroyd:1967:ASS

[Boo67a] J. Boothroyd. Algorithm 25: Sort a section of the elements of an array by determining the rank of each element. The Computer Journal, 10(3):308–309, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_10/Issue_03/tiff/308.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_

03/tiff/309.tif. See note [Hil69, Woo70g].

Boothroyd:1967:AOS

[Boo67b] J. Boothroyd. Algorithm 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 10(3):309–310, November 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_10/Issue_03/tiff/309.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_10/Issue_

03/tiff/310.tif. See note [Hil69, Woo70g].

Boothroyd:1974:CRC

[Boo74a] J. Boothroyd. Correspondence: On a radix conversion process. The Computer Journal, 17(1):95, February 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/

tiff/95.tif. See erratum [Boo74b] and correspondence [Ham74a].

Boothroyd:1974:ERC

[Boo74b] J. Boothroyd. Erratum: “On a radix conversion process”. The Computer Journal, 17(4):381, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/

tiff/381.tif. See [Boo74a, Ham74a].

Boothroyd:1978:CAM

[Boo78] J. Boothroyd. Certification of Algorithm 1: Mintree. The Computer Journal, 21(3):277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_21/Issue_03/tiff/277.tif. See

[Obr78a].

Boothroyd:1979:AP

[Boo79a] J. Boothroyd. Algorithm 6: PERM. The Computer Journal, 22(1): 88–89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/88.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/89.

tif. Reprinted from The Computer Bulletin, September 1964, p. 67.

Boothroyd:1979:AM

[Boo79b] J. Boothroyd. Algorithm 7: MINX. The Computer Journal, 22(1): 89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/89.tif. Reprinted from

The Computer Bulletin, September 1964, p. 67.

Boothroyd:1979:CAF

[Boo79c] J. Boothroyd. Certification of Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/

tiff/86.tif. Reprinted from The Computer Bulletin. See [PP79].

Boothroyd:1979:CAE

[Boo79d] J. Boothroyd. Certification of Algorithm 3: Easter. The Computer Journal, 22(1):87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_22/Issue_01/tiff/87.tif. See

[Hil79c].

Bowring:1973:DCU

[Bow73] C. A. Bowring. Discussion and correspondence: using macros to aid assembly-language teaching. The Computer Journal, 16(3):281– 283, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/281.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/160281.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/283.tif.

Bozyigit:1979:FRP

[BP79] M. Bozyigit and Y. Paker. A fixed routing problem in large and high connectivity networks. The Computer Journal, 22(3):246– 250, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/3/246.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/246.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/247.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/248.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/249.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/250.tif.

Boari:1974:PEP

[BPR74] M. Boari, P. Pellizzardi, and R. Rossi. Performance evaluation of a process control system by simulation model. The Computer Journal, 17(4):365–370, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/365.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/365.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/366.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/367.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/368.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/369.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/370.tif.

Brown:1972:DCD

[BPSS72] A. E. Brown, C. E. Phillips, J. S. Scandale, and D. P. Sparrow. A data collection and display system for a large-scale simulation. The Computer Journal, 15(2):105–108, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/105.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150105.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/105.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/106.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/107.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/108.

tif. See [Tem72b].

Bradly:1972:CHL

[Bra72] H. W. Bradly. Correspondence: On high level languages. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/

tiff/258.tif. See [Pal72b].

Brandon:1974:IUS

[Bra74] D. M. Brandon, Jr. The implementation and use of sparse matrix techniques in general simulation programs. The Computer Journal, 17(2):165–171, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/165.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/165.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_17/Issue_02/tiff/167.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/168.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/169.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/170.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/171.tif.

Bradley:1975:CAS

[Bra75] John Bradley. Correspondence: On “An algorithm for the solution of constrained generalised polynomial programming problems”. The Computer Journal, 18(3):285–286, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/285.3.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/285.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/286.tif.

See [BJS74].

Brady:1976:PAS

[Bra76] J. M. Brady. A programming approach to some concepts and results in the theory of computation. The Computer Journal, 19(3): 234–237, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/3/234.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/234.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/235.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/237.tif.

Brady:1977:HPR

[Bra77] J. M. Brady. Hints on proofs by recursion induction. The Computer Journal, 20(4):353–355, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/4/353.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/353.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/354.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/355.tif.

Brady:1978:BR

[Bra78a] J. M. Brady. Book review. The Computer Journal, 21(4):384, ????

1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
4/384.full.pdf+html.

Brady:1978:BRB

[Bra78b] J. M. Brady. Book review: Structural Pattern Recognition, by T. Pavlidis, 1977. The Computer Journal, 21(4):384, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/384.tif.

Bramer:1978:BR

[Bra78c] M. A. Bramer. Book review. The Computer Journal, 21(1):14, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/14.2.full.pdf+html.

Brady:1979:BRa

[Bra79a] J. M. Brady. Book review. The Computer Journal, 22(3):261, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/261.2.full.pdf+html.

Brady:1979:BRb

[Bra79b] J. M. Brady. Book review. The Computer Journal, 22(4):356, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/356.1.full.pdf+html.

Bramer:1979:COS

[Bra79c] M. A. Bramer. Correct and optimal strategies in game playing programs. The Computer Journal, 23(4):347–352, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/347.full.pdf+html.

Brent:1971:AGC

[Bre71] R. P. Brent. An algorithm with guaranteed convergence for finding a zero of a function. The Computer Journal, 14(4):422–425, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/4/422.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_04/140422.sgm.abs.html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/422.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/423.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/424.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/425.tif.

Barrodale:1970:CBA

[BRH70] I. Barrodale, F. D. K. Roberts, and C. R. Hunt. Computing best l_{p} approximations by functions nonlinear in one parameter. The Computer Journal, 13(4):382–386, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/382.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130382.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/382.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/383.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/384.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/385.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/386.tif.

Briggs:1979:RRT

[Bri79] T. Briggs. A review of road traffic simulation methods using a general purpose digital computer. The Computer Journal, 22(1):11– 16, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/1/11.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/13.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/16.tif.

Broyden:1969:ASN

[Bro69] C. G. Broyden. Algorithm 44: Solution of nonlinear simultaneous equations. The Computer Journal, 12(4):406–408, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_04/tiff/406.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_12/Issue_04/tiff/407.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/tiff/408.

tif. See note [Fie70, Red71b].

Broyden:1970:BR

[Bro70] C. G. Broyden. Book review. The Computer Journal, 13(4): 358, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/358.full.pdf+html.

Brons:1972:C

[Bro72a] K. A. Brons. Correspondence. The Computer Journal, 15(2):112, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/2/112.full.pdf+html.

Broyden:1972:LSS

[Bro72b] C. G. Broyden. Large sparse sets of linear equations. The Computer Journal, 15(1):70, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/1/70.full.pdf+html.

Brown:1973:DCP

[Bro73a] D. C. Brown. Discussion and correspondence: the project, and the future of Computing Science courses. The Computer Journal, 16(4):380–382, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/4/380.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/
160380.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/380.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/381.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_

04/tiff/382.tif. See correspondence [Fis74b, Bel74].

Brown:1973:PUO

[Bro73b] P. J. Brown. A practical usage of outer and inner syntax. The Computer Journal, 16(2):122–123, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/122.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/160122.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/122.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/123.tif.

Brown:1973:UPL

[Bro73c] P. J. Brown. UNRAVEL — a programming language to put intelligence into dumps. The Computer Journal, 16(1):10–12, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/1/10.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160010.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/12.tif.

Broadbent:1974:MSA

[Bro74a] J. K. Broadbent. Microprogramming and system architecture. The Computer Journal, 17(1):2–8, February 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/1/2.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/3.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/6.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/7.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/8.tif. See correspondence [WS74].

Broderick:1974:BRa

[Bro74b] W. R. Broderick. Book review. The Computer Journal, 17(1):88, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/88.2.full.pdf+html.

Broderick:1974:BRc

[Bro74c] W. R. Broderick. Book review. The Computer Journal, 17(1):92, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/92.2.full.pdf+html.

Broderick:1974:BRB

[Bro74d] W. R. Broderick. Book reviews: Spline Analysis, by Martin H. Schultz, 1972; Introduction to Programming, by A. Sharp; Algol Programming, by R. W. Sharp. The Computer Journal, 17(1): 92, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/92.tif.

Brooke:1974:BR

[Bro74e] N. M. Brooke. Book review. The Computer Journal, 17(2):171, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/171.full.pdf+html.

Brown:1974:BDE

[Bro74f] D. A. H. Brown. Biquinary decimal error detection codes with one, two and three check digits. The Computer Journal, 17(3): 201–204, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/201.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/201.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/202.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/203.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/204.tif.

Brown:1974:BR

[Bro74g] D. A. H. Brown. Book review. The Computer Journal, 17(3):210, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/210.full.pdf+html.

Brown:1974:CCR

[Bro74h] D. A. H. Brown. Correspondence: Cyclic redundancy checking by program. The Computer Journal, 17(1):94–95, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/94.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/95.tif.

Brown:1974:SEC

[Bro74i] D. A. H. Brown. Some error correcting codes for certain transposition and transcription errors in decimal integers. The Computer Journal, 17(1):9–12, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/1/9.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/12.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/9.tif.

Brown:1975:BRa

[Bro75a] D. A. H. Brown. Book review. The Computer Journal, 18(3):249, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/249.full.pdf+html.

Brown:1975:BRb

[Bro75b] P. J. Brown. Book review. The Computer Journal, 18(4):354, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
4/354.1.full.pdf+html.

Brokate:1976:C

[Bro76a] K. Brokate. Correspondence. The Computer Journal, 19(4):383, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/383.2.full.pdf+html.

Brown:1976:BR

[Bro76b] D. A. H. Brown. Book review. The Computer Journal, 19(4):332, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/332.3.full.pdf+html.

Brookes:1977:BR

[Bro77] B. C. Brookes. Book review. The Computer Journal, 20(4):358, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/358.3.full.pdf+html.

Brown:1979:AAU

[Bro79a] A. Brown. Algorithm 106: An algorithm for updating an ordered list on a VDU. The Computer Journal, 22(2):188–189, May 1979.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_02/tiff/188.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_02/tiff/189.tif.

Brown:1979:SMV

[Bro79b] P. J. Brown. Software methods for virtual storage of executable code. The Computer Journal, 22(1):50–52, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/50.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/51.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/52.tif.

Boggs:1971:NIO

[BS71] R. A. C. Boggs and F. J. Smith. A note on the integration of ordinary differential equations using Chebyshev series. The Computer Journal, 14(3):270–271, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/270.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140270.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/270.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/271.tif.

Bordier:1973:NSC

[BS73a] J. Bordier and H. Saya. A necessary and sufficient condition for a power language to be LL(k). The Computer Journal, 16(4):351– 356, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/351.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160351.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/352.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/353.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/355.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/356.tif.

Burstall:1973:BRB

[BS73b] R. M. Burstall and P. A. Samet. Book reviews: Representation and Meaning, by Herbert A. Simon and Laurent Siklossy, 1972; A Collection of Programming Problems and Techniques, by H. A. Maurer and M. R. Williams, 1972. The Computer Journal, 16(3): 253, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/253.tif.

Bezdek:1975:NCM

[BS75] R. H. Bezdek and A. H. Sameh. A note on computational methods for input-output econometric models. The Computer Journal, 18(3): 248–249, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/248.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/248.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/249.tif.

Burdekin:1979:CCE

[BS79] R. Burdekin and R. S. Scowen. Computers and the construction of economic models. The Computer Journal, 22 (1):3–10, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/1/3.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/3.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/5.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/6.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/7.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/8.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/9.tif.

Buckley-Sharp:1970:CPB

[BSH70] M. D. Buckley-Sharp and F. T. C. Harris. A computer program for banking multiple choice questions. The Computer Journal, 13(3): 230–236, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/230.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_13/Issue_03/130230.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/230.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/231.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/232.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/233.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/234.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/235.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/236.tif.

Baldwin:1971:MAC

[BT71] J. C. Baldwin and R. E. Thomas. Multi-access on the Chilton Atlas. The Computer Journal, 14(2):119–122, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/119.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140119.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/119.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/120.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/121.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/122.
tif.

Buckle:1976:C

[Buc76] J. K. Buckle. Correspondence. The Computer Journal, 19(4):381, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/381.2.full.pdf+html.

Bull:1972:DDB

[Bul72] G. M. Bull. Dynamic debugging in BASIC. The Computer Journal, 15(1):21–24, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/1/21.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/150021.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/21.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/22.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_01/tiff/24.tif. See correspon-

dence [Sil72].

Bull:1973:BR

[Bul73] G. M. Bull. Book review. The Computer Journal, 16(4):374, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/374.2.full.pdf+html.

Bumstead:1978:BR

[Bum78] D. Bumstead. Book review. The Computer Journal, 21(1):78, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/78.full.pdf+html.

Bumstead:1979:BR

[Bum79] D. Bumstead. Book review. The Computer Journal, 22(2):154, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/154.full.pdf+html.

Burstall:1970:BRa

[Bur70a] R. M. Burstall. Book review. The Computer Journal, 13(2):130, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/130.full.pdf+html.

Burstall:1970:BRb

[Bur70b] R. M. Burstall. Book review. The Computer Journal, 13(2):177, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/177.full.pdf+html.

Burstall:1971:BR

[Bur71] R. M. Burstall. Book review. The Computer Journal, 14(4):381, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/381.full.pdf+html.

Burgess:1972:CTE

[Bur72] C. J. Burgess. Compile-time error diagnostics in syntax-directed compilers. The Computer Journal, 15(4):302–307, November 1972.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/302.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150302.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/302.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/304.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/305.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/306.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/307.tif. See correspon-

dence [Mux73].

Burstall:1973:BRa

[Bur73a] R. M. Burstall. Book review. The Computer Journal, 16(2):105, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/105.2.full.pdf+html.

Burstall:1973:BRc

[Bur73b] R. M. Burstall. Book review. The Computer Journal, 16(3):253, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/3/253.1.full.pdf+html.

Burkhard:1975:FTQ

[Bur75a] W. A. Burkhard. Full table quadratic quotient searching. The Computer Journal, 18(2):161–163, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/161.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/163.tif.

Burkhard:1975:NTT

[Bur75b] W. A. Burkhard. Nonrecursive traversals of trees. The Computer Journal, 18(3):227–230, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/3/227.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/227.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_18/Issue_03/tiff/228.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/229.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/230.tif.

Burstall:1975:BR

[Bur75c] R. M. Burstall. Book review. The Computer Journal, 18(3):197, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/197.full.pdf+html.

Burton:1979:FPS

[Bur79] W. Burton. A FORTRAN preprocessor to support encapsulated data abstraction definitions. The Computer Journal, 22(4):307– 312, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/4/307.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/307.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/308.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/309.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/310.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/311.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/312.tif.

Bromley:1974:CTL

[BVH74] A. G. Bromley, Z. G. Vranesic, and V. C. Hamacher. Correspondence: Ternary logic in parallel multipliers. The Computer Journal, 17(2): 189–190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/17/2/189.2.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_17/Issue_02/tiff/189.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/190.tif. See [VH72].

Bernal:1970:NTB

[BW70a] M. J. M. Bernal and J. R. Whiteman. Numerical treatment of biharmonic boundary value problems with re-entrant boundaries. The Computer Journal, 13(1):87–91, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
13/1/87.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_01/130087.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/87.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/89.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/90.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/91.tif.

Boulton:1970:PNC

[BW70b] D. M. Boulton and C. S. Wallace. A program for numerical classification. The Computer Journal, 13(1):63–69, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/1/63.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130063.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/63.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/64.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/65.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/66.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/67.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/68.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/69.tif.

Beaubien:1971:IFD

[BW71] M. J. Beaubien and A. Wexler. Iterative, finite difference solution of interior eigenvalues and eigenfunctions of Laplace’s operator. The Computer Journal, 14(3):263–269, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/263.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140263.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/265.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/268.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/269.
tif.

Boulton:1973:IMH

[BW73a] D. M. Boulton and C. S. Wallace. An information measure for hierarchic classification. The Computer Journal, 16(3):254–261, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/254.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/160254.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/254.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/255.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/256.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/257.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/258.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/259.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/261.tif.

Boulton:1973:ORA

[BW73b] D. M. Boulton and C. S. Wallace. Occupancy of a rectangular array. The Computer Journal, 16(1):57–63, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
16/1/57.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160057.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/57.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/59.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/60.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/62.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/63.tif.

Boulton:1975:IMS

[BW75a] D. M. Boulton and C. S. Wallace. An information measure for single link classification. The Computer Journal, 18(3):236–238, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/236.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/236.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/237.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/238.tif.

Buckholtz:1975:APR

[BW75b] T. J. Buckholtz and C. S. Wetherell. Algorithm 89: a program to referee Kriegspiel and chess. The Computer Journal, 18(2):177–183, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_18/Issue_02/tiff/177.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/178.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/179.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/180.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/181.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/182.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/183.tif.

Bochmann:1978:CWS

[BW78] Gregor V. Bochmann and P. Ward. Compiler writing system for attribute grammars. The Computer Journal, 21(2):144–148, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/144.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/144.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/145.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/147.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/148.tif.

See also: publication 199, Dépt. d’Informatique, University de Montréal (July 1975).

Barton:1971:ASS

[BWZ71] David Barton, I. M. Willers, and R. V. M. Zahar. The automatic solution of systems of ordinary differential equations by the method of Taylor series. The Computer Journal, 14(3):243–248, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/243.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140243.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/243.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/245.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/246.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/248.tif.

Ben-Yaacov:1979:ICD

[BY79] G. Z. Ben-Yaacov. Interactive computation and data management for power system studies. The Computer Journal, 22(1):76–79, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/76.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/76.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/77.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/79.tif.

Cairns:1976:BR

[Cai76] D. J. Cairns. Book review. The Computer Journal, 19(1):20, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/20.full.pdf+html.

Campbell:1971:NOF

[Cam71] J. A. Campbell. A note on an optimal-fit method for dynamic allocation of storage. The Computer Journal, 14(1):7–9, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/7.full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/140007.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/7.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/8.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/9.tif.

Campbell:1973:ESA

[Cam73] D. V. A. Campbell. Errata: “A simple analysis of the n^{th} order polyphase sort”. The Computer Journal, 16(3):244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/244.tif. See [Mac73].

Campbell:1975:C

[Cam75] J. L. Campbell. Correspondence. The Computer Journal, 18(2): 190–191, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/190.3.full.pdf+html.

Campbell:1978:CDF

[Cam78] W. R. Campbell. A compiler definition facility based on the syntactic macro. The Computer Journal, 21(1):35–41, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/35.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/35.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/36.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/37.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/38.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/39.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/40.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/41.tif.

Carrick:1970:CIL

[Car70] D. A. Carrick. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 13(2):220–221, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/220.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/220.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/221.tif.

See [Kin69, Kin71b].

Cash:1975:NSL

[Cas75] J. R. Cash. The numerical solution of linear stiff boundary value problems. The Computer Journal, 18(4):371–372, November 1975.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/371.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/371.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/372.tif.

Castell:1976:C

[Cas76] S. P. Castell. Correspondence. The Computer Journal, 19(4):382, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/382.1.full.pdf+html.

Coles:1976:ACP

[CBB^{+}76] E. C. Coles, L. J. Beilin, C. J. Bulpitt, C. T. Dollery, B. F. Johnson, A. D. Munro-Faure, and S. C. Turner. An approach to computer processing of clinical data. The Computer Journal, 19(2): 104–109, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/104.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/104.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/105.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/106.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/107.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/108.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/109.tif.

Cohen:1974:NDF

[CC74] J. Cohen and Eileen Carton. Non-deterministic Fortran. The Computer Journal, 17(1):44–51, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/44.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/45.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/47.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/48.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_

01/tiff/50.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/51.tif.

Chaplin:1973:GRB

[CCH73] R. I. Chaplin, R. E. Crosbie, and J. L. Hay. A graphical representation of the Backus-Naur form. The Computer Journal, 16(1): 28–29, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/28.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/160028.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/28.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/29.tif.

Clare:1972:IVL

[CCL72] A. C. Clare, Elizabeth M. Cook, and M. F. Lynch. The identification of variable-length, equifrequent character strings in a natural language data base. The Computer Journal, 15(3):259– 262, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/259.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/150259.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/260.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/261.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/262.tif.

Carpenter:1977:OTM

[CD77] B. E. Carpenter and R. W. Doran. The other Turing machine. The Computer Journal, 20(3):269–279, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/269.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/269.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/270.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/271.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/272.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/273.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/274.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_20/Issue_03/tiff/275.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/276.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/277.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/278.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/279.tif.

Hodges page 318 note 6.1.

Chau:1974:IIG

[CDZ74] A. Y. C. Chau, B. W. Davies, and B. Zacharov. Island — an interactive graphics system for mathematical analysis. The Computer Journal, 17(2):104–112, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/104.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/105.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/106.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/107.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/108.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/109.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/110.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/111.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/112.tif.

Coulouris:1972:TCA

[CEM72] G. F. Coulouris, J. M. Evans, and R. W. Mitchell. Towards contentaddressing in data bases. The Computer Journal, 15(2):95–98, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/2/95.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/150095.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/95.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_02/tiff/96.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/tiff/97.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
tiff/98.tif.

Chawla:1970:EEG

[Cha70] M. M. Chawla. Estimation of errors of Gauss–Chebyshev quadratures. The Computer Journal, 13(1):107–109, February 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/1/107.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/130107.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/107.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/108.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/109.tif.

Chambers:1971:ARF

[Cha71] J. M. Chambers. Another round of Fortran. The Computer Journal, 14(3):312–314, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/312.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140312.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/312.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/313.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_

03/tiff/314.tif. See correspondence [Dew72c, Fla72b, Cha72b].

Chambers:1972:C

[Cha72a] J. M. Chambers. Correspondence. The Computer Journal, 15(4):381, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/4/381.full.pdf+html.

Chambers:1972:CAR

[Cha72b] J. M. Chambers. Correspondence: On “Another round of FOR-TRAN”. The Computer Journal, 15(4):381, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/150381.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_04/tiff/381.tif. See

[Cha71, Dew72c, Fla72b].

Chance:1973:BR

[Cha73a] E. M. Chance. Book review. The Computer Journal, 16(3):265, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/3/265.2.full.pdf+html.

Chance:1973:BRB

[Cha73b] E. M. Chance. Book review: FORTRAN Techniques, by A. Colin Day, 1972. The Computer Journal, 16(3):265, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/265.tif.

Chambers:1975:ATP

[Cha75] A. D. Chambers. Audit test packs and computer audit programs. The Computer Journal, 18(2):98–101, May 1975. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/2/98.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/100.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/101.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/98.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/99.tif.

Chantler:1977:BRa

[Cha77a] A. Chantler. Book review. The Computer Journal, 20(2):169, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/169.3.full.pdf+html.

Chantler:1977:BRb

[Cha77b] A. Chantler. Book review. The Computer Journal, 20(2):169, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/169.4.full.pdf+html.

Chambers:1978:CFA

[Cha78a] A. D. Chambers. Computer fraud and abuse. The Computer Journal, 21(3):194–198, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/3/194.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/194.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/195.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/196.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/197.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/198.tif.

Chantler:1978:BR

[Cha78b] A. E. Chantler. Book review. The Computer Journal, 21(4):358, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/4/358.2.full.pdf+html.

Chantler:1979:BRa

[Cha79a] A. Chantler. Book review. The Computer Journal, 22(1):36, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/36.1.full.pdf+html.

Chantler:1979:BRb

[Cha79b] A. Chantler. Book review. The Computer Journal, 22(1):36, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/36.2.full.pdf+html.

Cherry:1971:BR

[Che71] E. C. Cherry. Book review. The Computer Journal, 14(2):144, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
2/144.full.pdf+html.

Coulson:1973:ASP

[CHH73] V. G. Coulson, I. D. Hill, and V. F. Hillier. Algol survey program. The Computer Journal, 16(1):13–15, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
16/1/13.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160013.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/13.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/15.tif.

Christofides:1971:ACN

[Chr71a] N. Christofides. An algorithm for the chromatic number of a graph. The Computer Journal, 14(1):38–39, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
14/1/38.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_01/140038.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/38.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/39.tif.

Christofides:1971:ZOP

[Chr71b] N. Christofides. Zero-one programming using non-binary treesearch. The Computer Journal, 14(4):418–421, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/418.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140418.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/418.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/419.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/420.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/421.
tif.

Christiansen:1973:ASS

[Chr73a] Jan Christiansen. Algorithm 77: Solving a system of simultaneous ordinary differential equations of the first order using a method for automatic step change. The Computer Journal, 16(2):187–188, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_16/Issue_02/tiff/187.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/188.tif.

Christofides:1973:LSP

[Chr73b] N. Christofides. Large scheduling problems with bivalent costs. The Computer Journal, 16(3):262–264, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/262.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160262.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/262.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/263.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/264.tif.

Clint:1970:EEE

[CJ70] M. Clint and A. Jennings. The evaluation of eigenvalues and eigenvectors of real symmetric matrices by simultaneous iteration. The Computer Journal, 13(1):76–80, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
13/1/76.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130076.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/76.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/78.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/79.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/80.tif. See erratum [Sea71].

Cash:1978:CAS

[CL78] J. R. Cash and C. B. Liem. On the computational aspects of semiimplicit Runge–Kutta methods. The Computer Journal, 21(4):363– 365, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/363.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/363.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/364.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/365.tif.

Clarke:1976:BRa

[Cla76a] P. A. Clarke. Book review. The Computer Journal, 19(4):352, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/352.1.full.pdf+html.

Clarke:1976:BRb

[Cla76b] P. A. Clarke. Book review. The Computer Journal, 19(4):352, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/352.2.full.pdf+html.

Clarkson:1976:BR

[Cla76c] B. L. Clarkson. Book review. The Computer Journal, 19(3):288, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/288.1.full.pdf+html.

Clapson:1977:TPS

[Cla77a] P. J. Clapson. Toward a performance science: a comparative analysis of computing systems. The Computer Journal, 20(4):308– 315, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/308.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/308.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/309.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/311.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/312.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/314.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/315.tif.

Clasen:1977:NUC

[Cla77b] R. J. Clasen. A note on the use of the conjugate gradient method in the solution of a large system of sparse equations. The Computer Journal, 20(2):185–186, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/185.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/185.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/186.tif.

Clark:1978:SSA

[Cla78] I. A. Clark. Strema: Specifying application processes using streams. The Computer Journal, 21(1):25–30, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
21/1/25.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/27.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/30.tif.

Clarke:1979:BR

[Cla79] P. A. Clarke. Book review. The Computer Journal, 22(4):306, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/306.1.full.pdf+html.

Cleave:1975:BR

[Cle75] J. P. Cleave. Book review. The Computer Journal, 18(2):147, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/147.full.pdf+html.

Cleave:1977:BR

[Cle77] J. P. Cleave. Book review. The Computer Journal, 20(2):101, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/101.1.full.pdf+html.

Clowes:1977:BR

[Clo77] M. Clowes. Book review. The Computer Journal, 20(4):345, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/345.2.full.pdf+html.

Clowes:1979:RCM

[Clo79] M. Clowes. Review: Cad meets ai. The Computer Journal, 22(4):322, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/322.1.full.pdf+html.

Cluley:1977:BRa

[Clu77a] J. C. Cluley. Book review. The Computer Journal, 20(2):150, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/150.2.full.pdf+html.

Cluley:1977:BRb

[Clu77b] J. C. Cluley. Book review. The Computer Journal, 20(4):301, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/301.3.full.pdf+html.

Cash:1979:IAN

[CM79] J. R. Cash and J. C. P. Miller. On an iterative approach to the numerical solution of difference schemes. The Computer Journal, 22 (2):184–187, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/2/184.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/184.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/185.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/186.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/187.tif.

Capon:1972:MCT

[CMRW72] P. C. Capon, D. Morris, J. S. Rohl, and I. R. Wilson. The MU5 compiler target language and autocode. The Computer Journal, 15 (2):109–112, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/109.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/150109.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/110.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/111.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/112.tif.

Cunningham:1972:EHG

[CO72] Kathrine M. Cunningham and J. C. Ogilvie. Evaluation of hierarchical grouping techniques; a preliminary study. The Computer Journal, 15(3):209–213, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/3/209.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150209.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/209.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/210.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/211.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/212.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/213.tif.

Casey:1976:RAS

[CO76] R. G. Casey and I. M. Osman. Replacement algorithms for storage management in relational data bases. The Computer Journal, 19(4):306–314, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/4/306.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/306.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/307.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/308.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/309.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/310.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/311.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/312.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/314.tif.

Coates:1974:DMC

[Coa74] H. N. Coates. Discussion: Management in the computer business. The Computer Journal, 17(3):279–282, August 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/279.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/279.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/280.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/281.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/282.tif.

Coates:1977:BR

[Coa77] M. P. Coates. Book review. The Computer Journal, 20(4):329, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/329.2.full.pdf+html.

Coan:1978:CTP

[Coa78] D. R. A. Coan. Current trends in programming methodology and standardised development of computer software. The Computer Journal, 21(2):160, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/160.full.pdf+html.

Cohn:1970:BR

[Coh70] P. M. Cohn. Book review. The Computer Journal, 13(3):308, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
3/308.full.pdf+html.

Colin:1970:DTF

[Col70] A. J. T. Colin. DOLPHIN — a text filing system for University use. The Computer Journal, 13(2):136–141, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/136.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130136.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/136.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/137.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/138.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/139.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/140.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/141.tif.

Coleman:1975:DIT

[Col75] P. W. F. Coleman. Discussion: Integrated training for the blind programmer. The Computer Journal, 18(2):186–187, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/186.full.pdf+html.

Colin:1976:PSP

[Col76] A. J. T. Colin. The principles of systems programming. The Computer Journal, 19(2):159, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/159.full.pdf+html.

Colin:1977:BRa

[Col77a] A. J. T. Colin. Book review. The Computer Journal, 20(1):77, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/77.2.full.pdf+html.

Colin:1977:BRb

[Col77b] A. J. T. Colin. Book review. The Computer Journal, 20(3):212, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/3/212.1.full.pdf+html.

Coles:1978:BR

[Col78] E. C. Coles. Book review. The Computer Journal, 21(3):285, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/285.3.full.pdf+html.

Comerford:1975:LSMa

[Com75a] P. J. Comerford. Link systems for multi-computer control of a large process: Part 1. The Computer Journal, 18(2):144–147, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/144.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/144.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/145.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/147.tif.

Comerford:1975:LSMb

[Com75b] P. J. Comerford. Link systems for multi-computer control of a large process: Part 2: Fault detection. The Computer Journal, 18(3): 220–222, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/220.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/220.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/221.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/222.tif.

Conolly:1976:BR

[Con76] B. W. Conolly. Book review. The Computer Journal, 19(3):267, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/267.full.pdf+html.

Conolly:1977:BRb

[Con77a] B. Conolly. Book review. The Computer Journal, 20(4):295, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/295.2.full.pdf+html.

Conolly:1977:BRa

[Con77b] B. W. Conolly. Book review. The Computer Journal, 20(1):67, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/67.3.full.pdf+html.

Cooper:1971:BR

[Coo71] D. C. Cooper. Book review. The Computer Journal, 14(4):343, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/343.2.full.pdf+html.

Cook:1972:C

[Coo72a] D. B. Cook. Correspondence. The Computer Journal, 15(4):361, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/4/361.1.full.pdf+html.

Cook:1972:CPF

[Coo72b] D. B. Cook. Correspondence: ‘Packing’ in FORTRAN. The Computer Journal, 15(4):361, November 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
tiff/361.tif.

Cooper:1972:AIH

[Coo72c] D. C. Cooper. Artificial intelligence and heuristic programming. The Computer Journal, 15(2):169, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/169.full.pdf+html.

Cooper:1972:ECA

[Coo72d] D. C. Cooper. Essays on cellular automata. The Computer Journal, 15(1):83, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/83.full.pdf+html.

Cooper:1972:C

[Coo72e] W. Cooper. Correspondence. The Computer Journal, 15(3):246, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/246.full.pdf+html.

Cooper:1972:CQI

[Coo72f] W. Cooper. Correspondence: On “A quasi-intrinsic scheme for passing a smooth curve through a discrete set of points”. The Computer Journal, 15(3):246, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150246.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/246.tif. See [McC70b, Ecc71b,

Smi71, Sab72].

Cooper:1975:TDM

[Coo75] B. Cooper. Time-domain matrix evaluation of a digital servomechanism using semi-hybrid computational techniques. The Computer Journal, 18(2):169–172, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/169.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/169.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/170.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/171.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/172.tif.

Cooper:1976:BR

[Coo76a] D. C. Cooper. Book review. The Computer Journal, 19(1):24, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/24.2.full.pdf+html.

Cooper:1976:DOC

[Coo76b] D. C. Cooper. The design of an optimizing compiler. The Computer Journal, 19(2):150a–150, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/150-a.full.pdf+html.

Corner:1979:BRB

[Cor79] Lewis Corner. Book reviews: Model Building in Mathematical Programming, by H. P. Williams, 1978; Computer Simulation and Modeling: An Introduction, by R. S. Lehman. The Computer Journal, 22(1):2, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/2.tif.

Cox:1970:BTC

[Cox70] M. G. Cox. A bracketing technique for computing a zero of a function. The Computer Journal, 13(1):101–102, February 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/1/101.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/130101.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/101.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/102.tif. See note [SC71].

Cox:1971:AAC

[Cox71] M. G. Cox. An algorithm for approximating convex functions by means of first degree splines. The Computer Journal, 14(3):272– 275, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/3/272.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/140272.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/273.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/274.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/275.tif.

Cranley:1971:ANE

[CP71] R. Cranley and T. N. L. Patterson. On the automatic numerical evaluation of definite integrals. The Computer Journal, 14(2): 189–198, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/189.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140189.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/189.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/190.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_02/tiff/191.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/192.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/193.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/194.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/195.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/196.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/197.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_

02/tiff/198.tif. See correspondence [Val71b].

Chattergy:1979:DFC

[CP79] R. Chattergy and U. W. Pooch. A distributed function computer with dedicated processors. The Computer Journal, 22(1):37–40, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/37.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/37.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/40.tif.

Chung-Phillips:1975:NDD

[CPR75] Alice Chung-Phillips and R. W. Rosen. A note on dynamic data storage in Fortran IV. The Computer Journal, 18(4):342–343, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/4/342.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/342.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/343.tif.

Craggs:1971:BR

[Cra71] J. W. Craggs. Book review. The Computer Journal, 14(2):122, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
2/122.full.pdf+html.

Crane:1972:ACP

[Cra72] C. M. Crane. Algorithm 75: Contour plotting for functions specified at nodal points of an irregular mesh based on an arbitrary two-parameter coordinate system. The Computer Journal, 15(4):382–384, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_15/Issue_04/tiff/382.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/383.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/384.tif.

Cross:1974:NSD

[Cro74] M. J. Cross. Nimbus 4 SCR data handling. The Computer Journal, 17 (2):172–180, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/172.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/172.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/173.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/174.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/175.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/176.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/177.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/178.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/179.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/180.tif.

Crombie:1975:CCH

[Cro75] D. B. Crombie. Correspondence: Comparison of hybrid and digital computation performances for distillation simulation. The Computer Journal, 18(3):285, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/18/3/285.2.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_03/tiff/285.tif.

Crossley:1976:C

[Cro76a] A. D. Crossley. Correspondence. The Computer Journal, 19(1):95, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/95.2.full.pdf+html.

Croston:1976:C

[Cro76b] J. K. Croston. Correspondence. The Computer Journal, 19(4):381, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/381.1.full.pdf+html.

Crocker:1977:BR

[Cro77] P. Crocker. Book review. The Computer Journal, 20(3):212, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/212.2.full.pdf+html.

Cruise:1978:BR

[Cru78] A. M. Cruise. Book review. The Computer Journal, 21(3):223, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/223.2.full.pdf+html.

Cardenas:1975:SDS

[CS75] A. F. Cárdenas and R. W. Seeley. A simple data structure for interactive graphic design/drafting. The Computer Journal, 18(1): 30–33, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/18/1/30.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/31.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/32.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/33.tif.

Cardenas:1977:DCT

[CS77] Alfonso F. Cardenas and James P. Sagamang. Doubly-chained tree data base organisation — analysis and design strategies. The Computer Journal, 20(1):15–26, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/1/15.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/16.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/19.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/21.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_20/Issue_01/tiff/22.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/24.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/25.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/26.tif.

Coates:1974:CSB

[CT74] H. N. Coates and The Editor. Correspondence: The silent business user. The Computer Journal, 17(1):96, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/96.tif. See [TT73].

Cuff:1972:CCF

[Cuf72] R. N. Cuff. A conversational compiler for full PL/I. The Computer Journal, 15(2):99–104, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/2/99.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/150099.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/100.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/101.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/103.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/104.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/99.tif.

Cunningham:1977:BR

[Cun77] A. Cunningham. Book review. The Computer Journal, 20(1):55, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/55.2.full.pdf+html.

Curran:1978:BRB

[Cur78] M. P. J. Curran. Book review: Computer Methods for Mathematical Computations, by G. Forsythe, M. Malcolm, and C. Moler, 1977; 259 pages. (Prentice/Hall, £12.80). The Computer Journal, 21(4):336, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/4/336.2.full.pdf+html.

Cole:1970:IAJ

[CW70] A. J. Cole and D. Wishart. An improved algorithm for the Jardine-Sibson method of generating overlapping clusters. The Computer Journal, 13(2):156–163, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/2/156.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/
130156.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/156.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/157.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/158.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/159.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/160.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/161.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/162.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/163.tif.

Coggan:1971:LSE

[CW71] G. C. Coggan and J. A. Wilson. On-line state estimation with a small computer. The Computer Journal, 14(1):61–64, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/61.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140061.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/61.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/62.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/63.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/64.tif.

Curnow:1976:SB

[CW76] H. J. Curnow and B. A. Wichmann. A synthetic benchmark. The Computer Journal, 19(1):43–49, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
19/1/43.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/43.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/44.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/45.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_19/Issue_01/tiff/46.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/47.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/48.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/49.tif.

Delves:1977:AFG

[DA77] L. M. Delves and L. F. Abd. El Al. Algorithm 97: The fast Galerkin algorithm for the solution of linear Fredholm equations. The Computer Journal, 20(4):374–376, November 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/374.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/376.tif.

Dakin:1973:C

[Dak73a] R. J. Dakin. Correspondence. The Computer Journal, 16(1):18, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/18.1.full.pdf+html.

Dakin:1973:CIF

[Dak73b] R. J. Dakin. Correspondence: Impressions following the BCS Conference on Computer Performance. The Computer Journal, 16(1): 18, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/18.tif. See correspon-

dence [Gla73].

Dakin:1975:GCL

[Dak75] R. J. Dakin. A general control language: language structure and translation. The Computer Journal, 18(4):324–332, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/324.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/324.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/325.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/326.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/327.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/328.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/329.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/330.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/331.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/332.tif.

Davenport:1976:DI

[Dav76] R. A. Davenport. Database integrity. The Computer Journal, 19 (2):110–116, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/110.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/110.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/111.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/112.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/113.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/114.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/115.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/116.tif.

Davies:1977:BR

[Dav77] G. Davies. Book review. The Computer Journal, 20(1):14, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/14.2.full.pdf+html.

Davenport:1978:DCD

[Dav78] R. A. Davenport. Distributed or centralised data base. The Computer Journal, 21(1):7–14, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/1/7.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_01/tiff/12.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/
tiff/13.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_01/tiff/7.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/

tiff/8.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/9.tif.

Dawson:1973:CIC

[Daw73] J. L. Dawson. Combining interpretive code with machine code. The Computer Journal, 16(3):216–219, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/216.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160216.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/216.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/217.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/218.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/219.

tif. See correspondence [Mac74b, Bem74].

Dawkins:1974:CBC

[Daw74] B. W. Dawkins. Correspondence: The Computer Journal. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/17/2/190.2.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_02/tiff/190.tif.

Day:1970:USS

[Day70] A. C. Day. The use of symbol-state tables. The Computer Journal, 13(4):332–339, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/4/332.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/
130332.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/332.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/333.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/335.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/336.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/338.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/339.tif.

See note [Lan71b].

Day:1973:BRa

[Day73a] A. C. Day. Book review. The Computer Journal, 16(1):68, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/68.full.pdf+html.

Day:1973:BRb

[Day73b] A. C. Day. Book review. The Computer Journal, 16(4):330, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/330.full.pdf+html.

Day:1975:BR

[Day75] A. C. Day. Book review. The Computer Journal, 18(2):130, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/130.2.full.pdf+html.

Day:1976:ABB

[Day76] A. C. Day. Algorithm 91: Balancing a binary tree. The Computer Journal, 19(4):360–361, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/360.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/361.tif.

Day:1977:AGP

[Day77] A. C. Day. Algorithm 95: Generating a parity testing table. The Computer Journal, 20(3):280, August 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/280.tif.

Day:1978:ASS

[Day78a] A. C. Day. Algorithms supplement: Statement of policy. The Computer Journal, 21(4):377, November 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/377.tif.

Day:1978:BR

[Day78b] C. Day. Book review. The Computer Journal, 21(1):85, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/85.1.full.pdf+html.

Day:1979:BR

[Day79] A. C. Day. Book review. The Computer Journal, 22(4):356, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/356.2.full.pdf+html.

Day:1976:DCP

[DCHR76] A. C. Day, P. A. Clarke, D. Hill, and J. K. Reid. Discussion and correspondence: the proposed new standard for Fortran: a critical examination. The Computer Journal, 19(3):268–271, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/268.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/268.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/269.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/271.tif.

deChampeaux:1977:OTB

[dCS77] D. de Champeaux and L. Sint. An optimality theorem for a bidirectional heuristic search algorithm. The Computer Journal, 20 (2):148–150, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/148.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/148.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/149.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/150.tif.

Dawkins:1970:CUF

[DD70] B. W. Dawkins and H. C. Dawkins. Correspondence: On “A University faculty timetable”. The Computer Journal, 13 (1):121–122, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_01/tiff/121.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/122.tif. See [Alm69].

Dearnley:1974:MSO

[Dea74a] P. A. Dearnley. A model of a self-organising data management system. The Computer Journal, 17(1):13–16, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/13.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/13.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/tiff/14.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_01/tiff/15.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/16.tif.

Dearnley:1974:OMS

[Dea74b] P. A. Dearnley. The operation of a model self organising data management system. The Computer Journal, 17(3):205–210, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/205.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/205.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/206.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/208.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/209.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/210.tif.

Dearnley:1976:IDR

[Dea76] P. A. Dearnley. An investigation into database resilience. The Computer Journal, 19(2):117–121, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/117.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/117.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/118.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/119.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/120.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/121.tif.

Dearnley:1978:MDS

[Dea78] P. Dearnley. Monitoring database system performance. The Computer Journal, 21(1):15–19, February 1978. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/1/15.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/16.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/17.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/19.tif.

Dearnley:1979:ALM

[Dea79] P. A. Dearnley. Application level microcode to speed data base management. The Computer Journal, 22(3):200–202, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/200.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/200.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/201.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/202.tif.

Defays:1977:EAC

[Def77] D. Defays. An efficient algorithm for a complete link method. The Computer Journal, 20(4):364–366, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/364.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/364.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/365.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/366.tif.

Delves:1977:ALE

[Del77] L. M. Delves. Algorithm 96: a linear equation solver for Galerkin and least squares methods. The Computer Journal, 20 (4):371–374, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_20/Issue_04/tiff/371.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/372.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/373.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/374.tif.

Denham:1975:BRa

[Den75a] M. J. Denham. Book review. The Computer Journal, 18(4):323, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/323.2.full.pdf+html.

Denham:1975:BRb

[Den75b] M. J. Denham. Book review. The Computer Journal, 18(4):341, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/341.1.full.pdf+html.

Denvir:1975:C

[Den75c] B. T. Denvir. Correspondence. The Computer Journal, 18(4):380, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/380.1.full.pdf+html.

Deransart:1979:PSA

[Der79] Pierre Deransart. Proof by semantic attributes of a Lisp compiler. The Computer Journal, 22(3):240–245, August 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/240.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/240.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/241.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/242.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/243.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/244.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/245.tif.

Deutsch:1969:CCL

[Deu69] E. S. Deutsch. Correspondence: Comments on a line thinning algorithm. The Computer Journal, 12(4):412, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/12/
4/412.2.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/412.tif. See correspon-

dence [Fra70].

Dewar:1971:API

[Dew71] J. K. S. Dewar. Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 14(4):447–450, November 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
tiff/447.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/448.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_04/tiff/449.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
tiff/450.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/451.tif. See note [Dew72a, Dew74].

Dewar:1972:NAP

[Dew72a] J. K. S. Dewar. Note on Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 15(3):285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_03/tiff/285.tif. See [Dew71, Dew74].

Dewar:1972:C

[Dew72b] R. B. K. Dewar. Correspondence. The Computer Journal, 15(1): 7, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/7.full.pdf+html.

Dewar:1972:CAR

[Dew72c] R. B. K. Dewar. Correspondence: On “Another round of FOR-TRAN”. The Computer Journal, 15(1):7, February 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/150007.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/7.tif. See

[Cha71, Fla72b, Cha72b].

Dewar:1974:NAP

[Dew74] J. K. S. Dewar. Note on Algorithm 71: Procedures for interval arithmetic. The Computer Journal, 17(4):379–380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_04/tiff/380.tif. See [Dew71,

Dew72a].

Dewis:1975:BR

[Dew75] I. G. Dewis. Book review. The Computer Journal, 18(2):130, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/130.1.full.pdf+html.

DiManzo:1979:LOP

[DFO79] M. Di Manzo, A. L. Frisiani, and G. Olimpo. Loop optimisation for parallel processing. The Computer Journal, 22(3):234– 239, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/3/234.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/234.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/235.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/237.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/238.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/239.tif.

DeMorgan:1976:SAR

[DHW76] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A supplement to the Algol 60 revised report. The Computer Journal, 19(3): 276–288, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/3/276.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/276.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/277.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/278.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/279.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/280.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/281.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/282.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/283.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/285.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/286.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_19/Issue_03/tiff/287.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/288.tif.

DeMorgan:1978:MAS

[DHW78] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. Modified Algol 60 and the step-until element. The Computer Journal, 21(3): 282–285, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/282.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/282.tif.

Dwyer:1978:CSP

[DI78] B. Dwyer and J. Inglis. Correspondence: On “Structured programming and input statements”. The Computer Journal, 21(2):190–191, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_21/Issue_02/tiff/190.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_

02/tiff/191.tif. See [Ing76, Hop78].

Dickinson:1975:BR

[Dic75] C. J. Dickinson. Book review. The Computer Journal, 18(1):33, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/33.3.full.pdf+html.

Dirac:1970:BR

[Dir70] G. A. Dirac. Book review. The Computer Journal, 13(4):412, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/4/412.full.pdf+html.

Dives:1974:BR

[Div74] D. F. Dives. Book review. The Computer Journal, 17(1):8, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/8.2.full.pdf+html.

Dvorak:1974:DGG

[DK74] S. Dvorák and B. Kropác. Decomposition of the Gozinto’s graph with the use of a nesting store. The Computer Journal, 17(3): 245–248, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/245.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/245.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/247.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/248.tif.

Devillers:1976:IPF

[DL76] R. Devillers and G. Louchard. Improvement of parallelism in a finite buffer sharing policy. The Computer Journal, 19(3):238– 245, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/3/238.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/238.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/239.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/240.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/241.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/242.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/244.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/245.tif.

Dos:1972:ODU

[DO72] K. Dos and H. Otto. Optimal dynamic use of memory for PL/1 object programs in a real memory environment. The Computer Journal, 15(1):18–20, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/1/18.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/150018.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/18.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/19.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/20.tif.

Donovan:1978:NPV

[Don78] J. J. Donovan. A note on a performance of VM/370 in the integration of models and data bases. The Computer Journal, 21(1): 20–24, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/21/1/20.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/20.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/21.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/22.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/23.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/24.tif.

Dorling:1976:EDH

[Dor76] E. B. Dorling. Environmental data handling. The Computer Journal, 19(2):172b–172, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/172-b.full.pdf+html.

Downing:1971:CCC

[Dow71] J. Downing. A computer controlled circulation system. The Computer Journal, 14(1):2–6, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
140002.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_01/tiff/3.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_01/tiff/6.tif.

Dakin:1973:MCA

[DP73] R. J. Dakin and P. C. Poole. A mixed code approach. The Computer Journal, 16(3):219–222, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/219.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160219.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/219.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/221.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/222.

tif. See correspondence [Mac74b, Bem74].

Dance:1976:ALD

[DP76] D. L. Dance and U. W. Pooch. An adaptive on line data compression system. The Computer Journal, 19(3):216–224, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/216.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/216.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/217.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/218.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/219.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/220.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/221.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/222.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/223.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/224.tif.

deRoever:1977:BR

[dR77] W. P. de Roever. Book review. The Computer Journal, 20(2):136, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/136.1.full.pdf+html.

dInverno:1974:CFS

[dRC74] R. A. d’Inverno and R. A. Russell-Clark. CLAM — its function, structure and implementation. The Computer Journal, 17(3):229– 233, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/229.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/229.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/230.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/231.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/232.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/233.tif.

Duff:1972:PLM

[Duf72] M. J. B. Duff. Picture language machines. The Computer Journal, 15(1):79, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/79.full.pdf+html.

Duff:1973:BR

[Duf73] M. J. B. Duff. Book review. The Computer Journal, 16(4):307, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/4/307.1.full.pdf+html.

Duff:1974:BR

[Duf74] M. J. B. Duff. Book review. The Computer Journal, 17(1):58, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/58.full.pdf+html.

Duff:1978:BR

[Duf78] M. J. B. Duff. Book review. The Computer Journal, 21(3):198, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/198.3.full.pdf+html.

Duff:1979:BR

[Duf79] M. J. B. Duff. Book review. The Computer Journal, 22(1):16, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/16.1.full.pdf+html.

Dunham:1973:WAA

[Dun73] C. B. Dunham. When absolute approximations are optimal in function subroutines. The Computer Journal, 16(3):265, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/265.1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160265.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/265.
tif.

Dunham:1978:CCR

[Dun78] C. B. Dunham. Correspondence: Checking reproduction of published programs. The Computer Journal, 21(4):383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/383.tif. See correspondence [Jac79b].

Duncan:1979:DCQ

[Dun79a] G. Duncan. Discussion and correspondence Q charts — a method of specification. The Computer Journal, 22(4):380–382, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/380.full.pdf+html.

Duncan:1979:QCM

[Dun79b] G. Duncan. Q charts — a method of specification. The Computer Journal, 22(4):380–381, November 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/380.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/381.tif.

Durbin:1974:NIL

[Dur74] F. Durbin. Numerical inversion of Laplace transforms: an efficient improvement to Dubner and Abate’s method. The Computer Journal, 17(4):371–376, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/371.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/371.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/372.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/373.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/374.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/375.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/376.tif.

Dunham:1973:CDM

[DW73] C. B. Dunham and G. A. Watson. Correspondence: Degeneracy in the matrix of partial derivatives. The Computer Journal, 16(3): 284, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/284.tif. See [OW71].

Duff:1977:CLA

[DW77] M. J. B. Duff and D. M. Watson. The cellular logic array image processor. The Computer Journal, 20(1):68–72, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/68.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/68.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/69.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/71.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/72.
tif.

Duncan:1975:PNL

[DZW75] F. G. Duncan, D. Zissos, and Maureen Walls. A postfix notation for logic circuits. The Computer Journal, 18(1):63–69, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/63.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/63.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/64.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/65.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/66.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/67.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/69.tif.

Evans:1970:AAS

[EA70] D. J. Evans and L. V. Atkinson. Algorithm 51: An algorithm for the solution of general three term linear systems. The Computer Journal, 13(3):323–324, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_03/tiff/323.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/324.tif.

Earnshaw:1975:AVP

[Ear75] R. A. Earnshaw. Is APL a viable programming language? The Computer Journal, 18(4):318–323, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/318.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/318.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/319.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/320.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/321.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/322.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/323.tif.

Easteal:1975:BR

[Eas75] C. D. Easteal. Book review. The Computer Journal, 18(3):205, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/205.2.full.pdf+html.

Eason:1976:UNC

[Eas76] K. D. Eason. Understanding the naive computer user. The Computer Journal, 19(1):3–7, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/1/3.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/3.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/4.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/5.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/6.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/7.tif.

Eaton:1972:BR

[Eat72] J. R. Eaton. Book review. The Computer Journal, 15(3):194, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/194.full.pdf+html.

Eaton:1974:PAC

[Eat74] J. R. Eaton. A practical approach to computer simulation in business. The Computer Journal, 17(4):354, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/4/354.full.pdf+html.

Ecclestone:1971:C

[Ecc71a] M. J. Ecclestone. Correspondence. The Computer Journal, 14(4):351, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/351.1.full.pdf+html.

Ecclestone:1971:CQI

[Ecc71b] M. J. Ecclestone. Correspondence: On “A quasi-intrinsic scheme for passing a smooth curve through a discrete set of points”. The Computer Journal, 14(4):351, November 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140351.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/351.tif. See [McC70b, Smi71,

Sab72, Coo72f].

Ecker:1974:PSQ

[Eck74] A. Ecker. The period of search for the quadratic and related hash methods. The Computer Journal, 17(4):340–343, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/340.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/340.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/341.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/342.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/343.tif.

Edmonds:1973:BR

[Edm73] E. A. Edmonds. Book review. The Computer Journal, 16(2):117, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/117.full.pdf+html.

Edmonds:1974:PH

[Edm74] E. A. Edmonds. Problems of heuristics. The Computer Journal, 17(4):336a–336, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/336-a.full.pdf+html.

Edwards:1972:LBM

[Edw72] C. R. Edwards. The logic of Boolean matrices. The Computer Journal, 15(3):247–253, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/3/247.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150247.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/247.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/248.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/249.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/250.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/251.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/253.tif.

Edwards:1974:DSH

[Edw74] D. B. G. Edwards. Digital systems: Hardware organization and design. The Computer Journal, 17(4):364a–364, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/364-a.
full.pdf+html.

Edwards:1977:BRb

[Edw77a] B. G. Edwards. Book review. The Computer Journal, 20(3):231, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/3/231.1.full.pdf+html.

Edwards:1977:CBS

[Edw77b] B. J. Edwards. Choice of block sizes for magnetic tape files. The Computer Journal, 20(1):10–14, February 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/1/10.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/10.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/11.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/12.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/13.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/14.tif.

Edwards:1977:BRa

[Edw77c] R. Edwards. Book review. The Computer Journal, 20(1):36, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/36.1.full.pdf+html.

Edwards:1978:SID

[Edw78] C. R. Edwards. Some improved designs for the Digital Summation Threshold Logic (DSTL) gate. The Computer Journal, 21(1):73– 78, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/21/1/73.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/73.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/75.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/76.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/77.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/78.tif.

Edwards:1979:BR

[Edw79] D. B. G. Edwards. Book review. The Computer Journal, 22(4):312, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/312.2.full.pdf+html.

Evans:1976:SCB

[EH76] D. J. Evans and M. Hatzopoulos. The solution of certain banded systems of linear equations using the folding algorithm. The Computer Journal, 19(2):184–187, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/184.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/184.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/185.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/186.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/187.tif.

Eves:1978:IMT

[EH78] J. M. Eves and S. L. Hurst. Improvements in multioutput threshold-logic gates. The Computer Journal, 21(1):79–85, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/79.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/79.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/82.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/83.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/85.tif.

Ehrman:1972:C

[Ehr72a] J. R. Ehrman. Correspondence. The Computer Journal, 15(4):332, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/4/332.full.pdf+html.

Ehrman:1972:CSE

[Ehr72b] J. R. Ehrman. Correspondence: On “Suggested Extensions to FOR-TRAN IV”. The Computer Journal, 15(4):332, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/150332.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_04/tiff/332.tif. See

[Fin72a].

Ellis:1976:C

[Ell76] I. R. Ellis. Correspondence. The Computer Journal, 19(4):383, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/383.1.full.pdf+html.

Evans:1978:CCM

[ELR78] R. V. Evans, G. S. Lockington, and T. N. Reid. A compiler compiler and methodology for problem oriented language compiler implementors. The Computer Journal, 21(2):117–121, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/117.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/117.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/118.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/119.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/120.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/121.tif.

Elsworth:1979:CIL

[Els79] E. F. Elsworth. Compilation via an intermediate language. The Computer Journal, 22(3):226–233, August 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/226.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/226.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/227.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/228.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/229.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/230.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/231.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/232.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/233.tif.

Evangelisti:1971:GMU

[EM71] C. J. Evangelisti and S. P. Morse. Graphical modelling using contextually implied functions. The Computer Journal, 14(4):382– 389, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/4/382.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_04/140382.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/382.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/383.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/384.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/385.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/386.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/387.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/388.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/389.tif.

Emery:1975:BR

[Eme75] G. Emery. Book review. The Computer Journal, 18(2):107, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/107.2.full.pdf+html.

Emery:1976:BR

[Eme76] G. Emery. Book review. The Computer Journal, 19(3):206, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/206.2.full.pdf+html.

Emerson:1978:BRB

[Eme78] J. A. Emerson. Book review: Advanced ANS COBOL with structured programming and High Level COBOL Programming. The Computer Journal, 21(2):116, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/116.full.pdf+html.

Emerson:1979:BR

[Eme79] J. Emerson. Book review. The Computer Journal, 22(3):219, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/219.2.full.pdf+html.

Elcock:1972:DDC

[EMM72] E. W. Elcock, J. J. McGregor, and A. M. Murray. Data directed control and operating systems. The Computer Journal, 15 (2):125–129, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/125.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/150125.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/125.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/126.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/127.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/128.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/129.tif.

Evershed:1971:HLL

[ER71] D. G. Evershed and G. E. Rippon. High level languages for low level users. The Computer Journal, 14(1):87–90, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/87.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140087.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/87.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/89.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/

tiff/90.tif. See correspondence [Bac71b, Wal71b, FW72, Hil72b].

Eriksen:1974:DBC

[Eri74] Sven Eriksen. The data base concept. The Computer Journal, 8(1): ??, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Ewen-Smith:1975:CCD

[ES75] B. M. Ewen-Smith. Correspondence: Check digits and error correction. The Computer Journal, 18(3):284, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/284.1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/284.tif. See [And74, And75b,

How75b].

Evans:1977:ACB

[ES77] D. J. Evans and S. A. Smith. Algorithm 99: On the construction of balanced binary trees for parallel processing. The Computer Journal, 20(4):378–379, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_20/Issue_04/tiff/378.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/379.tif.

Edwards:1975:CSC

[ET75] J. P. N. Edwards and The Editor. Correspondence: Saving CPU time in scientific computation. The Computer Journal, 18(3):285, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/285.1.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_03/tiff/285.tif. See [HV74,

Art75, Hat76].

Etherton:1971:DSN

[Eth71] M. Etherton. Data structures for a network design system. The Computer Journal, 14(4):366–374, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/366.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140366.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/367.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_04/tiff/368.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/369.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/371.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/374.tif.

Evans:1970:BR

[Eva70] D. J. Evans. Book review. The Computer Journal, 13(4):381, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/4/381.full.pdf+html.

Evans:1971:AAS

[Eva71] D. J. Evans. Algorithm 69: An algorithm for the solution of symmetric general three term linear systems. The Computer Journal, 14(4):444–445, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_04/tiff/444.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/445.tif.

Evans:1972:ASC

[Eva72] D. J. Evans. An algorithm for the solution of certain tridiagonal systems of linear equations. The Computer Journal, 15(4):356– 359, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/356.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_04/150356.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/356.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/357.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/358.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/359.tif.

Evans:1975:RAD

[Eva75] D. J. Evans. A recursive algorithm for determining the eigenvalues of a quindiagonal matrix. The Computer Journal, 18(1):70–73, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/70.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/70.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/73.tif.

Evans:1977:UFM

[Eva77] D. J. Evans. On the use of fast methods for solving boundary value problems. The Computer Journal, 20(2):181–184, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/181.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/181.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/182.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/184.tif.

Evans:1978:BRB

[Eva78] M. Evans. Book review: Computer Operating Systems, by D. W. Barron; 1977; 135 pages (Chapman and Hall), £2.95. The Computer Journal, 21(4):336, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/4/336.3.full.pdf+html.

Eve:1971:BR

[Eve71] J. Eve. Book review. The Computer Journal, 14(1):20, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/20.full.pdf+html.

Eve:1974:BR

[Eve74] J. Eve. Book review. The Computer Journal, 17(1):12, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/12.full.pdf+html.

Farmakides:1973:LAC

[F^{+}73] John B. Farmakides et al. Legal aspects of computerized information systems. The Computer Journal, 7(1):??, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Farr:1978:AGA

[Far78] G. L. Farr. Algorithm 104: Generation of the alphabetically ordered sequences from a set of N distinct objects. The Computer Journal, 21 (4):373, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_04/tiff/373.tif.

Fleck:1975:AMN

[FB75] R. Fleck and J. Bailey. Algorithm 87: Minimum of a nonlinear function by the application of the geometric programming technique. The Computer Journal, 18(1):86–89, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_01/tiff/87.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/88.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/89.tif.

Fellgett:1969:CWA

[Fel69] Peter Fellgett. Correspondence: What is an analyst? The Computer Journal, 12(1):104, February 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_12/Issue_01/tiff/104.tif. See

correspondence [Fel70, AT70].

Fellgett:1970:CWA

[Fel70] P. B. Fellgett. Correspondence: What is an analyst? The Computer Journal, 13(1):118–119, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/119.tif. See [Fel69, AT70].

Fellgett:1971:BR

[Fel71] P. B. Fellgett. Book review. The Computer Journal, 14(4):403, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/403.full.pdf+html.

Fellgett:1978:BRa

[Fel78a] P. B. Fellgett. Book review. The Computer Journal, 21(3):286, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/286.1.full.pdf+html.

Fellgett:1978:BRb

[Fel78b] P. B. Fellgett. Book review. The Computer Journal, 21(3):286, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/286.2.full.pdf+html.

Fenton:1974:MS

[Fen74] J. S. Fenton. Memoryless subsystems. The Computer Journal, 17 (2):143–147, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/143.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/143.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/144.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/145.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/146.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/147.tif.

Ferriss:1974:BR

[Fer74] D. H. Ferriss. Book review. The Computer Journal, 17(1):63, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/63.2.full.pdf+html.

Forbes:1977:PAS

[FG77] K. Forbes and A. W. Goldsworthy. Prescheduling algorithm — scheduling a suitable mix prior to processing. The Computer Journal, 20(1):27–29, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/1/27.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/27.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/28.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/29.tif.

Fletcher:1971:CLB

[FGH71] R. Fletcher, J. A. Grant, and M. D. Hebden. The calculation of linear best L_{p} approximations. The Computer Journal, 14(3):276– 279, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/3/276.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/140276.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/276.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/277.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/278.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/279.tif.

Foley:1971:PRP

[FH71] M. Foley and C. A. R. Hoare. Proof of a recursive program: Quicksort. The Computer Journal, 14(4):391–395, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/391.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140391.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/391.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/392.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/393.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/394.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/395.tif.

Ford:1977:LLA

[FH77] W. S. Ford and V. C. Hamacher. Low level architecture features for supporting process communication. The Computer Journal, 20 (2):156–162, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/156.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/156.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/158.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/159.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/160.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/162.tif.

Fielding:1970:NAS

[Fie70] K. Fielding. Note on Algorithm 44: Solution of nonlinear simultaneous equations. The Computer Journal, 13(2):219, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_02/tiff/219.tif. See [Bro69, Red71b].

Fike:1975:PGM

[Fik75] C. T. Fike. A permutation generation method. The Computer Journal, 18(1):21–22, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/21.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/21.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/22.tif.

Finn:1972:CSE

[Fin72a] A. J. Finn. Correspondence: Suggested extensions to FORTRAN IV. The Computer Journal, 15(1):12, February 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
150012.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/12.tif. See correspondence [Mee72b,

Hal72b, Ehr72b].

Finn:1972:SEF

[Fin72b] A. J. Finn. Suggested extension to Fortran IV. The Computer Journal, 15(1):12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/1/12.full.pdf+html.

Finn:1974:C

[Fin74] A. J. Finn. Correspondence. The Computer Journal, 17(3):287, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/287.2.full.pdf+html.

Fisher:1974:C

[Fis74a] D. L. Fisher. Correspondence. The Computer Journal, 17(2):190, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/190.3.full.pdf+html.

Fisher:1974:CPF

[Fis74b] D. L. Fisher. Correspondence: On “The project, and the future of Computing Science courses”. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/190.tif. See [Bro73a, Bel74].

Fisher:1974:CES

[Fis74c] G. A. Fisher. Correspondence: On “Editorial: the silent business user”. The Computer Journal, 17(1):95–96, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/95.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/96.tif. See

[Gea73, TT73].

Fisher:1976:BPT

[Fis76] D. L. Fisher. Blind programmers — their manager’s experience. The Computer Journal, 19(2):189–192, May 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/189.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/189.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/190.tif.

Fisher:1979:BR

[Fis79a] D. L. Fisher. Book review. The Computer Journal, 22(3):266, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/266.2.full.pdf+html.

Fisher:1979:RCM

[Fis79b] D. L. Fisher. Review: Cad meets ai. The Computer Journal, 22 (4):322, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/4/322.2.full.pdf+html.

Fitch:1973:AS

[Fit73] J. P. Fitch. On algebraic simplification. The Computer Journal, 16(1):23–27, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/1/23.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
160023.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/23.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/tiff/24.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/tiff/25.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/26.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/tiff/27.tif.

Fitzpatrick:1974:BR

[Fit74] I. J. Fitzpatrick. Book review. The Computer Journal, 17(3):260, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/260.1.full.pdf+html.

Ferrie:1976:ESP

[FKLM76] J. Ferrie, C. Kaiser, D. Lanciaux, and B. Martin. An extensible structure for protected systems’ design. The Computer Journal, 19(4):315–321, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/4/315.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/315.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/317.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/318.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/319.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/320.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/321.tif.

Fenner:1979:BTR

[FL79] T. I. Fenner and G. Loizou. A binary tree representation and related algorithms for generating integer partitions. The Computer Journal, 23(4):332–337, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/4/332.full.pdf+html.

Flavell:1972:C

[Fla72a] A. J. Flavell. Correspondence. The Computer Journal, 15(1): 92, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/92.full.pdf+html.

Flavell:1972:CAR

[Fla72b] A. J. Flavell. Correspondence: On “Another round of FORTRAN”. The Computer Journal, 15(1):92, February 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
150092.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/92.tif. See [Cha71, Dew72c,

Hal72b, Cha72b].

Flavell:1975:BR

[Fla75] R. B. Flavell. Book review. The Computer Journal, 18(4):354, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
4/354.2.full.pdf+html.

Flavell:1977:BR

[Fla77] R. Flavell. Book review. The Computer Journal, 20(2):140, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/140.1.full.pdf+html.

Fletcher:1970:NAV

[Fle70] R. Fletcher. A new approach to variable metric algorithms. The Computer Journal, 13(3):317–322, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/317.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130317.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/317.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/318.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/319.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/320.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/321.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/322.tif.

Fletcher:1973:BRa

[Fle73a] R. Fletcher. Book review. The Computer Journal, 16(4):314, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/314.1.full.pdf+html.

Fletcher:1973:BRb

[Fle73b] R. Fletcher. Book review. The Computer Journal, 16(4):314, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/314.2.full.pdf+html.

Fleck:1975:CEL

[Fle75] Robert A. Fleck, Jr. Computer efficiency and linear programming. The Computer Journal, 18(2):115–116, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/115.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/115.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/116.tif.

Florentin:1970:BRa

[Flo70a] J. J. Florentin. Book review. The Computer Journal, 13(1):27, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/27.full.pdf+html.

Florentin:1970:BRb

[Flo70b] J. J. Florentin. Book review. The Computer Journal, 13(4): 400, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/400.full.pdf+html.

Florentin:1973:BRa

[Flo73a] J. J. Florentin. Book review. The Computer Journal, 16(1):63, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/63.full.pdf+html.

Florentin:1973:BRb

[Flo73b] J. J. Florentin. Book review. The Computer Journal, 16(4):374, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/4/374.1.full.pdf+html.

Florentin:1974:BRa

[Flo74a] J. J. Florentin. Book review. The Computer Journal, 17(2):120, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/120.1.full.pdf+html.

Florentin:1974:BRb

[Flo74b] J. J. Florentin. Book review. The Computer Journal, 17(2):180, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/180.full.pdf+html.

Florentin:1974:CAD

[Flo74c] J. J. Florentin. Consistency auditing of databases. The Computer Journal, 17(1):52–58, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/52.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/52.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/53.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/54.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/55.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/56.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/57.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/58.tif.

Florentin:1976:DBR

[Flo76] J. J. Florentin. Data base representations of application models. The Computer Journal, 19(1):13–16, February 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/1/13.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/13.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/14.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/15.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/16.tif.

Florentin:1979:BRa

[Flo79a] J. J. Florentin. Book review. The Computer Journal, 22(1):16, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/16.2.full.pdf+html.

Florentin:1979:BRb

[Flo79b] J. J. Florentin. Book review. The Computer Journal, 22(3):239, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/3/239.3.full.pdf+html.

Flower:1979:BR

[Flo79c] M. Flower. Book review. The Computer Journal, 22(2):104, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/104.2.full.pdf+html.

Fitch:1978:NCG

[FN78] J. P. Fitch and A. C. Norman. A note on compacting garbage collection. The Computer Journal, 21(1):31–34, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
21/1/31.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/32.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/34.tif.

Forrest:1972:IIA

[For72] A. R. Forrest. Interactive interpolation and approximation by Bézier polynomials. The Computer Journal, 15(1):71–79, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/71.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150071.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/71.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/73.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/74.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/76.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/77.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/78.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/79.tif.

Ford:1975:APC

[For75] W. S. Ford. Algorithm 90: Periodic cubic spline interpolation with equidistant nodes. The Computer Journal, 18(2): 183–184, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_18/Issue_02/tiff/183.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/184.tif.

Foster:1971:UCP

[Fos71] C. C. Foster. Uncoupling central processor and storage device speeds. The Computer Journal, 14(1):45–48, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
14/1/45.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140045.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/45.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/46.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/47.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/48.tif.

Fowell:1975:C

[Fow75] E. J. C. Fowell. Correspondence. The Computer Journal, 18(2):191, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/191.1.full.pdf+html.

Fox:1974:BRI

[Fox74] L. Fox. Book review: Introduction to Matrix Computations, G. W. Stewart, 1973; 441 pages (Academic Press). The Computer Journal, 17(3):200, March 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/200.full.pdf+html.

Fox:1976:BRC

[Fox76] L. Fox. Book review: The Chebyshev Polynomials, Theodore J. Rivlin, 1974; 186 pages (John Wiley). The Computer Journal, 19 (1):42, January 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/1/42.full.pdf+html.

Foxley:1978:HCP

[Fox78] Eric Foxley. A hybrid computer performance modelling system. The Computer Journal, 21(3):205–209, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/205.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/205.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/206.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/208.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/209.tif.

Feng:1979:SNS

[FR79] D. Y. Feng and R. F. Riesenfeld. Some new surface forms for computer aided geometric design. The Computer Journal, 23(4):324–331, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/324.full.pdf+html.

Fraser:1970:CFC

[Fra70] J. G. Fraser. Correspondence: Further comments on a line-thinning scheme. The Computer Journal, 13(2):221–222, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/221.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/221.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/222.tif.

See correspondence [Deu69].

Frank:1976:JCM

[Fra76] G. R. Frank. Job control in the MU5 operating system. The Computer Journal, 19(2):139–143, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/139.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/139.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/140.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/141.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/142.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/143.tif.

Freeman:1974:C

[Fre74a] K. Freeman. Correspondence. The Computer Journal, 17(2):189, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/189.1.full.pdf+html.

Freeman:1974:CH

[Fre74b] K. Freeman. Correspondence: On “Himmelbett”. The Computer Journal, 17(2):189, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_17/Issue_02/tiff/189.tif. See

[Bel73, Str74b].

Freeman:1978:BR

[Fre78] P. R. Freeman. Book review. The Computer Journal, 21(1):72, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/72.full.pdf+html.

Fryer:1971:BR

[Fry71a] J. G. Fryer. Book review. The Computer Journal, 14(1):107, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/107.full.pdf+html.

Fryer:1971:BRB

[Fry71b] J. G. Fryer. Book review: Rank Order Probabilities: Two Sample Normal Shift Alternatives, by Roy C. Milton, 1970. The Computer Journal, 14(1):107, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_01/tiff/107.tif.

Florentin:1975:SSR

[FS75] J. J. Florentin and A. J. Sammes. Systems with state reset. The Computer Journal, 18(2):135–139, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/135.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/135.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/136.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/137.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_18/Issue_02/tiff/138.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/139.tif.

Fairthorne:1973:CWU

[FT73] Robert A. Fairthorne and The Editor. Correspondence: On word use. The Computer Journal, 16(3):286, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/286.tif. See [WNT73].

Fuller:1976:PPR

[Ful76] A. T. Fuller. The period of pseudo-random numbers generated by Lehmer’s congruential method. The Computer Journal, 19(2): 173–177, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/173.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/173.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/176.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/177.tif.

Fellgett:1972:CWO

[FW72] P. B. Fellgett and M. R. Williams. Correspondence: The wonderland of operating systems. The Computer Journal, 15(2):131, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/131.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/tiff/131.tif. See

note [Bac71b, ER71, Mee71, Wal71b, Hil72b].

Fyfe:1970:UCS

[Fyf70] D. J. Fyfe. The use of cubic splines in the solution of certain fourth order boundary value problems. The Computer Journal, 13 (2):204–205, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/2/204.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_02/130204.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/205.tif.

Gale:1971:C

[Gal71a] J. W. Gale. Correspondence. The Computer Journal, 14(4):360, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/360.full.pdf+html.

Gale:1971:CFE

[Gal71b] J. W. Gale. Correspondence: On the February 1971 editorial. The Computer Journal, 14(4):360, November 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140360.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/360.tif.

Gallaher:1973:CDG

[Gal73] L. J. Gallaher. Correspondence: On Dijkstra and the GOTO statement. The Computer Journal, 16(3):284–285, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/285.tif.

Garside:1973:BR

[Gar73] M. J. Garside. Book review. The Computer Journal, 16(2):123, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/123.2.full.pdf+html.

Garg:1978:ACB

[Gar78] V. K. Garg. Algorithm 103: Computing the Bessel functions Y_{n}(x + iy) and K_{n}(x + iy). The Computer Journal, 21 (3):272–276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_21/Issue_03/tiff/272.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/273.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/274.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/275.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/276.tif.

Gayler:1972:C

[Gay72] D. R. Gayler. Correspondence. The Computer Journal, 15(4):307, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/4/307.full.pdf+html.

Gough:1971:NAP

[GE71] N. E. Gough and J. B. A. Epton. Note on Algorithm 48: Procedure for the evaluation of an integral occurring in the mean square response analysis of linear systems. The Computer Journal, 14 (2):215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/215.tif. See [Lit70].

Gearing:1970:SAT

[Gea70] Harold Gearing. Sidney Albert Tasker — an appreciation. The Computer Journal, 13(4):331, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/4/331.full.pdf+html.

Gearing:1973:ESB

[Gea73] H. W. G. Gearing. Editorial: the silent business user. The Computer Journal, 16(3):194–195, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/3/194.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
160194.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/194.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/195.tif.

See correspondence [Fis74c].

George:1970:BR

[Geo70] F. H. George. Book review. The Computer Journal, 13(3):247, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
3/247.full.pdf+html.

George:1978:BR

[Geo78] F. H. George. Book review. The Computer Journal, 21(3):242, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/242.1.full.pdf+html.

Gertler:1970:HLP

[Ger70] J. Gertler. High-level programming for process control. The Computer Journal, 13(1):70–75, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/70.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130070.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/70.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/71.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/72.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/73.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/74.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/75.tif.

Grant:1973:SPE

[GH73] J. A. Grant and G. D. Hitchins. The solution of polynomial equations in interval arithmetic. The Computer Journal, 16(1):69–72, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/1/69.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160069.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/69.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/70.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/71.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/72.tif.

Grant:1975:TAS

[GH75] J. A. Grant and G. D. Hitchins. Two algorithms for the solution of polynomial equations to limiting machine precision. The Computer Journal, 18(3):258–264, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/258.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/258.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/259.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/260.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/261.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/262.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/263.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/264.tif.

Ghandour:1972:FSA

[Gha72] Z. J. Ghandour. Formal systems and analysis of context sensitive languages. The Computer Journal, 15(3):229–237, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/229.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150229.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/229.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/230.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/231.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/232.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/233.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/234.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/235.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/237.tif.

Giles:1973:IPW

[Gil73] P. Giles. Implementation problems — why are they seldom aired in the Journal? The Computer Journal, 16(3):283, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/283.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/tiff/283.tif.

Gildersleeves:1978:ORV

[Gil78a] E. Gildersleeves. Online review, vol. 1 no. 1. The Computer Journal, 21(2):143a–143, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/143-a.full.pdf+html.

Giles:1978:BR

[Gil78b] P. Giles. Book review. The Computer Journal, 21(4):372, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/372.1.
full.pdf+html.

Giles:1979:BR

[Gil79] P. Giles. Book review. The Computer Journal, 22(4):382, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/382.full.pdf+html.

Girling:1973:BR

[Gir73] B. Girling. Book review. The Computer Journal, 16(3):215, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/215.1.full.pdf+html.

Girling:1976:BR

[Gir76] B. Girling. Book review. The Computer Journal, 19(4):321, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/321.1.full.pdf+html.

Green:1973:NTF

[GK73] D. H. Green and R. G. Kelsch. Nonlinear ternary feedback shift registers. The Computer Journal, 16(4):360–367, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/4/360.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/160360.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/360.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/361.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/362.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/363.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/364.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/365.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/366.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/367.tif.

Gudes:1979:SML

[GKS79] E. Gudes, H. S. Koch, and F. A. Stahl. Security in a multi-level structured model of a data base. The Computer Journal, 22(4):303– 306, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/4/303.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/303.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/304.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/305.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/306.tif.

Garwick:1970:CEI

[GL70] Jan V. Garwick and J. N. Lyness. Correspondence: On “The effect of inadequate convergence criteria in automatic routines”. The Computer Journal, 13(1):122, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/122.tif. See [Lyn69].

Glading:1973:CDL

[Gla73] D. G. Glading. Correspondence: On Dakin’s letter. The Computer Journal, 16(3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_16/Issue_03/tiff/287.tif. See

[Dak73b].

Glennie:1979:BRa

[Gle79a] A. E. Glennie. Book review. The Computer Journal, 23(4):352, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/352.2.full.pdf+html.

Glennie:1979:BRb

[Gle79b] A. E. Glennie. Book review. The Computer Journal, 23(4):368, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/368.2.full.pdf+html.

Gill:1972:AIU

[GM72] P. E. Gill and G. F. Miller. An algorithm for the integration of unequally spaced data. The Computer Journal, 15(1):80–83, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/80.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150080.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/

tiff/80.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/81.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/82.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/83.tif.

Goble:1975:FTR

[Gob75] C. E. Goble. A free-text retrieval system using hash codes. The Computer Journal, 18(1):18–20, February 1975. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/18.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/18.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/19.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/20.tif.

Goddu:1975:APO

[God75] F. J. B. Goddu. The application of precedence operations to interactive Canadian income tax calculations. The Computer Journal, 18(4):344–348, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/344.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/344.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/346.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/347.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/348.tif.

Goldsworthy:1973:CPR

[Gol73] A. W. Goldsworthy. Computers and privacy — A review of the Younger Comm. Report. The Computer Journal, 5(1):??, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Goldsack:1975:BR

[Gol75] S. J. Goldsack. Book review. The Computer Journal, 18(4):323, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/323.1.full.pdf+html.

Goldsack:1977:BRa

[Gol77a] S. J. Goldsack. Book review. The Computer Journal, 20(1):14, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/14.1.full.pdf+html.

Goldsack:1977:BRb

[Gol77b] S. J. Goldsack. Book review. The Computer Journal, 20(1):44, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/44.2.full.pdf+html.

Goldsack:1977:BRc

[Gol77c] S. J. Goldsack. Book review. The Computer Journal, 20(1):44, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/44.3.full.pdf+html.

Golden:1978:CAS

[Gol78a] Bruce L. Golden. Correspondence: On “An algorithm for the solution of the postman’s problem”. The Computer Journal, 21(4): 383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_04/tiff/383.tif. See [Bar78a].

Goldsack:1978:BR

[Gol78b] S. J. Goldsack. Book review. The Computer Journal, 21(1):85, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/85.2.full.pdf+html.

Goldsack:1978:CSD

[Gol78c] S. J. Goldsack. Correspondence: Semantic definition of linear control algorithms. The Computer Journal, 21(4):381– 382, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_21/Issue_04/tiff/381.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/382.tif.

Good:1970:BR

[Goo70] I. J. Good. Book review. The Computer Journal, 13(3):236, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
3/236.full.pdf+html.

Goodwin:1974:BR

[Goo74] G. C. Goodwin. Book reviews. The Computer Journal, 17(2):147, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/147.1.full.pdf+html.

Goodwin:1975:AIC

[Goo75] D. T. Goodwin. An algorithm for inverting certain translators of context-free languages. The Computer Journal, 18(4):349–354, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/4/349.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/349.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/350.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/351.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/352.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/353.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/354.tif.

Goodwin:1977:CUO

[Goo77] D. T. Goodwin. Conditions for underflow and overflow of an arithmetic stack. The Computer Journal, 20(1):56–62, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/56.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/56.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/57.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/60.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/62.tif.

Gordon:1978:IGM

[Gor78] M. Gordon. An interactive graphical minicomputer system for the management and exploration of clinical renal data. The Computer Journal, 21(1):91–93, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/1/91.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
01/tiff/91.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_01/tiff/92.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_01/tiff/93.tif.

Gordon:1979:PLI

[Gor79] M. J. C. Gordon. On the power of list iteration. The Computer Journal, 22(4):376–379, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/4/376.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/376.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/378.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/379.tif.

Gourlay:1970:BRa

[Gou70a] A. R. Gourlay. Book review. The Computer Journal, 13(1):54, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/54.full.pdf+html.

Gourlay:1970:BRb

[Gou70b] A. R. Gourlay. Book review. The Computer Journal, 13(3):242, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/3/242.full.pdf+html.

Gourlay:1970:GEH

[Gou70c] A. R. Gourlay. Generalisation of elementary Hermitian matrices. The Computer Journal, 13(4):411–412, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/411.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130411.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/411.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/412.tif.

Gay:1973:IDS

[GP73] B. Gay and S. G. Payne. Interactive digital simulation on a small computer. The Computer Journal, 16(2):118–121, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/118.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/160118.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/118.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/120.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/121.
tif.

Ghoshdastidar:1975:SES

[GR75] D. Ghoshdastidar and M. K. Roy. A study on the evaluation of Shell’s sorting technique. The Computer Journal, 18(3):234– 235, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/234.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/234.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/235.tif.

Grad:1971:MB

[Gra71] J. Grad. Matrix balancing. The Computer Journal, 14(3):280– 284, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/3/280.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/140280.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/280.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/281.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/282.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/283.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/284.tif.

Greenspan:1970:DNS

[Gre70] D. Greenspan. Discrete, nonlinear string vibrations. The Computer Journal, 13(2):195–201, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/2/195.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/
130195.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/195.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/196.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/197.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/198.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/199.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/200.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/201.tif.

Griffiths:1973:DOS

[Gri73] J. G. Griffiths. Drawing opaque solids using an incremental plotter. The Computer Journal, 16(1):35–38, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
16/1/35.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160035.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/35.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/37.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/38.tif.

Grimsditch:1976:C

[Gri76] D. S. Grimsditch. Correspondence. The Computer Journal, 19(4): 383–384, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/383.3.full.pdf+html.

Grimsdale:1979:BRa

[Gri79a] R. L. Grimsdale. Book review. The Computer Journal, 22(1):16, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/16.3.full.pdf+html.

Grimsdale:1979:BRb

[Gri79b] R. L. Grimsdale. Book review. The Computer Journal, 22(1):49, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/49.1.full.pdf+html.

Grossman:1971:PCF

[Gro71] M. Grossman. Parametric curve fitting. The Computer Journal, 14 (2):169–172, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/169.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140169.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/169.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/170.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/171.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/172.tif.

Grover:1973:MOI

[Gro73] D. J. Grover. Message orientated interactive graphics. The Computer Journal, 16(1):39–45, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/1/39.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160039.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/39.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/40.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/41.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/43.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/45.tif.

Girling:1973:BRB

[GS73] B. Girling and D. A. Sturt. Book reviews: Computer Simulation of Dynamic Systems, by R. J. Kochenburger, 1972; Electronic Data Processing and Computers for Commercial Students, by E. A. Bird, 1972. The Computer Journal, 16(3):215, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/215.tif.

Green:1978:CDT

[GS78] P. J. Green and R. Sibson. Computing Dirichlet tessellations in the plane. The Computer Journal, 21(2):168–173, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/168.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/168.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/169.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/171.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/172.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/173.tif.

Green:1979:CCH

[GS79] P. J. Green and B. W. Silverman. Constructing the convex hull of a set of points in the plane. The Computer Journal, 22(3):262– 266, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/3/262.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/262.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/263.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/265.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/266.tif.

Gayler:1972:CEP

[GT72] D. R. Gayler and The Editor. Correspondence: On editorial policy. The Computer Journal, 15(4):307, November 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150307.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/307.tif. See [Wal71b, MT72,

Bar73a, KTW^{+}73].

Gunton:1970:RPD

[Gun70] A. Gunton. Recovery procedures for direct access commercial systems. The Computer Journal, 13(2):123–126, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/123.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130123.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

13/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/125.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/126.
tif.

Gurney:1975:BR

[Gur75] I. D. C. Gurney. Book review. The Computer Journal, 18(1):20, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/20.1.full.pdf+html.

Gurd:1978:BR

[Gur78] J. R. Gurd. Book review. The Computer Journal, 21(3):218, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/218.1.full.pdf+html.

Gibbons:1973:CFC

[GW73] Charles J. Gibbons and M. Wells. Correspondence: On “File Compression Using Variable Length Encodings”. The Computer Journal, 16(3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/287.tif. See [Wel72].

Grad:1972:ALS

[GZ72] J. Grad and E. Zakrajsek. LR algorithm with Laguerre shift for symmetric tridiagonal matrices. The Computer Journal, 15(3):268– 270, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/268.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/150268.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/268.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/269.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/270.tif.

Hadjidimos:1971:OEA

[Had71] A. Hadjidimos. Optimum extrapolated ADI iterative difference schemes for the solution of Laplace’s equation in three space variables. The Computer Journal, 14(2):179–183, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/179.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_02/140179.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/179.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/181.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/182.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/183.tif.

Haddon:1975:BR

[Had75] E. W. Haddon. Book review. The Computer Journal, 18(1):69, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/69.full.pdf+html.

Haine:1978:ADS

[Hai78] P. M. Haine. Analysis and design of sequential digital systems. The Computer Journal, 21(2):177a–177, May 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/2/177-a.
full.pdf+html.

Hale:1971:BR

[Hal71] V. W. D. Hale. Book review. The Computer Journal, 14(2):153, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/2/153.full.pdf+html.

Hall:1972:C

[Hal72a] D. T. Hall. Correspondence. The Computer Journal, 15(3):273, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/273.full.pdf+html.

Hall:1972:CFC

[Hal72b] D. T. Hall. Correspondence: On FORTRAN complainants. The Computer Journal, 15(3):273, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150273.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/273.tif. See [Hil72b, Fin72a,

Fla72b, Pal72b].

Halfpenny:1974:BR

[Hal74] D. H. Halfpenny. Book review. The Computer Journal, 17(3):252, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/252.2.full.pdf+html.

Hammersley:1970:ASPa

[Ham70a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(1):110, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/110.tif.

Hammersley:1970:ASPb

[Ham70b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(2):206–207, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/206.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/207.tif.

Hammersley:1970:ASPc

[Ham70c] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(3):323, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_03/tiff/323.tif.

Hammersley:1970:ASPd

[Ham70d] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 13(4):421, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_04/tiff/421.tif.

Hammersley:1971:ASPa

[Ham71a] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 14(1):103, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/103.tif.

Hammersley:1971:ASPb

[Ham71b] P. Hammersley. Algorithms supplement: Previously published algorithms. The Computer Journal, 14(2):207, May 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/
tiff/207.tif.

Hammersley:1973:E

[Ham73] P. Hammersley. Editorial. The Computer Journal, 16(1):2–3, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/tiff/3.tif. See

correspondence [TT73].

HamiltonKelly:1974:CRC

[Ham74a] B. Hamilton Kelly. Correspondence: On radix conversion. The Computer Journal, 17(4):382, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/

tiff/382.tif. See [Sod53, Boo74a, Boo74b].

Hammersley:1974:SIP

[Ham74b] P. Hammersley. The skyline of information processing. The Computer Journal, 17(4):370a–370, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/370-a.full.pdf+html.

Hammersley:1976:CPM

[Ham76] J. M. Hammersley. Combinatorial programming: Methods and applications. The Computer Journal, 19(2):116, May 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/2/116.full.pdf+html.

Hamlet:1977:TPF

[Ham77] R. G. Hamlet. Testing programs with finite sets of data. The Computer Journal, 20(3):232–237, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/232.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/232.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/233.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/234.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/235.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/236.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/237.tif.

Hanson:1970:AEB

[Han70] Richard J. Hanson. Automatic error bounds for real roots of polynomials having interval coefficients. The Computer Journal, 13(3): 284–288, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/284.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_03/130284.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/284.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/285.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/286.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/287.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/288.tif.

Hanson:1978:FS

[Han78] D. R. Hanson. Filters in SL5. The Computer Journal, 21(2): 134–143, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/134.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/134.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/135.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/136.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/137.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/140.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/141.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/142.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/143.tif.

Harrop:1970:BAR

[Har70] M. Harrop. A basic approach to remote access. The Computer Journal, 13(2):131–135, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/2/131.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/
130131.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/131.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/132.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/133.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/134.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/135.tif.

Hartley:1976:TPA

[Har76] P. J. Hartley. Tensor product approximations to data defined on rectangular meshes in N-space. The Computer Journal, 19(4):348– 352, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/348.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/348.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/349.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/350.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/351.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/352.tif.

Harrington:1979:NSI

[Har79] Steven J. Harrington. A new symbolic integration system in RE-DUCE. The Computer Journal, 22(2):127–131, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/127.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/127.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/128.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/129.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/130.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/131.tif.

Haskell:1974:SPR

[Has74] R. Haskell. Symmetrical precedence relations on general phrase structure grammars. The Computer Journal, 17(3):234–241, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/234.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/234.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/235.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/237.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/238.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/239.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/240.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/241.tif.

Haskell:1975:EIC

[Has75] R. Haskell. Efficient implementation of a class of recursively defined functions. The Computer Journal, 18(1):23–29, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/23.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/23.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/24.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/26.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/27.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/28.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/29.tif.

Hastings:1976:C

[Has76] J. T. Hastings. Correspondence. The Computer Journal, 19(4):380, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/380.1.full.pdf+html.

Hatter:1976:PSM

[Hat76] D. J. Hatter. On processing sparse matrices. The Computer Journal, 19(3):273–274, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_19/Issue_02/tiff/273.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/274.tif. See [ET75].

Hayes:1977:DCN

[Hay77] P. J. Hayes. Discussion and correspondence: a note on the Towers of Hanoi problem. The Computer Journal, 20(3):282–285, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/282.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/283.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/284.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/285.tif.  See correspondence

[Wod78].

Hayden:1979:CND

[Hay79] Jane Hayden. Correspondence: Nonmorphological data collection. The Computer Journal, 22(3):286–287, August 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/286.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/287.tif. See [Tur78].

Hazlewood:1978:AAC

[Haz78] L. Hazlewood. Algorithm 101: An algorithm for cautious adaptive quadrature. The Computer Journal, 21(2):180–183, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_02/tiff/180.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_02/tiff/181.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/tiff/182.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_02/tiff/183.tif.

Haarhoff:1970:NMO

[HB70] P. C. Haarhoff and J. D. Buys. A new method for the optimization of a nonlinear function subject to nonlinear constraints. The Computer Journal, 13(2):178–184, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/178.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130178.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/178.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

13/Issue_02/tiff/179.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/180.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/181.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/182.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/183.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/184.
tif.

Honig:1978:TUA

[HC78] William L. Honig and C. Robert Carlson. Toward an understanding of (actual) data structures. The Computer Journal, 21(2): 98–104, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/21/2/98.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/100.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/102.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/103.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
02/tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_02/tiff/98.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_02/tiff/99.tif.

Hay:1974:IRS

[HCC74] J. L. Hay, R. E. Crosbie, and R. I. Chaplin. Integration routines for systems with discontinuities. The Computer Journal, 17(3):275– 278, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/275.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/275.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/276.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/277.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/278.tif.

Hanson:1971:CAS

[HD71] R. J. Hanson and P. Dyer. A computational algorithm for sequential estimation. The Computer Journal, 14(3):285–290, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/285.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_03/140285.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/285.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/286.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/287.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/288.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/289.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/290.tif.

Hopgood:1972:QHM

[HD72] F. R. A. Hopgood and J. Davenport. The quadratic hash method when the table size is a power of 2. The Computer Journal, 15(4):314– 315, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/314.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_04/150314.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/314.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/315.tif.  See correspondence

[PH73].

Heath:1973:BR

[Hea73] F. G. Heath. Book review. The Computer Journal, 16(3):204, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/204.1.full.pdf+html.

Healy:1974:C

[Hea74] M. J. R. Healy. Correspondence. The Computer Journal, 17(3):287, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/287.1.full.pdf+html.

Heath:1976:BRa

[Hea76a] F. G. Heath. Book review. The Computer Journal, 19(1):49, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/49.full.pdf+html.

Heath:1976:BRb

[Hea76b] F. G. Heath. Book review. The Computer Journal, 19(3):233, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/233.1.full.pdf+html.

Healy:1977:BR

[Hea77a] M. J. R. Healy. Book review. The Computer Journal, 20(4):349, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/349.2.full.pdf+html.

Heath:1977:BR

[Hea77b] F. G. Heath. Book review. The Computer Journal, 20(2):136, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/136.2.full.pdf+html.

Hebditch:1979:BR

[Heb79] D. Hebditch. Book review. The Computer Journal, 22(4):312, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/312.1.full.pdf+html.

Hennell:1978:ETN

[Hen78] M. A. Hennell. An experimental testbed for numerical software. The Computer Journal, 21(4):333–336, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/333.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/334.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/335.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/336.tif.

Houston:1971:IBT

[HG71] G. Houston and R. Gillespie. Integration of batch and timesharing services. The Computer Journal, 14(4):348–351, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/348.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140348.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/348.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/349.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_04/tiff/350.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/351.
tif.

HughesJones:1975:AGC

[HGMM75] R. Hughes Jones, P. R. Graves-Morris, and G. J. Makinson. Algorithm 85: Generation of Chisholm approximants. The Computer Journal, 18(1):81–83, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/81.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_01/tiff/83.tif.

Hittos:1976:OMM

[HH76] R. L. Hittos and D. S. Henderson. Optimisation of main memory size versus speed. The Computer Journal, 19(1):40–42, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/40.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/41.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/42.tif.

Hennell:1979:ETN

[HH79a] M. A. Hennell and D. Hedley. An experimental testbed for numerical software. II. ALGOL 68. The Computer Journal, 22(1):53–56, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/53.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/53.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/54.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/55.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/56.tif.

Howkins:1979:TMP

[HH79b] T. J. Howkins and M. T. Harandi. Towards more portable Cobol. The Computer Journal, 22(4):290–295, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/290.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_22/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/293.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/294.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/295.tif.

Higley:1976:TCT

[Hig76a] C. J. Higley. Type checking in a typeless language. The Computer Journal, 19(2):166–169, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/166.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/166.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/167.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/168.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/169.tif.

Higman:1976:POV

[Hig76b] B. Higman. The place of own variables in programming language theory. The Computer Journal, 19(3):225–228, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/225.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/225.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/226.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/227.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/228.tif.

Higman:1977:BRa

[Hig77a] B. Higman. Book review. The Computer Journal, 20(2):180, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/180.1.full.pdf+html.

Higman:1977:BRb

[Hig77b] B. Higman. Book review. The Computer Journal, 20(4):345, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/345.1.full.pdf+html.

Hill:1969:NAS

[Hil69] I. D. Hill. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 12(4): 409, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_04/tiff/409.tif. See [Boo67a,

Boo67b, Woo70g].

Hill:1971:FFA

[Hil71] I. D. Hill. Faults in functions, in Algol and Fortran. The Computer Journal, 14(3):315–316, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/315.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140315.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/315.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/316.tif.

See notes [SS72b].

Hill:1972:C

[Hil72a] I. D. Hill. Correspondence. The Computer Journal, 15(1):17, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/17.full.pdf+html.

Hill:1972:CHL

[Hil72b] I. D. Hill. Correspondence: On “High level languages for low level users”. The Computer Journal, 15(1):17, February 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/150017.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/17.tif. See

[Bac71b, ER71, Wal71b, FW72, Hal72b].

Hill:1973:BR

[Hil73a] I. D. Hill. Book review. The Computer Journal, 16(1):34, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/34.full.pdf+html.

Hill:1973:C

[Hil73b] I. D. Hill. Correspondence. The Computer Journal, 16(2):190, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/190.2.full.pdf+html.

Hill:1974:BR

[Hil74] I. D. Hill. Book review. The Computer Journal, 17(1):79, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/79.full.pdf+html.

Hill:1975:AEM

[Hil75] I. D. Hill. Algorithm 88: An electoral method. The Computer Journal, 18(1):89–90, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/89.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/90.tif.

Hilden:1976:GDI

[Hil76] J. Hilden. Guidelines for the design of interactive systems. The Computer Journal, 19(2):144–150, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/144.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/144.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/145.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/146.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/147.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/149.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/150.tif.

Hill:1977:AHD

[Hil77] I. D. Hill. Algorithm 94: Handling dates on small computers. The Computer Journal, 20(3):280, August 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/280.tif.

Hill:1979:CPP

[Hil79a] G. W. Hill. Cyclic properties of pseudo-random sequences of Mersenne prime residues. The Computer Journal, 22(1):80–85, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/1/80.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/80.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/81.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/82.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/83.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/84.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/85.tif.

Hill:1979:AGS

[Hil79b] I. D. Hill. Algorithm 109: Golden section search. The Computer Journal, 22(4):383–384, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_22/Issue_03/tiff/383.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/384.tif.

Hill:1979:AE

[Hil79c] I. D. Hill. Algorithm 3: Easter. The Computer Journal, 22(1): 86–87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/86.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/87.

tif. Reprinted from The Computer Bulletin. See [Boo79d, Ran79].

Hill:1980:CJS

[Hil80] I. D. Hill. Correspondence: Jumping to some purpose. The Computer Journal, 23(1):94, February 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_23/Issue_01/tiff/94.tif. See

[ASG79, Rob80, Tay80, MR82].

Hinchcliff:1979:BR

[Hin79] F. E. Hinchcliff. Book review. The Computer Journal, 23(4):323, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/323.1.full.pdf+html.

Hiorns:1970:BR

[Hio70] R. W. Hiorns. Book review. The Computer Journal, 13(1):109, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/109.full.pdf+html.

Hiorns:1978:BRa

[Hio78a] R. W. Hiorns. Book review. The Computer Journal, 21(3):233, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/233.1.full.pdf+html.

Hiorns:1978:BRb

[Hio78b] R. W. Hiorns. Book review. The Computer Journal, 21(3):245, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/245.1.full.pdf+html.

Hitchin:1978:CMF

[Hit78] D. Hitchin. Correspondence: On “Modifications to the Forsythe– Moler algorithm for solving linear algebraic systems”. The Computer Journal, 21(4):384, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_21/Issue_04/tiff/384.tif. See

[SG78, LS78].

Hoskins:1972:APC

[HK72] W. D. Hoskins and P. R. King. Algorithm 73: Periodic cubic spline interpolation using parametric splines. The Computer Journal, 15(3):282–283, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_15/Issue_03/tiff/282.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/283.tif.

Higginson:1973:CCR

[HK73] P. L. Higginson and P. T. Kirstein. On the computation of cyclic redundancy checks by program. The Computer Journal, 16(1):19– 22, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/19.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/160019.sgm.abs.

html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/21.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/22.tif.

Hoskins:1972:AIU

[HKA72] W. D. Hoskins, P. R. King, and T. H. Andres. Algorithm 74: Interpolation using periodic splines of odd order with equi-distant knots. The Computer Journal, 15(3):283–285, August 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
tiff/283.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_03/tiff/284.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/tiff/285.tif. See

note [Hos74].

Hawley:1975:DCC

[HKT75] D. A. Hawley, J. S. Knowles, and E. E. Tozer. Database consistency and the CODASYL DBTG proposals. The Computer Journal, 18(3): 206–212, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/206.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/206.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/207.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/208.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/209.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/210.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/211.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/212.tif.

Hammersley:1970:C

[HL70a] P. Hammersley and J. Larmouth. Correspondence. The Computer Journal, 13(2):220, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/2/220.1.full.pdf+html.

Hammersley:1970:CTF

[HL70b] P. Hammersley and J. Larmouth. Correspondence: Towards Fortran VI? The Computer Journal, 13(2):220, May 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/

tiff/220.tif. See [Red71b, Wel70].

Henderson:1976:EOP

[HL76a] D. S. Henderson and M. R. Levy. An extended operator precedence parsing algorithm. The Computer Journal, 19(3):229–233, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/3/229.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/229.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/230.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_03/tiff/231.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_03/tiff/232.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_03/
tiff/233.tif.

Henshell:1976:PDW

[HL76b] R. D. Henshell and H. C. Lauer. Ph.D. work in Computer Science and Professional Standards. The Computer Journal, 19 (3):274–275, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_19/Issue_02/tiff/274.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/275.tif. See [Lau75].

Humphrey:1970:MIR

[HM70] A. L. Humphrey and W. G. Munro. Management information retrieval. The Computer Journal, 13(2):127–130, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/127.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130127.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/127.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/129.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/130.
tif.

Hoskins:1977:ANS

[HM77] W. D. Hoskins and G. E. McMaster. Algorithm 98: a note on the solution of certain tridiagonal systems of linear equations. The Computer Journal, 20(4):376–378, November 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/376.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/377.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/378.tif.

Haley:1979:SPR

[HN79] J. R. Haley and A. S. Noetzell. A study of processing repetition. The Computer Journal, 22(4):352–356, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/352.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/352.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/353.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/354.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/355.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/356.tif.

Hill:1978:RAM

[HO78] I. D. Hill and Alex K. Obruča. Remarks on Algorithm 1: Mintree. The Computer Journal, 21(3):277, August 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/

tiff/277.tif. See [Obr78a].

Hoare:1970:BR

[Hoa70] C. A. R. Hoare. Book review. The Computer Journal, 13(2):170, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/170.full.pdf+html.

Hoare:1972:PSP

[Hoa72] C. A. R. Hoare. Proof of a structured program: The sieve of Eratosthenes. The Computer Journal, 15(4):321–325, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/321.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150321.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/321.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/322.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_04/tiff/323.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/324.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/325.tif.

Hoare:1973:SPS

[Hoa73] C. A. R. Hoare. A structured paging system. The Computer Journal, 16(3):209–215, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/3/209.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
160209.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/209.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/210.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/211.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/212.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/213.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/214.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/215.tif.

Hoare:1976:CST

[Hoa76] C. A. R. Hoare. Computer science and technology and their application. The Computer Journal, 19(2):172, May 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/2/172.full.pdf+html.

Hoare:1977:BR

[Hoa77] C. A. R. Hoare. Book review. The Computer Journal, 20(2):180, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/180.2.full.pdf+html.

Holmes:1975:IND

[Hol75] W. N. Holmes. Identification number design. The Computer Journal, 18(2):102–107, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/102.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/102.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/103.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/104.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/105.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/106.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/107.tif.

Holstein:1977:BR

[Hol77] R. Holstein. Book review. The Computer Journal, 20(4):349, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/349.1.full.pdf+html.

Holmes:1978:APL

[Hol78] W. N. Holmes. Is APL a programming language? The Computer Journal, 21(2):128–131, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/128.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/128.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/129.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/130.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/131.tif.

Honey:1971:C

[Hon71a] D. W. Honey. Correspondence. The Computer Journal, 14(4):443, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/443.full.pdf+html.

Honey:1971:CCD

[Hon71b] D. W. Honey. Correspondence: Calculation of a double-length square root from a double length number using single precision techniques. The Computer Journal, 14(4):443, November 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140443.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/443.tif.

Honey:1974:QMT

[Hon74] D. W. Honey. The ‘Queensway’ Mersey tunnel entrance control project. The Computer Journal, 17(4):290–294, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/290.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/293.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/294.tif.

Hope:1971:HOS

[Hop71] E. L. V. Hope. A high-order solution to the characteristic initial value problem in hyperbolic partial differential equations. The Computer Journal, 14(1):65–72, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
14/1/65.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140065.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/65.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/67.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/68.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/70.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/71.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/72.tif.

Hope:1978:CSP

[Hop78] Eric L. V. Hope. Correspondence: Structured programming input statements. The Computer Journal, 21(4):382, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/382.tif. See [Ing76, DI78].

Hoskins:1970:AIQ

[Hos70] W. D. Hoskins. Algorithm 62: Interpolating quintic splines on equidistant knots. The Computer Journal, 13(4):437–438, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_04/tiff/437.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/tiff/438.tif.

Hoshino:1971:DSC

[Hos71] S. Hoshino. On Davies, Swann and Campey minimisation process. The Computer Journal, 14(4):426–427, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/426.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140426.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/426.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/427.tif.

Hoskins:1974:NAI

[Hos74] W. D. Hoskins. Note on Algorithm 74: Interpolation using periodic splines of odd order with equi-distant knots. The Computer Journal, 17(1):91, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_01/tiff/91.tif. See [HKA72].

Housden:1971:PSF

[Hou71] R. J. W. Housden. Phrase structures in Fortran. The Computer Journal, 14(3):224–228, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/224.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140224.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/224.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/225.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/226.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/227.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/228.tif.

Housden:1975:SCT

[Hou75] R. J. W. Housden. On string concepts and their implementation. The Computer Journal, 18(2):150–156, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/150.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/150.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/151.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/152.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/153.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_18/Issue_02/tiff/154.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/156.tif.

Howard:1975:BR

[How75a] B. V. Howard. Book review. The Computer Journal, 18(4):297, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/297.full.pdf+html.

Howell:1975:CCD

[How75b] K. M. Howell. Correspondence: Check digits and error correction. The Computer Journal, 18(3):283–284, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/283.2.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/283.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/284.tif.

See [And74, And75b, ES75].

Howarth:1976:BRa

[How76a] D. J. Howarth. Book review. The Computer Journal, 19(1):24, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/24.1.full.pdf+html.

Howarth:1976:BRb

[How76b] D. J. Howarth. Book review. The Computer Journal, 19(3):245, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/245.1.full.pdf+html.

Hill:1967:NAT

[HP67] I. D. Hill and M. C. Pike. Note on algorithm 4: TWO BY TWO. The Computer Journal, 9(4):416, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_09/Issue_04/tiff/416.tif. See [HP79].

Haddon:1971:ALS

[HP71] E. W. Haddon and L. G. Proll. An Algol line-syntax checker. The Computer Journal, 14(2):128–132, May 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/2/128.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/140128.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/128.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/129.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/131.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/132.
tif.

Huxtable:1977:HSI

[HP77] D. H. R. Huxtable and J. M. M. Pinkerton. The hardware/software interface of the ICL 2900 range of computers. The Computer Journal, 20(4):290–295, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/290.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/293.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/294.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/295.tif.

Hill:1979:AT

[HP79] I. D. Hill and M. C. Pike. Algorithm 4: TWOBYTWO. The Computer Journal, 22(1):87–88, February 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/
tiff/87.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_01/tiff/88.tif. Reprinted from The Computer

Bulletin, September 1964, p. 67. See note [HP67].

Haring:1979:USJ

[HPLG79] G. Haring, R. Posch, C. Leonhardt, and G. Gell. The use of a synthetic jobstream in performance evaluation. The Computer Journal, 22(3):209–219, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/3/209.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/

tiff/209.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/210.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/211.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/213.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/214.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/215.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/216.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/217.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/218.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/219.tif.

Hext:1970:SAD

[HR70] J. B. Hext and P. S. Roberts. Syntax analysis by Domölki’s algorithm. The Computer Journal, 13(3):263–271, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/263.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130263.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/265.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/268.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/269.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/271.tif.

Heath:1973:BRB

[HR73] F. G. Heath and C. M. Reeves. Book reviews: Computer Logic Design, by M. Morris Mano, 1972; Logic and Logic Design, by B. Girling and H. G. Moring; An Introduction to Job Control Language, by Ronald J. DeMasi and Francis J. Longo, 1972. The Computer Journal, 16(3):204, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_16/Issue_03/tiff/204.tif.

Heindel:1974:ARP

[HR74] L. E. Heindel and J. T. Roberto. ARPL — A Retrieval Process Language. The Computer Journal, 17(2):113–116, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/113.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/113.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/114.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/115.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/116.tif.

Higgins:1971:DAA

[HS71] L. D. Higgins and Francis J. Smith. Disc access algorithms. The Computer Journal, 14(3):249–253, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/249.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140249.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/249.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/251.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/252.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/253.tif.

Hill:1975:C

[HT75] G. Hill and H. R. A. Townsend. Correspondence. The Computer Journal, 18(1):94, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/1/94.full.pdf+html.

Haskell:1978:CDS

[HT78] R. Haskell and The Editor. Correspondence: Dissemination of specialised information. The Computer Journal, 21(4):382, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/382.tif.

Hunter:1970:CME

[Hun70] D. B. Hunter. Clenshaw’s method for evaluating certain finite series. The Computer Journal, 13(4):378–381, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/378.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130378.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/378.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/380.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/381.
tif.

Hunt:1972:NSF

[Hun72] L. B. Hunt. Numerical smoothing and filtering in N dimensions. The Computer Journal, 15(1):58–65, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
15/1/58.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150058.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/58.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/59.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/60.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/61.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/62.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/63.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/64.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/65.tif.

Hunter:1975:QMP

[Hun75] G. Hunter. A quantitative measure of precision. The Computer Journal, 18(3):231–233, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/3/231.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/231.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/232.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/233.tif.

Hunter:1977:BR

[Hun77] R. B. Hunter. Book review. The Computer Journal, 20(3):256, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/256.2.full.pdf+html.

Huntly:1979:BR

[Hun79] R. A. Huntly. Book review. The Computer Journal, 22(2):145, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/145.2.full.pdf+html.

Hurst:1973:ACP

[Hur73] S. L. Hurst. The application of Chow parameters and Rademacher-Walsh matrices in the synthesis of binary functions. The Computer Journal, 16(2):165–173, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/165.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160165.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/165.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/166.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/167.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/168.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/169.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/170.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/171.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/172.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/173.tif.

Hurst:1979:ECS

[Hur79] S. L. Hurst. An engineering consideration of spectral transforms for ternary logic synthesis. The Computer Journal, 22(2):173– 183, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/2/173.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/173.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/176.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/177.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_22/Issue_02/tiff/178.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/179.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/180.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/181.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/182.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/183.tif.

Huzan:1979:MIS

[Huz79] Eva Huzan. Modelling of indexed sequential files: Monitoring disc transfers. The Computer Journal, 22(1):22–27, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/22.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/22.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/23.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/24.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/25.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/26.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/27.tif.

Hathaway:1974:CEM

[HV74] P. J. Hathaway and D. Van Vliet. Correspondence: An efficient method for reading and writing subscripted arrays using FOR-TRAN IV. The Computer Journal, 17(2):190–191, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/190.4.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/190.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/191.tif.

See correspondence [Art75, ET75].

Haentjens:1978:LAT

[HV78] R. Haentjens and E. Van de Vel. Lexical analysis and the transportation of programs. The Computer Journal, 21(2):190, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/190.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/190.tif.

Hill:1974:NAC

[HW74] I. D. Hill and R. W. M. Wedderburn. Note on Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 17(4):380, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/380.tif. See [Van73, Par74].

Haseman:1977:AAP

[HW77] W. D. Haseman and A. B. Whinston. Automatic application program interface to a data base. The Computer Journal, 20(3):222– 226, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/222.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/222.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/223.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/225.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/226.tif.

Hynds:1973:BR

[Hyn73] R. J. Hynds. Book reviews. The Computer Journal, 16(4):356, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/356.1.full.pdf+html.

Hyslop:1972:NAO

[Hys72] J. Hyslop. A note on the accuracy of optimisation techniques. The Computer Journal, 15(2):140, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/140.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150140.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/140.
tif.

Ibbett:1972:MIP

[Ibb72] R. N. Ibbett. The MU5 instruction pipeline. The Computer Journal, 15(1):42–50, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/1/42.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/150042.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/43.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/44.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/46.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/49.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/50.tif.

Inglis:1973:FBL

[ID73] J. Inglis and E. G. Dee. Flexibility of block-length for magnetic files. The Computer Journal, 16(4):303–307, November 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/4/303.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/160303.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/303.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/304.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/305.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/306.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/307.tif.

Ibbett:1977:MNS

[IH77] R. N. Ibbett and M. A. Husband. The MU5 name store. The Computer Journal, 20(3):227–231, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/227.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/227.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/228.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/229.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/230.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/231.tif.

Inglis:1974:BRB

[IL74] J. Inglis and S. H. Lavington. Book reviews: The Analysis of Information Systems, by C. T. Meadow, 1973; Introduction to Digital Computer Design, by D. J. Wooloms, 1973. The Computer Journal, 17(2):192, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_02/tiff/192.tif.

Inglis:1974:IIM

[Ing74] J. Inglis. Inverted indexes and multi-list structures. The Computer Journal, 17(1):59–63, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/59.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/59.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/60.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/62.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/63.tif.

Inglis:1975:BR

[Ing75] J. Inglis. Book review. The Computer Journal, 18(1):48, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/48.1.full.pdf+html.

Inglis:1976:SPI

[Ing76] J. Inglis. Structured programming and input statements. The Computer Journal, 19(2):188–189, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/188.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/188.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/189.tif. See correspondence [DI78,

Hop78].

Inglis:1977:BRa

[Ing77a] J. Inglis. Book review. The Computer Journal, 20(4):301, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/301.1.
full.pdf+html.

Inglis:1977:BRb

[Ing77b] J. Inglis. Book review. The Computer Journal, 20(4):301, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/301.2.
full.pdf+html.

Ibbett:1973:CCS

[IWT73] R. N. Ibbett, E. T. Warburton, and The Editor. Correspondence: The computing speed of a new machine. The Computer Journal, 16 (3):287, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/287.tif.

Jackson:1971:PCC

[Jac71] P. A. Jackson. PRBS cross-correlation measurements by hybrid computational techniques. The Computer Journal, 14(1):49–54, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/49.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140049.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/49.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/51.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/52.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/54.tif.

Jackson:1973:CFA

[Jac73] D. M. Jackson. Closed form approximations for random errors in distance functions. The Computer Journal, 16(1):52–56, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/1/52.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160052.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/52.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/54.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/55.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/56.tif.

Jackson:1974:BR

[Jac74] M. I. Jackson. Book review. The Computer Journal, 17(3):248, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/248.full.pdf+html.

Jackson:1977:BR

[Jac77] M. A. Jackson. Book review. The Computer Journal, 20(3):262, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/3/262.full.pdf+html.

Jacobs:1979:BR

[Jac79a] D. Jacobs. Book review. The Computer Journal, 22(3):233, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/233.2.full.pdf+html.

Jacobs:1979:CVS

[Jac79b] David Jacobs. Correspondence: Validating software. The Computer Journal, 22(3):287, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_22/Issue_03/tiff/287.tif. See

[Dun78].

Janacek:1977:BR

[Jan77] G. J. Janacek. Book review. The Computer Journal, 20(3):231, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/231.2.full.pdf+html.

Jardine:1970:DCA

[Jar70] N. Jardine. Discussion and correspondence: Algorithms, methods and models in the simplification of complex data. The Computer Journal, 13(1):116–117, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/1/116.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
130116.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/116.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_01/tiff/117.tif.

Jarvis:1975:MFM

[Jar75] J. E. Jarvis. The many faces of Multics. The Computer Journal, 18(1):2–6, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/2.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/3.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/5.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/6.tif.

Jorre:1973:DCUa

[JBD73a] R. Jorré, G. Boyfield, and J. Darkin. Data communication user research in the Post Office: Part 1. The Computer Journal, 16(3): 196–204, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/196.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/160196.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/197.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/198.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/199.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/200.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/201.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/202.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/203.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/204.tif.

Jorre:1973:DCUb

[JBD73b] R. Jorré, G. Boyfield, and J. Darkin. Data communication user research in the Post Office: Part 2. The Computer Journal, 16(4):292– 302, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/292.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160292.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/293.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_16/Issue_04/tiff/294.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/295.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/296.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/297.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/298.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/299.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/300.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/302.tif.

Jelinek:1979:ATP

[Jel79] J. Jelínek. An algebraic theory for parallel processor design. The Computer Journal, 22(4):363–375, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/363.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/363.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/364.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/365.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/366.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/367.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/370.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/372.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/373.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/374.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/375.tif.

Jennings:1971:ASV

[Jen71] A. Jennings. Algorithm 70: Solution of variable bandwidth positive definite simultaneous equations. The Computer Journal, 14(4): 446, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/446.tif.

Jenkins:1974:MDP

[Jen74] J. O. Jenkins. Modern data processing for management: a basic systems approach. The Computer Journal, 17(4):364b–364, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/4/364-b.full.pdf+html.

Jones:1974:PMR

[JO74] B. E. Jones and M. A. Ould. The patient medical record as a database. The Computer Journal, 17(4):295–301, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/295.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/295.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/296.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/297.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/298.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/299.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/301.tif.

Joffe:1976:C

[Jof76] B. Joffe. Correspondence. The Computer Journal, 19(1):95, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/95.1.full.pdf+html.

Johnson:1979:BR

[Joh79] D. B. Johnson. Book review. The Computer Journal, 22(4):306, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/306.2.full.pdf+html.

Jones:1970:SNA

[Jon70] A. Jones. Spiral — A new algorithm for non-linear parameter estimation using least squares. The Computer Journal, 13(3):301– 308, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/301.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_03/130301.sgm.abs.html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/302.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/303.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/304.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/305.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/306.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/307.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/308.tif.

Jones:1971:C

[Jon71a] A. H. Jones. Correspondence. The Computer Journal, 14(4):337, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/337.full.pdf+html.

Jones:1971:CQW

[Jon71b] A. J. Jones. Correspondence: On quality of writing. The Computer Journal, 14(4):337, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140337.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/337.tif. See [All71b].

Jones:1971:NAH

[Jon71c] C. B. Jones. A new approach to the ‘hidden line’ problem. The Computer Journal, 14(3):232–237, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/232.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140232.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/232.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/233.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/234.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/235.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/236.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/237.tif.

Jones:1972:AMI

[Jon72] H. S. P. Jones. Algorithm 72: Multiple integer arithmetic procedures in ALGOL. The Computer Journal, 15(3):281–282, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/281.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/282.tif.

Jones:1979:NDP

[Jon79] Susan Jones. A network display program. The Computer Journal, 22 (2):98–104, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/2/98.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/100.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/102.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/103.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/98.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/99.tif.

Jordan:1973:NLU

[Jor73] C. R. Jordan. A note on LISP universal S-functions. The Computer Journal, 16(2):124–125, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/124.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160124.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/124.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/125.tif.

Joseph:1970:APP

[Jos70] M. Joseph. An analysis of paging and program behaviour. The Computer Journal, 13(1):48–54, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/48.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130048.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/49.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/52.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/54.tif.

James:1976:TIC

[JP76] E. B. James and D. P. Partridge. Tolerance to inaccuracy in computer programs. The Computer Journal, 19(3):207–212, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/207.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/207.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/208.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/209.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/210.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/211.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/212.tif.

Jardine:1971:CMA

[JS71] N. Jardine and R. Sibson. Choice of methods for automatic classification. The Computer Journal, 14(4):404–406, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/404.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140404.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/404.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/405.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/406.tif.

Jackson:1972:SPN

[JW72] D. M. Jackson and L. J. White. Stability problems in nonstatistical classification theory. The Computer Journal, 15(3):214– 221, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/214.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/150214.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_

03/tiff/214.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/215.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/216.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/217.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/218.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/219.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/220.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/221.tif.

Jackson:1974:NFR

[JW74] D. M. Jackson and W. W. Wadge. Normal form reduction of probabilistic computations in non-parametric classification. The Computer Journal, 17(3):249–252, August 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/249.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/249.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/250.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/252.tif.

Kaiser:1972:JMP

[Kai72] H. F. Kaiser. The JK method: a procedure for finding the eigenvectors and eigenvalues of a real symmetric matrix. The Computer Journal, 15(3):271–273, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/271.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150271.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/271.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/272.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/273.tif.

Kammerer:1977:ER

[Kam77] P. Kammerer. Excluding regions. The Computer Journal, 20(2): 128–131, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/128.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/128.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/129.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/130.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/131.tif.

Keir:1979:CVA

[Kei79] K. M. Keir. Correspondence: Vector approximation to curves. The Computer Journal, 22(4):381, November 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/381.tif.

Kemp:1975:BR

[Kem75] R. H. Kemp. Book review. The Computer Journal, 18(3):264, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/264.1.full.pdf+html.

Kerridge:1976:PDS

[Ker76] J. M. Kerridge. PATCOSY — a database system for the National Health Service. The Computer Journal, 19(2):98– 103, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/2/98.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/100.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/101.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/102.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/103.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/98.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/99.tif.

Kaposi:1979:C

[KGC79] A. A. Kaposi, D. F. Gillies, and D. F. Cowell. Correspondence. The Computer Journal, 22(2):190–191, ???? 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/190.1.
full.pdf+html.

Khabaza:1975:BRa

[Kha75a] I. M. Khabaza. Book review. The Computer Journal, 18(1):42, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/42.2.full.pdf+html.

Khabaza:1975:BRb

[Kha75b] I. M. Khabaza. Book review. The Computer Journal, 18(2):149, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/149.full.pdf+html.

Khabaza:1976:BR

[Kha76] I. M. Khabaza. Book review. The Computer Journal, 19(1):39, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/39.full.pdf+html.

Khabaza:1977:BR

[Kha77] I. M. Khabaza. Book review. The Computer Journal, 20(3):256, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/3/256.1.full.pdf+html.

Khabaza:1978:BR

[Kha78] I. M. Khabaza. Book review. The Computer Journal, 21(3):218, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/218.2.full.pdf+html.

Khabaza:1979:BR

[Kha79] I. M. Khabaza. Book review. The Computer Journal, 22(2):104, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/2/104.1.full.pdf+html.

King:1969:ILE

[Kin69] P. J. H. King. The interpretation of limited entry decision table format and relationships among conditions. The Computer Journal, 12(4):320–326, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/4/320.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
120320.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/320.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/321.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/323.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/324.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/325.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/326.tif.  See correspondence

[Car70, Kin71b].

King:1971:C

[Kin71a] P. J. H. King. Correspondence. The Computer Journal, 14(1):54, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/54.full.pdf+html.

King:1971:CIL

[Kin71b] P. J. H. King. Correspondence: Interpretation of limited entry decision table format. The Computer Journal, 14(1):54, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/140054.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/54.tif. See

[Car70, Kin69].

King:1975:BR

[Kin75] P. J. H. King. Book review. The Computer Journal, 18(2):139, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/139.full.pdf+html.

Kirstein:1973:BR

[Kir73] P. T. Kirstein. Book review. The Computer Journal, 16(2):134, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/134.2.full.pdf+html.

Kirstein:1974:BRa

[Kir74a] P. T. Kirstein. Book review. The Computer Journal, 17(2):120, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/120.2.full.pdf+html.

Kirstein:1974:BRb

[Kir74b] P. T. Kirstein. Book review. The Computer Journal, 17(3):228, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/3/228.1.full.pdf+html.

Kirstein:1975:BR

[Kir75] P. T. Kirstein. Book review. The Computer Journal, 18(1):48, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/48.2.full.pdf+html.

Kirstein:1979:BR

[Kir79] P. T. Kirstein. Book review. The Computer Journal, 22(4):375, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/375.full.pdf+html.

Kitney:1974:BRB

[Kit74] R. I. Kitney. Book review: Computing Principles and Techniques, by B. L. Vickery, 1979. The Computer Journal, 17(4):383, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/383.tif.

Kitney:1977:BRa

[Kit77a] R. I. Kitney. Book review. The Computer Journal, 20(2):169, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/169.1.full.pdf+html.

Kitney:1977:BRb

[Kit77b] R. I. Kitney. Book review. The Computer Journal, 20(2):169, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/169.2.full.pdf+html.

Kitney:1979:BR

[Kit79] R. I. Kitney. Book review. The Computer Journal, 23(4):383, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/383.2.full.pdf+html.

King:1975:CDT

[KJ75] P. J. H. King and R. G. Johnson. The conversion of decision tables to sequential testing procedures. The Computer Journal, 18(4):298– 306, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/4/298.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/298.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/300.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/301.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/304.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/305.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/306.tif.

Kobayashi:1979:SMC

[KK79] Y. Kobayashi and T. L. Kunii. A simulation of a minicomputer controlled system and its use as a debugging tool. The Computer Journal, 22(1):41–49, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/22/1/41.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/41.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/42.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/43.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/44.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/45.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/46.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/47.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/48.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/49.tif.

Kovacs:1971:NAM

[KL71] Zs. Kovács and Shirley A. Lill. Note on Algorithm 46: a modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 14(2):214–215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_02/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_

02/tiff/215.tif. See [Lil70, Lil71].

Kaletzky:1977:POL

[KL77] A. Kaletzky and D. W. Lewin. Problem orientated language for logic design. The Computer Journal, 20(2):163–169, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/163.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/163.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/164.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/165.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/166.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/167.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/168.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/169.tif.

Kline:1975:CES

[Kli75] John V. Kline. Correspondence: On ETAOIN SHRDLU. The Computer Journal, 18(3):286, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/18/3/286.2.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_03/tiff/286.tif.

Kardasz:1974:SBS

[KM74] J. H. Kardasz and G. Molnar. A Simula-based structure oriented language for the dynamic simulation of chemical plants. The Computer Journal, 17(1):28–37, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/28.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/28.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/29.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/30.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/31.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/32.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/33.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/34.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/35.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/36.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/37.tif.

Knott:1975:HF

[Kno75a] Gary D. Knott. Hashing functions. The Computer Journal, 18(3): 265–278, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/3/265.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/265.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/266.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/268.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/269.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/270.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/271.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/272.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/273.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/274.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/275.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/277.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/

tiff/278.tif. Also published in/as: Stanford University Report, 1975. Section 3, “A history of hashing schemes”, and the lengthy bibliography, are recommended and useful resources.

Knowles:1975:BR

[Kno75b] J. S. Knowles. Book review. The Computer Journal, 18(3):222, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/222.3.full.pdf+html.

Knowles:1977:BR

[Kno77] J. S. Knowles. Book review. The Computer Journal, 20(1):26, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/26.full.pdf+html.

Kollias:1978:EST

[Kol78] J. G. Kollias. An estimate of seek time for batched searching of random or index sequential structured files. The Computer Journal, 21 (2):132–133, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/132.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/132.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/133.tif.

Komor:1974:NLF

[Kom74] T. Komor. A note on left factored languages. The Computer Journal, 17(3):242–244, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/3/242.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/242.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/243.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/244.tif.

Korn:1979:ADC

[KP79] G. A. Korn and O. A. Palusinski. Application of a digital computer simulation language to control hybrid computer Monte Carlo studies. The Computer Journal, 22(2):136–141, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/136.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/136.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/137.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/139.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/140.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/141.tif.

Kral:1971:SPS

[Krá71] J. Král. Some properties of the scatter storage technique with linear probing. The Computer Journal, 14(2):145–149, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/145.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140145.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/147.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/149.tif.

Kramer:1977:C

[Kra77] J. Kramer. Correspondence. The Computer Journal, 20(4):382–383, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/382.full.pdf+html.

Kohavi:1973:MDE

[KRK73] Z. Kohavi, J. A. Rivierre, and I. Kohavi. Machine distinguishing experiments. The Computer Journal, 16(2):141–147, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/141.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/160141.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/141.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/142.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/143.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/144.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/145.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/146.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/147.
tif.

Kronsjo:1970:AF

[Kro70a] T. O. M. Kronsjö. Algorithm 57: Find. The Computer Journal, 13(4):425–426, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_04/tiff/425.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/426.tif.

Kronsjo:1970:AIPa

[Kro70b] T. O. M. Kronsjö. Algorithm 58: an illustrative primal simplex linear program. The Computer Journal, 13(4):426– 428, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_04/tiff/426.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/427.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/428.tif. See note [Kro71].

Kronsjo:1970:AIPb

[Kro70c] T. O. M. Kronsjö. Algorithm 59: An illustrative primal simplex linear program using multipliers. The Computer Journal, 13 (4):428–429, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_04/tiff/428.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_

04/tiff/429.tif. See note [Kro71].

Kronsjo:1970:AIPc

[Kro70d] T. O. M. Kronsjö. Algorithm 60: An illustrative primal simplex linear program using multipliers and single dimensional arrays. The Computer Journal, 13(4):429–430, November 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/
tiff/429.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_04/tiff/430.tif. See note [Kro71].

Kronsjo:1970:AIS

[Kro70e] T. O. M. Kronsjö. Algorithm 61: An illustrative self-dual parametric simplex linear program using multipliers. The Computer Journal, 13(4):430–437, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_04/tiff/430.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/431.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/432.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/433.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/434.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/435.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/436.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/437.tif. See [Kro71].

Kronsjo:1971:NA

[Kro71] T. O. M. Kronsjö. Note on Algorithms 58 to 61. The Computer Journal, 14(2):215, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_02/tiff/215.tif. See

[Kro70b, Kro70c, Kro70d, Kro70e].

Keedy:1979:ISS

[KRR79] J. L. Keedy, K. Ramamohanarao, and J. Rosenberg. On implementing semaphores with sets. The Computer Journal, 22(2): 146–150, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/2/146.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/146.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/147.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/148.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/149.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/150.tif.

Knibb:1971:SPP

[KS71] D. Knibb and R. E. Scraton. On the solution of parabolic partial differential equations in Chebyshev series. The Computer Journal, 14(4):428–432, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/4/428.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140428.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/428.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/429.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/430.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/431.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/432.tif.

Kollias:1977:IPI

[KSD77] J. G. Kollias, P. M. Stocker, and P. A. Dearnley. Improving the performance of an intelligent data management system.

The Computer Journal, 20(4):302–307, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/302.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/302.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/303.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/304.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/305.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/306.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/307.tif.

Kirby:1973:CLB

[KTW^{+}73] G. H. Kirby, The Editor, R. Wane, I. D. Hill, and Brian Meek. Correspondence: Letters to The Computer Journal. The Computer Journal, 16(2):189–190, May 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:

//comjnl.oxfordjournals.org/content/16/2/189.2.full.pdf+
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/189.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/190.tif. See [Wal71b,

GT72, MT72].

Khalil:1974:LSN

[KU74] H. M. Khalil and D. L. Ulery. LINEAL: a system for numerical linear algebra. The Computer Journal, 17(3):267–274, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/267.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/267.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/268.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/269.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/270.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/271.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/273.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/274.tif.

Lambert:1976:CSO

[Lam76] J. D. Lambert. Computer solution of ordinary differential equations. The Computer Journal, 19(2):155, May 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/2/155.full.pdf+html.

Lansdown:1971:C

[Lan71a] R. John Lansdown. Correspondence. The Computer Journal, 14 (2):155, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/155.1.full.pdf+html.

Lansdowne:1971:CUS

[Lan71b] R. John Lansdowne. Correspondence: On “The use of symbolstate tables”. The Computer Journal, 14(2):155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_02/tiff/155.tif. See [Day70].

Land:1974:BR

[Lan74a] F. F. Land. Book review. The Computer Journal, 17(2):191, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/191.2.full.pdf+html.

Land:1974:FAI

[Lan74b] F. F. Land. Functional analysis of information processing. The Computer Journal, 17(4):376, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/376.full.pdf+html.

Lang:1974:BR

[Lan74c] C. A. Lang. Book review. The Computer Journal, 17(1):63, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/63.1.full.pdf+html.

Lance:1976:C

[Lan76a] G. N. Lance. Correspondence. The Computer Journal, 19(4):380–381, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/380.2.full.pdf+html.

Land:1976:ESG

[Lan76b] F. F. Land. Evaluation of systems goals in determining a design strategy for a computer based information system. The Computer Journal, 19(4):290–294, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/4/290.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/290.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/291.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/292.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/293.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/294.tif.

Laski:1973:BR

[Las73] J. G. Laski. Book review. The Computer Journal, 16(1):38, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
1/38.full.pdf+html.

Last:1979:BR

[Las79] R. W. Last. Book review. The Computer Journal, 22(4):351, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/351.full.pdf+html.

Lauer:1975:DPD

[Lau75] H. C. Lauer. Discussion on Ph.D. thesis proposals in computing science. The Computer Journal, 18(3):279–281, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/279.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/279.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/280.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_

03/tiff/281.tif. See note [HL76b].

Laver:1970:JRD

[Lav70] F. J. M. Laver. JANUS — reflections on DATAFAIR 69. The Computer Journal, 13(1):2–6, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/1/2.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
130002.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/2.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_01/tiff/3.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/4.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/5.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_01/tiff/6.tif.

Laver:1974:BR

[Lav74] F. J. M. Laver. Book review. The Computer Journal, 17(2):142, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/142.1.full.pdf+html.

Lavington:1978:BR

[Lav78] S. Lavington. Book review. The Computer Journal, 21(3):198, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/198.1.full.pdf+html.

Lim:1976:PVN

[LB76] R. T. C. Lim and J. D. Blackburn. On program verification in numerical methods. The Computer Journal, 19(3):272–273, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/272.tif. See [Smi72].

Learner:1975:C

[Lea75] A. Learner. Correspondence. The Computer Journal, 18(2):188, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/188.2.full.pdf+html.

Leavenworth:1977:NPD

[Lea77] B. M. Leavenworth. Non-procedural data processing. The Computer Journal, 20(1):6–9, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/1/6.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/
tiff/6.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/7.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/8.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/9.
tif.

Lea:1978:TCA

[Lea78] R. M. Lea. Text compression with an associative parallel processor. The Computer Journal, 21(1):45–56, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
21/1/45.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/46.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/49.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/50.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/52.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/54.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/55.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/56.tif.

Lea:1979:BR

[Lea79] R. M. Lea. Book review. The Computer Journal, 22(4):327, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/327.2.full.pdf+html.

Lee:1972:FDB

[Lee72] J. A. N. Lee. The formal definition of the BASIC language. The Computer Journal, 15(1):37–41, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
15/1/37.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150037.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/37.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/38.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/39.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/

tiff/40.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/41.tif. See correspondence [Bar72b].

Leech:1974:CSA

[Lee74] J. Leech. Correspondence: On the spelling of ALGORITHM. The Computer Journal, 17(4):383, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/

tiff/383.tif. See [Mar74d, Mul75].

Leech:1975:BR

[Lee75] J. Leech. Book review. The Computer Journal, 18(4):365, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/365.1.
full.pdf+html.

Lee:1978:UCS

[Lee78a] B. S. Lee. Use of case studies in teaching data processing. The Computer Journal, 21(4):366–372, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/366.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/366.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/367.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/370.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/372.tif.

Lee:1978:RRB

[Lee78b] P. A. Lee. A reconsideration of the recovery block scheme. The Computer Journal, 21(4):306–310, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/306.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/306.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/307.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/308.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_21/Issue_04/tiff/309.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/310.tif.

Lee:1979:BRa

[Lee79a] P. A. Lee. Book review. The Computer Journal, 22(1):21, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/21.1.full.pdf+html.

Lee:1979:BRb

[Lee79b] P. A. Lee. Book review. The Computer Journal, 22(2):150, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/150.2.
full.pdf+html.

Lee:1979:BRc

[Lee79c] P. A. Lee. Book review. The Computer Journal, 22(3):239, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/239.1.
full.pdf+html.

Lehman:1973:CUC

[Leh73] M. M. Lehman. Computer usage control. The Computer Journal, 16 (2):106–110, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/106.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160106.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/106.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/107.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/108.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/109.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/110.tif.

Lehman:1974:DCS

[Leh74] M. M. Lehman. Digital computer systems principles. The Computer Journal, 17(4):364, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/364.full.pdf+html.

Lehman:1975:BR

[Leh75] M. M. Lehman. Book review. The Computer Journal, 18(3):242, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/242.1.full.pdf+html.

Lehman:1976:BR

[Leh76] M. M. Lehman. Book review. The Computer Journal, 19(3):253, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/253.1.full.pdf+html.

Lehman:1977:BR

[Leh77] M. M. Lehman. Book review. The Computer Journal, 20(4):349, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/349.3.full.pdf+html.

Lemos:1979:DCM

[Lem79] R. S. Lemos. Discussion and correspondence: Measuring students’ attitudes towards structured walk-throughs. The Computer Journal, 23 (4):377–383, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/4/377.full.pdf+html.

Lewin:1971:AAA

[Lew71a] D. W. Lewin. Advanced aspects of asynchronous logic design. The Computer Journal, 14(3):254–259, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/254.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140254.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/256.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/257.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/258.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/259.tif.

Lewis:1971:SAI

[Lew71b] R. P. W. Lewis. The selection of ADI iteration parameters by numerical experiment for the solution of Poisson’s equation over a circular area. The Computer Journal, 14(1):73–74, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/73.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140073.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/73.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/74.tif.

Lewis:1972:CIN

[Lew72a] B. A. Lewis. Correspondence: On “Improved numerical methods for Volterra integral equations of the first kind”. The Computer Journal, 15(2):137, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150137.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/137.tif. See [AW71].

Lewis:1972:INM

[Lew72b] B. A. Lewis. Improved numerical methods for Volterra integral equations of the first kind. The Computer Journal, 15(2):137, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
2/137.full.pdf+html.

Lewin:1973:BR

[Lew73] D. W. Lewin. Book review. The Computer Journal, 16(3):222, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/222.2.full.pdf+html.

Lewin:1976:BR

[Lew76] D. Lewin. Book review. The Computer Journal, 19(4):314, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/314.4.
full.pdf+html.

Lewins:1978:SS

[Lew78] J. Lewins. Simulation of systems. The Computer Journal, 21(2):167, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/2/167.full.pdf+html.

Lewins:1979:BR

[Lew79] J. Lewins. Book review. The Computer Journal, 22(3):261, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/261.1.full.pdf+html.

Liddell:1974:BRa

[Lid74a] H. M. Liddell. Book review. The Computer Journal, 17(1):27, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/27.1.full.pdf+html.

Liddell:1974:BRb

[Lid74b] H. M. Liddell. Book review. The Computer Journal, 17(3):204, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/204.1.full.pdf+html.

Liggins:1973:AAC

[Lig73] D. Liggins. Algorithm 79: an algorithm for the capacitated transportation problem. The Computer Journal, 16(3):276–278, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/tiff/277.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/278.tif.

Lill:1970:AMD

[Lil70] Shirley A. Lill. Algorithm 46: a modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 13(1):111– 113, 106, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_01/tiff/111.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/113.tif. See note [Lil71, KL71].

Lill:1971:NAM

[Lil71] Shirley A. Lill. Note on Algorithm 46: a modified Davidon method for finding the minimum of a function using difference approximation for derivatives. The Computer Journal, 14(1):106, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/106.tif. See [Lil70, KL71].

Lim:1974:PAP

[Lim74] A. L. Lim. A PL/I approach to programming languages. The Computer Journal, 17(4):312, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/312.full.pdf+html.

Lim:1975:BR

[Lim75] A. L. Lim. Book review. The Computer Journal, 18(1):42, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/42.1.full.pdf+html.

Lindsey:1972:AFT

[Lin72a] C. H. Lindsey. Algol 68 with fewer tears. The Computer Journal, 15 (2):176–188, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/176.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/150176.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/176.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/177.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/178.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/179.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/180.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/181.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/182.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/183.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/184.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/185.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/186.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/187.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/188.tif.

Ling:1972:TCC

[Lin72b] R. F. Ling. On the theory and construction of k-clusters. The Computer Journal, 15(4):326–332, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/326.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150326.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/326.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/327.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/328.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/329.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/330.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/331.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/332.
tif.

Lindley:1973:BR

[Lin73] D. V. Lindley. Book review. The Computer Journal, 16(4):307, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/307.2.full.pdf+html.

Lipsky:1979:STS

[Lip79] L. Lipsky. A study of time sharing systems considered as queueing networks of exponential servers. The Computer Journal, 23(4): 290–297, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/23/4/290.full.pdf+html.

Lister:1974:VSP

[Lis74] A. M. Lister. Validation of systems of parallel processes. The Computer Journal, 17(2):148–151, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/148.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/148.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/150.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/151.tif.

Littlewood:1970:APE

[Lit70] P. G. Littlewood. Algorithm 48: Procedure for the evaluation of an integral occurring in the mean square response analysis of linear systems. The Computer Journal, 13(2): 207–208, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_02/tiff/206.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/207.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/208.tif. See note [GE71].

Lucas:1976:SPE

[LK76] H. C. Lucas, Jr. and R. B. Kaplan. A structured programming experiment. The Computer Journal, 19(2):136–138, May 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/136.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/136.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/137.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/138.tif.

Lyness:1977:TCA

[LK77] J. N. Lyness and J. J. Kaganove. A technique for comparing automatic quadrature routines. The Computer Journal, 20(2):170– 177, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/170.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/170.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/171.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/172.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/173.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/176.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/177.tif.

Lavington:1971:EPU

[LKK71] S. H. Lavington, D. J. Kinniment, and A. E. Knowles. An experimental paging unit. The Computer Journal, 14(1):55–60, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/55.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140055.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/55.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/56.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/57.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/58.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/59.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/60.tif.

Learner:1970:NTC

[LL70] A. Learner and A. L. Lim. A note on transforming context-free grammars to Wirth–Weber precedence form. The Computer Journal, 13 (2):142–144, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/2/142.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_02/130142.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/142.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/143.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/144.tif.

Lim:1975:TSP

[LL75] A. L. Lim and G. R. Lewis. Towards structured programs in APL. The Computer Journal, 18(2):140–143, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/140.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/140.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/141.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/142.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/143.tif.

Llewellyn:1973:DEF

[Lle73] J. A. Llewellyn. The deadly embrace — a finite state model approach. The Computer Journal, 16(3):223–225, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/223.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160223.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/225.tif.

Lloyd:1970:BR

[Llo70] E. Lloyd. Book review. The Computer Journal, 13(2):144, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
2/144.full.pdf+html.

Lloyd:1974:BR

[Llo74a] C. Lloyd. Book review. The Computer Journal, 17(2):186, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/186.full.pdf+html.

Lloyd:1974:BRB

[Llo74b] Cliff Lloyd. Book review: Structured programming, by O.-J. Dahl, E. W. Dijkstra and C. A. R. Hoare. The Computer Journal, 17 (2):186, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_02/tiff/186.tif.

Leaman:1973:DTP

[LLR73] R. J. Leaman, M. H. Lloyd, and C. S. Repton. The development and testing of a processor self-test program. The Computer Journal, 16(4):308–314, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/4/308.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/
160308.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/308.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/309.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/310.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/311.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/312.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/313.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/314.tif.

Levy:1977:AAP

[LM77a] L. S. Levy and R. Melville. The algebraic anatomy of programs. The Computer Journal, 20(4):340–345, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/340.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/340.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/341.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/342.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/343.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/344.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/345.tif.

Lowden:1977:SCD

[LM77b] B. G. T. Lowden and I. R. MacCallum. Some consequences of deferred binding in COBOL. The Computer Journal, 20(1):51–55, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/51.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/51.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/52.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/53.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/54.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/55.
tif.

Loetzsch:1979:SIS

[LMP79] J. Loetzsch, M. Machura, and W. Paetzold. Sample implementation of a special purpose language by the DEPOT system. The Computer Journal, 22(4):357–362, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/357.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/357.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/358.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/359.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_22/Issue_04/tiff/360.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/361.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/362.tif.

Lozier:1973:BCP

[LMS73] D. W. Lozier, L. C. Maximon, and W. L. Sadowski. A bit comparison program for algorithm testing. The Computer Journal, 16 (2):111–117, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/111.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160111.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/111.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/112.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/113.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/115.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/116.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/117.tif.

Leipala:1979:PSS

[LN79] T. Leipälä and O. Nevalainen. A plotter sequencing system. The Computer Journal, 22(4):313–316, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/313.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/313.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/314.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/315.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/316.tif.

Lanzarone:1975:PCR

[LO75] G. A. Lanzarone and M. Ornaghi. Program construction by refinements preserving correctness. The Computer Journal, 18(1):55– 62, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/18/1/55.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/55.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/56.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_18/Issue_01/tiff/57.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/58.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/59.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/60.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/61.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/62.tif.

LoDato:1970:PCC

[LoD70] V. A. LoDato. The permutation of a certain class of matrices. The Computer Journal, 13(4):405–410, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/405.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130405.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/405.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/406.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/407.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/408.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/409.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/410.tif.

LoDato:1971:MDP

[LoD71] V. A. LoDato. The minimisation of distance in placement algorithms. The Computer Journal, 14(3):260–262, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/260.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140260.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/260.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/261.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/262.tif.

Loizou:1972:QCJ

[Loi72] G. Loizou. On the quadratic convergence of the Jacobi method for normal matrices. The Computer Journal, 15(3):274–276, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/3/274.full.pdf+html;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_15/Issue_03/150274.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/274.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/275.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/276.tif.

Loretan:1976:TPE

[Lor76] R. P. Lorétan. Timing problems in electronic lockout chains. The Computer Journal, 19(3):250–253, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/250.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/250.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/251.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/252.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/253.tif.

Losty:1977:BRa

[Los77a] P. A. Losty. Book review. The Computer Journal, 20(3):226, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/226.1.full.pdf+html.

Losty:1977:BRb

[Los77b] P. A. Losty. Book review. The Computer Journal, 20(3):226, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/226.2.full.pdf+html.

Lowden:1977:DCN

[Low77a] B. G. T. Lowden. Discussion and correspondence: a note on the oscillating sort. The Computer Journal, 20(1):92–96, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/92.full.pdf+html.

Lowden:1977:NOS

[Low77b] B. G. T. Lowden. A note on the oscillating sort. The Computer Journal, 20(1):92, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_20/Issue_01/tiff/92.tif.

Lucas:1973:MSE

[LP73a] H. C. Lucas, Jr. and L. Presser. A method of software evaluation: the case of programming language translators. The Computer Journal, 16(3):226–231, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/3/226.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
160226.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/226.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/227.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/228.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/229.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/230.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/231.tif.

Lynch:1973:PSP

[LP73b] M. F. Lynch and J. H. Petrie. A program suite for the production of articulated subject indexes. The Computer Journal, 16(1): 46–51, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/46.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/160046.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/46.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/47.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/48.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/49.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/51.
tif.

Lubran:1972:SOL

[LR72] J. F. Lubran and J. D. Roberts. Some observations on ‘least time to go’ scheduling. The Computer Journal, 15(1):32–36, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/32.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150032.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/32.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_15/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_01/tiff/34.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/35.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/36.tif.

Lewis:1978:TPR

[LR78] B. A. Lewis and J. S. Robinson. Triangulation of planar regions with applications. The Computer Journal, 21(4):324–332, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/324.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/324.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/325.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/326.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/327.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/328.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/329.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/330.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/331.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/332.tif.

Little:1978:CMF

[LS78] W. Little and D. A. Smith. Correspondence: On “Modifications to the Forsythe–Moler algorithm for solving linear algebraic systems”. The Computer Journal, 21(4):383, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/383.tif. See [SG78, Hit78].

Leonard:1971:NFO

[LT71] A. J. Leonard and M. E. Tribe. Notable features of Orion. The Computer Journal, 14(4):344–347, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/344.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140344.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/344.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_14/Issue_04/tiff/346.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/347.
tif.

Lotfi:1979:MNN

[LT79] Z. M. Lotfi and A. J. Tosser. Minimising the NAND-NOR-XOR network of modulo 2 sum of Boolean products. The Computer Journal, 22(3):251–255, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/3/251.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/251.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/253.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/254.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/255.tif.

Lunnon:1969:PSP

[Lun69] W. F. Lunnon. A postage stamp problem. The Computer Journal, 12(4):377–380, November 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/4/377.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_04/
120377.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/377.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_04/tiff/378.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
04/tiff/379.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_04/tiff/380.tif.  See correspondence

[Sel72b].

Lunnon:1971:MDM

[Lun71] W. F. Lunnon. Multi-dimensional map-folding. The Computer Journal, 14(1):75–80, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/1/75.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/140075.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/75.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/76.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/78.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/80.tif.

Lunnon:1975:CMP

[Lun75] W. F. Lunnon. Counting multidimensional polyominoes. The Computer Journal, 18(4):366–367, November 1975. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/4/366.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/366.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/367.tif.

Lazos:1979:MEB

[LV79] C. Lazos and A. Vafiadis. A method to estimate the I/O buffer size in a computer system. The Computer Journal, 22(4):323– 327, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/22/4/323.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/323.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/324.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_04/tiff/325.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_04/tiff/326.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_04/
tiff/327.tif.

Lance:1971:NND

[LW71] G. N. Lance and W. T. Williams. A note on a new divisive classificatory program for mixed data. The Computer Journal, 14(2): 154–155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/154.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140154.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/155.tif.

Lazos:1979:ICU

[LY79] C. Lazos and J. R. Yandle. Improving CPU utilisation in a multiprogramming system. The Computer Journal, 22(3):203–205, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/

content/22/3/203.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/203.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/204.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/205.tif.

Lyness:1969:EIC

[Lyn69] J. N. Lyness. The effect of inadequate convergence criteria in automatic routines. The Computer Journal, 12(3):279–281, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/12/3/279.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/120279.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/tiff/279.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
12/Issue_03/tiff/280.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_12/Issue_03/tiff/281.tif. See correspon-

dence [GL70].

Lynn:1972:RDF

[Lyn72] P. A. Lynn. Recursive digital filters with linear-phase characteristics. The Computer Journal, 15(4):337–342, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/337.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150337.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/337.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/338.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/339.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/342.tif.

MacCallum:1970:SIG

[Mac70a] K. J. MacCallum. Surfaces for interactive graphical design. The Computer Journal, 13(4):352–358, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/352.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130352.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/352.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/353.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/354.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/355.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/356.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/357.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/358.
tif.

Macleod:1970:SFI

[Mac70b] I. A. Macleod. SP/1 — a Fortran integrated string processor. The Computer Journal, 13(3):255–260, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/255.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130255.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/255.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/256.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/257.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/258.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/259.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/260.tif.

MacCallum:1971:C

[Mac71a] K. J. MacCallum. Correspondence. The Computer Journal, 14(1):9, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/9.full.pdf+html.

Macleod:1971:MFM

[Mac71b] I. A. Macleod. Mp/1 — a Fortran macroprocessor. The Computer Journal, 14(3):229–231, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/229.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140229.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/229.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/230.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/231.tif.

MacCallum:1973:SAO

[Mac73] I. R. MacCallum. A simple analysis of the n^{th} order polyphase sort. The Computer Journal, 16(1):16–18, February 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/1/16.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160016.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/16.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/17.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/18.tif. See errata [Cam73].

MacCallum:1974:C

[Mac74a] K. J. MacCallum. Correspondence. The Computer Journal, 17 (2):188, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/188.1.full.pdf+html.

MacCallum:1974:CMI

[Mac74b] K. J. MacCallum. Correspondence: On mixing interpretive and machine code. The Computer Journal, 17(2):188, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/188.tif. See [Daw73, DP73, Bem74].

Maes:1978:RPS

[Mae78] R. Maes. On the representation of program structures by decision tables: a critical assessment. The Computer Journal, 21(4):290– 295, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/290.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/290.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/291.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/292.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/293.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/294.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/295.tif.

Mallach:1977:SST

[Mal77] E. G. Mallach. Scatter storage techniques: a uniform viewpoint and a method for reducing retrieval times. The Computer Journal, 20 (2):137–140, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/137.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/137.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/138.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/139.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/140.tif.

Man:1971:HOM

[Man71a] F. T. Man. A high-order method of solution for the Lyapunov matrix equation. The Computer Journal, 14(3):291–292, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/291.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140291.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/291.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/292.tif.

Manning:1971:AWM

[Man71b] J. R. Manning. Algorithm 68: White to move and mate in n moves. The Computer Journal, 14(2):209–213, May 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/
tiff/209.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_02/tiff/210.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/tiff/211.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/
tiff/212.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_02/tiff/213.tif.

Manning:1974:CCS

[Man74] J. R. Manning. Continuity conditions for spline curves. The Computer Journal, 17(2):181–186, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/181.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/181.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/182.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/183.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/184.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_17/Issue_02/tiff/185.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/186.tif. See

correspondence [San75].

Marovac:1973:SDD

[Mar73] N. Marovac. A single data-display structure: a new view on interactive computer graphics in CAD. The Computer Journal, 16 (2):152–156, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/152.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160152.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/152.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/153.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/154.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/155.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/156.tif.

Marovac:1974:BR

[Mar74a] N. Marovac. Book review. The Computer Journal, 17(3):244, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/244.1.full.pdf+html.

Marovac:1974:MDG

[Mar74b] N. Marovac. A method for defining general networks for CAD, using interactive computer graphics. The Computer Journal, 17(4):332– 336, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/332.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/332.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/333.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/335.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/336.tif.

Marriott:1974:C

[Mar74c] H. A. Marriott. Correspondence. The Computer Journal, 17(2):187, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/187.1.full.pdf+html.

Marriott:1974:CSA

[Mar74d] H. A. Marriott. Correspondence: On the spelling of ALGO-RITHM. The Computer Journal, 17(2):187, May 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/

tiff/187.tif. See [Lee74, Mul75].

Marjasoo:1976:MHS

[Mar76] O. Marjasoo. A methodology for hardware and software availability analysis. The Computer Journal, 19(2):132–135, May 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/132.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/133.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/134.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/135.tif.

Marovac:1979:DCM

[Mar79a] N. Marovac. Discussion and correspondence: Microcomputers in computer science programmes. The Computer Journal, 22(1):91– 94, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/1/91.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/91.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/92.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/93.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/94.tif.

Marriott:1979:BR

[Mar79b] I. Marriott. Book review. The Computer Journal, 22(4):327, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/327.1.full.pdf+html.

Maude:1973:IMG

[Mau73] A. D. Maude. Interpolation — mainly for graph plotters. The Computer Journal, 16(1):64–65, February 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/1/64.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160064.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/64.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/65.
tif.

Mayers:1974:BR

[May74] D. F. Mayers. Book review. The Computer Journal, 17(2):164, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/164.1.full.pdf+html.

Mayne:1976:BR

[May76] D. Q. Mayne. Book review. The Computer Journal, 19(3):228, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/228.1.full.pdf+html.

May:1979:BR

[May79] D. May. Book review. The Computer Journal, 23(4):359, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/23/4/359.1.
full.pdf+html.

Mazlack:1978:PSS

[Maz78] L. J. Mazlack. Predicting student success in an introductory programming course. The Computer Journal, 21(4):380–381, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/380.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/380.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/381.tif.

McCallien:1970:SRD

[McC70a] C. W. J. McCallien. The solution of reactor diffusion problems. The Computer Journal, 13(4):369–377, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/369.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130369.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/369.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/370.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_04/tiff/371.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/372.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/373.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/374.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/375.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/377.tif.

McConalogue:1970:QIS

[McC70b] D. J. McConalogue. A quasi-intrinsic scheme for passing a smooth curve through a discrete set of points. The Computer Journal, 13(4):392–396, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/4/392.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/
130392.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/392.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/393.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/394.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/395.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/396.tif.

See correspondence [Ecc71b, Smi71, Sab72, Coo72f].

McConalogue:1971:AAF

[McC71a] D. J. McConalogue. Algorithm 66: An automatic French-curve procedure for use with an incremental plotter. The Computer Journal, 14(2):207–209, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_02/tiff/207.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/208.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/209.tif. See note [McC72].

McConalogue:1971:AAI

[McC71b] D. J. McConalogue. Algorithm 67: An axis invariant procedure to provide slopes at points by local three-point parametric curve fitting. The Computer Journal, 14(2):209, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_02/tiff/209.tif. See note [SM71].

McConalogue:1972:NAA

[McC72] D. J. McConalogue. Note on Algorithm 66: An automatic Frenchcurve procedure for use with an incremental plotter. The Computer Journal, 15(3):285, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_15/Issue_03/tiff/285.tif. See

[McC71a].

McConalogue:1978:ACI

[McC78] D. J. McConalogue. Algorithm 102: Convolution integrals involving probability distribution functions. The Computer Journal, 21(3):270–272, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_21/Issue_03/tiff/270.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/271.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/272.tif.

McDonell:1977:III

[McD77] K. J. McDonell. An inverted index implementation. The Computer Journal, 20(2):116–123, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/2/116.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/116.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/117.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/118.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/119.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/120.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/121.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/122.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/123.tif.

McGettrick:1977:TMC

[McG77] A. D. McGettrick. Teaching mathematics by computer. The Computer Journal, 20(3):263–268, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/3/263.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/263.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_20/Issue_03/tiff/265.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/266.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/268.tif.

McKay:1979:BR

[McK79] M. McKay. Book review. The Computer Journal, 22(1):62, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/62.3.full.pdf+html.

McLain:1974:DCA

[McL74] D. H. McLain. Drawing contours from arbitrary data points. The Computer Journal, 17(4):318–324, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/318.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/318.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/319.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/320.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/321.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/322.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/323.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/324.tif.

McLain:1976:TDI

[McL76] D. H. McLain. Two dimensional interpolation from random data. The Computer Journal, 19(2):178–181, May 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/178.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/178.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/179.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/180.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/181.tif.

McLain:1978:AVA

[McL78] D. H. McLain. Algorithm 100: Vector approximation to curves. The Computer Journal, 21(2):178–180, May 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/178.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/179.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/180.tif.

Matthews:1971:CMN

[MD71] A. Matthews and D. Davies. A comparison of modified Newton methods for unconstrained optimisation. The Computer Journal, 14(3): 293–294, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/3/293.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_03/140293.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/293.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/294.tif.

Morris:1972:SMO

[MDFS72] D. Morris, G. D. Detlefsen, G. R. Frank, and T. J. Sweeney. The structure of the MU5 operating system. The Computer Journal, 15 (2):113–116, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/113.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_02/150113.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/113.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/115.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/116.tif.

Meadows:1973:AFM

[Mea73] N. G. Meadows. Automatic frequency and Mikhailov locus plotting for systems with distributed lags: a new analogue computer technique. The Computer Journal, 16(3):266–268, August 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/266.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/160266.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_03/tiff/267.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/268.tif.

Meditch:1970:NMD

[Med70] J. S. Meditch. Newton’s method in discrete-time nonlinear data smoothing. The Computer Journal, 13(4):387–391, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/387.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130387.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/387.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/388.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/389.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/390.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/391.tif.

Meek:1971:C

[Mee71] B. L. Meek. Correspondence. The Computer Journal, 14(4):332, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/332.full.pdf+html.

Meek:1972:C

[Mee72a] B. L. Meek. Correspondence. The Computer Journal, 15(3):267, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/267.full.pdf+html.

Meek:1972:CSE

[Mee72b] B. L. Meek. Correspondence: On “Suggested Extensions to FOR-TRAN IV”. The Computer Journal, 15(3):267, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/150267.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_15/Issue_03/tiff/267.tif. See

[Fin72a].

Meek:1973:BR

[Mee73] B. L. Meek. Book review. The Computer Journal, 16(4):379, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/379.2.full.pdf+html.

Merret:1970:C

[Mer70a] T. H. Merret. Correspondence. The Computer Journal, 13(2):221, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/221.1.full.pdf+html.

Merret:1970:CCP

[Mer70b] T. H. Merret. Correspondence: Chain processing in PL/1. The Computer Journal, 13(2):221, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_13/Issue_02/tiff/221.tif.

Mercer:1975:BRa

[Mer75a] C. A. Mercer. Book review. The Computer Journal, 18(2):134, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/134.full.pdf+html.

Mercer:1975:BRb

[Mer75b] C. A. Mercer. Book review. The Computer Journal, 18(3):226, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/226.full.pdf+html.

Mercer:1975:BRc

[Mer75c] C. A. Mercer. Book review. The Computer Journal, 18(3):230, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/230.1.full.pdf+html.

Macnaghten:1977:FFT

[MH77] A. M. Macnaghten and C. A. R. Hoare. Fast Fourier transform free from tears. The Computer Journal, 20(1):78–83, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/78.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/80.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/82.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/83.tif.

Michie:1971:HS

[Mic71] D. Michie. Heuristic search. The Computer Journal, 14(1):96– 102, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/14/1/96.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_01/140096.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/100.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/101.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/102.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/96.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/97.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/98.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/99.tif.

Michie:1979:ES

[Mic79] D. Michie. Expert systems. The Computer Journal, 23(4):369–376, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/369.full.pdf+html.

Middleton:1976:UMI

[Mid76] A. G. Middleton. On the use of macros for iteration. The Computer Journal, 19(2):170–172, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/170.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/170.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/172.tif.

Middleton:1977:CTF

[Mid77] A. G. Middleton. A case for type and form flow analysis. The Computer Journal, 20(3):238–241, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/238.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/238.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/239.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/241.tif.

Millenson:1970:LLS

[Mil70] J. R. Millenson. Language and list structure of a compiler for experimental control. The Computer Journal, 13(4):340–343, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/340.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/130340.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/340.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/341.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/342.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/343.tif.

Miller:1979:BR

[Mil79] G. F. Miller. Book review. The Computer Journal, 22(2):163, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/163.1.full.pdf+html.

Mitchell:1970:BR

[Mit70] A. R. Mitchell. Book review. The Computer Journal, 13(2):126, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/2/126.full.pdf+html.

Mitchem:1976:VAE

[Mit76] John Mitchem. On various algorithms for estimating the chromatic number of a graph. The Computer Journal, 19(2):182–183, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/182.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/182.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/183.tif.

Mitra:1977:BR

[Mit77] G. Mitra. Book review. The Computer Journal, 20(3):256, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/256.3.
full.pdf+html.

Mayne:1975:ICF

[MJ75] A. Mayne and E. B. James. Information compression by factorising common strings. The Computer Journal, 18(2):157–160, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/157.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/157.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/158.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/160.tif.

Morris:1971:F

[MKL71] D. Morris, T. G. Kennedy, and L. Last. Flocoder. The Computer Journal, 14(3):221–223, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/221.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140221.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/221.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/222.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/223.tif.

McKee:1970:ADM

[MM70] S. McKee and A. R. Mitchell. Alternating direction methods for parabolic equations in two space dimensions with a mixed derivative. The Computer Journal, 13(1):81–86, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
13/1/81.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130081.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/81.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/82.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/83.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/84.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/85.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/86.tif.

McKee:1971:ADM

[MM71] S. McKee and A. R. Mitchell. Alternating direction methods for parabolic equations in three space dimensions with mixed derivatives. The Computer Journal, 14(3):295–300, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/295.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140295.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/295.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/296.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/297.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/298.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/299.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/300.tif.

Mottl:1978:DPS

[MM78] J. Mottl and L. Mottlová. Description of a program for solving problems of nonlinear programming. The Computer Journal, 21(3): 263–269, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/263.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/263.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/264.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/265.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/266.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/267.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/268.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/269.tif.

Mojena:1977:HGM

[Moj77] R. Mojena. Hierarchical grouping methods and stopping rules: An evaluation. The Computer Journal, 20(4):359–363, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/359.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/359.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/360.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/361.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_20/Issue_04/tiff/362.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/363.tif.

Molnar:1971:SSE

[Mol71] G. Molnar. SEL — a self-extensible programming language. The Computer Journal, 14(3):238–242, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/238.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140238.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/238.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/239.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/240.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/241.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/242.tif.

Moon:1972:C

[Moo72] G. Moon. Correspondence. The Computer Journal, 15(2):124, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
2/124.full.pdf+html.

Moore:1979:DTJ

[Moo79] L. Moore. Design for a transportable job organisation language. The Computer Journal, 22(4):296–302, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/296.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/296.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/297.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/298.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/299.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/300.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/301.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/302.tif.

Morton:1975:BR

[Mor75] K. W. Morton. Book review. The Computer Journal, 18(3):264, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/3/264.2.full.pdf+html.

Morice:1976:BR

[Mor76] P. B. Morice. Book review. The Computer Journal, 19(4):332, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/332.1.full.pdf+html.

Morton:1978:RRN

[Mor78] K. W. Morton. Review: Recoding of natural language for economy of transmission or storage. The Computer Journal, 21(1):44, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/44.full.pdf+html.

Morris:1979:BR

[Mor79] D. Morris. Book review. The Computer Journal, 22(2):150, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/150.1.full.pdf+html.

Mottl:1979:DPNa

[Mot79a] J. Mottl. Description of a program for nonlinear programming. The Computer Journal, 22(3):256–261, August 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/256.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/256.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/257.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/258.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/259.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/260.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/261.tif.

Mottl:1979:DPNb

[Mot79b] J. Mottl. Description of a program for nonlinear programming: special problem of bivalent nonlinear programming. The Computer Journal, 23(4):360–368, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/4/360.full.pdf+html.

Moylan:1978:BR

[Moy78] P. J. Moylan. Book review. The Computer Journal, 21(3):242, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/242.2.full.pdf+html.

Mavaddat:1979:DSF

[MP79] F. Mavaddat and B. Parhami. A data structure for family relations. The Computer Journal, 22(2):110–113, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/110.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/110.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/111.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/112.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/113.tif.

McBride:1976:MOW

[MR76a] W. E. McBride and A. K. Rigler. A modification of the Osborne and Watson algorithm for nonlinear minimax approximation. The Computer Journal, 19(1):79–81, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
19/1/79.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/79.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/80.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/81.tif.

Mohn:1976:IPA

[MR76b] Karl-Heinz Mohn and Roger V. Roman. An interactive polynomial approximation algorithm. The Computer Journal, 19(1):74– 78, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/1/74.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/74.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/75.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/76.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/77.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/78.tif.

Mamrak:1977:ASE

[MR77] S. Mamrak and J. M. Randal. An analysis of a software engineering failure. The Computer Journal, 20(4):316–320, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/316.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/316.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/317.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/319.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/320.tif.

Missala:1982:CJS

[MR82] Marek Missala and Piotr Rudnicki. Correspondence: Jumping to some purpose. The Computer Journal, 25(2):286, May 1982. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_25/Issue_02/tiff/286.tif. See [ASG79, Hil80, Rob80,

Tay80].

Millard:1975:SES

[MRW75] G. E. Millard, D. J. Rees, and H. Whitfield. The standard EMAS subsystem. The Computer Journal, 18(3):213–219, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/213.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/215.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/216.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/217.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/218.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/219.tif.

Murtagh:1970:CEQ

[MS70] B. A. Murtagh and R. W. H. Sargent. Computational experience with quadratically convergent minimisation methods. The Computer Journal, 13(2):185–194, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/2/185.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/
130185.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/185.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/186.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/188.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/189.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/190.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/191.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_02/tiff/192.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
02/tiff/193.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/tiff/194.tif.

Mullin:1974:OBF

[MS74] J. K. Mullin and G. F. Stuart. Optimum blocking factors for a mix of sequentially processed jobs. The Computer Journal, 17(3): 224–228, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/3/224.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/224.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/225.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/226.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/227.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/228.tif.

McLain:1970:MMA

[MT70] T. G. McLain and A. R. Trice. The MINIMOP multi-access operating system. The Computer Journal, 13(3):237–242, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/237.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130237.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/237.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/238.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/239.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/240.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

13/Issue_03/tiff/241.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/242.tif.

Moon:1972:CEP

[MT72] G. Moon and The Editor. Correspondence: On editorial policy. The Computer Journal, 15(2):124, May 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150124.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/124.tif. See [Wal71b, GT72,

KTW^{+}73].

Mullins:1975:COW

[Mul75] Roy S. Mullins. Correspondence: On the origin of the word ‘algorithm’. The Computer Journal, 18(3):286, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/286.1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/286.tif. See [Mar74d, Lee74].

Muxworthy:1970:NAC

[Mux70] D. T. Muxworthy. Note on Algorithm 47: a clustering algorithm. The Computer Journal, 13(2):219, May 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/

tiff/219.tif. See [Sal70c, vR70a, vR70c].

Muxworthy:1973:CCT

[Mux73] D. T. Muxworthy. Correspondence: On “Compile-Time Error Diagnostics in Syntax-Directed Compilers”. The Computer Journal, 16 (3):285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/285.tif. See [Bur72].

Muxworthy:1974:BR

[Mux74] D. T. Muxworthy. Book review. The Computer Journal, 17(2):123, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/123.1.full.pdf+html.

Muxworthy:1975:C

[Mux75] D. T. Muxworthy. Correspondence. The Computer Journal, 18 (4):381, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/4/381.1.full.pdf+html.

Muxworthy:1976:BR

[Mux76] D. T. Muxworthy. Book review. The Computer Journal, 19(3):212, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/212.1.full.pdf+html.

Muxworthy:1977:BRa

[Mux77a] D. T. Muxworthy. Book review. The Computer Journal, 20(2):123, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/123.3.full.pdf+html.

Muxworthy:1977:BRb

[Mux77b] D. T. Muxworthy. Book review. The Computer Journal, 20(2):131, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/2/131.3.full.pdf+html.

Muxworthy:1978:BR

[Mux78] D. T. Muxworthy. Book review. The Computer Journal, 21(4):354, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/4/354.2.full.pdf+html.

Morris:1970:SPG

[MWC70] D. Morris, I. R. Wilson, and P. C. Capon. A system program generator. The Computer Journal, 13(3):248–254, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/248.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130248.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/248.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/249.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/250.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/251.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/252.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/253.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/254.
tif.

Nash:1975:OST

[Nas75] J. C. Nash. A one-sided transformation method for the singular value decomposition and algebraic eigenproblem. The Computer Journal, 18(1):74–76, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/74.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/74.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/75.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/76.tif. Cited

in Åke Björck’s bibliography on least squares, which is available by

anonymous ftp from math.liu.se in pub/references.

Naur:1979:CID

[Nau79a] P. Naur. Critique of the ISO data processing vocabulary. The Computer Journal, 22(3):286, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/286.tif.

Naur:1979:DCC

[Nau79b] P. Naur. Discussion and correspondence: critique of the ISO data processing vocabulary. The Computer Journal, 22(3):286–288, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/286.full.pdf+html.

Ng:1976:RTI

[NC76] T. S. Ng and A. Cantoni. Run time interaction with Fortran using mixed code. The Computer Journal, 19(1):91–92, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/91.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/91.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/92.
tif.

Neal:1975:C

[Nea75] L. R. Neal. Correspondence. The Computer Journal, 18(1):93, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/93.2.full.pdf+html.

Needham:1974:BR

[Nee74] R. Needham. Book review. The Computer Journal, 17(2):191, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/191.1.full.pdf+html.

Nelson:1979:CSC

[Nel79] Sallie S. Nelson. CONSIM: a study of control issues in conversational simulation. The Computer Journal, 22(2):119–126, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/119.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/119.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/120.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/121.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/122.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/123.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/124.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/125.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/126.tif.

Napper:1976:AUE

[NF76] R. B. E. Napper and R. N. Fisher. ALEC — a user extensible scientific programming language. The Computer Journal, 19(1):25– 31, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/1/25.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/25.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/26.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/27.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/28.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/29.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/30.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/31.tif.

Nievergelt:1970:BST

[NI70] J. Nievergelt and M. I. Irland. Bounce-and-skip. A technique for directing the flow of control in programs. The Computer Journal, 13(3): 261–262, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/3/261.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_03/130261.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/261.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/262.tif. See note [Osb71b].

Needham:1974:BRB

[NLW74] R. Needham, F. F. Land, and M. Wells. Book reviews: Computer Data Security, by H. Katzan, Jr., 1974; Information Systems in Management, by R. J. Radford, 1973; Data Transmission, by M. D. Bacon and G. M. Bull, 1973. The Computer Journal, 17 (2):191, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_02/tiff/191.tif.

Norman:1976:ESI

[Nor76] A. C. Norman. Expanding the solutions of implicit sets of ordinary differential equations in power series. The Computer Journal, 19(1): 63–68, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/1/63.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/63.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/64.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/65.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/67.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/68.tif.

Norris:1978:DCS

[Nor78] G. Norris. Discussion and correspondence summary of the report: ‘the computer board: policies past, present and future?’. The Computer Journal, 21(1):90–91, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/1/90.full.pdf+html.

Notley:1970:CRL

[Not70] M. G. Notley. The cumulative recurrence library. The Computer Journal, 13(1):14–19, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/14.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130014.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/14.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/15.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/16.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/19.tif.

Nessett:1979:AAD

[NR79] D. M. Nessett and O. W. Rechard. Allocation algorithms for dynamically microprogrammable multiprocessor systems. The Computer Journal, 22(2):155–163, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/2/155.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/155.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/156.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/157.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/158.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/159.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/160.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/161.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/163.tif.

Nicklas:1977:ISI

[NS77] B. M. Nicklas and G. Schlageter. Index structuring in inverted data bases by TRIES. The Computer Journal, 20(4):321–324, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/321.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/321.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/

tiff/322.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/323.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/324.tif.

Nevalainen:1979:PQA

[NT79] O. Nevalainen and J. Teuhola. Priority queue administration by sublist index. The Computer Journal, 22(3):220–225, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/220.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/220.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/221.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/222.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/223.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/224.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/225.tif.

Nudds:1977:DMM

[Nud77] D. Nudds. The design of the MAX macroprocessor. The Computer Journal, 20(1):30–36, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/1/30.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/30.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/31.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/32.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/33.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_01/tiff/34.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_01/tiff/35.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
01/tiff/36.tif.

Nutt:1978:BR

[Nut78] W. Owen Nutt. Book review. The Computer Journal, 21(1):6, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/6.full.pdf+html.

Nevalainen:1977:DBF

[NV77] O. Nevalainen and M. Vesterinen. Determining blocking factors for sequential files by heuristic methods. The Computer Journal, 20(3):

245–247, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/245.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/245.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/246.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/247.tif.

Needham:1974:DPM

[NW74] R. M. Needham and M. V. Wilkes. Domains of protection and the management of processes. The Computer Journal, 17(2):117– 120, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/117.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/117.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/120.tif.

OBeirne:1970:CPW

[O’B70] T. H. O’Beirne. Computer programs which play music with microtones. The Computer Journal, 13(4):344–349, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/344.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130344.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/344.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/345.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/346.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/347.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/348.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/349.tif.

Osborn:1975:SPS

[OBLP75] R. A. Osborn, W. P. Bain, T. Lloyd, and P. H. Perring. SPG — a programming system for commercial transaction processing. The Computer Journal, 18(4):290–297, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/290.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_18/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/293.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/294.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/295.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/296.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/297.tif.

Obruca:1978:AM

[Obr78a] A. Obruča. Algorithm 1: Mintree. The Computer Journal, 21(3): 276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_03/tiff/276.tif. Reprinted from

The Computer Bulletin, September 1964, p. 67. See notes [Boo78, HO78, Obr78b, Pit78, RT78, WT79].

Obruca:1978:RAM

[Obr78b] Alex K. Obruča. Remark on Algorithm 1: Mintree. The Computer Journal, 21(3):277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_21/Issue_03/tiff/277.tif. See

[Obr78a].

OConnell:1979:BR

[O’C79] S. O’Connell. Book review. The Computer Journal, 22(3):219, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/219.1.full.pdf+html.

Osborne:1978:ISP

[OL78] M. Osborne and F. Little. An implementation of structured programming in APL. The Computer Journal, 21(4):311–315, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/311.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/311.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/312.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/313.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_21/Issue_04/tiff/314.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/315.tif.

Oliver:1971:EDI

[Oli71] J. Oliver. The evaluation of definite integrals using high-order formulae. The Computer Journal, 14(3):301–306, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/301.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140301.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/301.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/302.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/303.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/304.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/305.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/306.tif.

Oliver:1972:DAC

[Oli72] J. Oliver. A doubly-adaptive Clenshaw–Curtis quadrature method. The Computer Journal, 15(2):141–147, May 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/2/141.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/150141.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/141.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/142.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/143.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/144.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/147.tif.

Olle:1975:BR

[Oll75] T. W. Olle. Book review. The Computer Journal, 18(3):251, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/251.full.pdf+html.

Oliver:1979:PEA

[ON79] J. R. Oliver and R. S. Newton. Practical experience with Algol 68-Rt. The Computer Journal, 22(2):114–118, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/114.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/114.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/115.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/116.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/117.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/118.tif.

Oliveira-Pinto:1971:ATC

[OP71] F. Oliveira-Pinto. Algorithm 69: Trigonometric curve fitting to equally or unequally spaced data. The Computer Journal, 14(2):213–214, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_02/tiff/213.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/214.tif.

Oliveira-Pinto:1973:GCP

[OP73a] F. Oliveira-Pinto. Generalised Chebyshev polynomials and their use in numerical approximation. The Computer Journal, 16(4):375– 379, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/375.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160375.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/375.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/376.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/377.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/378.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/379.tif.

Oliveira-Pinto:1973:STA

[OP73b] F. Oliveira-Pinto. Simultaneous trigonometric approximation of the function and its first derivative. The Computer Journal, 16(1):73– 76, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/73.full.pdf+html;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_16/Issue_01/160073.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/73.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/74.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/75.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/76.tif.

Oliveira-Pinto:1975:UDO

[OP75] F. Oliveira-Pinto. The use of derivatives in the optimisation of the truncation error of polynomial approximations. The Computer Journal, 18(1):77–80, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/77.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/77.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/78.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/79.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/80.tif.

ORegan:1970:SSA

[O’R70] P. G. O’Regan. Step size adjustment at discontinuities for fourth order Runge–Kutta methods. The Computer Journal, 13(4):401– 404, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/401.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/130401.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/401.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/402.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/403.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_

04/tiff/404.tif. See correction [Whi72b].

Ortony:1971:SSV

[Ort71] A. Ortony. A system for stereo viewing. The Computer Journal, 14 (2):140–144, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/140.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140140.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/140.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_02/tiff/141.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/142.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/144.tif.

Ord-Smith:1970:GPS

[OS70] R. J. Ord-Smith. Generation of permutation sequences: part 1. The Computer Journal, 13(2):152–155, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/152.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130152.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/152.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/153.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/154.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/155.
tif.

Ord-Smith:1971:BR

[OS71a] R. J. Ord-Smith. Book review. The Computer Journal, 14(2):139, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/2/139.full.pdf+html.

Ord-Smith:1971:GPS

[OS71b] R. J. Ord-Smith. Generation of permutation sequences: Part 2. The Computer Journal, 14(2):136–139, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/136.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140136.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/136.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/137.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/138.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/139.
tif.

Ord-Smith:1975:BR

[OS75] R. J. Ord-Smith. Book review. The Computer Journal, 18(2):130, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/130.3.full.pdf+html.

Osborne:1971:C

[Osb71a] G. W. Osborne. Correspondence. The Computer Journal, 14(2):155, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/2/155.2.full.pdf+html.

Osborne:1971:CBS

[Osb71b] G. W. Osborne. Correspondence: On “Bounce-and-skip. A technique for directing the flow of control in programs”. The Computer Journal, 14(2):155, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/155.tif. See [NI70].

Osterby:1979:AES

[Øst79] Ole Østerby. Algorithm 108: Efficient solution of tridiagonal linear systems. The Computer Journal, 22(3):283–285, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_03/tiff/283.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_03/tiff/284.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/tiff/285.
tif.

Overholt:1967:NAF

[Ove67] K. J. Overholt. Note on Algorithm 2: Fibonacci search. The Computer Journal, 9(4):414, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_04/
tiff/414.tif. See [PP79].

Osborne:1971:ADN

[OW71] M. R. Osborne and G. A. Watson. On an algorithm for discrete nonlinear L_{1} approximation. The Computer Journal, 14(2): 184–188, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/2/184.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_02/140184.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/184.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/185.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_14/Issue_02/tiff/186.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/187.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/188.tif.  See correspondence

[DW73].

Pacifico:1977:BR

[Pac77] R. E. Pacifico. Book review. The Computer Journal, 20(2):101, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
2/101.2.full.pdf+html.

Padgett:1973:TDD

[Pad73] M. I. Padgett. Tree driven data input and its validation. The Computer Journal, 16(4):315–321, November 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/4/315.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/160315.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/317.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/318.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/319.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/320.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/321.
tif.

Page:1971:SGO

[Pag71] E. S. Page. Systematic generation of ordered sequences using recurrence relations. The Computer Journal, 14(2):150–153, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/150.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140150.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/150.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/151.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/152.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/153.
tif.

Page:1973:EJW

[Pag73] E. S. Page. Editorial: the journal — why and what for. The Computer Journal, 16(4):290, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/4/290.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/
160290.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/290.tif.

Page:1974:CMD

[Pag74] E. S. Page. Concepts and methods in discrete event digital simulation. The Computer Journal, 17(4):343, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/4/343.full.pdf+html.

Pagan:1976:IOD

[Pag76a] Frank G. Pagan. On interpreter-oriented definitions of programming languages. The Computer Journal, 19(2):151–155, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/2/151.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/151.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/152.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
02/tiff/153.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_02/tiff/154.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_02/tiff/155.tif.

Page:1976:DSC

[Pag76b] E. S. Page. Distribution sampling for computer simulation. The Computer Journal, 19(2):109, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/109.full.pdf+html.

Pagan:1979:AMD

[Pag79a] F. G. Pagan. Algol 68 as a metalanguage for denotational semantics. The Computer Journal, 22(1):63–66, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
22/1/63.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/63.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/64.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/65.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/66.tif.

Page:1979:BR

[Pag79b] I. P. Page. Book review. The Computer Journal, 22(4):345, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/345.1.full.pdf+html.

Paine:1978:BR

[Pai78] R. M. Paine. Book review. The Computer Journal, 21(3):209, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/209.1.full.pdf+html.

Paker:1973:BR

[Pak73] W. Paker. Book review. The Computer Journal, 16(2):123, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/123.1.full.pdf+html.

Palme:1972:C

[Pal72a] J. Palme. Correspondence. The Computer Journal, 15(1):4, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/1/4.full.pdf+html.

Palme:1972:CHL

[Pal72b] J. Palme. Correspondence: On high level languages. The Computer Journal, 15(1):4–4, 36–36, February 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/36.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/4.tif. See correspondence [Bra72,

Hal72b].

Palme:1973:CSN

[Pal73] J. Palme. Correspondence: SIMPL/I — a new simulation language from IBM. The Computer Journal, 16(2):189, May 1973. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/189.1.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/189.tif.

Pankhurst:1970:CPG

[Pan70] R. J. Pankhurst. A computer program for generating diagnostic keys. The Computer Journal, 13(2):145–151, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/145.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130145.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/145.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/146.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/147.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/150.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/151.
tif.

Parker:1974:NAC

[Par74] Frank Parker. Note on Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 17(4):380, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/

tiff/380.tif. See [Van73, HW74].

Parish:1975:AQT

[Par75a] C. C. M. Parish. On the application of queueing theory to analysing on-line computing systems. The Computer Journal, 18(2):117– 121, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/117.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/117.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/118.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/119.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/120.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/121.tif.

Parkin:1975:BR

[Par75b] A. Parkin. Book review. The Computer Journal, 18(1):33, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/33.1.full.pdf+html.

Parkin:1975:BCF

[Par75c] A. Parkin. Bringing cost into file design decisions. The Computer Journal, 18(3):198–199, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/3/198.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/198.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/199.tif.

Parkin:1975:POA

[Par75d] A. Parkin. Programming objectives: an analysis. The Computer Journal, 18(4):307–311, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/307.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/307.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/309.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/310.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/311.tif.

Partridge:1975:DDW

[Par75e] D. Partridge. A dynamic database which automatically removes unwanted generalisation for the efficient analysis of language features that exhibit a disparate frequency distribution. The Computer Journal, 18(1):43–48, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/43.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/43.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/44.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/45.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/46.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/47.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/48.tif.

Parkin:1976:C

[Par76] A. Parkin. Correspondence. The Computer Journal, 19(1):94–95, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/94.2.full.pdf+html.

Parkin:1977:C

[Par77a] A. Parkin. Correspondence. The Computer Journal, 20(4):381, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/381.full.pdf+html.

Parkin:1977:POD

[Par77b] A. Parkin. The probable outcomes of a data processing project. The Computer Journal, 20(2):98–101, May 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/2/98.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/100.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/101.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/98.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/99.tif.

Parkin:1977:BR

[Par77c] J. E. Parkin. Book review. The Computer Journal, 20(4):358, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/358.1.full.pdf+html.

Parkin:1978:CMC

[Par78] A. Parkin. Critique on multi-criteria assessment applied to alternative data processing systems. The Computer Journal, 21(2): 188–190, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/188.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/188.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/189.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/190.tif.

Park:1979:BR

[Par79] D. Park. Book review. The Computer Journal, 22(3):233, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/233.1.
full.pdf+html.

Paul:1979:BR

[Pau79] R. J. Paul. Book review. The Computer Journal, 22(1):62, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/62.2.full.pdf+html.

Presser:1974:ECD

[PB74] L. Presser and J. Benson. Evaluation of compiler diagnostics. The Computer Journal, 17(2):121–123, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/121.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/121.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/122.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/123.tif.

Patel:1976:ASB

[PB76] M. R. Patel and K. H. Bennett. Analysis of speed of a binary multiplier using a variable number of shifts per cycle. The Computer Journal, 19(3):254–257, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/254.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/254.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/255.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/256.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/257.tif.

Patel:1978:ASB

[PB78] M. R. Patel and K. H. Bennett. Analysis of speed of a binary divider using a variable number of shifts per cycle. The Computer Journal, 21(3):246–252, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/246.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/246.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/247.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/248.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/249.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/250.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/252.tif.

Parkin:1978:EEB

[PC78] A. Parkin and R. B. Coats. EDSIM-event based discrete simulation using general purpose languages such as Fortran. The Computer Journal, 21(2):122–127, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/122.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/122.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/123.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/124.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/127.tif.

Price:1979:AWS

[PD79] W. L. Price and M. Dowson. Algorithm 107: a weighted simplex procedure for the solution of simultaneous nonlinear equations. The Computer Journal, 22(3):282, August 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/

tiff/282.tif. This algorithm is published under the Journal’s agreement with the Institute of Mathematics and its Applications, and is associated with the authors’ article in the JIMA, Volume 24, Number 1, August 1979.

Parsons:1974:DML

[PDY74] R. G. Parsons, A. G. Dale, and C. V. Yurkanan. Data manipulation language requirements for data base management systems. The Computer Journal, 17(2):99–103, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/

17/2/99.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_02/tiff/100.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/tiff/101.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
17/Issue_02/tiff/102.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_02/tiff/103.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/tiff/99.
tif.

Peckham:1970:NMM

[Pec70] G. Peckham. A new method for minimising a sum of squares without calculating gradients. The Computer Journal, 13(4):418– 420, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/418.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/130418.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/418.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/419.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/420.tif.

Pengilly:1976:ASD

[Pen76] P. J. Pengilly. An approach to systems design. The Computer Journal, 19(1):8–12, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/1/8.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_01/tiff/12.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/8.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/9.tif.

Pengilly:1977:ASD

[Pen77] P. J. Pengilly. An approach to systems design. II. the popularisation of PROPLAN. The Computer Journal, 20(4):296–297, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/296.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/296.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/297.tif.

Perlis:1976:ACS

[Per76] Donald Perlis. An application of compiler simulation at the source language level. The Computer Journal, 19(1):90, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/90.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/90.tif.

Pezarro:1976:NEH

[Pez76] M. T. Pezarro. A note on estimating hit ratios for direct-access storage devices. The Computer Journal, 19(3):271–272, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/271.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/271.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/272.tif.

Piper:1974:ARP

[PFW74] L. R. Piper, I. R. Franklin, and N. H. Westwood. Algorithm 83: a rejection procedure for sampling from the distribution of chi-squared (χ^{2}) or the sampling variance (s^{2}). The Computer Journal, 17(4): 378, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_17/Issue_04/tiff/378.tif.

Pawson:1973:CHT

[PH73] A. J. D. Pawson and F. R. A. Hopgood. Correspondence: Hashing techniques for table searching. The Computer Journal, 16(3): 285, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_03/tiff/285.tif. See [HD72].

Phillips:1970:ANI

[Phi70a] G. M. Phillips. Analysis of numerical iterative methods for solving integral and integrodifferential equations. The Computer Journal, 13(3):297–300, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/3/297.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/
130297.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/297.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_03/tiff/298.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
03/tiff/299.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/tiff/300.tif.

Phillips:1970:BR

[Phi70b] G. M. Phillips. Book review. The Computer Journal, 13(1):24, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/24.full.pdf+html.

Phillips:1974:PIF

[Phi74a] D. A. Phillips. A preliminary investigation of function optimisation by a combination of methods. The Computer Journal, 17(1):75– 79, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/17/1/75.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/75.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/76.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/77.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/78.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/79.tif.

Phillips:1974:TNA

[Phi74b] G. M. Phillips. Topics in numerical analysis. The Computer Journal, 17(4):370, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/370.full.pdf+html.

Phillips:1976:C

[Phi76] B. P. Phillips. Correspondence. The Computer Journal, 19(1):93, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/93.1.full.pdf+html.

Pike:1967:NAF

[PHJ67] M. C. Pike, I. D. Hill, and F. D. James. Note on Algorithm 2: Fibonacci search, and on Algorithm 7: MINX. The Computer Journal, 9(4):416–417, February 1967. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_09/Issue_04/tiff/416.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_09/Issue_
04/tiff/417.tif. See [PP79].

Piessens:1971:SAG

[Pie71] R. Piessens. Some aspects of Gaussian quadrature formulae for the numerical inversion of the Laplace transform. The Computer Journal, 14(4):433–436, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/4/433.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140433.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/433.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/434.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/435.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/436.tif.

Pierce:1974:SLD

[Pie74] R. H. Pierce. Source language debugging on a small computer. The Computer Journal, 17(4):313–317, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/313.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/313.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/314.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/315.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/316.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/317.tif.

Pike:1979:AR

[Pik79] M. C. Pike. Algorithm 5: RANCOMB. The Computer Journal, 22 (1):88, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/88.tif. Reprinted from

The Computer Bulletin, September 1964, p. 67.

Piney:1977:GIS

[Pin77] C. Piney. A generalised interactive system for the acquisition of structured data. The Computer Journal, 20(1):37–44, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/37.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_20/Issue_01/tiff/37.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/41.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/43.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/44.
tif.

Pitcher:1974:CDR

[Pit74] H. H. W. Pitcher. Correspondence: Definition of random. The Computer Journal, 17(4):381, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/381.tif. See [SL73].

Pitman:1978:NAM

[Pit78] Anthony E. N. T. Pitman. Note on Algorithm 1: Mintree. The Computer Journal, 21(3):276, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/

tiff/276.tif. See [Obr78a].

Pai:1974:ASL

[PL74] D. Pai and D. Lewin. Analysis of sequential logic circuits. The Computer Journal, 17(1):64–68, February 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/64.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/64.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/65.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/66.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/67.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/68.tif.

Powell:1972:RCI

[PM72] D. R. Powell and J. R. Macdonald. A rapidly convergent iterative method for the solution of the generalised nonlinear least squares problem. The Computer Journal, 15(2):148–155, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/148.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150148.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/148.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/149.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/150.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/151.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/152.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/153.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/154.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/155.tif.

Potter:1978:BR

[Pot78] C. Potter. Book review. The Computer Journal, 21(3):233, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/233.3.full.pdf+html.

Powell:1973:BR

[Pow73] M. J. D. Powell. Book review. The Computer Journal, 16(1):45, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/1/45.full.pdf+html.

Powell:1974:BR

[Pow74] M. J. D. Powell. Book review. The Computer Journal, 17(1):92, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/1/92.1.full.pdf+html.

Powell:1976:BR

[Pow76] M. J. D. Powell. Book review. The Computer Journal, 19(3):245, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/245.2.full.pdf+html.

Pike:1979:AFS

[PP79] M. C. Pike and J. Pixner. Algorithm 2: Fibonacci search. The Computer Journal, 22(1):86, February 1979. CODEN CMPJA6.

ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/

tiff/86.tif. Reprinted from The Computer Bulletin, vol. 8, p. 147, 1965. See [Boo79c, Ove67, PHJ67].

Pierce:1977:TDC

[PR77] R. H. Pierce and J. Rowell. A transformation-directed compiling system. The Computer Journal, 20(2):109–115, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/109.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/109.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/110.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/111.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/112.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/113.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/114.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/115.tif.

Price:1970:BR

[Pri70] V. E. Price. Book review. The Computer Journal, 13(3):288, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
3/288.full.pdf+html.

Price:1977:BR

[Pri77a] V. E. Price. Book review. The Computer Journal, 20(1):55, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/55.3.full.pdf+html.

Price:1977:CRS

[Pri77b] W. L. Price. A controlled random search procedure for global optimisation. The Computer Journal, 20(4):367–370, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/367.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/367.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/368.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_

04/tiff/369.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/370.tif.

Price:1978:SSD

[Pri78] W. L. Price. Simulation studies of data communication networks operating in datagram mode. The Computer Journal, 21(3):219– 223, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/219.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/219.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/220.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/221.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/222.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/223.tif.

Proll:1974:BR

[Pro74a] L. G. Proll. Book review. The Computer Journal, 17(1):8, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/8.1.full.pdf+html.

Proskurowski:1974:OMA

[Pro74b] W. Proskurowski. Ordering method to accelerate the solution of mate-in-two chess problems by computer. The Computer Journal, 17(1):80–81, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/80.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/80.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/81.tif.

Poole:1974:CSO

[PS74] F. Poole and P. M. Stocker. Correspondence: On “Self-organising data management systems”. The Computer Journal, 17(1):95, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/95.tif. See [SD73].

Peters:1970:LSP

[PW70] G. Peters and J. H. Wilkinson. The least squares problem and pseudoinverses. The Computer Journal, 13(3):309–316, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/309.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130309.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/309.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/310.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/311.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/312.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/313.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/314.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/315.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/316.tif.

Page:1974:IAS

[PW74a] E. S. Page and L. Waller. An interactive approach for the solution of a class of discrete optimisation problems. The Computer Journal, 17(1):69–74, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/69.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/69.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/70.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/71.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/72.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/73.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/74.tif.

Proll:1974:COD

[PW74b] L. G. Proll and F. P. Wyman. Correspondence: Observations on a decision rule for binary programming. The Computer Journal, 17(2): 187–188, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/17/2/187.2.full.pdf+html;   http://www3.oup.
co.uk/computer_journal/hdb/Volume_17/Issue_02/tiff/187.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/188.tif. See [Wym73].

Quilliam:1979:BR

[Qui79] H. Quilliam. Book review. The Computer Journal, 23(4):316, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/316.1.full.pdf+html.

Radford:1973:BR

[Rad73] A. S. Radford. Book review. The Computer Journal, 16(2):134, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/134.1.full.pdf+html.

Radford:1975:BRa

[Rad75a] A. S. Radford. Book review. The Computer Journal, 18(2):111, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/111.1.full.pdf+html.

Radford:1975:BRb

[Rad75b] A. S. Radford. Book review. The Computer Journal, 18(2):111, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/111.2.full.pdf+html.

Radford:1975:BRc

[Rad75c] A. S. Radford. Book review. The Computer Journal, 18(2):111, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/111.3.full.pdf+html.

Radford:1975:BRd

[Rad75d] A. S. Radford. Book review. The Computer Journal, 18(2):111, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/111.4.full.pdf+html.

Radford:1976:BR

[Rad76] A. S. Radford. Book review. The Computer Journal, 19(3):228, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/228.2.full.pdf+html.

Radford:1979:BR

[Rad79] A. S. Radford. Book review. The Computer Journal, 22(4):345, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/345.2.full.pdf+html.

Ramsay:1970:FGM

[Ram70] J. O. Ramsay. A family of gradient methods for optimization. The Computer Journal, 13(4):413–417, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/413.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130413.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/413.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/414.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/415.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/416.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/417.tif.

Randell:1971:BR

[Ran71a] B. Randell. Book review. The Computer Journal, 14(1):90, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/90.full.pdf+html.

Randell:1971:LAM

[Ran71b] B. Randell. Ludgate’s analytical machine of 1909. The Computer Journal, 14(3):317–326, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/317.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140317.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/317.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/318.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/319.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/320.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/321.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/323.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/324.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/325.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_14/Issue_03/tiff/326.tif.  See correspondence

[WR72].

Randall:1974:IDS

[Ran74] F. E. Randall. Interrogating date-sensitive files. The Computer Journal, 17(4):302–305, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/302.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/304.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/305.tif.

Randle:1979:CAE

[Ran79] J. Randle. Certification of Algorithm 3: Easter. The Computer Journal, 22(1):87, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_22/Issue_01/tiff/87.tif. See

[Hil79c].

Raymont:1979:BR

[Ray79] P. G. Raymont. Book review. The Computer Journal, 22(4):331, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/4/331.2.full.pdf+html.

Rohl:1970:LPF

[RC70] J. S. Rohl and G. Cordingley. List processing facilities in Atlas Autocode. The Computer Journal, 13(1):20–24, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/1/20.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130020.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/20.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/21.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/22.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/23.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/24.tif.

Reader:1970:RDM

[Rea70] A. V. Reader. Random digits by a mincing process. The Computer Journal, 13(1):118, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/1/118.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
130118.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/118.tif.

Redish:1971:C

[Red71a] K. A. Redish. Correspondence. The Computer Journal, 14(1):64, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/64.full.pdf+html.

Redish:1971:CNA

[Red71b] K. A. Redish. Correspondence: (note on) Algorithm 44. The Computer Journal, 14(1):64, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
140064.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/64.tif. See [Bro69, Fie70, HL70b].

Reeves:1973:BR

[Ree73] C. M. Reeves. Book review. The Computer Journal, 16(3):204, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/204.2.full.pdf+html.

Reeves:1974:BR

[Ree74] C. M. Reeves. Book review. The Computer Journal, 17(2):164, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/164.2.full.pdf+html.

Rees:1975:ED

[Ree75] D. J. Rees. The EMAS director. The Computer Journal, 18(2): 122–130, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/122.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/122.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/

tiff/123.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/124.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/125.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/126.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/127.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/128.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/129.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/130.tif.

Reeves:1976:BR

[Ree76a] C. M. Reeves. Book review. The Computer Journal, 19(4):332, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/332.2.full.pdf+html.

Reeves:1976:BRB

[Ree76b] C. M. Reeves. Book review: Computer Science: Programming in FORTRAN with WATFOR WATFIV. The Computer Journal, 19 (2):172a–172, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/172-a.full.pdf+html.

Reeves:1979:AS

[Ree79a] C. M. Reeves. Algorithm 8: SUD. The Computer Journal, 22(1): 89, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/89.tif. Reprinted from

The Computer Bulletin, September 1964, p. 67.

Reeves:1979:AB

[Ree79b] C. M. Reeves. Algorithm 9: Boyserf. The Computer Journal, 22(1): 89–90, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/90.

tif. Reprinted from The Computer Bulletin, September 1964, p. 67.

Reeves:1979:FSDa

[Ree79c] C. M. Reeves. Free store distribution under random fit allocation: Part 1. The Computer Journal, 22(4):346–351, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/346.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/346.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/347.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/348.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/349.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/350.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/351.tif.

Reeves:1979:FSDb

[Ree79d] C. M. Reeves. Free store distribution under random fit allocation: part 2. The Computer Journal, 23(4):298–306, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/298.full.pdf+html.

Reid:1971:BR

[Rei71a] J. K. Reid. Book review. The Computer Journal, 14(1):44, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/44.full.pdf+html.

Reid:1971:NAP

[Rei71b] J. K. Reid. A note on the approximation of plane regions. The Computer Journal, 14(3):307–308, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/307.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140307.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/307.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/308.tif.

Reid:1973:BR

[Rei73] J. K. Reid. Book review. The Computer Journal, 16(3):222, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/222.1.full.pdf+html.

Repko:1972:ACD

[Rep72] Marya Goldman Repko. Animated cartoon design with a CRT. The Computer Journal, 15(4):316–320, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/316.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150316.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/316.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/317.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/318.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/319.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/320.tif.

Reynolds:1971:CPI

[Rey71a] C. F. Reynolds. CODIL, part 1. The importance of flexibility. The Computer Journal, 14(3):217–220, August 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/3/217.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/140217.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/217.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_03/tiff/218.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_03/tiff/219.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/tiff/220.
tif.

Reynolds:1971:CPC

[Rey71b] C. F. Reynolds. CODIL, part 2. The CODIL language and its interpreter. The Computer Journal, 14(4):327–332, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/327.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140327.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/327.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/328.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/329.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/330.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/331.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/332.tif.

Reynolds:1977:CSH

[Rey77] R. A. Reynolds. Character string handling in Fortran. The Computer Journal, 20(4):325–329, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/4/325.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/325.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/326.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/327.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/328.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/329.tif.

Ricci:1971:ARH

[Ric71] A. Ricci. An algorithm for the removal of hidden lines in 3D scenes. The Computer Journal, 14(4):375–377, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/375.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140375.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/375.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/376.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/377.tif.

Ricci:1973:CGC

[Ric73] A. Ricci. A constructive geometry for computer graphics. The Computer Journal, 16(2):157–160, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/157.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160157.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/157.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/158.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/159.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/160.tif.

Richards:1974:CMS

[Ric74] B. Richards. Computers in medicine series. The Computer Journal, 17(4):305, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/305.full.pdf+html.

Richards:1976:MIM

[Ric76] B. Richards. Methoden der informatik in der medizin. The Computer Journal, 19(2):165, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/165.full.pdf+html.

Richards:1978:BR

[Ric78] B. Richards. Book review. The Computer Journal, 21(3):269, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/269.full.pdf+html.

Riley:1977:BR

[Ril77] N. Riley. Book review. The Computer Journal, 20(1):77, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/77.1.full.pdf+html.

Riley:1978:BR

[Ril78] N. Riley. Book review. The Computer Journal, 21(3):209, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/209.2.
full.pdf+html.

Riley:1979:BR

[Ril79] N. Riley. Book review. The Computer Journal, 22(1):16, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/16.4.full.pdf+html.

Ringland:1975:SFV

[Rin75] Gill Ringland. Software functional variability. The Computer Journal, 18(4):312–317, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/312.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/312.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/313.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/314.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_04/
tiff/315.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_04/tiff/316.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_04/tiff/317.tif.

Rivett:1979:BR

[Riv79] P. Rivett. Book review. The Computer Journal, 22(2):109, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/109.1.
full.pdf+html.

Rohl:1972:NCA

[RL72] J. S. Rohl and J. A. Linn. A note on compiling arithmetic expressions. The Computer Journal, 15(1):13–14, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
15/1/13.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/150013.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
tiff/13.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_15/Issue_01/tiff/14.tif.

Reid:1973:BRB

[RL73] J. K. Reid and D. W. Lewin. Book reviews: Sparse Matrices and their Applications, edited by D. L. Rose and R. A. Willoughby, 1972; Introduction to Digital Computer Technology, by Louis Nashelsky, 1973. The Computer Journal, 16(3):222, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/222.tif.

Rokne:1975:ACI

[RL75] J. Rokne and P. Lancaster. Algorithm 86: Complex interval arithmetic. The Computer Journal, 18(1):83–85, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_01/tiff/84.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/85.tif.

Romao:1978:AMN

[RM78] J. C. Romão and R. Vilela Mendes. Algorithms for multidimensional numerical integration with singularities. The Computer Journal, 21(4):355–358, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/4/355.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/

tiff/355.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/356.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/357.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/358.tif.

Robinson:1975:DIB

[Rob75] K. A. Robinson. Database — the ideas behind the ideas. The Computer Journal, 18(1):7–11, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/1/7.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_01/tiff/7.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/
tiff/8.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_01/tiff/9.tif.

Robson:1977:WCF

[Rob77] J. M. Robson. Worst case fragmentation of first fit and best fit storage allocation strategies. The Computer Journal, 20(3):242– 244, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/242.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/242.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/243.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/244.tif.

Robinson:1979:EED

[Rob79] K. A. Robinson. An entity/event data modelling method. The Computer Journal, 22(3):270–279, August 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/3/270.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/270.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/271.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/272.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/273.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/274.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/275.tif;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_22/Issue_03/tiff/276.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_03/tiff/277.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
03/tiff/278.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_03/tiff/279.tif. See discussion [Ano79l].

Robinson:1980:CJS

[Rob80] G. L. Robinson. Correspondence: Jumping to some purpose. The Computer Journal, 23(3):288, August 1980. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_03/

tiff/288.tif. See [ASG79, Hil80, Tay80, MR82].

Rodabaugh:1970:SC

[Rod70] D. J. Rodabaugh. On stable correctors. The Computer Journal, 13(1):98–100, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/13/1/98.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
130098.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/tiff/100.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_01/tiff/98.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/99.tif.

Rogers:1972:CNT

[Rog72] C. A. Rogers. Computers in number theory. The Computer Journal, 15(2):175, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/2/175.full.pdf+html.

Rogers:1974:BR

[Rog74] C. A. Rogers. Book review. The Computer Journal, 17(3):260, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/260.2.full.pdf+html.

Rogers:1975:BR

[Rog75] M. H. Rogers. Book review. The Computer Journal, 18(3):222, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/222.1.full.pdf+html.

Rogers:1976:BR

[Rog76] A. Rogers. Book review. The Computer Journal, 19(3):212, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/212.2.full.pdf+html.

Rogers:1979:BRa

[Rog79a] A. Rogers. Book review. The Computer Journal, 22(1):40, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/40.full.pdf+html.

Rogers:1979:BRb

[Rog79b] M. H. Rogers. Book review. The Computer Journal, 22(3):205, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/205.1.full.pdf+html.

Rohlf:1973:AHC

[Roh73] F. James Rohlf. Algorithm 76: Hierarchical clustering using the minimum spanning tree. The Computer Journal, 16(1):93–95, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/93.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/94.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/95.tif.

Rohlf:1974:ADP

[Roh74] R. James Rohlf. Algorithm 81: Dendrogram plot. The Computer Journal, 17(1):89–91, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/89.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/90.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/91.tif.

Rohlf:1975:NAC

[Roh75] F. J. Rohlf. A new approach to the computation of the Jardine-Sibson B_{k} clusters. The Computer Journal, 18(2):164–168, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/164.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/164.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/165.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/167.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/168.tif.

Rohl:1976:PIF

[Roh76] J. S. Rohl. Programming improvements to Fike’s algorithm for generating permutations. The Computer Journal, 19(2):156–159, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/2/156.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/156.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/158.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/159.tif.

Rohl:1978:GPC

[Roh78] J. S. Rohl. Generating permutations by choosing. The Computer Journal, 21(4):302–305, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/4/302.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/304.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/305.tif.

Rohl:1979:C

[Roh79] J. S. Rohl. Correspondence. The Computer Journal, 22(2):191, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/191.full.pdf+html.

Ross:1970:BSD

[Ros70] H. McGregor Ross. The British Standard Data Code and how to exploit it. The Computer Journal, 13(3):223–229, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/223.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130223.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/223.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/224.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/225.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/226.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/227.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/228.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/229.
tif.

Roy:1978:EPA

[Roy78] M. K. Roy. Evaluation of permutation algorithms. The Computer Journal, 21(4):296–301, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/4/296.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/296.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/297.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/298.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/300.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/301.tif.

Robinson:1976:EAF

[RT76] S. K. Robinson and I. S. Torsun. An empirical analysis of Fortran programs. The Computer Journal, 19(1):56–62, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/56.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/56.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/57.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/58.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/59.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/60.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/61.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/62.tif.

Rodgers:1978:CAM

[RT78] M. H. Rodgers and R. H. Thomason. Certification of Algorithm 1: Mintree. The Computer Journal, 21(3):276–277, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_03/tiff/277.tif. See [Obr78a].

Rozenberg:1974:GMP

[RW74] G. Rozenberg and D. Wood. Generative models for parallel processes. The Computer Journal, 17(4):344–348, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/344.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/344.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/345.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/346.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/347.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/348.tif.

Rzveski:1979:BR

[Rzv79] G. Rzveski. Book review. The Computer Journal, 23(4):316, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/316.2.full.pdf+html.

Schneck:1973:FFO

[SA73] P. B. Schneck and Ellinor Angel. A Fortran to Fortran optimising compiler. The Computer Journal, 16(4):322–330, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/4/322.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/160322.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/322.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/323.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/324.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/325.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/326.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/327.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/328.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/329.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/330.tif.

Sandys:1979:CPP

[SA79] M. A. Sandys and Aleph Null. Correspondence: Point in polygon. The Computer Journal, 22(3):287–288, August 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/287.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/288.tif.

Sabin:1972:CCS

[Sab72] M. A. Sabin. Correspondence: Comments on some algorithms for the representation of curves by straight line segments. The Computer Journal, 15(2):104, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/104.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150104.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/104.

tif. See [Ecc71b, McC70b, Smi71, Coo72f].

Sabin:1976:C

[Sab76a] M. A. Sabin. Correspondence. The Computer Journal, 19(1):93–94, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/93.3.full.pdf+html.

Sabin:1976:MDI

[Sab76b] M. A. Sabin. A method for displaying the intersection curve of two quadric surfaces. The Computer Journal, 19(4):336–338, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/336.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/336.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/337.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/338.tif.

Scowen:1971:SPW

[SAHS71] R. S. Scowen, D. Allin, A. L. Hillman, and M. Shimell. SOAP — a program which documents and edits ALGOL 60 programs. The Computer Journal, 14(2):133–135, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/133.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140133.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/133.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/134.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/135.tif.

Sale:1970:AAS

[Sal70a] A. H. J. Sale. Algorithm 54: Approximation of straight lines. The Computer Journal, 13(4):422–424, November 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/
tiff/422.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_04/tiff/423.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/tiff/424.tif.

Sale:1970:NAI

[Sal70b] A. H. J. Sale. Note on Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 13(1):115, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/115.tif. See [Spä69].

Sale:1970:NAC

[Sal70c] A. H. J. Sale. Note on Algorithm 47: a clustering algorithm. The Computer Journal, 13(3):326, August 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/

tiff/326.tif. See [vR70a, vR70c, Mux70].

Sale:1971:AIC

[Sal71a] A. H. J. Sale. Algorithm 65: An improved clustering algorithm. The Computer Journal, 14(1):104–106, February 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/104.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_14/Issue_01/tiff/105.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/106.tif.

Sale:1971:CFS

[Sal71b] A. H. J. Sale. The classification of Fortran statements. The Computer Journal, 14(1):10–12, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/1/10.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/140010.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/10.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/11.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/12.tif.

Salzer:1972:LIC

[Sal72] H. E. Salzer. Lagrangian interpolation at the Chebyshev points x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n; some unnoted advantages. The Computer Journal, 15(2):156–159, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/156.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150156.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/156.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_02/tiff/157.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
02/tiff/158.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/159.tif.  See correspondence

[Sal73b].

Salzer:1973:C

[Sal73a] H. E. Salzer. Correspondence. The Computer Journal, 16(2):190, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/190.3.full.pdf+html.

Salzer:1973:CLI

[Sal73b] H. E. Salzer. Correspondence: On “Lagrangian interpolation at the Chebyshev points x_{n,ν} ≡ cos(νπ/n), ν = 0(1)n”. The Computer Journal, 16(2):190, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_16/Issue_02/tiff/190.tif. See

[Sal72].

Salomon:1974:EAW

[Sal74] Kenneth B. Salomon. An efficient algorithm which determines the output from a sequential machine for regular inputs. The Computer Journal, 17(4):349–354, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/349.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/349.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/350.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/351.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/352.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/353.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/354.tif.

Samet:1972:IN

[Sam72] P. A. Samet. Insight, not numbers. The Computer Journal, 15(1): 88–92, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/1/88.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/150088.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/88.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/89.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/90.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/91.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/92.tif.

Samet:1973:BRa

[Sam73a] P. A. Samet. Book review. The Computer Journal, 16(2):134, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/134.3.full.pdf+html.

Samet:1973:BRb

[Sam73b] P. A. Samet. Book review. The Computer Journal, 16(3):253, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/253.2.full.pdf+html.

Samet:1973:BRc

[Sam73c] P. A. Samet. Book review. The Computer Journal, 16(3):261, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/261.2.full.pdf+html.

Samet:1974:BRa

[Sam74a] P. A. Samet. Book review. The Computer Journal, 17(2):112, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/112.2.full.pdf+html.

Samet:1974:BRb

[Sam74b] P. A. Samet. Book review. The Computer Journal, 17(3):228, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/228.2.full.pdf+html.

Samson:1976:TOA

[Sam76] W. B. Samson. Testing overflow algorithms for a table of variable size. The Computer Journal, 19(1):92, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
19/1/92.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/92.tif.

Sandys:1975:CCC

[San75] M. A. Sandys. Correspondence: On “Continuity Conditions for Spline Curves”. The Computer Journal, 18(3):282–283, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/282.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/282.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/283.tif.

See [Man74].

Saunders:1972:DAC

[Sau72] R. P. Saunders. A description and analysis of character maps. The Computer Journal, 15(2):160–169, May 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/15/2/160.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_

02/150160.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/160.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/161.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/162.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/163.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/164.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/165.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/166.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/167.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/169.tif.

Savage:1969:JOT

[Sav69] M. J. Savage. Junction optimisation technique. The Computer Journal, 12(3):268–272, August 1969. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/12/3/268.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
120268.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/268.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/269.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_
03/tiff/270.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_12/Issue_03/tiff/271.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_12/Issue_03/tiff/272.tif.

See erratum [Sav70].

Savage:1970:EJO

[Sav70] M. J. Savage. Erratum: “Junction optimisation technique”. The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/120.tif. See [Sav69].

Silvey:1970:HCT

[SB70] T. I. Silvey and J. R. Barker. Hybrid computing techniques for solving parabolic and hyperbolic partial differential equations. The Computer Journal, 13(2):164–170, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/164.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130164.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/164.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/165.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/166.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/167.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/168.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_02/tiff/169.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/170.
tif.

Scraton:1972:C

[SBS72] R. E. Scraton, H. W. Bradly, and David Silber. Correspondence. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/3/258.1.full.pdf+html.

Squire:1971:CBT

[SC71] William Squire and M. G. Cox. Correspondence: On “A bracketing technique for computing a zero of a function”. The Computer Journal, 14(1):72, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/72.tif. See [Cox70].

Schlageter:1976:PLV

[Sch76] G. Schlageter. The problem of lock by value in large data bases. The Computer Journal, 19(1):17–20, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
19/1/17.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/20.tif.

Scheuermann:1977:MIS

[Sch77a] P. Scheuermann. Modelling the information space in physical storage at different levels of detail. The Computer Journal, 20(3):248– 256, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/

content/20/3/248.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/248.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/249.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/250.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/251.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/252.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/253.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/254.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/255.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/256.tif.

Schwarz:1977:GCT

[Sch77b] Jerald Sydney Schwarz. Generic commands — A tool for partial correctness formalisms. The Computer Journal, 20(2):151– 155, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/151.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/151.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/152.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/153.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/154.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/155.tif.

Scott:1970:NSU

[Sco70] Melvin R. Scott. Numerical solution of unstable initial value problems by invariant imbedding. The Computer Journal, 13(4):397– 400, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/397.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/130397.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/397.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/398.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/399.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/400.tif.

Scott:1974:CDC

[Sco74] Jean V. Scott. Computer design of cascade logic. The Computer Journal, 17(3):253–260, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/3/253.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/253.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/255.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/256.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/257.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/258.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/259.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/260.tif.

Scraton:1970:MIC

[Scr70] R. E. Scraton. A method for improving the convergence of Chebyshev series. The Computer Journal, 13(2):202–203, May 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/2/202.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_02/130202.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/tiff/202.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_02/tiff/203.tif.

Scraton:1972:CGE

[Scr72] R. E. Scraton. Correspondence: The generalised Euler transformation. The Computer Journal, 15(3):258, August 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
tiff/258.tif. See [Wyn71].

Stocker:1973:SOD

[SD73] P. M. Stocker and P. A. Dearnley. Self-organising data management systems. The Computer Journal, 16(2):100–105, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/2/100.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/160100.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/100.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/101.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_16/Issue_02/tiff/102.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_02/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_02/tiff/105.tif. See correspon-

dence [PS74].

Samson:1978:STU

[SD78] W. B. Samson and R. H. Davis. Search times using hash tables for records with non-unique keys. The Computer Journal, 21(3): 210–214, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/3/210.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/210.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/211.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/212.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/213.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/214.tif.

Seaks:1971:EEE

[Sea71] Terry G. Seaks. Erratum: The evaluation of eigenvalues and eigenvectors of real symmetric matrices by simultaneous iteration. The Computer Journal, 14(4):351, November 1971. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140351.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/351.tif. See [CJ70].

Seldon:1972:C

[Sel72a] J. L. Seldon. Correspondence. The Computer Journal, 15(4):361, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/4/361.2.full.pdf+html.

Seldon:1972:CPS

[Sel72b] J. L. Seldon. Correspondence: The postage stamp problem. The Computer Journal, 15(4):361, November 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
tiff/361.tif. See [Lun69].

Setchfield:1979:BR

[Set79] K. Setchfield. Book review. The Computer Journal, 22(1):10, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/10.full.pdf+html.

Smith:1978:MFM

[SG78] D. A. Smith and Pamela S. Guipre. Modifications to the Forsythe– Moler algorithm for solving linear algebraic systems. The Computer Journal, 21(2):174–177, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/174.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/174.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/175.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/176.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/

tiff/177.tif. See correspondence [LS78, Hit78].

Samet:1972:CDL

[SH72a] P. A. Samet and D. W. Honey. Calculation of a double-length square root from double-length number using single precision techniques. The Computer Journal, 15(2):116, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/116.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/116.tif.

Smith:1972:ROS

[SH72b] J. L. M. Smith and T. S. Holden. Restart of an operating system having a permanent file structure. The Computer Journal, 15(1):25– 31, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/1/25.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/150025.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/25.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/26.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/27.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/28.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/29.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/30.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/31.tif.

Shain:1972:ATR

[Sha72a] M. Shain. Acceptance trials for a real time system. The Computer Journal, 15(4):298–301, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/298.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150298.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/298.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/299.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/300.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/301.tif.

Shanks:1972:RTS

[Sha72b] J. A. Shanks. Romberg tables for singular integrands. The Computer Journal, 15(4):360–361, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/360.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150360.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/360.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/361.tif.

Shackleton:1977:BR

[Sha77a] J. A. Shackleton. Book review. The Computer Journal, 20(4):329, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/329.1.full.pdf+html.

Shaw:1977:BR

[Sha77b] A. Shaw. Book review. The Computer Journal, 20(3):201, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/201.full.pdf+html.

Shepherd:1971:ASP

[She71] R. F. Shepherd. Algorithms supplement: Previously published algorithms. The Computer Journal, 14(4):444, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/444.tif.

Shepherd:1972:ASP

[She72a] R. F. Shepherd. Algorithms supplement: Previously published algorithms. The Computer Journal, 15(3):280–281, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/280.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/281.tif.

Shepherd:1972:ASS

[She72b] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 15(4):384, November 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
tiff/384.tif.

Shepherd:1973:ASSa

[She73a] R. F. Shepherd. Algorithms supplement — statement of policy. The Computer Journal, 16(1):95, February 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/95.tif.

Shepherd:1973:ASSb

[She73b] R. F. Shepherd. Algorithms supplement — statement of policy. The Computer Journal, 16(3):279–280, August 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/279.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/280.tif.

Shepherd:1974:ASS

[She74] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 17(1):91–92, February 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/91.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/92.tif.

Shepherdson:1977:R

[She77] J. C. Shepherdson. Review. The Computer Journal, 20(3):244, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
3/244.full.pdf+html.

Shepherd:1979:ASS

[She79] R. F. Shepherd. Algorithms supplement: Statement of policy. The Computer Journal, 22(1):90, February 1979. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/
tiff/90.tif.

Shrivastava:1975:VCP

[Shr75] S. K. Shrivastava. A view of concurrent process synchronisation. The Computer Journal, 18(4):375–379, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/375.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/375.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/376.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/377.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/378.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/379.tif.

Sibson:1971:SOP

[Sib71] R. Sibson. Some observations on a paper by Lance and Williams. The Computer Journal, 14(2):156–157, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/156.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140156.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/156.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/157.tif.

Sibson:1973:SOE

[Sib73] R. Sibson. SLINK: an optimally efficient algorithm for the singlelink cluster method. The Computer Journal, 16(1):30–34, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/1/30.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/160030.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/30.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_16/Issue_01/tiff/31.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_01/tiff/32.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
tiff/33.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/34.tif.

Sibson:1978:LET

[Sib78] R. Sibson. Locally equiangular triangulations. The Computer Journal, 21(3):243–245, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/3/243.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/243.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/244.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/245.tif.

Silber:1972:CDD

[Sil72] David Silber. Correspondence: On “Dynamic debugging in BA-SIC”. The Computer Journal, 15(3):258, August 1972. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
tiff/258.tif. See [Bul72].

Silberschatz:1979:SPC

[Sil79] A. Silberschatz. On the safety of the I/O primitive in Concurrent PASCAL. The Computer Journal, 22(2):142–145, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/142.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/142.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/143.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/144.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/145.tif.

Simpson:1974:DM

[Sim74] D. Simpson. Documentation manual. The Computer Journal, 17(4): 348, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/348.full.pdf+html.

Simpson:1979:BRa

[Sim79a] D. Simpson. Book review. The Computer Journal, 22(3):239, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/239.2.full.pdf+html.

Simpson:1979:BRb

[Sim79b] D. Simpson. Book review. The Computer Journal, 22(3):245, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/245.2.full.pdf+html.

Simpson:1979:PSM

[SJ79] H. R. Simpson and K. Jackson. Process synchronisation in MAS-COT. The Computer Journal, 22(4):332–345, November 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/332.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/332.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/333.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/334.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/335.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/336.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/337.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/338.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/339.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/340.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/341.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/342.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/344.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/345.tif.

Sethi:1975:C

[SK75] A. S. Sethi and M. S. Krishnamoorthy. Correspondence. The Computer Journal, 18(4):380–381, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/4/380.2.full.pdf+html.

Shacham:1976:DMS

[SK76] M. Shacham and E. Kehat. A direct method for the solution of large sparse systems of linear equations. The Computer Journal, 19(4):353– 359, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/353.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/353.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/354.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/355.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/356.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/357.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/358.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/359.tif.

Smith:1978:BR

[SK78] J. Smith and F. E. Knowles. Book review. The Computer Journal, 21(4):358, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/358.1.full.pdf+html.

Spyropoulos:1973:DAN

[SKY73] K. Spyropoulos, E. Kiountouzis, and A. Young. Discrete approximation in the L_{1} norm. The Computer Journal, 16(2):180– 186, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/180.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160180.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/180.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/181.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/182.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/183.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/184.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/185.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/186.tif.

Stentiford:1973:EAC

[SL73] F. W. M. Stentiford and D. W. Lewin. An evolutionary approach to the concept of randomness. The Computer Journal, 16(2):

148–151, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/148.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160148.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/148.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/149.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/150.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_

02/tiff/151.tif. See correspondence [Pit74].

Swift:1978:CCM

[SL78] A. Swift and G. R. Lindfield. Comparison of a continuation method with Brent’s method for the numerical solution of a single nonlinear equation. The Computer Journal, 21(4):359–362, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/359.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/359.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/360.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/361.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/362.tif.

Slater:1970:BR

[Sla70] L. J. Slater. Book review. The Computer Journal, 13(4):386, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/4/386.full.pdf+html.

Slater:1971:BR

[Sla71a] L. J. Slater. Book review. The Computer Journal, 14(4):436, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/436.full.pdf+html.

Slater:1971:BRB

[Sla71b] L. J. Slater. Book review: Mathematical Model Building in Economics and Industry (Second Series), by M. G. Kendall (editor), 1970. The Computer Journal, 14(4):436, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/436.tif.

Sloman:1978:BR

[Slo78] A. Sloman. Book review. The Computer Journal, 21(1):34, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/34.1.full.pdf+html.

Sabin:1971:NAA

[SM71] M. A. Sabin and D. J. McConalogue. Note on Algorithm 67: An axis invariant procedure to provide slopes at points by local three-point parametric curve fitting. The Computer Journal, 14(4): 451, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/451.tif. See [McC71b].

Small:1978:BRB

[Sma78a] R. E. Small. Book review: On-Line Data Bases: Infotech State of the Art Report, 1: Analysis and Bibliography, 2: Invited Papers (Infotech, £110). The Computer Journal, 21(4):336, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/336.1.
full.pdf+html.

Small:1978:CPS

[Sma78b] R. E. Small. Computational probability and simulation and mathematical modelling and digital simulation for engineers and scientists. The Computer Journal, 21(2):121, May 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/2/121.full.pdf+html.

Smillie:1970:AIS

[Smi70a] K. W. Smillie. Algorithm 49: Indexing subarrays in multidimensional arrays. The Computer Journal, 13(2):208, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_02/tiff/208.tif.

Smith:1970:DSR

[Smi70b] B. Smith. Developments in SPECOL — a retrieval language for the non-programmer. The Computer Journal, 13(1):10–13, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/1/10.full.pdf+html;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_13/Issue_01/130010.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/10.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/11.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/12.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/13.tif.

Smith:1971:DEH

[Smi71] L. B. Smith. Drawing ellipses, hyperbolas or parabolas with a fixed number of points and maximum inscribed area. The Computer Journal, 14(1):81–86, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/1/81.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/140081.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/81.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/82.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/tiff/84.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/tiff/85.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

14/Issue_01/tiff/86.tif. See [McC70b, Ecc71b, Sab72, Coo72f].

Smith:1972:PVP

[Smi72] J. Meredith Smith. Proof and validation of program correctness. The Computer Journal, 15(2):130–131, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/130.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150130.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/130.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/131.tif. See note [LB76].

Smith:1973:BR

[Smi73] C. A. B. Smith. Book review. The Computer Journal, 16(4):314, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/4/314.3.full.pdf+html.

Smith:1975:BR

[Smi75] A. B. Smith. Book review. The Computer Journal, 18(4):341, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
4/341.2.full.pdf+html.

Smith:1976:BR

[Smi76] A. B. Smith. Book review. The Computer Journal, 19(3):233, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/233.2.full.pdf+html.

Smith:1977:BR

[Smi77] C. A. B. Smith. Book review. The Computer Journal, 20(4):363, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/363.2.full.pdf+html.

Smith:1979:BR

[Smi79] J. M. Smith. Book review. The Computer Journal, 23(4):368, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/368.1.full.pdf+html.

Slagle:1975:ATP

[SN75] J. R. Slagle and L. M. Norton. Automated theorem-proving for the theories of partial and total ordering. The Computer Journal, 18(1): 49–54, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/18/1/49.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/49.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/50.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/51.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/52.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/53.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/54.tif.

Snell:1970:SCS

[Sne70] J. Snell. The solution in Chebyshev series of systems of linear differential equations with general boundary conditions.

The Computer Journal, 13(1):103–106, February 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/1/103.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_01/130103.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/103.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/104.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/105.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/106.
tif.

Sakaki:1976:AFL

[SNH76] T. Sakaki, K. Nakashima, and Y. Hattori. Algorithms for finding in the lump both bounds of the chromatic number of a graph. The Computer Journal, 19(4):329–332, November 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/329.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/329.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/330.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/331.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/332.tif.

Snow:1978:BR

[Sno78] C. R. Snow. Book review. The Computer Journal, 21(4):372, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
4/372.2.full.pdf+html.

Soden:1953:NMR

[Sod53] Walter Soden. Note: a method of radix conversion. Mathematical Tables and Other Aids to Computation, 7(44):273–274, October 1953. CODEN MTTCAS. ISSN 0891-6837. URL http://www.jstor.org/

stable/2002843.

Southwell:1976:FDN

[Sou76] W. H. Southwell. Fitting data to nonlinear functions with uncertainties in all measurement variables. The Computer Journal, 19(1): 69–73, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/19/1/69.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/69.tif;

http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/70.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_01/tiff/71.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_01/tiff/72.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
01/tiff/73.tif.

Soule:1977:NNT

[Sou77] S. Soule. A note on the nonrecursive traversal of binary trees. The Computer Journal, 20(4):350–352, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/350.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/350.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/351.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/352.tif.

Spath:1969:AIC

[Spä69] H. Späth. Algorithm 42: Interpolation by certain quintic splines. The Computer Journal, 12(3):292–293, August 1969. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_12/Issue_03/
tiff/292.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_12/Issue_03/tiff/293.tif. See note [Sal70b].

Squire:1971:C

[Squ71] William Squire. Correspondence. The Computer Journal, 14(1):72, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/72.full.pdf+html.

Smyth:1974:SSH

[SR74] W. F. Smyth and E. Radaceanu. A storage scheme for hierarchic structures. The Computer Journal, 17(2):152–156, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/152.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/152.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/153.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/154.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/155.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/156.tif.

Siskind:1976:STD

[SR76] V. Siskind and J. Rosenhead. Seek times for disc file processing: Some results from probability theory. The Computer Journal, 19(4):301– 305, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/301.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/301.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/302.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/303.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/304.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/
tiff/305.tif.

Schlechtendahi:1972:C

[SS72a] E. G. Schlechtendahi and U. Schumann. Correspondence. The Computer Journal, 15(2):147, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/2/147.full.pdf+html.

Schlechtendahl:1972:CRF

[SS72b] E. G. Schlechtendahl and U. Schumann. Correspondence: Remarks on “Faults in functions, in ALGOL and FORTRAN”. The Computer Journal, 15(2):147, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/
150147.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/147.tif. See [Hil71].

Smith:1972:NSD

[SS72c] I. R. Smith and L. A. Snider. Numerical solution of the differential equations of a synchronous generator. The Computer Journal, 15(4):350–355, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/350.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150350.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/350.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/351.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/352.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/353.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/354.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/355.tif.

Stoy:1972:OEOa

[SS72d] J. E. Stoy and C. Strachey. OS6 — an experimental operating system for a small computer. Part 1: general principles and structure. The Computer Journal, 15(2):117–124, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/117.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150117.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/117.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/118.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/119.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/120.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/121.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/122.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/123.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/124.tif.

Stoy:1972:OEOb

[SS72e] J. E. Stoy and C. Strachey. OS6 — an experimental operating system for a small computer. Part 2: input/output and filing system. The Computer Journal, 15(3):195–203, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/195.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150195.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/195.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/197.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/198.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/200.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/201.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/202.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/203.tif.

Smillie:1975:CDT

[SS75] K. W. Smillie and M. J. R. Shave. Converting decision tables to computer programs. The Computer Journal, 18(2):108–111, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/108.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/108.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/109.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
02/tiff/110.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/111.tif.

Schmidt:1979:TCA

[SS79] G. Schmidt and T. Ströhlein. Timetable construction — an annotated bibliography. The Computer Journal, 23(4):307–316, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/23/
4/307.full.pdf+html.

Stark:1971:LA

[Sta71] R. H. Stark. A language for algorithms. The Computer Journal, 14(1):40–44, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/1/40.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
140040.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/40.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_01/tiff/41.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
01/tiff/42.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_01/tiff/43.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_01/tiff/44.tif.

Stacey:1974:FIC

[Sta74] G. M. Stacey. A Fortran interface to the CODASYL Database Task Group specifications. The Computer Journal, 17(2):124– 129, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/2/124.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/124.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/

computer_journal/hdb/Volume_17/Issue_02/tiff/127.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/128.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/129.tif.

Stamper:1975:C

[Sta75] R. K. Stamper. Correspondence. The Computer Journal, 18(1):93, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/93.1.full.pdf+html.

Stamper:1977:BR

[Sta77a] R. K. Stamper. Book review. The Computer Journal, 20(4):363, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/363.1.full.pdf+html.

Stamper:1977:LPS

[Sta77b] R. K. Stamper. The LEGOL 1 prototype system and language. The Computer Journal, 20(2):102–108, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/102.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/102.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/103.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/104.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/105.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/106.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/107.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/108.tif.

Stephens:1974:ILC

[Ste74] P. D. Stephens. The IMP language and compiler. The Computer Journal, 17(3):216–223, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/3/216.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/216.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/217.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/218.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/

tiff/219.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/220.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_03/tiff/221.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_03/
tiff/222.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_03/tiff/223.tif.

Stephenson:1975:BR

[Ste75] J. Stephenson. Book review. The Computer Journal, 18(2):107, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
2/107.1.full.pdf+html.

Stecher:1977:PIS

[Ste77a] P. Stecher. Proposal for an interface system between the business and data processing systems. The Computer Journal, 20(3):194– 201, August 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/3/194.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/194.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/195.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/197.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/198.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_03/tiff/199.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_03/tiff/200.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_03/
tiff/201.tif.

Steele:1977:CNA

[Ste77b] K. A. Steele. CPM network analysis with a storage tube terminal. The Computer Journal, 20(4):330–334, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/330.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/330.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/331.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/332.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/334.tif.

Storey:1972:OFB

[Sto72] S. H. Storey. Optimal fixed batch sizes for real-time inquiry systems. The Computer Journal, 15(2):132–133, May 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/132.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150132.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/132.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/133.tif.

Stojak:1973:HDC

[Sto73] P. F. Stojak. Hybrid and digital computation results in multicomponent distillation simulation. The Computer Journal, 16(4):368– 374, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/368.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160368.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/368.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/369.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/370.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/371.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/372.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/373.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/374.tif.

Stross:1972:ODD

[Str72] C. O. M. Stross. Operation of a disc data base. The Computer Journal, 15(4):290–297, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/4/290.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150290.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/290.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/291.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
04/tiff/292.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/293.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/294.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_

04/tiff/295.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/296.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_04/tiff/297.tif.

Strawson:1974:C

[Str74a] J. Strawson. Correspondence. The Computer Journal, 17(2):190, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
17/2/190.1.full.pdf+html.

Strawson:1974:CH

[Str74b] Joy Strawson. Correspondence: On “Himmelbett”. The Computer Journal, 17(2):190, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_17/Issue_02/tiff/190.tif. See

[Bel73, Fre74b].

Sturman:1971:FPM

[Stu71] J. N. Sturman. The functional partitioning of memory — its application to interactive computing. The Computer Journal, 14(4):352– 360, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/14/4/352.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_04/140352.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/352.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/353.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/354.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/355.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/356.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/357.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/358.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/359.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/360.tif.

Sturt:1973:BR

[Stu73] D. A. Sturt. Book review. The Computer Journal, 16(3):215, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
3/215.2.full.pdf+html.

Sufrin:1978:CC

[Suf78] B. A. Sufrin. Compiler construction. The Computer Journal, 21 (2):143, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/143.full.pdf+html.

Sufrin:1979:BR

[Suf79] B. Sufrin. Book review. The Computer Journal, 22(2):150, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/150.3.
full.pdf+html.

Sumner:1976:BR

[Sum76] F. H. Sumner. Book review. The Computer Journal, 19(4):325, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/325.full.pdf+html.

Sumner:1978:CAS

[Sum78] F. H. Sumner. Computer architecture, second edition. The Computer Journal, 21(2):133, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/2/133.full.pdf+html.

Sutcliffe:1976:ADC

[Sut76a] D. C. Sutcliffe. An algorithm for drawing the curve F (X, Y) = 0. The Computer Journal, 19(3):246–249, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/246.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/246.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/247.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/248.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/249.tif.

Sutcliffe:1976:RCA

[Sut76b] D. C. Sutcliffe. A remark on a contouring algorithm. The Computer Journal, 19(4):333–335, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/4/333.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_04/

tiff/333.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_04/tiff/334.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_04/tiff/335.tif.

Szwarcfiter:1978:SPT

[SW78] J. L. Szwarcfiter and L. B. Wilson. Some properties of ternary trees. The Computer Journal, 21(1):66–72, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
21/1/66.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/66.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/67.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/68.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/69.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/70.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/71.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/72.tif.

Spengler:1979:ADR

[SW79] C. B. Spengler and J. B. Westwood. Assessing data retrieval capabilities: a case study. The Computer Journal, 22(1):17–21, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/17.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/17.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/18.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/19.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/20.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/21.
tif.

Smith:1972:GRE

[SY72] L. W. Smith and S. S. Yau. Generation of regular expressions for automata by the integral of regular expressions. The Computer Journal, 15(3):222–228, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/3/222.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/
150222.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/222.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/223.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/224.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/225.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_03/tiff/226.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
03/tiff/227.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/tiff/228.tif.

Symons:1971:CAF

[Sym71] C. R. Symons. A cost accounting formula for multiprogramming computers. The Computer Journal, 14(1):13–20, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/13.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140013.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/13.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/14.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/15.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/16.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/17.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/18.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/19.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/20.tif.

Shaftel:1974:ACR

[SZ74] T. L. Shaftel and R. W. Zmud. Allocation of computer resources through flexible pricing. The Computer Journal, 17(4):306–312, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/17/4/306.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/306.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/307.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/308.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/309.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/310.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/311.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_04/tiff/312.tif.

Szanser:1973:BTE

[Sza73] A. J. Szanser. Bracketing technique in elastic matching. The Computer Journal, 16(2):132–134, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/132.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160132.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/132.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/133.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/134.tif.

Tanenbaum:1978:CPA

[Tan78] A. S. Tanenbaum. A comparison of Pascal and Algol 68. The Computer Journal, 21(4):316–323, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/316.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/316.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/317.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/318.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/319.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/320.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/321.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/322.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/323.tif.

Taylor:1970:BVP

[Tay70a] J. C. Taylor. Boundary value problems for singular ordinary differential equations. A widely applicable method for calculating eigenvalues. The Computer Journal, 13(3):291–296, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/291.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130291.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/291.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/292.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/293.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/294.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

13/Issue_03/tiff/295.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/296.tif.

Taylor:1970:SDF

[Tay70b] P. J. Taylor. The stability of the Du Fort–Frankel method for the diffusion equation with boundary conditions involving space derivatives. The Computer Journal, 13(1):92–97, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
13/1/92.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/130092.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/92.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/93.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/94.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/95.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/96.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/97.tif.

Taylor:1973:BRa

[Tay73a] F. E. Taylor. Book review. The Computer Journal, 16(4):321, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/321.2.full.pdf+html.

Taylor:1973:BRb

[Tay73b] F. E. Taylor. Book reviews. The Computer Journal, 16(4):346, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/346.full.pdf+html.

Taylor:1976:BR

[Tay76a] P. J. Taylor. Book review. The Computer Journal, 19(1):55, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
1/55.full.pdf+html.

Taylor:1976:FIO

[Tay76b] R. Taylor. Fast input/output of variable-length arrays in FOR-TRAN IV. The Computer Journal, 19(3):273, August 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/273.tif.

Taylor:1980:CJS

[Tay80] N. B. Taylor. Correspondence: Jumping to some purpose. The Computer Journal, 23(2):189–190, May 1980. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_23/Issue_02/
tiff/189.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_23/Issue_02/tiff/190.tif. See [ASG79, Hil80, Rob80,

MR82].

Tomlin:1971:BRB

[TCB71] J. A. Tomlin, D. C. Cooper, and D. A. Bell. Book reviews: Integer and Dynamic Programming, by J. Abadie (editor), 1970; Algorithms and Recursive Functions, by A. L. Mal’cev, 1970; Information Theory for Systems Engineers (Econometrics and Operations Research XVII), by L. P. Hyvärinen, 1970. The Computer Journal, 14(4): 343, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/343.tif.

Tosser:1977:AON

[TD77] A. Tosser and D. Dubus. Algebraic optimisation of NAND-NOR switching circuits. The Computer Journal, 20(1):73–77, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/73.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/73.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/74.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/75.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/76.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/77.
tif.

Templeman:1972:C

[Tem72a] J. A. Templeman. Correspondence. The Computer Journal, 15 (4):359, ???? 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/4/359.full.pdf+html.

Templeman:1972:CDC

[Tem72b] J. A. Templeman. Correspondence: On “A data collection and display system for a large-scale simulation”. The Computer Journal, 15(4):359, November 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/
150359.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/tiff/359.tif. See [BPSS72].

Templeman:1975:C

[Tem75] J. A. Templeman. Correspondence. The Computer Journal, 18 (2):189, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/189.1.full.pdf+html.

Terrine:1973:CGP

[Ter73] G. Terrine. Coordinate grammars and parsers. The Computer Journal, 16(3):232–244, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/3/232.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
160232.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/232.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/233.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/234.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/235.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/236.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/237.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/238.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/239.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/240.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/241.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/242.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/243.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/244.tif.

Teskey:1977:DCK

[Tes77] N. Teskey. Discussion and correspondence: KATE: a macro-processor for extending command languages. The Computer Journal, 20(2):

187–189, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/187.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/187.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/188.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/189.tif.

Tewarson:1972:MSL

[Tew72] R. P. Tewarson. On minimax solutions of linear equations. The Computer Journal, 15(3):277–279, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/277.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150277.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/277.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/278.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/279.tif.

Thornton:1972:RLQ

[Tho72] B. S. Thornton. Reduction and linearisation of quadratic constrained optimisation in linked systems. The Computer Journal, 15(1):84– 87, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/1/84.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/150084.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/84.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/85.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/86.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/87.tif.

Thomas:1973:BR

[Tho73] M. Thomas. Book review. The Computer Journal, 16(4):321, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/321.1.full.pdf+html.

Thomas:1974:BRa

[Tho74a] A. J. Thomas. Book review. The Computer Journal, 17(1):63, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/63.3.full.pdf+html.

Thomas:1974:BRc

[Tho74b] M. Thomas. Book review. The Computer Journal, 17(2):112, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/112.1.full.pdf+html.

Thomas:1974:BRb

[Tho74c] M. C. Thomas. Book review. The Computer Journal, 17(1):74, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/74.1.full.pdf+html.

Thomas:1979:BRa

[Tho79a] A. J. Thomas. Book review. The Computer Journal, 22(1):27, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/27.2.full.pdf+html.

Thomas:1979:BRb

[Tho79b] A. J. Thomas. Book review. The Computer Journal, 22(4):331, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/331.1.full.pdf+html.

Thomas:1979:BRc

[Tho79c] M. Thomas. Book review. The Computer Journal, 22(4):362, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/362.2.full.pdf+html.

Thomson:1979:BR

[Tho79d] A. Thomson. Book review. The Computer Journal, 22(3):250, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/250.full.pdf+html.

Tabata:1975:ETI

[TI75] T. Tabata and R. Ito. Effective treatment of the interpolation factor in Marquardt’s nonlinear least-squares fit algorithm. The Computer Journal, 18(3):250–251, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/250.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/250.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/251.tif.

Tomlin:1971:BR

[Tom71] J. A. Tomlin. Book review. The Computer Journal, 14(4):343, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/343.1.full.pdf+html.

Tomlinson:1975:C

[Tom75] J. Tomlinson. Correspondence. The Computer Journal, 18(2):190, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/190.1.full.pdf+html.

Townsend:1977:BR

[Tow77a] H. R. A. Townsend. Book review. The Computer Journal, 20(3):231, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/3/231.3.full.pdf+html.

Townsend:1977:SPA

[Tow77b] H. R. A. Townsend. SST — a program for the automatic generation of symbol state tables. The Computer Journal, 20(2): 124–127, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/2/124.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/124.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_02/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_02/tiff/126.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_02/tiff/127.tif.

Townsend:1978:AE

[Tow78] H. R. A. Townsend. Analysing electroencephalograms. The Computer Journal, 21(3):259–262, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/3/259.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/

tiff/259.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/260.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/261.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/262.tif.

Treble:1970:CCC

[Tre70] D. P. Treble. Computers and composition in change ringing. The Computer Journal, 13(4):350–351, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/350.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/130350.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_04/tiff/350.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/351.tif.

Triance:1974:DCH

[Tri74] J. M. Triance. Discussion and correspondence: Handling records with a variable structure in COBOL. The Computer Journal, 17(1):93– 94, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/17/1/93.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/93.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/94.tif.

Triance:1976:SCS

[Tri76] J. M. Triance. The significance of the 1974 COBOL standard. The Computer Journal, 19(4):295–300, November 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/295.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/295.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/296.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/297.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/298.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/299.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/300.tif.

Triance:1978:DCS

[Tri78] J. M. Triance. Discussion and correspondence: a study of COBOL portability. The Computer Journal, 21(3):278–281, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/278.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/278.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/279.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/280.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/281.tif.

Trapp:1975:SNV

[TS75] G. E. Trapp, Jr. and W. Squire. Solving nonlinear Vandermonde systems. The Computer Journal, 18(4):373–374, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/373.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/373.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/374.tif.

Tyler:1973:CE

[TT73] C. R. Tyler and The Editor. Correspondence: Editorial. The Computer Journal, 16(3):286–287, August 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/286.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/287.tif. See [Ham73, CT74, Fis74c].

Turner:1975:GCS

[Tur75] L. F. Turner. The ‘on-ground’ compression of satellite data. The Computer Journal, 18(3):243–247, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/243.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/243.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/244.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/246.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/247.tif.

Turski:1978:MND

[Tur78] W. M. Turski. A model of nonmorphological data collection based on sentential calculus. The Computer Journal, 21(4):337– 340, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/4/337.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/337.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/338.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/339.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/340.tif. See

note [Hay79].

Turner:1979:CDW

[Tur79] J. C. Turner. The conditional distribution of waiting time given queue length in a computer system. The Computer Journal, 22(1): 57–62, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/22/1/57.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/57.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/58.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/59.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_01/tiff/60.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
01/tiff/61.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_01/tiff/62.tif.

Thrower:1975:C

[TW75] C. A. Thrower and C. A. W. Wells. Correspondence. The Computer Journal, 18(2):188–189, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/188.3.full.pdf+html.

Ting:1977:MRS

[TW77] T. C. Ting and Y. W. Wang. Multiway replacement selection sort with dynamic reservoir. The Computer Journal, 20(4):298– 301, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/298.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/298.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/299.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_20/Issue_04/tiff/300.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/301.tif.

Twizell:1975:NSS

[Twi75] E. H. Twizell. The numerical solution of second order hyperbolic partial differential equations with unequally spaced initial conditions. The Computer Journal, 18(3):252–257, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/252.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/252.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/253.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/254.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/255.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/256.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/257.tif.

Tzafestas:1973:SEA

[Tza73] S. G. Tzafestas. State estimation algorithms for non-linear stochastic sequential machines. The Computer Journal, 16(3):245–253, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-

2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/3/245.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/160245.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/245.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/246.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/247.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/248.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/249.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/250.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
03/tiff/251.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_03/tiff/252.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_03/tiff/253.tif.

Ulery:1974:SLO

[UK74] Dana L. Ulery and H. M. Khalil. A survey of language-oriented systems for numerical linear algebra. The Computer Journal, 17(1): 82–88, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/17/1/82.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/82.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/83.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/84.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/85.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/86.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/87.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/88.tif.

Ullmann:1973:BR

[Ull73a] J. R. Ullmann. Book review. The Computer Journal, 16(3):225, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/3/225.full.pdf+html.

Ullmann:1973:BRB

[Ull73b] J. R. Ullmann. Book review: Introduction to Mathematical Techniques in Pattern Recognition, by Harry C. Andrews, 1972. The Computer Journal, 16(3):225, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/225.tif.

Ullmann:1977:BGT

[Ull77] J. R. Ullmann. A binary n-gram technique for automatic correction of substitution, deletion, insertion and reversal errors in words. The Computer Journal, 20(2):141–147, May 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/141.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/141.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/142.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/143.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/144.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/145.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/146.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/147.tif.

Vajda:1970:BR

[Vaj70] S. Vajda. Book review. The Computer Journal, 13(4):412a–412, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/412-a.full.pdf+html.

Vajda:1976:BRa

[Vaj76a] S. Vajda. Book review. The Computer Journal, 19(4):305, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/305.1.
full.pdf+html.

Vajda:1976:BRb

[Vaj76b] S. Vajda. Book review. The Computer Journal, 19(4):305, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/305.2.
full.pdf+html.

Valentine:1971:C

[Val71a] S. H. Valentine. Correspondence. The Computer Journal, 14(4):365, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/365.1.full.pdf+html.

Valentine:1971:CAN

[Val71b] S. H. Valentine. Correspondence: “On the automatic numerical evaluation of definite integrals”. The Computer Journal, 14(4): 365, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/365.tif. See [CP71].

Valentine:1974:CNA

[Val74] S. H. Valentine. Comparative notes on Algol 68 and PL/I. The Computer Journal, 17(4):325–331, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/325.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/325.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/326.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/327.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/328.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/329.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/330.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/331.tif.

VandeVel:1970:NAP

[Van70] H. Van de Vel. A note on the automatic pretreatment of polynomials. The Computer Journal, 13(3):289–290, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/289.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130289.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/289.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/290.tif.

VanHoa:1973:ACP

[Van73] Tran Van Hoa. Algorithm 78: Counting preferential votes in multi-member constituencies using absolute majority criteria. The Computer Journal, 16(3):273–276, August 1973. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/
tiff/273.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/274.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_03/tiff/275.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_03/

tiff/276.tif. See note [HW74, Par74].

Varney:1976:TUO

[Var76a] R. C. Varney. Toward the understandability of an operating system. The Computer Journal, 19(3):213–215, August 1976. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/3/213.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_03/tiff/213.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_03/tiff/214.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
03/tiff/215.tif.

Varol:1976:SRC

[Var76b] Y. L. Varol. Some remarks on computer acquisition. The Computer Journal, 19(2):127–131, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/127.full.pdf+html;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/127.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/128.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/129.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/130.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/131.tif.

Varol:1977:GPR

[Var77] Y. L. Varol. On the generation of the pseudo-remainder in polynomial division. The Computer Journal, 20(2):178–180, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/2/178.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/178.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/179.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/180.tif.

Varol:1979:SAP

[Var79] Y. L. Varol. On a subset of all the permutations of n marks. The Computer Journal, 23(4):344–346, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/23/4/344.full.pdf+html.

Varney:1973:SFO

[VG73] R. C. Varney and M. H. Gotterer. The structural foundation for an operating system. The Computer Journal, 16(4):357–359, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/4/357.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_04/160357.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
04/tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/tiff/358.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_04/tiff/359.tif.

Vranesic:1972:TLP

[VH72] Z. G. Vranesic and V. C. Hamacher. Ternary logic in parallel multipliers. The Computer Journal, 15(3):254–258, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/254.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_15/Issue_03/150254.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/254.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/255.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/256.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/257.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

15/Issue_03/tiff/258.tif. See correspondence [BVH74].

Vickers:1979:BR

[Vic79] T. Vickers. Book review. The Computer Journal, 22(3):245, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/245.1.full.pdf+html.

Vit:1975:RIE

[Vit75] K. Vit. Rational interpolation and extrapolation for SUMT. The Computer Journal, 18(4):368–370, November 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/368.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/368.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/369.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/370.tif.

vanOosterom:1978:THT

[vO78] A. van Oosterom. Triangulating the human torso. The Computer Journal, 21(3):253–258, August 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/3/253.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/253.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/254.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/255.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_03/
tiff/256.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_03/tiff/257.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_03/tiff/258.tif.

Vorvoreanu:1975:C

[Vor75] J. C. Vorvoreanu. Correspondence. The Computer Journal, 18(2): 189–190, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/189.2.full.pdf+html.

vanRijsbergen:1970:ACA

[vR70a] C. J. van Rijsbergen. Algorithm 47: a clustering algorithm. The Computer Journal, 13(1):113–115, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/
tiff/113.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_01/tiff/114.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_01/tiff/115.tif. See

notes [vR70c, Mux70, Sal70c].

vanRijsbergen:1970:AFH

[vR70b] C. J. van Rijsbergen. Algorithm 52: a fast hierarchic clustering algorithm. The Computer Journal, 13(3):324–326, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/324.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/325.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/326.
tif.

vanRijsbergen:1970:NAC

[vR70c] C. J. van Rijsbergen. Note on Algorithm 47: a clustering algorithm. The Computer Journal, 13(2):219, May 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_02/

tiff/219.tif. See [vR70a, Mux70, Sal70c].

vanRijsbergen:1971:AIS

[vR71] C. J. van Rijsbergen. An algorithm for information structuring and retrieval. The Computer Journal, 14(4):407–412, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/407.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140407.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/407.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/408.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/409.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/410.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

14/Issue_04/tiff/411.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/412.tif.

Waite:1970:BMP

[Wai70] W. M. Waite. Building a mobile programming system. The Computer Journal, 13(1):28–31, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/28.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130028.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/31.tif.

Walwyn:1971:C

[Wal71a] P. R. Walwyn. Correspondence. The Computer Journal, 14(4):425, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/425.full.pdf+html.

Walwyn:1971:CHL

[Wal71b] P. R. Walwyn. Correspondence: On “High level languages for low level users”. The Computer Journal, 14(4):425, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/140425.sgm.abs.html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_04/tiff/425.tif. See

[ER71] and correspondence [Bac71b, FW72, MT72, GT72, Hil72b, KTW^{+}73].

Walker:1973:BRa

[Wal73a] B. S. Walker. Book review. The Computer Journal, 16(2):125, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/125.full.pdf+html.

Walker:1973:BRb

[Wal73b] B. S. Walker. Book review. The Computer Journal, 16(4):379, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
4/379.1.full.pdf+html.

Walsh:1973:BR

[Wal73c] J. Walsh. Book review. The Computer Journal, 16(3):261, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/3/261.1.
full.pdf+html.

Walker:1974:BR

[Wal74a] B. S. Walker. Book review. The Computer Journal, 17(2):123, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
2/123.2.full.pdf+html.

Walker:1974:ICD

[Wal74b] B. S. Walker. Integrated circuits in digital electronics. The Computer Journal, 17(4):354a–354, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/4/354-a.full.pdf+html.

Walker:1975:BR

[Wal75a] B. S. Walker. Book review. The Computer Journal, 18(3):242, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/242.2.full.pdf+html.

Wallis:1975:C

[Wal75b] S. E. Wallis. Correspondence. The Computer Journal, 18(2):188, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/2/188.1.full.pdf+html.

Walker:1976:BRa

[Wal76a] B. S. Walker. Book review. The Computer Journal, 19(3):201, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
3/201.full.pdf+html.

Walker:1976:BRb

[Wal76b] B. S. Walker. Book review. The Computer Journal, 19(4):314, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
4/314.1.full.pdf+html.

Walker:1976:MFA

[Wal76c] B. S. Walker. Microcomputers: Fundamentals and applications; minicomputer evaluation and selection; minicomputers: in industrial process control. The Computer Journal, 19(2):121, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/19/
2/121.full.pdf+html.

Walters:1976:BTC

[Wal76d] R. E. Walters. Benchmark techniques: a constructive approach. The Computer Journal, 19(1):50–55, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
19/1/50.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/50.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/51.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/52.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/53.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/54.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/55.tif.

Wallach:1977:SAC

[Wal77] Y. Wallach. Scheduling algorithms for concurrent execution. The Computer Journal, 20(2):132–136, May 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/20/2/132.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/133.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_02/tiff/134.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
02/tiff/135.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_02/tiff/136.tif.

Wallis:1978:BR

[Wal78] P. Wallis. Book review. The Computer Journal, 21(3):214, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/214.1.
full.pdf+html.

Wallis:1979:BRa

[Wal79a] P. Wallis. Book review. The Computer Journal, 22(1):27, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/27.1.full.pdf+html.

Wallis:1979:BRb

[Wal79b] P. Wallis. Book review. The Computer Journal, 22(2):163, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/163.2.
full.pdf+html.

Wallis:1979:BRc

[Wal79c] P. J. L. Wallis. Book review. The Computer Journal, 23(4):337, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/337.full.pdf+html.

Wane:1973:C

[Wan73] R. Wane. Correspondence. The Computer Journal, 16(2):190, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/16/
2/190.1.full.pdf+html.

Ward:1974:VDM

[War74] R. G. Ward. A variable delay method for improving recognition of parallel processable code in computer programs. The Computer Journal, 17(2):157–164, May 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/2/157.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/157.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/158.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/159.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/160.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_02/tiff/162.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_02/
tiff/163.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_02/tiff/164.tif.

Waters:1971:BSP

[Wat71a] S. J. Waters. Blocking sequentially processed magnetic files. The Computer Journal, 14(2):109–112, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/14/2/109.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/140109.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/109.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/110.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/111.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/112.tif.

Watt:1971:BR

[Wat71b] J. M. Watt. Book review. The Computer Journal, 14(4):412, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/412.1.full.pdf+html.

Waters:1972:EFD

[Wat72a] S. J. Waters. Erratum: “File design fallacies”. The Computer Journal, 15(3):258, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_15/Issue_03/tiff/258.tif. See

[Wat72b].

Waters:1972:FDF

[Wat72b] S. J. Waters. File design fallacies. The Computer Journal, 15(1):1–4, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/1/1.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
150001.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/1.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/tiff/2.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/tiff/3.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/4.tif. See erratum [Wat72a].

Watt:1972:NIV

[Wat72c] J. M. Watt. Numerical initial value problems in ordinary differential equations. The Computer Journal, 15(2):155, May 1972.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
2/155.full.pdf+html.

Watson:1973:EBA

[Wat73] G. A. Watson. On estimating best approximations of functions defined by integral equations. The Computer Journal, 16(1):77– 80, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/16/1/77.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/160077.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/77.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/78.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/79.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/80.tif.

Waters:1974:BR

[Wat74a] S. Waters. Book review. The Computer Journal, 17(3):252, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/252.1.full.pdf+html.

Waters:1974:CAM

[Wat74b] S. J. Waters. Computer-aided methodology of computer systems design. The Computer Journal, 17(3):211–215, August 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/211.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/211.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/212.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/213.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/214.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/215.tif.

Waters:1974:MCS

[Wat74c] S. J. Waters. Methodology of computer systems design. The Computer Journal, 17(1):17–24, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/17/1/17.full.pdf+html;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/17.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/18.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/19.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/20.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/21.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_01/tiff/22.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
01/tiff/23.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_01/tiff/24.tif.

Waters:1975:ASI

[Wat75a] S. J. Waters. Analysis of self-indexing, disc files. The Computer Journal, 18(3):200–205, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/3/200.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/200.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/201.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/202.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_03/
tiff/203.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_03/tiff/204.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_03/tiff/205.tif.

Waters:1975:EMD

[Wat75b] S. J. Waters. Estimating magnetic disc seeks. The Computer Journal, 18(1):12–17, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/18/1/12.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/13.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/14.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
01/tiff/15.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_01/tiff/16.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_01/tiff/17.tif.

Waters:1976:CPA

[Wat76a] S. J. Waters. CAM 01: a precedence analyser. The Computer Journal, 19(2):122–126, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/122.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/122.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/123.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_02/tiff/124.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_02/
tiff/125.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_02/tiff/126.tif.

Waters:1976:HR

[Wat76b] S. J. Waters. Hit ratios. The Computer Journal, 19(1):21–24, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/1/21.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/21.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/tiff/22.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
19/Issue_01/tiff/23.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_19/Issue_01/tiff/24.tif.

Waters:1977:CSP

[Wat77a] S. J. Waters. CAM02: a structured precedence analyser. The Computer Journal, 20(1):2–5, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/
tiff/2.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/3.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/5.
tif.

Waters:1977:MVM

[Wat77b] S. J. Waters. Majority verdicts in multi-processing — any two from three. The Computer Journal, 20(3):207–212, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/207.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/207.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/208.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/209.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/210.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_20/Issue_03/tiff/211.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/212.tif.

Waters:1979:TCS

[Wat79] S. J. Waters. Towards comprehensive specifications. The Computer Journal, 22(3):195–199, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/3/195.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/195.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/196.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_03/tiff/197.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_03/
tiff/198.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/199.tif.

Wetherell:1972:DKV

[WBB72] C. S. Wetherell, T. J. Buckholtz, and K. S. Booth. A director for Kriegspiel, a variant of chess. The Computer Journal, 15(1):66– 70, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.
org/content/15/1/66.full.pdf+html;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/150066.sgm.abs.
html;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/66.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/67.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/tiff/68.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_01/tiff/70.tif.

Wiseman:1978:MGA

[WCH78] N. E. Wiseman, C. I. C. Campbell, and J. Harradine. On making graphic arts quality output by computer. The Computer Journal, 21(1):2–6, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/1/2.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/
tiff/2.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_01/tiff/3.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/4.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/5.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/6.tif.

Williams:1971:GSD

[WCL71] W. T. Williams, H. T. Clifford, and G. N. Lance. Group-size dependence: a rationale for choice between numerical classifications. The Computer Journal, 14(2):157–162, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/157.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140157.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/157.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/158.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/159.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/160.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/161.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/162.tif.

Wells:1974:DEP

[WD74] M. Wells and A. Denson. Direct execution of programming languages. The Computer Journal, 17(2):130–134, May 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/130.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/130.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/131.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
02/tiff/132.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_02/tiff/133.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_02/tiff/134.tif.

Wilton-Davies:1976:C

[WD76] C. C. Wilton-Davies. Correspondence. The Computer Journal, 19 (4):381, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/381.3.full.pdf+html.

Wichmann:1979:PCG

[WD79] B. A. Wichmann and J. Du Croz. A program to calculate the GAMM measure. The Computer Journal, 22(4):317–322, November 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/4/317.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/317.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/318.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/319.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_04/tiff/320.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_04/tiff/321.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
04/tiff/322.tif.

Webb:1970:BRB

[Web70] M. J. H. Webb. Book review: The Economics of Computers, by William F. Sharpe, 1969. The Computer Journal, 13(2):222, May 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_02/tiff/222.tif.

Webb:1972:CPP

[Web72] M. H. J. Webb. Computer procurement policies for universities and similar users. The Computer Journal, 15(2):170–175, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/2/170.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/150170.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/170.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/171.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/172.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_02/tiff/173.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_02/tiff/174.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_02/tiff/175.tif.

Wegbreit:1972:GCG

[Weg72] B. Wegbreit. A generalised compactifying garbage collector. The Computer Journal, 15(3):204–208, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/204.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150204.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/204.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

15/Issue_03/tiff/205.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/206.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/207.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/208.tif.

Wegbreit:1974:PCE

[Weg74] Ben Wegbreit. Procedure closure in EL1. The Computer Journal, 17(1):38–43, February 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/17/1/38.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/38.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/40.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_01/
tiff/41.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_01/tiff/42.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_17/Issue_01/tiff/43.tif.

Weir:1979:BR

[Wei79] J. A. S. Weir. Book review. The Computer Journal, 22(2):145, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
2/145.1.full.pdf+html.

Wells:1970:CTF

[Wel70] M. Wells. Correspondence: Towards Fortran VI? The Computer Journal, 13(1):120, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_13/Issue_01/tiff/120.tif. See

[HL70b].

Wells:1972:FCU

[Wel72] M. Wells. File compression using variable length encodings. The Computer Journal, 15(4):308–313, November 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/4/308.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_04/150308.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/308.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/309.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_15/Issue_04/tiff/310.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_04/tiff/311.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_04/tiff/312.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_04/tiff/313.tif. See correspon-

dence [GW73].

Welland:1973:BR

[Wel73] R. C. Welland. Book review. The Computer Journal, 16(2):131, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
16/2/131.full.pdf+html.

Wells:1974:BR

[Wel74] M. Wells. Book review. The Computer Journal, 17(2):191, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/191.3.
full.pdf+html.

Welland:1975:BR

[Wel75] R. C. Welland. Book review. The Computer Journal, 18(1):20, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/20.2.full.pdf+html.

Wellicome:1976:BR

[Wel76] J. F. Wellicome. Book review. The Computer Journal, 19(3):206, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/206.1.full.pdf+html.

Wells:1977:BR

[Wel77] M. Wells. Book review. The Computer Journal, 20(3):256, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/256.4.
full.pdf+html.

Wells:1979:BR

[Wel79] M. Wells. Book review. The Computer Journal, 22(3):202, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/202.full.pdf+html.

West:1971:BR

[Wes71] John C. West. Book review. The Computer Journal, 14(1):102, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
1/102.full.pdf+html.

Westcott:1974:BR

[Wes74] J. H. Westcott. Book review. The Computer Journal, 17(1):88, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/88.1.full.pdf+html.

Wexelblat:1976:C

[Wex76] R. L. Wexelblat. Correspondence. The Computer Journal, 19(4): 382–383, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/19/4/382.2.full.pdf+html.

Williams:1978:DCA

[WF78] M. H. Williams and E. V. C. Fielding. Discussion and correspondence: automatic generation of payroll programs. The Computer Journal, 21(4):378–380, November 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/21/4/378.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_04/
tiff/378.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_04/tiff/379.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_04/tiff/380.tif.

Wheeler:1978:RFA

[Whe78] D. J. Wheeler. Rational fault analysis. The Computer Journal, 21 (2):177, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/177.full.pdf+html.

Wheeler:1979:BR

[Whe79] D. J. Wheeler. Book review. The Computer Journal, 23(4):359, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
23/4/359.2.full.pdf+html.

Whiteley:1972:C

[Whi72a] E. Whiteley. Correspondence. The Computer Journal, 15(1): 24, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/15/1/24.full.pdf+html.

Whiteley:1972:CSS

[Whi72b] E. Whiteley. Correspondence: On “Step size adjustment at discontinuities for fourth order Runge–Kutta methods”. The Computer Journal, 15(1):24, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
150024.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/24.tif. See [O’R70].

Wells:1971:EOS

[WHM71] M. Wells, D. Holdsworth, and A. P. McCann. The Eldon 2 operating system for KDF9. The Computer Journal, 14(1):21–24, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/1/21.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_01/140021.sgm.abs.html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/21.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_01/tiff/22.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_14/Issue_01/tiff/23.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/24.tif.

Wichmann:1972:FAC

[Wic72] B. A. Wichmann. Five Algol compilers. The Computer Journal, 15(1):8–12, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/15/1/8.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_01/
150008.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/10.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/tiff/11.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_15/Issue_
01/tiff/12.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_01/tiff/8.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_15/Issue_01/tiff/9.tif.

Wichmann:1978:BR

[Wic78] B. A. Wichmann. Book review. The Computer Journal, 21(3):214, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/3/214.2.full.pdf+html.

Wiederhold:1979:TSS

[Wie79] J. F. A. Wiederhold. Tuning to service standards without monitors. The Computer Journal, 22(1):28–36, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
22/1/28.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/28.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/29.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/30.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/31.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/32.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/33.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/35.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/36.tif.

Wight:1975:EAP

[Wig75] A. S. Wight. The EMAS archiving program. The Computer Journal, 18(2):131–134, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/18/2/131.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/131.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/132.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/133.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/134.tif.

Wilkinson:1970:BRB

[Wil70] J. H. Wilkinson. Book review: A Collection of Matrices for Testing Computational Algorithms, by Robert T. Gregory, D. L. Karney, 1969; 154 pages. (John Wiley and Sons Ltd., £4.75). The Computer Journal, 13(4):391, November 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/4/387.
full.pdf;   http://comjnl.oxfordjournals.org/content/13/
4/391.full.pdf+html.

Wilkes:1972:PID

[Wil72] M. V. Wilkes. On preserving the integrity of data bases. The Computer Journal, 15(3):191–194, August 1972. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/191.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150191.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/191.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/192.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/193.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/194.
tif.

Wilkes:1973:DP

[Wil73] M. V. Wilkes. The dynamics of paging. The Computer Journal, 16(1):4–9, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/1/4.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/
160004.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/4.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/tiff/5.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/tiff/6.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_01/tiff/7.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_01/tiff/8.tif;
http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/tiff/9.tif.

Wilkes:1974:ACP

[Wil74] M. V. Wilkes. The art of computer programming, volume 3, sorting and searching. The Computer Journal, 17(4):324, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
4/324.full.pdf+html.

Wilbur:1975:BR

[Wil75] S. R. Wilbur. Book review. The Computer Journal, 18(1):54, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
1/54.full.pdf+html.

Wilbur:1976:LPI

[Wil76a] S. R. Wilbur. Lang-Pak — an interactive language design system. The Computer Journal, 19(2):150, May 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/2/150.full.pdf+html.

Willers:1976:SDG

[Wil76b] I. M. Willers. The syntax directed graph algorithm for the input of equations to the Taylor series system for solving ordinary differential equations. The Computer Journal, 19(4):344–347, November 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/19/4/344.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/344.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_19/Issue_04/tiff/345.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/346.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/347.tif.

Williams:1976:BRa

[Wil76c] I. R. Williams. Book review. The Computer Journal, 19(4):314, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/314.2.full.pdf+html.

Williams:1976:BRb

[Wil76d] I. R. Williams. Book review. The Computer Journal, 19(4):314, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/4/314.3.full.pdf+html.

Williams:1976:DE

[Wil76e] M. R. Williams. The Difference Engines. The Computer Journal, 19(1):82–89, February 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/19/1/82.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/82.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/83.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_01/tiff/84.tif;   http:

//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/85.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/86.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_19/Issue_01/tiff/87.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_01/
tiff/88.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_19/Issue_01/tiff/89.tif.

Wilkie:1977:TLE

[Wil77a] A. D. Wilkie. Table look-up (with examples in COBOL). The Computer Journal, 20(3):202–206, August 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/202.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/202.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/203.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/204.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/205.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/206.tif.

Williams:1977:GSF

[Wil77b] M. H. Williams. Generating structured flow diagrams: the nature of unstructuredness. The Computer Journal, 20(1):45–50, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/1/45.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/45.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/46.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/47.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/48.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/49.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
20/Issue_01/tiff/50.tif.

Williams:1977:C

[Wil77c] M. R. Williams. Correspondence. The Computer Journal, 20(4):380, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
20/4/380.1.full.pdf+html.

Willis:1977:RTH

[Wil77d] P. J. Willis. A real time hidden surface technique. The Computer Journal, 20(4):335–339, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/4/335.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/335.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/336.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/337.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/338.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/339.tif.

Wilson:1977:BSI

[Wil77e] J. M. Wilson. Boolean simplification and integer inequalities. The Computer Journal, 20(4):356–358, November 1977. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/4/356.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_04/tiff/356.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_04/tiff/357.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
04/tiff/358.tif.

Wilkes:1978:BR

[Wil78a] M. V. Wilkes. Book review. The Computer Journal, 21(3):245, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/245.2.full.pdf+html.

Williams:1978:CAH

[Wil78b] I. R. Williams. Computational analysis with the HP-25 pocket calculator. The Computer Journal, 21(2):148, May 1978. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/21/2/148.full.pdf+html.

Williams:1978:SSF

[Wil78c] M. H. Williams. Static semantic features of Algol60 and Basic. The Computer Journal, 21(3):234–242, August 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/3/234.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/

hdb/Volume_21/Issue_03/tiff/234.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/235.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/236.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/237.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/238.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/239.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_03/tiff/240.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_03/tiff/241.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
03/tiff/242.tif.

Williams:1978:BDH

[Wil78d] P. W. Williams. Balancing demand in a hierarchical structure. The Computer Journal, 21(4):341–346, November 1978. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/21/4/341.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/341.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/342.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/343.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_21/Issue_04/tiff/344.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_21/Issue_04/tiff/345.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_
04/tiff/346.tif.

Williams:1978:CCS

[Wil78e] P. W. Williams. Criteria for choosing subsets to obtain maximum relative entropy. The Computer Journal, 21(1):57–62, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/57.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/57.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/58.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/59.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/60.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/61.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/62.tif.

Williams:1978:RBE

[Wil78f] P. W. Williams. The relationship between entropy and file search times. The Computer Journal, 21(1):63–65, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/63.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/63.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/64.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/65.tif.

Williams:1979:CEO

[Wil79] M. H. Williams. Conditions for extended operator precedence parsing. The Computer Journal, 22(2):164–168, May 1979. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/22/2/164.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/164.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/165.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_
02/tiff/166.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_22/Issue_02/tiff/167.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_22/Issue_02/tiff/168.tif.

Winn:1975:DAE

[Win75] G. C. E. Winn. A digital approach to the efficient synthesis of threshold gates. The Computer Journal, 18(3):239–242, August 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/3/239.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/239.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_03/tiff/240.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
03/tiff/241.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_03/tiff/242.tif.

Wise:1974:BR

[Wis74] A. H. Wise. Book review. The Computer Journal, 17(3):244, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
3/244.2.full.pdf+html.

Wiseman:1975:BR

[Wis75] N. E. Wiseman. Book review. The Computer Journal, 18(4):359, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/359.full.pdf+html.

Wiseman:1978:BR

[Wis78] N. E. Wiseman. Book review. The Computer Journal, 21(1):62, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/62.full.pdf+html.

Witts:1979:BRa

[Wit79a] R. C. Witts. Book review. The Computer Journal, 22(1):49, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/49.2.full.pdf+html.

Witts:1979:BRb

[Wit79b] R. C. Witts. Book review. The Computer Journal, 22(3):205, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
3/205.2.full.pdf+html.

Wagle:1974:CPM

[WJ74] B. Wagle and P. M. Jenkins. Corporate planning, models and computer systems. The Computer Journal, 17(3):194–200, August 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/3/194.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/194.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/195.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/196.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/197.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_03/tiff/198.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
03/tiff/199.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_03/tiff/200.tif.

Williams:1977:DRU

[WK77] P. W. Williams and M. T. Khallaghi. Document retrieval using a substring index. The Computer Journal, 20(3):257–262, August 1977.

CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/20/3/257.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/257.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/258.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/259.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_20/Issue_03/tiff/260.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_20/Issue_03/tiff/261.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_
03/tiff/262.tif.

Willcox:1972:ACD

[WL72] W. R. Willcox and S. P. Lapage. Automatic construction of diagnostic tables. The Computer Journal, 15(3):263–267, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/15/3/263.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_03/150263.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/263.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/264.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_15/Issue_03/tiff/265.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_15/Issue_03/tiff/266.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
15/Issue_03/tiff/267.tif.

Williams:1971:CCC

[WLDC71] W. T. Williams, G. N. Lance, M. B. Dale, and H. T. Clifford. Controversy concerning the criteria for taxonometric strategies. The Computer Journal, 14(2):162–165, May 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/2/162.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/140162.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/162.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_02/tiff/163.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_02/tiff/164.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/tiff/165.
tif.

Wheeler:1973:C

[WN73] D. Wheeler and R. Needham. Correspondence. The Computer Journal, 16(1):18, ???? 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/1/18.2.full.pdf+html.

Wheeler:1973:CWU

[WNT73] D. Wheeler, R. Needham, and The Editor. Correspondence: On word use. The Computer Journal, 16(1):18, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_01/tiff/18.tif. See correspondence [FT73].

Williams:1975:CMI

[WO75] M. H. Williams and H. L. Ossher. A computer model for instructional purposes. The Computer Journal, 18(4):333–341, November 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/4/333.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/333.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/334.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/335.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/336.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/337.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/338.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_04/tiff/339.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_04/tiff/340.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_
04/tiff/341.tif.

Williams:1978:CUF

[WO78] M. H. Williams and H. L. Ossher. Conversion of unstructured flow diagrams to structured form. The Computer Journal, 21(2): 161–167, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/21/2/161.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/161.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/162.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/163.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/164.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_21/Issue_02/
tiff/165.tif;   http://www3.oup.co.uk/computer_journal/hdb/

Volume_21/Issue_02/tiff/166.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_21/Issue_02/tiff/167.tif.

Wodon:1978:CNT

[Wod78] P. Wodon. Correspondence: On “A note on the Towers of Hanoi problem”. The Computer Journal, 21(2):191, May 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_21/Issue_02/tiff/191.tif. See [Hay77].

Wolfe:1971:NSI

[Wol71] M. A. Wolfe. The numerical solution of implicit first order ordinary differential equations with initial conditions. The Computer Journal, 14(2):173–178, May 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/2/173.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_02/
140173.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/173.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/174.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/175.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_02/tiff/176.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_02/tiff/177.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
02/tiff/178.tif.

Wolff:1978:RNL

[Wol78] J. G. Wolff. Recoding of natural language for economy of transmission or storage. The Computer Journal, 21(1):42–44, February 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
21/1/42.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_21/Issue_01/tiff/42.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_21/Issue_01/tiff/43.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
21/Issue_01/tiff/44.tif.

Wood:1970:GNF

[Woo70a] D. Wood. A generalised normal form theorem for context-free grammars. The Computer Journal, 13(3):272–277, August 1970. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/13/3/272.

full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_03/130272.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/272.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/273.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/274.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/tiff/275.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_03/tiff/276.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_03/tiff/277.tif.

Wood:1970:TLF

[Woo70b] D. Wood. The theory of left factored languages: Part 2. The Computer Journal, 13(1):55–62, February 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/13/1/55.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/130055.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/55.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/56.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/57.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/58.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/59.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_01/tiff/60.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_01/tiff/61.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_13/Issue_01/tiff/62.
tif.

Woodall:1970:AIS

[Woo70c] A. D. Woodall. Algorithm 45: An internal sorting procedure using a two-way merge. The Computer Journal, 13(1): 110–111, February 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_01/tiff/110.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
01/tiff/111.tif.

Woodall:1970:AIM

[Woo70d] A. D. Woodall. Algorithm 55: An internal merge sort giving ranks of items. The Computer Journal, 13(4):424–425, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://www3.oup.co.uk/computer_journal/hdb/Volume_
13/Issue_04/tiff/424.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_13/Issue_04/tiff/425.tif.

Woodall:1970:ADC

[Woo70e] A. D. Woodall. Algorithm 56: To disentangle a chain. The Computer Journal, 13(4):425, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_13/Issue_04/tiff/425.tif.

Woodall:1970:NAL

[Woo70f] A. D. Woodall. Note on Algorithm 43: a listed radix sort. The Computer Journal, 13(3):326, August 1970. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_03/
tiff/326.tif.

Woodall:1970:NAS

[Woo70g] A. D. Woodall. Note on Algorithms 25, 26. 25: Sort a section of the elements of an array by determining the rank of each element. 26: Order the subscripts of an array section according to the magnitudes of the elements. The Computer Journal, 13(3):326, August 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_13/Issue_03/tiff/326.tif. See [Boo67a, Boo67b, Hil69].

Wood:1971:FNT

[Woo71a] D. Wood. A further note on top-down deterministic languages. The Computer Journal, 14(4):396–403, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/396.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140396.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/396.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/397.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/398.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/399.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/400.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/401.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/402.

tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/403.tif.

Woodall:1971:ART

[Woo71b] A. D. Woodall. Algorithm 63: a recursive tree sort. The Computer Journal, 14(1):103, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_01/tiff/103.tif.

Woodall:1971:ANR

[Woo71c] A. D. Woodall. Algorithm 64: a non-recursive tree sort. The Computer Journal, 14(1):104, February 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_01/
tiff/104.tif.

Wood:1974:BR

[Woo74a] P. A. Wood. Book review. The Computer Journal, 17(1):74, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/74.2.full.pdf+html.

Woods:1974:BR

[Woo74b] J. Woods. Book review. The Computer Journal, 17(2):142, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/2/142.2.
full.pdf+html.

Woodward:1976:BR

[Woo76] P. M. Woodward. Book review. The Computer Journal, 19(3):253, ???? 1976. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
19/3/253.2.full.pdf+html.

Wood:1977:TVP

[Woo77a] D. Wood. Two variations on N-parallel finite state generators. The Computer Journal, 20(1):63–67, February 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic).  URL http://comjnl.oxfordjournals.org/content/
20/1/63.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/63.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/64.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

20/Issue_01/tiff/65.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_20/Issue_01/tiff/66.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/tiff/67.
tif.

Woodall:1977:GPS

[Woo77b] A. D. Woodall. Generation of permutation sequences. The Computer Journal, 20(4):346–349, November 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/20/4/346.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/346.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_04/tiff/347.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_04/tiff/348.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_04/
tiff/349.tif.

Wood:1978:BRa

[Woo78a] B. Wood. Book review. The Computer Journal, 21(1):34, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/34.2.full.pdf+html.

Wood:1978:BRb

[Woo78b] B. M. Wood. Book review. The Computer Journal, 21(3):285, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
3/285.4.full.pdf+html.

White:1975:SLT

[WP75] J. R. White and L. Presser. A structured language for translator construction. The Computer Journal, 18(1):34–42, February 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/1/34.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/34.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/35.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/36.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_18/Issue_01/tiff/37.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/38.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/39.tif;   http://www3.oup.co.uk/computer_

journal/hdb/Volume_18/Issue_01/tiff/40.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_18/Issue_01/tiff/41.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
18/Issue_01/tiff/42.tif.

Williams:1972:CLA

[WR72] M. R. Williams and B. Randell. Correspondence: On “Ludgate’s analytical machine of 1909’. The Computer Journal, 15(2):131, May 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://www3.oup.co.uk/computer_journal/
hdb/Volume_15/Issue_02/tiff/131.tif. See [Ran71b]. From The

Times (London, Saturday, 21 October, 1871), under the heading ‘Deaths’: ‘On the 18th Oct, at 1, Dorset-street Manchester Square, Charles Babbage, Esq. in the 80th year of his age.’.

Wren:1971:BR

[Wre71a] Anthony Wren. Book review. The Computer Journal, 14(4):377, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
14/4/377.full.pdf+html.

Wren:1971:BRB

[Wre71b] Anthony Wren. Book review: Case Exercises in Operations Research, by J. J. C. Martin and R. A. Denison, 1971. The Computer Journal, 14(4):377, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_04/tiff/377.tif.

Wren:1972:PPS

[Wre72] Anthony Wren. Proceedings of the Princeton Symposium on Mathematical Programming. The Computer Journal, 15(1):87, February 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/15/
1/87.full.pdf+html.

Wright:1974:ASE

[Wri74] W. E. Wright. An axiomatic specification of Euclidean analysis. The Computer Journal, 17(4):355–364, November 1974. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/17/4/355.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/355.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/356.tif;

 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/357.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/358.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/359.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/360.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/361.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_17/Issue_04/tiff/362.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_
04/tiff/363.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_17/Issue_04/tiff/364.tif.

Walsh:1973:BRB

[WS73] Joan Walsh and P. A. Samet. Book reviews: Computer Applications of Numerical Methods, by Shan S. Kuo, 1972; Introduction to Computational Methods for Students of Calculus, by S. S. McNeary, 1972. The Computer Journal, 16(3):261, August 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_16/Issue_03/tiff/261.tif.

Whitby-Strevens:1974:CMS

[WS74] C. Whitby-Strevens. Correspondence: Microprogramming and systems architecture. The Computer Journal, 17(4):381, November 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/381.tif. See [Bro74a].

Walker:1979:AAM

[WT79] A. N. Walker and The Editor. Algorithms: On “Algorithm 1: Mintree”. The Computer Journal, 22(3):288, August 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_03/tiff/288.tif. See [Obr78a].

Waliuzzaman:1970:DMV

[WV70] K. M. Waliuzzaman and Z. G. Vranesic. On decomposition of multi-valued switching functions. The Computer Journal, 13(4):359– 362, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/13/4/359.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_13/Issue_04/130359.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_

04/tiff/359.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_13/Issue_04/tiff/360.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_13/Issue_04/tiff/361.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/362.tif.

Whitfield:1973:EEM

[WW73] H. Whitfield and A. S. Wight. EMAS — the Edinburgh Multi-Access System. The Computer Journal, 16(4):331–346, November 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/16/4/331.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_04/160331.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/331.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/332.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/333.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/334.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/335.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/336.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/337.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/338.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/339.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/340.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/341.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/342.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/343.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_04/tiff/344.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_04/tiff/345.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_16/Issue_04/tiff/346.
tif.

Walker:1976:LBB

[WW76] A. Walker and D. Wood. Locally balanced binary trees. The Computer Journal, 19(4):322–325, November 1976. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/19/4/322.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/322.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_19/Issue_04/tiff/323.tif;   http://www3.oup.co.

uk/computer_journal/hdb/Volume_19/Issue_04/tiff/324.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_19/Issue_
04/tiff/325.tif.

Williams:1979:CAA

[WW79] P. W. Williams and D. Woodhead. Computer assisted analysis of cryptic crosswords. The Computer Journal, 22(1):67–70, February 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
22/1/67.full.pdf+html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/67.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_22/Issue_01/tiff/68.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
22/Issue_01/tiff/69.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_22/Issue_01/tiff/70.tif.

Watt:1971:BRB

[WY71] J. M. Watt and J. Z. Young. Book reviews: Computation and Theory in Ordinary Differential Equations, by James W. Daniel and Ramon E. Moore, 1970; Progress in Brain Research, Vol. 33. Computers and Brains, by J. P. Schade and J. Smith, 1970. The Computer Journal, 14(4):412, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_14/Issue_04/tiff/412.tif.

Wyman:1973:BPD

[Wym73] F. P. Wyman. Binary programming: a decision rule for selecting optimal vs heuristic techniques. The Computer Journal, 16(2): 135–140, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/16/2/135.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_16/Issue_02/160135.sgm.abs.html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/135.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/136.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/137.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/138.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/139.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/140.tif.

See correspondence [PW74b].

Wynn:1971:NGE

[Wyn71] P. Wynn. A note on the generalised Euler transformation. The Computer Journal, 14(4):437–441, November 1971. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/14/4/437.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/140437.sgm.abs.html;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/437.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
14/Issue_04/tiff/438.tif;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_14/Issue_04/tiff/439.tif;   http://www3.
oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/tiff/440.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_

14/Issue_04/tiff/441.tif. See correspondence [Scr72].

Wyvill:1977:BR

[Wyv77] G. Wyvill. Book review. The Computer Journal, 20(4):358, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
4/358.2.full.pdf+html.

Young:1971:BRC

[YFM71] A. Young, L. Fox, and D. F. Mayers. Book reviews: Convexity and Optimization in Finite Dimensions I, by Josef Stoer and Christoph Witzgall, 1970; On Round-off Errors in Linear Programming, by H. Müller-Merbach, 1970; Approximation Theory, edited by A. Talbot, 1970. The Computer Journal, 14(4):390, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://www3.oup.co.uk/computer_journal/hdb/
Volume_14/Issue_04/tiff/390.tif.

Young:1970:BR

[You70] A. Young. Book review. The Computer Journal, 13(1):13, ???? 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/13/
1/13.full.pdf+html.

Young:1971:BR

[You71] J. Z. Young. Book review. The Computer Journal, 14(4):412, ???? 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/14/
4/412.2.full.pdf+html.

Young:1972:BR

[You72] A. Young. Book review. The Computer Journal, 15(3):228, August 1972. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
15/3/228.full.pdf+html.

Young:1974:BR

[You74] A. Young. Book review. The Computer Journal, 17(1):27, ???? 1974. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/17/
1/27.2.full.pdf+html.

Young:1975:BRa

[You75a] A. Young. Book review. The Computer Journal, 18(3):205, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/205.1.full.pdf+html.

Young:1975:BRb

[You75b] A. Young. Book review. The Computer Journal, 18(3):222, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/18/
3/222.2.full.pdf+html.

Young:1975:C

[You75c] K. J. Young. Correspondence. The Computer Journal, 18(4):381, ???? 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067

(electronic). URL http://comjnl.oxfordjournals.org/content/
18/4/381.2.full.pdf+html.

Young:1977:BRa

[You77a] A. Young. Book review. The Computer Journal, 20(1):44, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/44.1.full.pdf+html.

Young:1977:BRb

[You77b] A. Young. Book review. The Computer Journal, 20(1):55, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/55.1.full.pdf+html.

Young:1977:BRc

[You77c] A. Young. Book review. The Computer Journal, 20(1):77, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/20/
1/77.3.full.pdf+html.

Young:1978:BR

[You78] A. Young. Book review. The Computer Journal, 21(1):14, ???? 1978. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/21/
1/14.1.full.pdf+html.

Young:1979:BRa

[You79a] A. Young. Book review. The Computer Journal, 22(1):52, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
1/52.full.pdf+html.

Young:1979:BRb

[You79b] A. Young. Book review. The Computer Journal, 22(4):362, ???? 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

tronic). URL http://comjnl.oxfordjournals.org/content/22/
4/362.1.full.pdf+html.

Yau:1979:GST

[YT79] S. S. Yau and Y. S. Tang. Generation of shortest test sequences for detecting individual faults of sequential circuits. The Computer Journal, 22(2):169–172, May 1979. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/22/2/169.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/169.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_22/Issue_02/tiff/170.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_22/Issue_02/tiff/171.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_22/Issue_02/
tiff/172.tif.

Zafarullah:1971:SSI

[Zaf71] A. Zafarullah. Some stable implicit difference methods for heat equation with derivative boundary condition. The Computer Journal, 14(3):309–311, August 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/3/309.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_03/
140309.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_03/tiff/309.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_03/tiff/310.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
03/tiff/311.tif.

Zakrajsek:1977:AGE

[Zak77] E. Zakrajsek. Algorithm 93: Generalised eigenvalue problem. The Computer Journal, 20(1):86–91, February 1977. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/
tiff/86.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/87.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_01/tiff/88.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_20/Issue_01/
tiff/89.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_20/Issue_01/tiff/90.tif;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_20/Issue_01/tiff/91.tif.

Zambardino:1970:ADP

[Zam70] R. A. Zambardino. Algorithm 53: Decomposition of positive definite symmetric band matrices. The Computer Journal, 13 (4):421–422, November 1970. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.

uk/computer_journal/hdb/Volume_13/Issue_04/tiff/421.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_13/Issue_
04/tiff/422.tif.

Zambardino:1974:ASS

[Zam74] R. A. Zambardino. Algorithm 82: Solutions of systems of linear equations with partial pivoting and reduced storage requirements. The Computer Journal, 17(4):377–378, November 1974. CODEN CM-PJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http:/

/www3.oup.co.uk/computer_journal/hdb/Volume_17/Issue_04/
tiff/377.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_17/Issue_04/tiff/378.tif. The algorithm number was

omitted in the original journal issue.

Zaremba:1977:C

[Zar77] W. A. Zaremba. Correspondence. The Computer Journal, 20(4): 380–381, ???? 1977. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/20/4/380.2.full.pdf+html.

Zissos:1971:NNO

[ZD71] D. Zissos and F. G. Duncan. NOR and NAND operators in Boolean algebra applied to switching circuit design. The Computer Journal, 14(4):413–417, November 1971. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/14/4/413.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_04/
140413.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/413.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/414.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_14/Issue_
04/tiff/415.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_14/Issue_04/tiff/416.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_14/Issue_04/tiff/417.tif.

Zissos:1973:BM

[ZD73] D. Zissos and F. G. Duncan. Boolean minimisation. The Computer Journal, 16(2):174–179, May 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/16/2/174.full.pdf+html;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_02/
160174.sgm.abs.html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/174.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/175.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/176.tif;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_16/Issue_02/tiff/177.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_16/Issue_02/tiff/178.tif;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
02/tiff/179.tif.

Zakian:1973:NIL

[ZL73] V. Zakian and R. K. Littlewood. Numerical inversion of Laplace transforms by weighted least-squares approximation. The Computer Journal, 16(1):66–68, February 1973. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://

comjnl.oxfordjournals.org/content/16/1/66.full.pdf+html;
 http://www3.oup.co.uk/computer_journal/hdb/Volume_16/Issue_
01/160066.sgm.abs.html;   http://www3.oup.co.uk/computer_
journal/hdb/Volume_16/Issue_01/tiff/66.tif;   http://www3.

oup.co.uk/computer_journal/hdb/Volume_16/Issue_01/tiff/67.
tif;   http://www3.oup.co.uk/computer_journal/hdb/Volume_
16/Issue_01/tiff/68.tif.

Zmud:1975:ETE

[Zmu75a] R. W. Zmud. Erratum: “Two Enhancements to a Flexible Pricing Control System for Allocating Computer Resources”. The Computer Journal, 18(3):287, August 1975. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://www3.oup.co.uk/

computer_journal/hdb/Volume_18/Issue_03/tiff/287.tif. See

[Zmu75b].

Zmud:1975:TEF

[Zmu75b] R. W. Zmud. Two enhancements to a flexible pricing control system for allocating computer resources. The Computer Journal, 18 (2):112–114, May 1975. CODEN CMPJA6. ISSN 0010-4620 (print),

1460-2067 (electronic). URL http://comjnl.oxfordjournals.org/
content/18/2/112.full.pdf+html;   http://www3.oup.co.uk/
computer_journal/hdb/Volume_18/Issue_02/tiff/112.tif;   http:
//www3.oup.co.uk/computer_journal/hdb/Volume_18/Issue_02/
tiff/113.tif;   http://www3.oup.co.uk/computer_journal/hdb/
Volume_18/Issue_02/tiff/114.tif. See erratum [Zmu75a].

Zvegintzov:1975:PAD

[Zve75] N. Zvegintzov. Problems with analytic description of text formats. The Computer Journal, 18(2):148–149, May 1975. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

URL http://comjnl.oxfordjournals.org/content/18/2/148.
full.pdf+html;   http://www3.oup.co.uk/computer_journal/
hdb/Volume_18/Issue_02/tiff/148.tif;   http://www3.oup.co.
uk/computer_journal/hdb/Volume_18/Issue_02/tiff/149.tif.